A Complete Bibliography of Publications in the International Journal of Parallel Programming

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

05 August 2024 Version 2.43

Title word cross-reference

∗ [CS16]. + [HVF18, SBC17]. 0 [LS92]. 1 [LS92]. 2 [CTB14, ES11, IBA11]. 3 [BC15, HPVRPF15, HF14a, HF14b, JGM15, LLGC17, LHP^{+}17, SJKA99, fSxWC18, SBC17]. < [JS06a]. > [JS06a]. ^{(}R) [BKT08, SM09]. ^{T M} [BKT08]. i [TRD21]. K [LKS^{+}20, VVCA23]. kd [WR18]. l_{1} [GLLH17]. m [DPL86].

-D [ES11]. -Means [LKS^{+}20]. -Minimization [GLLH17]. -Tree [WR18]. -way [DPL86].

1000 [SSMO96]. 14 [HG18]. 16 [Swa88]. 18th [DB08].

2 [ELK18]. 2.0 [RSJ^{+}19]. 20 [TTF22]. 2000 [IPR^{+}05]. 2011 [MCE13]. 2013 [AG15]. 2014 [Bro15]. 2015 [PVG17]. 2019 [Kes20]. 2020 [RJO22]. 2022 [ORJ24]. 2DT [BARSW95]. 2DT-FP [BARSW95].

3 [EAK21]. 3.0 [KaM10, OP10].

512 [RSJ^{+}19].

95 [KaM10].

A. [Swa88]. Abingdon [AM95]. Above [LCT^{+}20]. abstraction [VR88]. Abstractions [ASS21, BCL14, LQWP10, YAI95]. Abstractive [KSF^{+}18]. Accelerated [BETK24, DMC^{+}18, HML^{+}20, KLK16, PES^{+}18, SF20, SBC17]. Accelerating [FFS18, FRT^{+}18, FJZ^{+}15, HF14a, HF14b, LLGC17, MAWD^{+}16, MNN22, PTdSF^{+}12, SCS23, WQZ^{+}24, XWH21, ZHF^{+}19, ZTY^{+}19]. Acceleration [BC10, MCFM12, STM15]. Accelerator [ALPS19, EK17, FVvL^{+}16, LWGZ18, SWG^{+}18, TTF22, YZZ^{+}19]. Accelerators [GP17, LCF21, SdLC21, ZJL22]. Access [JG97, Joh94, LMHW18, OOR13, RWMF24, ZK07]. Accesses [GV95, LPB13]. Accumulations [MM16]. Accumulative [IH04]. Accuracy [CEP97, KP04]. Accurate [PZL^{+}19, RGB^{+}08, TA99]. Accurately [BGdS09, Low00]. Achieving [AMP^{+}05, GAR^{+}16, GS90, Won02]. Acknowledgment [Nic14]. ACOTES [MAB^{+}11]. Action [WZG^{+}17]. Active [RLK20]. Activity [FR95]. Actor [ZLC^{+}19]. Actors [RTD20]. ACTS [DGMP09]. Acyclic [Hue97, ZLJA12]. Adaptability [SA10]. Adaptation [CCL12, SJT13]. Adapting [EFED05, JMSG02, PIP18]. Adaption [CLL21]. Adaptive [AO19, BBB^{+}17, BFRPVR^{+}15, CS20, DJS12, GLLH17, GRV^{+}17, GW19, GH89, HHW10, HP13, HR11, HTDL18, JLDF19, KSEG14, LLM^{+}12, LJE05, PSM97, RA09, SHC15, WR18, ZLD15]. AdaptiveLock [YLB19]. Address [SS01, TAY^{+}12, HR11]. Addressing [GG13]. Adjoining [PS92, PW92]. Adjustment [ZLC^{+}19]. ADL [PC13]. ADL-Based [PC13]. Admission [NYHA14]. Adoption [SdLC21]. Advance [SL14]. Advanced [AAN^{+}20, DLRS13, MAB^{+}11, LF15, NdMMW16]. Advantage [TKN^{+}08]. AES [XWH21]. Affine [Fea92b, KDV22, KP95, LM00, Mon97, Fea92a]. Affinity [GRC^{+}14]. After [AKBPV19]. Again [MP04]. Against [DDJ^{+}18, FDY^{+}19, GWHY19, MWES19]. Age [DKB^{+}09]. Agent [FLMR17a, FLMR17b, WLL^{+}08, STB^{+}18]. Agent-Based [FLMR17a, FLMR17b].

Agents [ES06]. Aggregation

[HHW20, LSA^{+}07, SBN03]. Aggressive [SK14]. Agnostic [AVM^{+}16, NAP02]. Agriculture [GKC22]. AI [GAG22]. AI-Assisted [GAG22]. Air [LCT^{+}20]. Aircraft [MSJ20]. ALE [HAA^{+}11]. Algebra [CCG^{+}14, CBR17, HKJ^{+}18, KTRZ^{+}17, MP04]. algebraic [SS89]. Algorithm [AFO^{+}08, AKT^{+}14, ASG20, BM09, CSCL20, CL96, Cra88, CDDM18, DMMS91, DWS16, DZW10, EKU22, FBV21, GF14, GKC22, HNC^{+}16, IP90, IKN00, JLDF19, JCW^{+}18, KBD03, LLM^{+}12, LMP98, LF15, LKS^{+}20, MMN15, MSJ20, MCT^{+}18, Mer86, MB12b, Moh19, MVD^{+}14, NFC^{+}09, NB15, NRR99, NRGB17, NdMMW16, PS92, RY20, RY22, RK92, RBR22, SI11, SF20, SWL05, Spr92, XZT20, ZQT20, ZTY^{+}19, EG86, FcF87, GZ87, GT86, Hua89, JGA^{+}88, LS92, Sch92, SRV88]. Algorithm-Based [NRR99]. Algorithmic [DMK21, DM17, EK17, GK18, HdMMK22, HK23, dMMHdLN21, SHK13, WE18, WMK19, WK20]. Algorithms [AT91, APR^{+}18, AMAH01, AK17, ABSSS19, AGT17, BR14b, CAT18, CAP88, Dam07, DPS90, DMC^{+}20, EO88, FG16, GM20, Ged13, GP17, GF14, HSXH19, IBA11, Iqb91, uHKAMFM16a, uHKAMFM16b, KPS14, LTSD15, Liv91, dMMKdLN22, ÖO07, PIP18, RG15, SH87, SS92, SAS18, SKAT91, SJC18, SR90, VVCA23, XWH21, Zey05, ZLD15, ZHF^{+}19, DPL86, ECSS88, HFM88, SDJS98, Swa88, Zha89]. Alias [LC11, WGW04]. Aligned [Dab21]. Alignment [FJZ^{+}15]. All-Parses [IP90]. All-Port [IBA11]. All-to-All [FPY08a]. Allgather [QA11]. Allocation [BE14, CND95, DLX^{+}17, DS20, LkCH94, NG92, PZL^{+}19, RY20, RY22, ZLD15, EO88, NP98]. Alloyed [LLSS03]. Allpairs [SFAG14]. Alltoall [QA11]. Alone [DJR16]. Alternate [PK20]. Alternative [KF99, FcF87]. Alternatives [Bel94, MB99, NPT86]. Altruism [LCL19].

Altruism-Based [LCL19]. AMAIX [ZJL22]. Amdahl [Ano87a, PM07]. AMR [NLRH07]. Analyses [CI96, GV95, SJW22]. Analysing [BDD^{+}18]. Analysis [ASS24, AK96, ABTZ00, AFO^{+}08, AW98, BEA^{+}19, BG96, BFRPVR^{+}15, CSC^{+}00, CSD21, CAZ02, CAT18, CPL^{+}10, Fea91, Gha19, GH96, HML^{+}20, Jak19, KP04, LT17, LCL19, LY95, LHF^{+}15, LWDL17, LHP^{+}17, LC11, MP91, MHL95, MP04, NSU22, NP19, PCJ20, PPEP08, RLEJ19, RRH03, Sca11, SSP^{+}96, SO89, US05, WGW04, dMP^{+}03, AD86, GTK^{+}88, NPD89, RS90, KR87]. Analytical [KWA^{+}10, NP19, VCP^{+}13, ZJG17, ZJL22]. Analytics [FJA^{+}18, FRT^{+}18, LWF^{+}19, TMXS24]. Analyze [ASW^{+}15, Dem11]. Analyzing [APR^{+}18, ALG^{+}95, DF98, FM09, HRH08, SD11]. AND-parallel [SRV88]. AND-Parallelism [SH96, BS89]. AND/OR [RK92]. Android [AER^{+}17]. Animation [BGMR11]. Announcement [Int98, Ano86d, Ano92]. Anomalies [Jan15]. ANSI [BG03]. Ant [ASG20, dMMHdLN21]. Anti [CDRV98]. Anti- [CDRV98]. Apache [ZLA21]. APCFS [KK11]. API [LCT^{+}20, TTF22]. APL [GS90]. app [DJR16]. Application [ACC^{+}02, API03, BGdS09, BS07, CZTM03, Dam07, EGK23, FJO^{+}16, HL21, HTDL18, JCH^{+}08, JAW17, KS97, Mat17, MP04, Moh19, PG07, PB04, RSK09, Sek09, SKG09, TOM^{+}11, VMS15, BH87, CRM92, WB87]. Application-Aware [JAW17]. Application-Dependent [VMS15]. Application-Level [HTDL18]. Application-Specific [API03, TOM^{+}11]. Applications [Ano16a, Ano18b, BEA^{+}19, BEJD21, BBR11b, BDD^{+}18, CY14, CR19, CBR17, CHCL14, CPT14, DPT17, DFH17, DS16, DGMP09, EWHS11, FM09, GHM14, GS11, GS13, GRC^{+}14, GGV17, Gre16, HK14, HMK09, HtBK^{+}10, HLK^{+}09, IPR^{+}05, KMjC02, KPRS96, KTBP18, LRG14, LLW^{+}17, LQWP10, LWLG11, MV17, Mar09, MAJD16, MG15, MCWK01, MANR09, Mis09, OK99, ÖA21, PPQV16, RLPN^{+}02, RSJ^{+}14, RGB^{+}08, SR15, SÜCV17, SSB^{+}17, SASH12, SBN03, TG21, TB23, TMHT96, WL16, WLL17, ZK07, ZZS^{+}19, ZD19, ZSH^{+}12, GKMB87, SDJS98, SS89]. applicative [Hun87]. Applied [BUMS02, KaM10, Lin91a]. Approach [AK90b, AVM^{+}16, BBB^{+}17, CHB06, DM17, FCZ16, FJA^{+}18, FBV21, FJO^{+}16, GAG22, GYL92, JQWG15, KK20, KSF^{+}18, KKKS24, KSA^{+}18, LTF^{+}12, LLL^{+}15, LCT^{+}20, MO91, NN95, OATGEL15a, PMV17, QZP15, STM15, VSDK09, qWlJzKhC17, WS08, WEJS94]. Approaches [BUMS02, JCH^{+}08, PCJ18, VRGC19]. Appropriate [Gen16]. Approximate [HZL16, Iqb91, TGT18, VCP^{+}16]. Arbitration [BS91]. ArchC [ARB^{+}05]. Architectural [LSHK09, NP01, SEP08, TCUV14, WGF^{+}16]. Architecture [AP86, ARB^{+}05, BGGT02, CHCL14, CFC^{+}19, CDC09, DB08, DLRS13, FCJV99, GWPV21, GL92, HTZ^{+}97, HL21, JLDS16, LHP^{+}17, MB12a, MB99, MSPR18, NdMMW16, NAP02, RD08, STF^{+}12, SJT13, TRD21, YS22, ZTY^{+}19, CB86, GKMB87]. Architecture-Agnostic [NAP02]. Architectures [Ano18b, Ano18a, BG96, BFG^{+}10, CPG01, CND95, CJA00, GBPK07, Ged13, GAG22, GGV17, HCEP98, HP13, LAD15, MCE13, MGJS15, Mis09, NFC^{+}09, NdMCdMMW16, PJS^{+}05, PMM^{+}18, PG16, PVG17, RJO22, RMH21, RSJ^{+}19, SJBV06, SMH21, TG21, TB23, TJY99, TF94, VHK^{+}18, ZLAV04, ZZS^{+}19, LRG^{+}91]. Area [RSP20, Roy10, SWZ^{+}15, WMN^{+}17]. Argument [ABASS12, NG92]. Argument-Fetching [NG92]. Arithmetic [ABASS12]. ARM [MGL^{+}17]. ARMv8 [CFC^{+}19, KHT21]. ARMv8-based [CFC^{+}19]. Armv8-M [KHT21]. Arnoldi [LEA15]. Array [AM04, BG96, CZ12, CI96, Fea91, GV95, GS06, GW19, GB20, SMM94, TG05]. Array-oriented [CZ12]. Arrays [EHKT07]. Arrival [FPY08b, QA11]. Art [KPS14, LHL^{+}16]. Artificial [CSCL20, GKC22]. ASIPs [ALTT17]. Aspect [KKSP18, KK20]. ASPmT [MWHS24]. ASPmT-Based [MWHS24]. Assembly [ABTZ00]. Assessing [EGK23, KKSP18]. Assessment [BKK20, BKK23, FJA^{+}18, Hal86, UWF^{+}20]. Assignment [CB01, Fos89]. Assimilation [XZT20]. Assisted [GRV^{+}17, GAG22, MMG04, RMG^{+}13, CMW^{+}94, LCF21]. Asteroid [RC16]. Astronomy [vNR11]. ASW [ZTY^{+}19]. Asynchronous [BBC07, CJS21, DF98, GSS10, GW19, PHS19, SAS18, CG94]. Atmospheric [SMSH13]. Atom [MARC24]. Atomic [SW16, Win89]. Atomics [HVF18]. Attack [ABSSS19, MWES19]. Attacks [AAI^{+}20a, AAI^{+}20b, GWHY19]. Attempting [GYL92]. attitude [WSC20]. Attribute [MO91]. Attributes [BDD^{+}18]. Auction [WWG^{+}19]. Auction-Based [WWG^{+}19]. Auto [CCG^{+}14, Ged13]. Auto-Tuning [CCG^{+}14, Ged13]. Automata [BR97, WSS18]. Automated [AZK^{+}18, BEJD21, JGP^{+}18, VNU19]. Automatic [AAB^{+}16, API03, ALG^{+}95, BG17, BGGT02, CZ12, CZTM03, Col95, CAZ02, EM14, FCRC16, GK94, GVB^{+}06, GRC^{+}14, GYP22, GMS00, HHC^{+}15, Jak19, JW16, KSTF24, LQWP10, PHS19, SRS06, SHK13, SSB^{+}17, SNS21, TFEK16, TG05, ZLC^{+}19, vdSGBW08, KMV87]. Automatically [DDJ^{+}18]. Autonomic [GGV17]. Autonomous [KK11]. Autotuning [BC15]. Avoidance [NBA13]. Avoider [YZZ20]. Avoiding [MMN15, SJBV06]. AVX [RSJ^{+}19]. AVX-512 [RSJ^{+}19]. Aware [AOAM21, AAB^{+}16, AVLV03, CTK^{+}11, CAK17, DCX^{+}17, FPCD14, GB20, HZZS20, HSM^{+}24, JQWG15, JAW17, JLDF19, JDF20, KHT21, LQWP10, LGY16, LHLT19, Mar17, MMD21, PS23, QA11, TMXS24, WTZ^{+}19, WTQ21, XLWX19, YHGW16, MEP07, YFC21]. Awareness [KAI20, RGB^{+}08]. Axiomatization [GM20]. axioms [FK87].

B [AP86, WZTH13]. B-Queue [WZTH13]. B-Spline [AP86]. B.E. [Sca11]. Back [LXL17]. Backtracking [BMA02, SRV88]. Backtracking-Based [BMA02]. Backup [XZX^{+}15]. BADCO [VMS15]. Balance [YHGW16]. balanced [DPL86]. Balancing [ASW^{+}15, AmWHM99, EWHS11, HR11, JK03, RLH14, RSJ^{+}14, SR15]. Bandwidth [FPY08a, KSEG14]. Bank [GG13]. bards [Par86a]. Barrier [GH89, HTK98, JHLM01, Liv91, Lub90, Bro86, HFM88]. Barriers [GE90, SMC94]. Based [AAN^{+}20, AAI^{+}20a, AAI^{+}20b, AOAM21, ABSSS19, AA15, BEA^{+}19, BMA02, BHL21, BKK20, BKK23, BEJD21, BDD^{+}18, CLJH16, CND95, CLL21, CDC09, CPMC96, DK16, DeB87, DM17, DGMP09, DWQ17, EVK22, FLMR17a, FLMR17b, FCZ16, FJA^{+}18, FR95, FJZ^{+}15, FC11, FPCD14, FCRC16, GBPK07, GMB06, GAG22, GGV17, GF14, GL18, GHR20, GL92, HZL16, HmWHR97, HF14a, HF14b, HHC^{+}15, JK12, JCW^{+}18, KBD03, KKMS99, KF99, KTF23, KGK20, KT01, KJPN10, LLM^{+}12, LLM16, LPF16, LJ09, LCL19, LLL^{+}15, LWP04, LWDL17, LCL17, LYG^{+}18, MLdlP02, MCFM12, MGL^{+}17, MPR^{+}05, MWHS24, NYHA14, NSU22, NRR99, NRGB17, OB13, OBB^{+}24, PDN21, PK20, PC13, QZP15, RSP20, RLH14, RMH21, RBR22, RSJ^{+}14, RSJ^{+}19, SAB11, SAI^{+}20, SZH18, SJC18, SS17, SÜCV17, SHZ^{+}14, fSxWC18, SW95, SWF^{+}17, SDL17, TSS99, TFEK16, TESK06, TG05]. Based [UKT00, US05, UWF^{+}20, VVCA23, WLL^{+}08, WL16, WHC^{+}17, WWG^{+}19, WHC^{+}24, WK20, YWW^{+}19, YHGW16, YLB19, YZZ20, ZLD15, ZQT20, ZLA21, ZWJK05, ZJG17, ZXY^{+}15, ZHF^{+}19, uRHH14, BBR11a, BC10, CFC^{+}19, KWA^{+}10, KSF^{+}18, KM86, LP94, LJ08, MB12b, OATGEL15a, Pan08, RD08, RK13, SL14, YFC21, Mil88]. Basis [DMC^{+}18, FT87, Sch92]. Bat [NRGB17]. Batch [QGT^{+}19]. be [DM87, BGMR11]. Bedding [WSC20]. Bee [CSCL20]. Before [LCT^{+}20]. Behavior [LGY16, TMHT96]. Behavioral [LCL19, TLSG05, VMS15]. Belief [MXP14, SF20]. Benchmark [SGJ^{+}03, AM95]. Benchmarking [Cza17]. Benchmarks [TSB03]. Berlekamp [Moh19]. Better [DCX^{+}17, Par86c]. Between [BS07, Cza17, AHKR01]. Bézier [PGLC^{+}18]. Bias [DKB^{+}09]. Bidomain [XOdFV^{+}09]. Big [APR^{+}18, DC20, DX14, HSXH19, KIT^{+}20, LT17, RSA^{+}18, WTZ^{+}19, WTQ21, YS22]. Billions [qWlJzKhC17]. Binaries [JMSG02]. Binary [ABSSS19, ABvK^{+}13, DPS90, HRC17, LSA^{+}07, LYG^{+}18, NRGB17, MA87]. Binding [Con88]. Bio [CSCL20]. Bio-Inspired [Mis09, CSCL20]. Bioinformatics [VRGC19]. Biological [LHP^{+}17, ECSS88]. Biotechnology [BR14a]. Bipartite [BM09]. BitTorrent [JJIL15]. Black [IS03]. Blackboard [Dav87]. Blade [SKG09]. Block [GRV^{+}17, HCEP98, IS03, LF15, Low00, MP95, SMDJ19, SSB21, TSS99]. Block-Based [TSS99]. Block-Structured [HCEP98, MP95]. Blocked [SMN09]. BlockGraphChi [SMDJ19]. Blocking [EKU22, BG17]. Blocks [CBR17, KTRZ^{+}17, TTMD23]. Blue [MSA^{+}07]. Boltzmann [SMN09]. Boost [HP13]. Boosting [KTBP18, LRG14]. both [LSM^{+}18]. Bottom [SKG09]. Bottom-Up [SKG09]. Bound [Hun91, JGA^{+}88].

Boundaries [SNB04]. Bounds [Hen21]. BPLG [LAD15]. Brain [NFC^{+}09]. Branch [BMA02, CHYP96, CEP97, Hun91, JSHP97, KMG01, LJ08, LLSS03, TF96, JGA^{+}88]. Branch-and-Bound [Hun91, JGA^{+}88]. Breadth [GAR^{+}16]. Breadth-First [GAR^{+}16]. Breakdown [LSHK09]. Brief [KPS14]. Bringing [GHM14]. Broadband [Gsc07]. Broadcast [FPY08a, LS05, LM00]. Broadcasting [BB90]. BSHIFT [YZZ^{+}19]. BSP [AGT17, FG16, GM20, HMF^{+}13, Jak19]. BSP-Why [FG16]. Buffer [YJY16]. Bug [WLWZ15, YZZ20]. Bugs [PCJ20]. Building [CBR17, DMC^{+}18, KTRZ^{+}17, TTMD23]. Burstiness [RNJ^{+}12]. Bus [GBPK07]. Butterfly [LAD15, Bro86]. Bypassing [MS99, OVA04]. Byte [KF99]. Byte-Codes [KF99]. Bzip2 [GHDF19].

C [BJM20a, BG03, CAZ02, GH96, HG18, HTZ^{+}97, LHP^{+}22, NLBB23, PB04, SSB^{+}17, SNS21, TTF22, YBRM14]. Cache [AOAM21, CMW90, FDY^{+}19, JQWG15, KKZN12, KSEG14, KD15, LTL15, LGY16, LJE05, NIO^{+}03, NBD98, PZL^{+}19, PMHC03, RLPN^{+}02, Roy10, SNB04, SS01, SS17, SJT13, TKN^{+}08, TGT18, TFMP97, YDV19, YBDJ17]. Cache-Coherent [SS01]. Cache-Incoherent [TGT18]. Cache-Integrated [KKZN12]. Caches [BBGM95, LVJ22, MKAP05, PO07, TFMP97, WMC98]. Caching [KVG18]. Cake [CHSC18]. Calculating [LBT17]. Calculation [Boh23]. Calculations [HdMMK22, HK23]. Calculus [HL21]. Call [JK12]. Calling [JK12]. can [DM87, NLBB23]. Cancellation [GN20]. Capabilities [OATGEL15b]. Capturing [FM09]. Car [RLK20]. Carbon [CDC09]. Cardiac [LLGC17, XOdFV^{+}09]. Carlo [BJM20b, PES^{+}18]. Cartesian [AKHD13]. CAS [MMG04]. CAS-DSM [MMG04].

Cascade [KTBP18]. Case [BKT08, CG94, CGPS18, CML04, DE00, LDHL05, SPS14, TESK06, KM86]. Categorization [LYL14]. Cattle [KSA^{+}18]. Caused [MKAP05]. CCAP [JQWG15]. ccNUMA [NP01]. CCRP [BHL21]. CDMA [GN20]. Celerity [TTF22]. Cell [WSO^{+}07, BGMR11, Gsc07, OOS^{+}08, Sca11, SKG09]. Cell/B.E. [Sca11]. Center [DFZ21]. Centers [AAN^{+}20, LVM16]. Central [FVvL^{+}16]. Centric [CM06, FPCD14, KP01]. Cetus [BML^{+}13, RMG^{+}13]. Challenges [Bel94]. Change [LFHAM19]. Channel [Gha19, GL92, HZZS20, WQJY17]. Channel-Aware [HZZS20]. Channels [FDY^{+}19, KLK16, YDV19]. Characteristics [SH96, Tic90]. Characterization [AVM^{+}16, GM20, YDV19]. Characterizing [BGdS09, CFX^{+}20]. Charismatic [LKS^{+}20]. Check [NRR99]. Checking [AKBPV19, Hen21, HMK09, TLSG05]. Checkpointing [LNP91, RMG^{+}13]. Chemistry [CGN^{+}09]. Chinese [FCZ16]. Chip [AO19, AOAM21, GRV^{+}17, GG13, Gsc07, JLDF19, JDF20, KKZN12, KSEG14, KT01, LS07, MVB^{+}06, OP12, OBB^{+}24, PM07, TGT18, TESK06, ZK07, ZGH^{+}15, ZC09, AH08]. Chip-Multi [AOAM21]. Chip-Multiprocessors [GRV^{+}17, TGT18]. Chips [NCR^{+}19]. choice [BS89]. Cholesky [GN89]. church [Ano86a]. Circuit [PMV17, WPC07]. circuits [BH87]. CISL [MPR^{+}05]. Cities [KIT^{+}20]. Clairvoyant [SY08]. Class [BEP13, MPR^{+}05, IPR^{+}05]. Class-Based [MPR^{+}05]. Classification [CHYP96, CS20, KTBP18, Mon97, NAS23, QZP15]. Cleaning [MCT^{+}18]. Clearance [GAK20]. Climate [HNC^{+}16, LHF^{+}15]. Cloaking [MS99]. CLOMP [BGdS09]. Closure [CAP88, KP95, KPRS96, VK88]. Cloud [AAI^{+}20a, AAI^{+}20b, CAK17, DS20, HZL16, HC17, JM20, JQWG15, KJHB14, RLH14, WQJY17, XZX^{+}15, XLWX19, uRHH14]. Cloud-Based [AAI^{+}20a, AAI^{+}20b]. Clouds [JAW17, LTF^{+}12, LCT^{+}20]. Cluster [CYS16, EAT14, ES11, FPCD14, LJ09, LTL15, LSYG15, MLdlP02, NIK00, SCB^{+}14, YS22]. Cluster-Based [FPCD14, LJ09]. Clustered [CPG01, GBPK07]. Clustering [ANS20, BABW14, CS20, CAP88, DMC91, FCZ16, LKS^{+}20]. Clusters [BEA^{+}19, BS03, BC15, DWQ17, EAK21, FPY08a, GCD^{+}03, GSY^{+}13, HC17, HOZ06, QA11, RPF18, TTF22, WK20]. CMP [DLX^{+}17, LTL15]. CMPs [BHJ06, FC11, KKZN12, LGY16]. CNN [SHS21]. CNNs [SWG^{+}18]. Co [GRAG00, MPR^{+}05, NB15]. Co-Generation [MPR^{+}05]. Co-operation [NB15]. Co-Scheduling [GRAG00]. Coarse [CSF^{+}20, NIO^{+}03, PSM97, SSM21, WW17, AD89]. Coarse-Grain [PSM97]. Coarse-Grained [CSF^{+}20, SSM21, WW17]. Code [AKBPV19, ABTZ00, BTB^{+}13, CPG01, GBLG10, GK94, HBC23, HRC17, JS10, KaM10, KAMAMA17, LF15, LC11, MGW99, MCA98, MP04, NRB94, NLBB23, ÖO07, PB04, TFEK16, TF94, WNMW16, WK20]. Coded [WHC^{+}24]. Codes [CAZ02, ELGE17, HTK98, KF99, RMG^{+}13, SF20]. Coding [DLRS13, MB12b, SSEA14, YMW^{+}17]. Coflow [CLL21]. Coherence [CMW90, FC11, KSEG14, MPAG18, PMM^{+}18, SNB04, SMH21, YDV19, BCK98]. Coherence-Free [PMM^{+}18]. Coherent [SS01, TGT18]. Cohesion [KKSP18]. Collaborative [Gen16, JCW^{+}18, VSDK09, WLWZ15]. Collection [Cra88, AH86]. Collective [BG17, FPY08b, IBA11, KSTF24]. Collector [Fos89, LWLG11]. Colluding [AKA^{+}20]. Colony [ASG20, CSCL20, dMMHdLN21]. Combinatorial [MAT23]. Combining [ABASS12, GV95, GH89, HSCI^{+}16, LSM^{+}18, LLSS03, RK92, SMC94, WMC98]. Coming [LS07]. Commands [GYL92]. commentary [Lin88a]. Comments [Swa88]. Commercial [NYHA14, RLPN^{+}02]. committed [BS89]. committed-choice [BS89]. Common [BEJD21]. Communal [BKK20, BKK23]. communicating [Mai87, RS90]. Communication [AAN^{+}20, AH08, BG17, CTB14, GAR^{+}16, GL95, IBA11, IKN00, JQJ^{+}16, KHH08, KKZN12, KSTF24, KT01, KTT^{+}99, LM00, MMN15, MEP07, MO91, OPLS17, PSM97, RGB^{+}08, TOM^{+}11, TA99, WZTH13, MO90]. Communication-Avoiding [MMN15]. Communication-Driven [TOM^{+}11]. Communications [HZZS20, Mon97]. Compaction [DH00, KGK20, PYX17]. Compactors [ZC09]. Comparative [BFRPVR^{+}15, HPVRPF15, LMPS05, YS22]. Compare [FLD15, Sun11]. Compare-and-Swap [FLD15, Sun11]. Comparison [BS07, DMC^{+}20, HMF^{+}13, OP10, SS01, ECSS88, FT87, GE89, Hua89, Kas86]. Compatibility [CS97]. Competitive [Gen16]. Compilation [AVLV03, GBLG10, HmWHR97, JB98, KL00]. Compile [KCW^{+}05, vdSGBW08]. Compile-Time [KCW^{+}05]. Compile/Run [vdSGBW08]. Compile/Run-time [vdSGBW08]. Compiler [AZK^{+}18, ALPS19, BML^{+}13, BKT08, CGN^{+}09, CTK^{+}11, CP04, CFB94, CEH13, EM13, FKM^{+}11, GBC^{+}08, HTK98, JCD^{+}14, Ken94, KTT^{+}99, LCF21, LEL^{+}99, MMG04, MO91, MCA98, MAB^{+}11, PB04, RMG^{+}13, RBES00, SSP^{+}00, SBC17, SG00, TMHT96, TJY99, YZ13]. Compiler-Assisted [RMG^{+}13, LCF21]. Compiler-Generated [JCD^{+}14, MCA98]. Compiler-Parallelized [HTK98, TMHT96]. Compiler-Towards [SSP^{+}00]. Compilers [HML^{+}20, MPR^{+}05, ME15, SGK12]. Compiling [HTZ^{+}97]. Complementary [LkCH94]. Complete [BdS07]. Complex [AMP^{+}05, CHCL14, fS18]. Complexity [DFH17]. Component [EFED05, MLdlP02, fSxWC18]. Component-Based [MLdlP02]. Components [DKB^{+}09, DJR16]. Composability [CB19]. Composable [AMP^{+}05]. Composed [LWF^{+}19]. Composition [GVB^{+}06, GGV18, HHC^{+}15, RK13]. Compositional [EHKT07, TLSG05]. Comprehensive [OATGEL15a]. Compressed [KK11]. Compression [BABW14, HNC^{+}16, KKMS99, TSS99, VHK^{+}18]. Computation [BG17, BEP13, CJA00, Cza17, DMMS91, FLMR02, HSCI^{+}16, JAW17, KSBN22, LEA15, MCWK01, MNN22, NdMM09, PGLC^{+}18, Ric90, SSM21, Ski91, SSB21, KMV87, MA87]. Computation-Oriented [SSM21]. Computational [HLK^{+}09, LLL^{+}15]. Computations [HKJ^{+}18, IH04, JLV21, NST89, PMHC03, SBC17, VCP^{+}16, YH18, LRG^{+}91, SS89, TKM89, Wai87]. Compute [LSM^{+}18, SR15]. Compute-intensive [LSM^{+}18]. Computer [ACC^{+}02, Ano18a, DB08, DMC91, Kuc94, MCE13, MGJS15, MB12a, PVG17, RJO22]. Computers [Bel94, GHC^{+}17, HOZ06, MLdlP02, Ano87d, Gao86]. Computing [APR^{+}18, ACD^{+}16, Ano16b, Ano18b, Ano19, BE14, Car09, CTP13, CSF^{+}20, CGPS18, CSTGL03, DDD^{+}19, DFH17, Den94, DS20, FBV21, FKT12, GWYQ18, Gha19, HMF^{+}13, HdMMK22, HK23, HDK24, HLS15, HS16, JM20, KJHB14, LLGC17, LRG14, MB12a, OATGEL15b, OG11, PLN^{+}04, RLH14, SM09, SZ17, SWG^{+}18, TG21, TRL09, TAY^{+}12, TFPF18, VCP^{+}16, WTZ^{+}19, WTQ21, WSO^{+}07, WGW04, YS22, YFC21, Zha10, ZZS^{+}19, NK88, DB08].

Concatenation [Zey05]. Concept [KaM10]. Concurrency [BAF94, Gen16, PCJ20, SB90, VSH^{+}11, WLWZ15, YZZ20, AD86, CP88, DM87, Pra86]. concurrency/synchronization [AD86]. Concurrent [Ano16c, AR16, CHSC18, GMP89, LWDL17, PB01, SBC17, TSS86, VVCA23]. Condition [NBN^{+}15]. Conditional [LNG12]. Conditions [MJ02]. Conference [MCE13, PVG17, RJO22]. Confidence [KMG01]. Configurable [HL21, TRD21]. Conflict [CRM17, NBA13]. Conflicts [GG13, SD11, WS14]. Conformance [TLSG05]. Congestion [AAN^{+}20, JLDF19]. conjugate [SDJS98]. Connected [DMC91, DWS16]. Connection [GH96, PW92]. Connectionism [Ano87b]. Connections [ALTT17]. conquer [MFC21]. Conscious [ZK07]. Consecutive [AKA^{+}20]. Consensus [NAS23]. Considering [OPLS17, WMC98, XLWX19]. Consistency [BAP01, BBGM95, LNG12, SHZ^{+}91]. Consistent [KS90, SH87, Swa88]. Consolidation [LVM16, XLWX19]. Constant [LMP98, MWES19]. Constant-Time [MWES19]. Constrained [BABW14, SHS21, VCP^{+}16]. Constraint [MRLR16, JB98]. Constraints [AKD98, AF15, API03, BEJD21, BBR11b, RBES00, SWZ^{+}15]. Construct [Spr92, FcF87]. Constructing [DWQ17, KP05, DPL86]. Construction [BNWL90, CP04, WR18]. constructs [BCK98]. Consumption [RSP20]. Container [DST21]. Containers [DK16, WP00]. Contention [Dem11, JQWG15, SMC94, SAL16]. Contention-Aware [JQWG15]. Contention-Free [SMC94]. Context [CJA00, IP90, Lan90]. Context-free [IP90, Lan90]. Contexts [JMSG02]. Contraction [SSP^{+}96]. Control [AAN^{+}20, AmWHM99, FM09, Gen16, Kas86, KHT21, MCA98, MJ02, NYHA14, RLK20, SKA96, SB90, VSH^{+}11, YKM03, ZGH^{+}15, FK87]. Control-Flow [KHT21]. Controlled [DJS12]. Controller [OBB^{+}24]. Controllers [ANS^{+}12, MFGEL19]. controlling [VR88]. ControlPULP [OBB^{+}24]. Converging [BHL21]. Conversion [AmWHM99, SJBV06]. Convex [SS92, SSMO96]. Convoider [YZZ20]. Convolutional [HSM^{+}24, WZG^{+}17, YYYX20]. cooperating [NP98]. Cooperation [LCL19, PTdSF^{+}12]. Cooperative [ANS20, AATD20, NdMCdMMW16]. Coordinated [Gen16, LLM16]. Coordination [BCS^{+}09, CM06, GGV17, Sek09]. Coprocessor [BP17]. Coq [GHM14, LBT17]. Core [ABvK^{+}13, AML^{+}10, ABB^{+}10, BKT08, CFC^{+}19, GGV17, uHKAMFM16b, LHP^{+}17, LZ17, MFU21, OBB^{+}24, PMM^{+}18, RPF18, SSEA14, SHLJ17, SMDJ19, SASH12, SA10, VMS15, WZTH13, Zha10, vNR11, BEJD21, CZ12, Ged13, HML^{+}20, MXP14, NdMCdMMW16, QZP15, RC16, XWH21, uHKAMFM16a]. Core-to-Core [WZTH13]. Cores [Boh23, CTK^{+}11, HG18, RTD20]. Cores-Enabled [Boh23]. Correction [AAI^{+}20a, HTDL18, Moh19]. Correctness [HZL16, HMK09, SNB04]. Correlated [JW16]. Correlating [vNR11]. Coscheduling [FR95]. Cost [BDH^{+}14, Jak19, VNU19, WS15, XLWX19, YZZ^{+}19]. Cost-Directed [BDH^{+}14]. Costs [TB23]. Countering [AAI^{+}20b, AAI^{+}20a]. Coupled [RWMF24, SS01, ZTY^{+}19, Zey05]. Coverage [MAWD^{+}16]. Covering [JLMW15]. Covert [WQJY17, YDV19]. CovertInspector [WQJY17]. CPU [BC15, FFS18, GGV17, GHC^{+}17, LLGC17, LRG14, Moh19, NCR^{+}19, PTdSF^{+}12, SSB^{+}17, SBC17, WE18, ZTY^{+}19].

CPU/GPU [GGV17, SSB^{+}17]. Credit [BHL21, YHGW16]. Credit-Based [BHL21, YHGW16]. Critical [SK97]. Cross [AM95, DSR17, WCC16, ZJG17]. Cross-ISA [WCC16]. Cross-Platform [ZJG17]. Cross-Thread [DSR17]. Crypto [MWES19]. Cryptography [Dam07]. CSMqGraph [CSF^{+}20]. CSP [FcF87]. CTA [YHGW16]. Cuckoo [DS20]. CUDA [HLP11, HF14a, HF14b, KAMAMA17, LRG14, MGL^{+}17]. Curve [Bos12]. Custom [MCFM12]. Customization [GSY^{+}13]. Customized [ASG20]. Cycle [FCJV99, HZZS20, SAB11, TGT18, dMP^{+}03]. Cycle-Approximate [TGT18]. cyclic [JB98]. Czip [HNC^{+}16].

D [BC15, CTB14, ES11, HPVRPF15, HF14a, HF14b, IBA11, JGM15, LLGC17, LHP^{+}17, SJKA99, fSxWC18, SBC17]. D-Stacked [LHP^{+}17]. DAFT [ZLJA12]. Daily [Ano87c]. DancerFly [JDF20]. Data [APR^{+}18, AKHD13, ABTZ00, Ano16d, AMKE18, ALPS19, AJF16, ANS^{+}12, ALG^{+}95, BARSW95, BS03, BBGM95, BG96, BCL17, CFB94, CAK17, CAT18, DDD^{+}19, DM17, DC20, DTLW16, DMC^{+}20, DX14, DFZ21, DLX^{+}17, DJS12, EW96, EK17, ELGE16, FJA^{+}18, FRT^{+}18, FPCD14, GSP^{+}17, GG14, GL18, GV99, GYL92, GB20, HSCI^{+}16, HZZ^{+}19, HRH08, HHW20, HP13, HGT^{+}12, HTmG^{+}12, HNC^{+}16, IR19, KIT^{+}20, KTF23, KP01, KP04, LSA^{+}07, LTL15, LVM16, LT17, LSM^{+}18, LWLG11, LHF^{+}15, MXP14, MHL95, MCWK01, MTT15, NRR99, NP19, NAP02, NLRH07, OK99, PCJ18, PMHC03, RSA^{+}18, RG15, RS90, Ric90, RSJ^{+}14, SdLC21, SNB04, SS99, SL14, SQH92, SR04, SH15, SASH12, TESK06, TFMP97, TMXS24, VVCA23, WTZ^{+}19, WSC20, WTQ21, WHC^{+}24, WB87, WE18, WW17, XZT20, XH98, YAI95, YWW^{+}19, YS22, YLB19, ZLA21]. Data [vdSGBW08, CG94, Gao86, Kas86, Win89].

Data- [LSM^{+}18]. Data-Analytics [TMXS24]. Data-Centric [FPCD14, KP01]. Data-Driven [AMKE18, DTLW16, TESK06]. Data-Flow [Ano16d]. Data-Intensive [LWLG11]. Data-Locality [CAK17]. Data-Obtaining [XZT20]. Data-Parallel [AJF16, SQH92]. Data-Sharing [SNB04]. Database [SB90, STM15, VK88]. Databases [WZB^{+}92]. Datacenters [BHL21]. Dataflow [BS15, CGPS18, CZTM03, Fea91, FBGEL19, JCW^{+}18, LS98, NRB94, NG92, OGP^{+}16, TB23, WGF^{+}16, ZD19]. Dataflow-Based [JCW^{+}18]. Dataflow-Inspired [OGP^{+}16]. Datapath [SdLC21]. Datasets [MV17]. Daytime [fS18]. DCF [JCW^{+}18]. dCompaction [PYX17]. Deadlock [LWDL17]. Deanonymization [GWHY19]. Debugging [BBGM95]. Declarative [KTF23]. Decomposed [WEJS94]. Decomposition [ADC^{+}17, BUMS02, KKKS24, QZP15, YYYX20]. Decomposition-Based [QZP15]. Decoupled [ZLJA12]. Deductive [FG16]. Dedup [DCX^{+}17, GHDF19]. Deduplication [DCX^{+}17]. Deep [CLL21, FFS18, GVB^{+}06, HZZ^{+}19, MFC21, PVF21, QGT^{+}19, WQZ^{+}24, YZZ^{+}19, ZJL22]. DeeperThings [SHS21]. Defect [OB13]. Defend [FDY^{+}19]. Defense [YDV19]. Define [JACK20]. Defined [CSCL20, KWA^{+}10, KIT^{+}20, NSU22, RMH21, DM87]. definite [GHLN86]. Definition [OK99]. Degree [AK17]. Delay [BMA02, NST89]. Delayed [PYX17]. Delta [WHC^{+}24]. demand [JK86]. demand-driven [JK86]. Demonstrating [ACC^{+}02]. Denotational [Hud86]. Dense [KFC08, MVB^{+}06]. Dependence [ABTZ00, DV97, GV95, KP04, MHL95, Mon97, NP19, SW95, YAI95, WB87]. Dependences [CDRV98, DSR17]. Dependencies [CH95, JB98]. Dependency [MWES19, WTZ^{+}19]. Dependency-Aware [WTZ^{+}19]. Dependent [DFA^{+}09, VMS15]. Deployed [AAN^{+}20]. Deployment [PK20]. Depth [PTdSF^{+}12, ZJL22, KR87, RK87, SJW22]. Derivation [MLdlP02, SO89]. Derivative [Boh23]. Derived [NFC^{+}09]. Deriving [Wai87]. Description [ARB^{+}05, MPR^{+}05]. Descriptions [GmWHR98, KP05]. Design [ACD^{+}14, AML^{+}10, AR16, BC10, BDD^{+}18, BS91, CHB06, CDC09, CZTM03, DMK21, FVvL^{+}16, HLK^{+}09, KHH08, KWA^{+}10, KS97, LMHW18, MB99, MARC24, MWHS24, OGP^{+}16, PG07, PP10, RK13, SSP^{+}00, SY08, TLSG05, TKN^{+}08, WLL^{+}08]. Designing [BKK20, BKK23, SCB^{+}14]. Desktop [GS13]. Detailed [LLGC17]. Detect [DS16]. Detecting [AKA^{+}20, SW95]. Detection [CLJH16, CRM17, DV97, HPY01, HTDL18, Jan15, KSJ14, LFHAM19, MTT15, NSU22, NRGB17, fSxWC18, fS18, WLWZ15, YWW^{+}19, YLB19, ZQT20, Tho87]. Detection/Correction [HTDL18]. Determinism [SÜCV17]. Deterministic [EVK22, PTD^{+}06, ZC17]. Developing [CEH13, EHKT07, GHC^{+}17]. Development [Dam07, TCUV14, dMP^{+}03]. Device [GZJ18, MFGEL19]. Devices [AER^{+}17, AGPGF14, Dam07, JQJ^{+}16, KKKS24, SHS21]. DFA [KJHB14]. Diagnosis [GKC22]. Diagnostics [RC16]. Diagonalization [HOZ06]. Difference [Ger10]. Different [DMC^{+}20, GWPV21, JCH^{+}08]. Differentiated [AKT^{+}14]. Diffusion [HF14a, HF14b, SDJS98]. Digital [GP17, RSP20, SS92, ZC09]. Digital-Microfluidic [ZC09]. Dijkstra [FBV21]. Dimensional [BARSW95, EHKT07, HDK24, LPB13, Fea92a, RG18, WLL17]. dining [RB86]. Directed [BDH^{+}14, Hue97, Zha89]. Directives [AAB^{+}16, HSCI^{+}16]. Discontinuous [CF19]. Discovery [GHC^{+}17, KSTF24]. Discrete [Dem11, PPQV16, SSB21]. Disjoint [SH15]. Dispatcher [SY08]. Dispersion [RSV^{+}05]. Distance [BEP13, HGT^{+}12, KAI20]. Distributed [AKA^{+}20, BS03, BR14b, Boh23, CY14, CCL12, CR19, CHCL14, CSTGL03, DS97, DZW10, EK14, FSS06, FPCD14, FBGEL19, GHM14, GL95, HK23, HOZ06, JLJ^{+}18, JAW17, KTF23, LWL^{+}19, Lin91a, LHP^{+}22, MP91, MMG04, MBE03, MVD^{+}14, NIK00, OATGEL15b, OG11, PLN^{+}04, PS23, SNB04, SW16, SB91, SHS21, TTF^{+}08, TTMD23, qWlJzKhC17, WTL^{+}23, WQZ^{+}24, AH86, GS90, GT86, PW87, RB86, RS90, TKM89, Tho87, Sek09]. Distributed- [FBGEL19]. Distributed-Memory [BS03, TTMD23, qWlJzKhC17]. Distributed-Order [Boh23]. Distributing [HHW10]. Distribution [ALG^{+}95, HSCI^{+}16, NAP02, SSP^{+}96]. Distributions [AKHD13]. Divergent [LVJ22]. Divide [MFC21]. Divide-and-conquer [MFC21]. Divisible [RSJ^{+}14]. DMR [ZC17]. DNN [LCF21, WTL^{+}23]. Do [Kuc94]. DocChip [TRD21]. Document [LKS^{+}20, TRD21]. Domain [CB19, GF14, TFEK16, WL16, WK20, RK13]. Domain-Based [GF14]. Domain-Specific [WK20]. Dominance [Spr92]. Dominating [DWS16]. Double [KJPN10, LLM^{+}12]. Double-Precision [KJPN10]. Downsampling [LTSD15]. DRAM [SJW22, WMN^{+}17, ZLJ^{+}17]. DRAMSpec [WMN^{+}17]. DRAMSys4.0 [SJW22]. Driven [AMKE18, CPMC96, DTLW16, GRC^{+}14, RNJ^{+}12, TOM^{+}11, TESK06, VNU19, XH98, JK86, Kas86]. Driver [GZJ18]. Drives [YJY16]. DRL [CLL21]. Drone [MMD21]. dRuby [Sek09]. DS [GN20]. DSM [BAP01, MMG04, WLL^{+}08]. DSMs [HTK98, KVG18]. DSP [SHK13]. DSParLib [LHP^{+}22]. Dual [WS08]. Dual-thread [WS08]. Duo [BKT08].

Duplication [CKC22]. Duty [HZZS20]. DVFS [CKC22]. DynaCo [SMH21]. Dynamic [ABvK^{+}13, CPG01, CR19, CS97, CML04, EWHS11, GAG22, Hue97, JK12, JM20, JCD^{+}14, KCW^{+}05, LSA^{+}07, LTF^{+}12, LSYG15, LCL19, LGY16, LMPS05, Lys08, MRLR16, MTT15, NBA13, NLRH07, OVA04, PDN21, PD89, PCJ20, PO07, PVF21, RSP20, RMH21, RD08, RRH03, SSNS16, SR04, SMH21, SJT13, TCUV14]. Dynamically [CHPC96, GMB^{+}11]. Dynamics [ACC^{+}02]. DySHARQ [RMH21]. DyTO [JM20].

Eager [SAL16]. Early [PYC16, TA99]. EARTH [HTZ^{+}97, HMT^{+}96]. EARTH-MANNA [HMT^{+}96]. Eat [CHSC18]. Economics [LCL19, YBDJ17]. Ecosystem [RSA^{+}18]. Edge [KKKS24, MMD21, SHS21, YFC21, ZQT20]. Editor [EA09, MA10, SS10, BCL90, Ano00a, Ano14, Ano16a, Ano16b, Ano16d, Ano16c, Ano18b, Ano18a, Ano19, Ano20, Ano21a, Ano21b, Ayg03, Ban94, Ban04a, Ban04b, Car09, Fur95, Gau96, Giv07, Giv08, Int98, JS06a, JS06b, Joe99, Joe03, Kes20, McK07, Mis09, NL23, Ora03, Pan08, Seh98, Vei01, Vei02]. Editorial [Ano86b, AG15, Bro19, CTP13, CAT18, DPT17, FKT12, FH05, GGE19, Gha19, GK18, Gre16, HK14, HSXH19, JACK20, LFL^{+}17, LT17, MCE13, MGJS15, MGD^{+}14, OG11, PP10, PVG17, RJO22, SGK12, SS10, TG21, TFPF18, ZZS^{+}19]. Editors [SMM11, HF06, AM07b, CHS99, CmHS99, EmH97, FmH96, GSA08, GS05, HN94]. Effect [NPD89, BCK98]. Effective [CPMC96, HGT^{+}12]. Effectiveness [GYP22, MHL95, PYC16, SBN03]. Effects [HRH08, TF96]. Efficiency [BBB^{+}17, EGK23, KTBP18, STF^{+}12, SWZ^{+}15]. Efficient [ASS24, ABvK^{+}13, BR97, BEP13, BCL14, BFG^{+}10, CR19, CPT14, CL96, CKC22, EAT14, FPY08a, Fea92b, FVvL^{+}16, GSP^{+}17, GG14, GN20, GL18, GAK20, GS06, GRR98, GHC^{+}17, GmWHR98, HZZ^{+}19, IP90, IBA11, JGM15, KDV22, KP05, LNP91, LS05, LNG12, LWLG11, LMHW18, LWGZ18, NRR99, NdMMW16, QRW00, RPF18, RSP20, RLEJ19, Roy10, SRS06, SSNS16, SL14, SSP^{+}96, SS23, SO89, SKAT91, SHC15, SHZ^{+}14, SJT13, TTF^{+}08, WZTH13, WQZ^{+}24, XZX^{+}15, YJY16, YLB19, Fea92a, Hua89]. Efficiently [EGJS15, HR11, JMSG02]. Elastic [GG13, YBDJ17]. ElasticActor [ZLC^{+}19]. Electronics [GWPV21]. Element [MCT^{+}18, RG15]. Elements [qWlJzKhC17]. Eliminate [KTT^{+}99]. Eliminating [HTK98]. Elliptic [Bos12]. Embedded [Ano18a, Ano21b, AF15, CHB06, CFF^{+}06, DLRS13, DLX^{+}17, Giv07, Giv08, KTBP18, LMPS05, MSJ20, MCE13, MGJS15, MAB^{+}11, Pan08, PP10, PVG17, PO07, PPEP08, RJO22, SSM21, TLSG05, TFEK16, TGT18, US05]. Embedding [Li03, CSG89]. Emergency [GAK20]. Emerging [HP13, JACK20, TG21, ZZS^{+}19]. Empirical [CCG^{+}14, LDHL05, PMV17, SSMO96, YYYX20]. Employing [CS97, KKKS24]. Emulation [OBB^{+}24]. Emulator [WCC16]. Enable [HP13, ID08, TAY^{+}12]. Enabled [Boh23, FKM^{+}11, GZJ18, GSY^{+}13, JACK20, MMD21, SAI^{+}20, RA09]. Enabling [GZJ18, SdLC21, SMDJ19]. Encore [GTK^{+}88]. Encryption [AAI^{+}20b, KBD03, NdMMW16, AAI^{+}20a]. End [LSHK09]. End-to-End [LSHK09]. Endpoint [JLDF19]. Energy [AVLV03, CPT14, CKC22, EAT14, FVvL^{+}16, HYBA18, KAI20, LMHW18, Mar17, SSM21, SJT13, VCP^{+}16, XLWX19]. Energy-Aware [Mar17, XLWX19]. Energy-Constrained [VCP^{+}16]. Energy-Efficient [CKC22, EAT14, FVvL^{+}16, LMHW18, SJT13]. Engine [BC15, RLK20, Gsc07]. Engineering [CPT14, KaM10]. Engines [MCFM12]. Enhanced [ABASS12, FMSG17, GRAG00, RY20, RY22, RSJ^{+}19]. Enhancement [AMP01, CYS16, HML^{+}20, KP01, LCL17, SAI^{+}20]. Enhancing [ACC^{+}01, GYP22, MP95, SZH18]. Ensembles [ASW^{+}15]. Enterprise [LVM16]. Enumeration [AG98, GL18]. Environment [AFM^{+}06, AA15, BFG^{+}10, DMMP18, MFG^{+}08, QFRA19, SQH92, UWF^{+}20, vdSGBW08]. Environments [BCS^{+}09, BFRPVR^{+}15, Car09, CCL12, CAK17, GWPV21, HHW10, HdMMK22, HK23, HDK24, KJHB14, LLM16, PCJ20, TTF^{+}08, XLWX19, BCL90, Con88]. Epidemic [LEA15]. Equation [CTB14, ES11]. Equations [LM00, XOdFV^{+}09]. Equivalence [AKBPV19]. equivalences [Mai87]. Era [AATD20, ABB^{+}10, DX14]. Erasure [WHC^{+}24]. Erasure-Coded [WHC^{+}24]. Erlang [BDH^{+}14, STB^{+}18]. Erratum [Ano03, FLMR17a, HF14b, uHKAMFM16a]. Error [DFC^{+}07, Moh19, ÖA21, RLK20]. ESL [GHR20]. Estimating [DMC^{+}20, HGT^{+}12, KK20]. Estimation [DKB^{+}09, GHR20, KMG01, LPF16, LLL^{+}15, MVD^{+}14, TSS99, YYYX20]. Evaluating [AM95, BCK98, SCB^{+}14, TF96]. Evaluation [AMAH01, BML^{+}13, BS15, BEG^{+}10, CCL12, CDC09, DMC^{+}18, FC11, GBPK07, GWPV21, IPR^{+}05, JGP^{+}18, JCH^{+}08, KHH08, LCL17, ME15, NRB94, NP19, OATGEL15a, PVAE98, SSMO96, TSB03, CSG89, LAV98, VK88]. Even [DCX^{+}17]. Event [Dem11, PPQV16, RNJ^{+}12, WZG^{+}17]. Event-Driven [RNJ^{+}12]. Eviction [GSP^{+}17]. Evolution [ACD^{+}16]. Evolutionary [ACD^{+}16, HSM^{+}24, MWHS24, PB01, STB^{+}18]. Evolving [GKC22]. Exact [MAT23]. Example [SO89, Wai87]. Exascale [MAJD16].

Exceed [LS98]. Exception [FMSG17]. Exceptions [AHKR01]. Executable [LC11]. Execute [GYL92, BS89]. Executing [FCRC16]. Execution [AMKE18, BS15, BEJD21, BAF94, CHPC96, Col95, CSTGL03, CFF^{+}06, DJS12, EAT14, FM09, GS06, GL95, JSHP97, KLG08, KGK20, LLL^{+}15, LEG11, LCL17, Lys08, MFG^{+}08, OGP^{+}16, SNB04, SAS18, SB91, SBC17, TTF^{+}08, Tic90, TF96, VVCA23, WE18, Ali86, Gol88, Kas86, KM86, SRV88]. exemplified [Tho87]. Expansion [BCC00]. Experience [Hal86, HmWHR97, RMG^{+}13, SCB^{+}14]. Experiences [CEH13, NP98]. Experimental [AFM^{+}06, IPR^{+}05]. Experiments [Hun87, NPT86]. expert [KM86]. Explainable [HSM^{+}24]. Explicit [BP17, DMC^{+}18, Ger10]. Explicitly [LMP98]. Exploit [ADC^{+}17]. Exploiting [BS03, Gsc07, GL92, JG97, JLDS16, LS98, SASH12, VCP^{+}16, YDV19]. Exploration [CZTM03, KWA^{+}10, MSJ01, MWHS24, PMM^{+}18, SEP08, WMN^{+}17]. Exploring [AHKR01, FVvL^{+}16, PG07]. Exponentiations [NdMM09]. Expose [GV95]. Express [GZJ18, JQJ^{+}16]. Expression [AFO^{+}08, Sca11]. Extend [DFA^{+}09]. Extended [BG03, DDD^{+}19, Sch92, YAI95]. Extending [ABB^{+}10, ML15]. Extensibility [CB19]. Extensible [CP04, SHK13]. Extension [BG03, CFB94]. Extensions [API03, CZTM03, RSJ^{+}19, SG00, TMXS24]. external [CSF^{+}20]. Extracted [KP04]. Extracting [PJS^{+}05]. Extraction [AER^{+}17, JK12, LKS^{+}20].

Fabrics [GBC^{+}08]. Face [LYG^{+}18].

FACILE [GMP89]. Factor

[BTB^{+}13, MXP14]. Factorization [DZW10, LKS^{+}20, GN89]. Failure [GCD^{+}03, Mai87]. failures [TKM89]. Fair [Lin91a]. FAIRIO [AKT^{+}14]. Fairness [FK87]. False [GMB95, MWES19]. Family [PVAE98]. Farm [EK14, JBB21]. Fast [BC15, CS16, HNC^{+}16, JGP^{+}18, JLMW15, Joh94, Ken01, KT01, NIK00, RGB^{+}08, SMC94, SHLJ17, TA99, WZTH13, XZT20]. Fast-Fits [Joh94]. FastFlow [TTMD23, TRB^{+}24]. Fault [AKHD13, CJS21, DFZ21, EAT14, GWYQ18, GJR09, Gha19, HTDL18, LJ09, MEP07, NAS23, NRR99, WGF^{+}16, ZLJA12]. Fault-aware [MEP07]. Fault-Management [GJR09]. Fault-Model-Relevant [NAS23]. Fault-Tolerant [DFZ21, EAT14]. Faults [LG10]. Faulty [BB90]. FEADS [PG07]. Feature [CS20, SAS18]. Features [PYC16, Zha10]. Feel [NLBB23]. Fences [HVF18, LNG12]. Ferret [GHDF19]. Fetch [HCEP98, MSJ01]. Fetching [NG92]. FFNNs [SDH22]. Field [QZP15]. File [ALTT17, AVLV03, CND95, KK11, Mar09, ZLAV04]. Fill [BMA02]. Filter [HSM^{+}24]. Filtering [JCW^{+}18]. Filters [MKAP05, WdSAM^{+}17]. Find [CAP88, KP95]. Finding [KLK16]. Fine [BG96, CTK^{+}11, DV97, GL92, PSM97, SDH22, SZH18, WTQ21]. Fine- [PSM97]. Fine-Grain [BG96]. Fine-Grained [CTK^{+}11, GL92, SDH22, SZH18, WTQ21]. Finite [BR97, Ger10, MCT^{+}18, RG15]. First [GAR^{+}16, KS90, MKAP05, KR87, RK87]. First-Level [MKAP05]. Fish [WMK19]. Fits [Joh94]. Fix [HZZ^{+}19]. fixed [Ano86a]. Fixpoints [Ano87c]. Flat [FT87, TSS86]. Flexible [ELK18, KHH08, KKZN12]. Flight [SI11]. Float [HZZ^{+}19]. Float-Fix [HZZ^{+}19]. Flow [Ano16d, AmWHM99, BG96, FM09, GG14, GYL92, GV95, KHT21, KSBN22, KTF23, MCA98, MJ02, NRR99, SBN03, Gao86, RS90]. Flower [MSJ20]. Flows [YKM03]. Fly [JDF20, KSJ14]. Flying [LCT^{+}20]. FOG [SHLJ17]. FORAY [ID08]. Forecast [BBB^{+}17]. Forecasting [MMD21]. Forensics [ZXY^{+}15]. Forest [YWY^{+}18]. ForestGOMP [BFG^{+}10]. Foreword [BmH98, NS97a]. Fork95 [KS97]. Form [CB01, TG05]. Formal [BdS07, KP05, LMPS05, MP91]. Formalised [GGV18]. Format [ASS24]. Formats [Mar09]. Fortran [KaM10, NLBB23]. Fortress [ASS21]. Forwarding [CLJH16]. Forwarding-Based [CLJH16]. Four [TSS99]. Fourth [BP17]. Fourth-Order [BP17]. FP [BARSW95]. FPGA [BETK24, KJPN10, MCFM12, OBB^{+}24, SCS23]. FPGA-Accelerated [BETK24]. FPGA-Based [MCFM12, OBB^{+}24]. FPGAs [STM15, VNU19]. Fractal [MP04, SC88]. Fractional [Boh23, JLMW15]. Framework [ASW^{+}15, ASS21, AmWHM99, BKK20, BKK23, BFS05, CP04, CHB06, CB19, DKB^{+}09, EWHS11, EHKT07, FJA^{+}18, GWYQ18, GHR20, JK12, KHH08, KKSP18, LFHAM19, MGL^{+}17, PG07, SHLJ17, SW16, SBC17, SJW22, TLSG05, TRL09, VFIN12, YWW^{+}19, ZGH^{+}15, ACD^{+}14, LP94]. Frameworks [Ano19, DX14, OP10, WTZ^{+}19, WTQ21]. Free [AR16, FLD15, LFD17, MARC24, PMM^{+}18, SMC94, Sun11, WTL^{+}23, IP90, Lan90]. Friendly [HZZ^{+}19, OOR13]. Frontiers [Ano16b, CTP13, FKT12, TFPF18]. FSI [HAA^{+}11]. FT [CFX^{+}20]. FT-2000 [CFX^{+}20]. Full [AK90a, GZJ18, MVD^{+}14]. Fully [LF15, SHS21]. Functional [ADC^{+}17, ACC^{+}01, AJF16, BARSW95, BFS05, GMP89, GS06, Hud86, KH18, Mat17, PC13, Gol88, Wai87]. Functions [ACC^{+}01, CFF^{+}06, DMC^{+}18, SNS21]. Fusion [EM14, Ken01, LZ17]. Fuzzy [GE90, KK20].

Galerkin [CF19]. Games [CYS16].

Garbage [Cra88, Fos89, LWLG11, AH86]. Gateway [AML^{+}10]. Gaussian [MVB^{+}06].

GCC [FKM^{+}11]. GCD [ABSSS19]. GCM [GHM14, MSPR18]. Gemini [OXL^{+}17]. Gene [AFO^{+}08, MSA^{+}07]. Gene/L [MSA^{+}07]. General [DDJ^{+}18, IP90, IH04, WP00, SS89]. General-Purpose [WP00]. Generalization [PMV17, WW17]. Generalized [GYL92, FcF87]. Generate [MGW99, BS89]. generate-and-test [BS89]. Generated [JCD^{+}14, MCA98]. Generating [AK17, ALTT17]. Generation [BTB^{+}13, BEJD21, CL96, Dar05, JW16, MPR^{+}05, QRW00, SR90, SSB^{+}17, TFEK16, qWlJzKhC17, WK20]. Generator [CPL^{+}10, EVK22]. Generic [BJM20a, GJK^{+}05, GW19, MAT23, MCT^{+}18, SM16, ZJL22]. Genetic [AMAH01, BM09, GKC22, MB12b, SO89]. Genome [BEA^{+}19, OOR13]. geometric [SS89]. Ghost [MS11, KTRZ^{+}17]. Girth [WS15]. Given [AK17]. Glacial [AW98]. GLE [DCX^{+}17]. GLE-Dedup [DCX^{+}17]. Global [AH86, LLSS03, PPQV16, RBES00, TAY^{+}12]. Globally [DCX^{+}17, TV15]. Globally-Locally [DCX^{+}17]. GMM [fS18]. Good [YBDJ17]. GOP [SSEA14]. GPGPU [AAB^{+}16, BCL14, BCL17, CBR17, STF^{+}12, YZ13, YHGW16]. GPGPUs [LMHW18]. GPI [HKJ^{+}18]. GPS [HVF18]. GPU [BC15, Boh23, BC10, CDDM18, CTB14, DK16, DMMP18, DMC^{+}18, FRT^{+}18, FJZ^{+}15, GLLH17, GGV17, GG13, Hen21, KLK16, LVJ22, LRG14, LTF^{+}12, LLW^{+}17, LEG11, LAD15, LFHAM19, Moh19, MGL^{+}17, NCR^{+}19, OOR13, OATGEL15a, PTdSF^{+}12, PHS19, PES^{+}18, RSA^{+}18, SI11, SF20, SLZB13, SJC18, SSB^{+}17, SBC17, SFAG14, SK14, VVCA23, WdSAM^{+}17, WR18, WE18, WK20, ZYOY13, ZHF^{+}19, ZD19, ZTY^{+}19]. GPU-Accelerated [DMC^{+}18, SBC17]. GPU-Based [DK16, VVCA23, BC10, OATGEL15a]. GPU-Friendly [OOR13]. GPUs [GL18, HLP11, JLDS16, KSBN22, KGK20, KPS14, LS20, MAWD^{+}16, MS11, MNN22, QGT^{+}19]. Grabbing [Sun11]. gradient [SDJS98]. GrADS [BCC^{+}05]. Grain [BG96, DV97, NRB94, NIO^{+}03, PSM97]. Grained [CTK^{+}11, CSF^{+}20, GL92, SDH22, SSM21, SZH18, WTQ21, WW17, AD89]. Grammar [MO91]. Grammars [PW92]. Granularity [PSM97, ZLC^{+}19]. Graph [ASS24, BCL90, CBR17, CSF^{+}20, CZTM03, GAR^{+}16, GWHY19, GP94, HKJ^{+}18, HSXH19, JK12, KSF^{+}18, KTF23, PS23, SHLJ17, SMDJ19, SSP^{+}96, Spr92, TH17, WZB^{+}92, ZHF^{+}19, GZ87, HKJ^{+}18]. Graph-Based [KTF23]. Graphical [RG15]. Graphics [CPP^{+}12, JGM15, SAB11]. Graphs [DV97, Hue97, KPRS96, KSTF24, LPF16, MXP14, OP10, OB13, PVF21, Zha89]. Graphs* [EKU22]. GraphTango [ASS24]. Greedy [AT91, Ken01, Sun11]. Grid [BFRPVR^{+}15, MMD21, SASH12, WL16, AFM^{+}06, BBC07, BCC^{+}05, SR04]. Grid-Based [WL16]. GridFOR [WL16]. Grids [HP13, LLL^{+}15, JS06b]. Gröbner [Sch92]. Group [KSA^{+}18]. Groups [BBC07]. GrPPI [BJM20a]. Guaranteed [MEP07]. Guarded [GYL92]. Guards [GYL92]. Guest [AG15, Bro19, CTP13, CAT18, DPT17, EA09, FKT12, GGE19, GK18, Gre16, HK14, HF06, HSXH19, JACK20, LFL^{+}17, LT17, MCE13, MGJS15, MGD^{+}14, MA10, OG11, PP10, PVG17, RJO22, SMM11, SGK12, SS10, TG21, TFPF18, ZZS^{+}19, Ano00a, Ayg03, AM07b, Ban04a, Ban04b, Car09, EmH97, FmH96, Fur95, GSA08, Gau96, GS05, Giv07, Giv08, HN94, JS06a, JS06b, Joe99, Joe03, Kes20, McK07, Mis09, NL23, Ora03, Pan08, Seh98, Vei01, Vei02]. Guided [MTT15]. GVirtuS [MGL^{+}17].

H [Roy10]. H-NMRU [Roy10]. Hadoop [LSM^{+}18, Mat17, NRGB17, RSA^{+}18]. Halo [PHS19]. Handle [ELGE16]. Handling [DFC^{+}07, FMSG17, HHW20, IR19, RBES00]. Hard [FJO^{+}16]. Hardware [AVM^{+}16, CHSC18, CPMC96, GP17, GV99, HZZ^{+}19, HSM^{+}24, HL21, KT01, KTBP18, Lys08, MSA^{+}07, NdMM09, NdMMW16, OBB^{+}24, OXL^{+}17, OPLS17, PMM^{+}18, RMH21, SWZ^{+}15, SD11, SH15, STM15, TRD21, WS14, YDV19, ZLAV04, vNR11]. Hardware-Agnostic [AVM^{+}16]. Hardware-Aware [HSM^{+}24]. Hardware-Based [CPMC96, KT01]. Hardware-Efficiency [KTBP18]. Hardware-Friendly [HZZ^{+}19]. Hardware-In-The-Loop [OBB^{+}24]. Hardware-Managed [RMH21]. Hardware-Supported [SD11]. Hardware/Software [GV99, Lys08, OPLS17, SWZ^{+}15, STM15]. HARE [JLDF19]. Harsh [GWPV21]. Hash [AR16, CHSC18, LFD17]. Health [AAN^{+}20, GKC22]. Healthcare [DC20]. Heap [GH96, LLM16, AH86]. Heap-Based [LLM16]. Heat [LYG^{+}18]. Height [ABASS12]. Helper [ZGH^{+}15]. Helping [Sun11]. Henderson [Swa88]. Heterogeneity [TMXS24]. Heterogeneity-Aware [TMXS24]. Heterogeneous [AER^{+}17, ANS20, Ano21a, AMKE18, ABB^{+}10, BEA^{+}19, Bro15, Bro19, BJM20b, ELK18, EAK21, EVK22, EGK23, GAG22, GGV18, GMB^{+}11, GHR20, HdMMK22, HK23, HDK24, HtBK^{+}10, HHC^{+}15, KTRZ^{+}17, LLGC17, LSYG15, LS05, MMN15, Mar17, MFGEL19, NCR^{+}19, OATGEL15b, OP12, OPLS17, PGLC^{+}18, PHS19, PVF21, SSM21, SEP08, WLL17, XWH21]. Heuristics [KPS14, CSG89]. HEVC [WdSAM^{+}17]. HICOR [GK94]. Hierarchical [Bro15, GP94, MV17, NN95, PG16, SSMO96, WSS18]. Hierarchically [PPEP08]. Hierarchies [GVB^{+}06]. Hierarchy [MCWK01]. High [APR^{+}18, Ano16a, Ano19, ASG20, BE14, BETK24, BCS^{+}09, BCL17, BS07, Bro15, Bro19, Car09, DPT17, DFH17, DB08, DST21, EAK21, GWYQ18, GGE19, GBLG10, Gha19, GK18, GJK^{+}05, Gre16, GHDF19, GE90, HG18, HK14, Jan15, KP05, KTRZ^{+}17, KJPN10, LPB13, LQWP10, LWP04, MB12a, dMMHdLN21, MSPR18, NFC^{+}09, NSU22, NdMM09, NL23, OXL^{+}17, PGLC^{+}18, SH96, SAL16, SCB^{+}14, SS23, TFEK16, TTF22, TGT18, WCC16, WMN^{+}17, WGW04, WK20, YZ13, YBRM14, ZLA21, Ano21a, Kes20]. High-Level [Ano16a, Ano19, BETK24, Bro15, Bro19, DPT17, EAK21, GGE19, GK18, Gre16, GHDF19, HG18, Jan15, KP05, LQWP10, dMMHdLN21, SH96, SS23, WMN^{+}17, HK14, TTF22, Kes20, Ano21a]. High-Performance [APR^{+}18, Ano19, GWYQ18, Gha19, GJK^{+}05, LPB13, MB12a, NdMM09, PGLC^{+}18, WCC16, WGW04, WK20, YBRM14, ZLA21, DST21, OXL^{+}17]. High-Productivity [BCS^{+}09]. High-Scalability [BS07]. higher [NPD89]. higher-order [NPD89]. Highly [TAY^{+}12, XZX^{+}15]. Highly-Scalable [TAY^{+}12]. Historical [TRD21]. History [BEA^{+}19, CEP97, JLDF19, LJ08, LLSS03, uRHH14]. History-Aware [JLDF19]. History-Based [BEA^{+}19]. Hitachi [TSB03]. HitFlow [FBGEL19]. HLFET [PIP18]. HLPGPU [Bro15]. HLPP [Ano16a]. Home [WLL^{+}08]. Homogeneous [MMN15]. Homomorphisms [LBT17, RG18]. horizontally [CB86]. Hotspotting [Ano86c]. HP [IPR^{+}05]. HPC [CAK17, CAT18, EAK21, HLK^{+}09, JQJ^{+}16, JQWG15, LLM^{+}12, LFL^{+}17, MAT23, NAS23, OBB^{+}24, YS22]. HSDC [DFZ21]. HW [KBG^{+}08]. Hybrid [ASS24, AOAM21, ADC^{+}17, BC15, CTB14, Cza17, DMMP18, EK14, FBGEL19, HSCI^{+}16, HBC23, JQJ^{+}16, LFL^{+}17, LRG14, MMD21, RY20, RY22, RRH03, SR15, VSH^{+}11, WHC^{+}24, YWW^{+}19, YLB19, ZLJ^{+}17]. Hybridization [DS20]. Hydrodynamics [Zey05]. Hypercube [CSG89, DPS90, GE89, NK88, Wai87]. Hypercubes [BB90]. HyperFatTree [SWF^{+}17]. Hypergraph [CND95]. Hypergraph-Based [CND95]. Hypersequential [UKT00]. Hyperspectral [CS20, LFHAM19]. Hyperthreading [HRH08].

I/O [AKT^{+}14, CSF^{+}20, MG15]. ICCG [IS03]. IDE [HLK^{+}09]. Identification [BR14a, FR95, OP12, PYC16, WQJY17]. Identifying [DM20]. Identity [JGP^{+}18]. IEEE [RJO22]. If [AmWHM99]. If-Conversion [AmWHM99]. iGridEdgeDrone [MMD21]. II [Fea92b, KR87]. ILP [SKA96]. Image [AM95, KBD03, RSK09, TRD21, YWY^{+}18]. Imagery [CS20]. Images [DPS90, KSA^{+}18, LFHAM19, SSB21]. Immune [MB12b]. Immune-based [MB12b]. Impact [BE14, KLG08]. Imperative [GM20, Jak19]. Imperfectly [AMP01]. Imperfectly-Nested [AMP01]. Implementation [AM95, AML^{+}10, CGJK95, CDDM18, DMMP18, ES11, GP17, GH89, HAA^{+}11, JSS^{+}15, JLMW15, KS97, LS91, LWP04, MXP14, NdMMW16, NSS12, OGP^{+}16, OXL^{+}17, PB01, PC13, RG18, RSV^{+}05, SM16, Sek09, SKG09, SY08, WLL^{+}08, WPC07, WS15, YZ13, ZQT20, ACD^{+}14, GTK^{+}88, TSS86, RK87]. Implementations [AJF16, BS07, BEG^{+}10, DE00, HPVRPF15, MWES19, Moh19, NdMCdMMW16, TSS99]. Implemented [MLdlP02]. Implementing [BAP01, Mil88, SPS14, SFAG14]. Implications [NP01]. Implicitly [AHKR01, LEA15]. Important [Ano86d, Ano92]. Improve [CHPC96]. Improved [EKU22, KSF^{+}18, LYL14]. Improved/Optimized [EKU22]. Improving [CHYP96, CEP97, GSY^{+}13, JLJ^{+}18, JHLM01, LWL^{+}19, MCWK01, PJS^{+}05, PMV17, RTD20, RSJ^{+}14, SBN03, SA10, XH98]. In-Depth [ZJL22, SJW22]. In-Loop [WdSAM^{+}17]. In-Memory [WTZ^{+}19, WTQ21]. Inaccuracy [JJIL15]. Incoherent [TGT18]. Incorporating [AK96]. Increased [KP04]. Increasing [HCEP98]. Incremental [CP04, XZX^{+}15]. Independent [EW96, FSS06, Ken94, SH96]. Index [GFL00]. Indexes [YJY16]. Induced [LG10]. Industrial [BR14a, FJO^{+}16]. Inference [PVF21, SHS21]. Inferential [RKG04]. InfiniBand [LWP04, QA11]. Infinite [FLMR02, KPRS96]. Information [AFM^{+}06, BE14, NRR99]. Infrared [YWY^{+}18]. Infrastructure [BML^{+}13, CEH13, EM13, SLZB13, UWF^{+}20]. Infrequent [ASG20]. Inheritance [Tho87]. Initial [AW98, HmWHR97, TKM89]. Initializing [Hem89]. Inlining [GYP22, LkCH94]. Input [SLZB13]. Input-Sensitivity [SLZB13]. Inspired [KPS14, Mis09, OGP^{+}16, CSCL20]. Instability [DKB^{+}09]. Instability-Estimation [DKB^{+}09]. Installation [CCG^{+}14]. Instruction [AHKR01, API03, BMA02, BR97, CSC^{+}00, CZTM03, HCEP98, JLDS16, LZ17, MP95, MSJ01, NN95, OVA04, RD08, SBN03, Tou05, TF94, VHK^{+}18, CMW^{+}94, NP98]. Instruction-level [NN95]. Instruction-Set [API03]. Instrumentation [AVM^{+}16, LSA^{+}07]. Integrated [CPL^{+}10, FRT^{+}18, GV99, KKZN12, MFU21]. Integrating [DTLW16]. Integration [GMP89, LLM^{+}12, PSM97, dMP^{+}03]. Integrity [KHT21]. Intel [BKT08, BP17, Cza17, RSJ^{+}19]. Intelligence [GKC22, dMMKdLN22]. Intelligent [MMD21]. Intel(R) [BGGT02]. Intensive [DDD^{+}19, LWLG11, RSJ^{+}14, LSM^{+}18]. Inter [GAR^{+}16, KTT^{+}99]. Inter-Node [GAR^{+}16]. Inter-Processor [KTT^{+}99]. Interaction [AHKR01, FJA^{+}18, GGV18].

Interactions [MHCF98]. Interactive [SJKA99]. Interchangeably [DJR16]. Interconnect [GBPK07]. Interconnection [MANR09]. Interconnects [RA09]. Interesting [VRGC19]. Interface [DGMP09, GZJ18, HKJ^{+}18, HTDL18, KBG^{+}08]. Interfaces [KKZN12]. Interference [CEP97, TMXS24]. Interference- [TMXS24]. Intermediate [CFB94, GP94, GBC^{+}08]. Internal [FWH^{+}94]. Internat [Swa88]. International [Ano21b, DB08, MCE13, PVG17, RJO22, SS10]. Internet [HZZS20, JACK20, KIT^{+}20, KAI20, MMD21, PYC16, PCJ18, SAI^{+}20]. Interpolation [DMC^{+}18, DMC^{+}20]. Interpreter [KGK20]. Interprocedural [CAZ02, CI96, GH96, HPY01, LkCH94]. Interprocess [CMW90, MO91, MO90]. Interprocessor [CH95]. Interruptible [TB23]. Interval [RWMF24, US05]. Intra [BGGT02]. Intra-Register [BGGT02]. IntraModule [MO91]. Introducing [SFAG14]. Introduction [Ano00a, Ano00b, Ano00c, Ano01, Ayg03, AM07a, AM07b, Ban94, Ban04a, Ban04b, Car09, CHS99, CmHS99, DB08, EmH97, EA09, Evr00, FmH96, Fur95, GSA08, Gau96, Giv07, Giv08, HmWHR97, HF06, JS06a, JS06b, Joe99, Joe03, LY98a, LY98b, McK07, MPZ06, Mis09, MA10, Ora03, Pan08, Pin95, Pin99, SMM11, Seh98, Vei01, Vei02]. Introspection [WHC^{+}17]. Introspection-Based [WHC^{+}17]. Intrusion [NSU22, NRGB17, YWW^{+}19]. intrusive [ZXY^{+}15]. Invalidate [BAP01]. Invasive [SR15]. invented [Par86b]. Inverse [fSxWC18]. Inversion [ABSSS19, KMG01, MMN15, SMM94]. Investigating [MWHS24]. Investigation [SdLC21]. IOT [GAK20, AAN^{+}20, AATD20, DMC^{+}20, FJA^{+}18, UWF^{+}20]. IP [AML^{+}10, CSD21, LSHK09]. IP-PBX [AML^{+}10]. IP-PBX/VoIP [AML^{+}10].

Irregular [ACC^{+}01, GF14, LLW^{+}17,

MCWK01, NST89, TB23]. ISA

[MP95, WCC16]. Isomorphic [Ano87d]. Issue [Ano16b, Ano18b, Ano18a, Ano19, Ano21a, AM07b, Bro19, Car09, DB08, GSA08, Gha19, Giv07, Giv08, HSXH19, JACK20, MCE13, MGJS15, MB12a, Mis09, ORJ24, Pan08, PP10, PVG17, RJO22, SS10, SZ17, TFPF18, WNMW16, ZZS^{+}19, JS06b, TG21, Ano21b, BmH98]. Issues [Bel94, NS97a]. Itemset [ASG20]. Iteration [HF14a, HF14b]. Iterative [MS11, PDN21, Rau96, ZHF^{+}19]. Iterator [GS11].

J [Swa88]. Jacobi [HOZ06]. Jacobians [BUMS02]. Java [AHKR01, FSS06, JQJ^{+}16, JMSG02, KF99, SS23, WGW04, WP00]. Job [LLL^{+}15, NSS12, WW17]. Join [RK92, RBR22]. Joint [HOZ06]. journal [Ano86b]. JPEG [SEP08]. Just [SA19].

kD [STF^{+}12]. kD-tree [STF^{+}12]. Kernel [LYG^{+}18, NLBB23, VVCA23, ZYOY13]. Kernelized [WCC16]. Kernels [KDV22, SSB^{+}17, WSO^{+}07]. Key [LKS^{+}20, PZL^{+}19]. Keyword [SNS21]. knapsack [LS92]. KNMF [LKS^{+}20]. Kutta [BP17].

L [MSA^{+}07]. Lab [ZC09]. Lab-on-Chip [ZC09]. Labeling [SH87, Swa88]. LACross [ZJG17]. Lagrangian [RSV^{+}05]. LALP [MCFM12]. LALR [BNWL90]. Landing [MSJ20]. Landslide [WSC20]. Language [ARB^{+}05, BARSW95, BCL17, CFB94, FCZ16, Fos89, GS06, Hud86, KS97, MCFM12, MPR^{+}05, SM09, TFEK16, WL16, WK20]. Languages [Ano19, CK02, FMSG17, Lan90, PS92, NPD89]. Laplace [CTB14]. Large [Cza17, GL18, HC17, HR11, HKJ^{+}18, KKZN12, LTSD15, LSA^{+}07, LWGZ18, SGJ^{+}03, SWF^{+}17, WW17, XZT20, ZWJK05]. Large-Scale [HC17, KKZN12, LWGZ18, SWF^{+}17, WW17]. Latency [AK96, Bos12, HZL16, JG97, LSHK09, MEP07]. Lattice [HLP11, SMN09, SKG09]. Launcher [NLBB23]. law [Ano87a, PM07]. layer [OATGEL15b]. Layered [Tic90]. Layout [SASH12]. Lazy [CRM17]. LCS [GSP^{+}17]. LDPC [SF20]. LEACH [KAI20]. Leaks [JGP^{+}18]. Learning [CR19, CLL21, CDDM18, DS16, FFS18, FKM^{+}11, HBC23, MAWD^{+}16, ÖA21, PVF21, WQZ^{+}24, ZJG17, ZD19, ZJL22]. Learning-Based [ZJG17]. Leases [CM06]. least [Ano86a]. Left [MP04]. Legacy [JBB21]. Legal [KP95]. Length [EM14, VHK^{+}18]. Lessons [Hal86]. Level [AG06, Ano16a, Ano19, BETK24, BCL17, Bro15, Bro19, DPT17, EAK21, GGE19, GBLG10, GK18, Gre16, GHDF19, HG18, HTDL18, Jan15, JF21, KP05, LLW^{+}17, LQWP10, dMMHdLN21, MHCF98, MKAP05, NL23, SSP^{+}00, SSEA14, SH96, SS23, SÜCV17, SMM94, SASH12, Tou05, WMN^{+}17, XOdFV^{+}09, YWW^{+}19, ZLJ^{+}17, BC10, HK14, NN95, TTF22, WS08, Kes20, Ano21a]. Levels [Gsc07]. Leveraging [LTL15]. LH [CS16]. Libraries [GJK^{+}05]. Library [BBR11a, LCF21, LAD15, LHP^{+}22, MFGEL19, SÜCV17, YKLD17, YBRM14]. Life [YYYX20, Ano87c]. Lifetime [SZH18]. Light [CM06]. Light-Weight [CM06]. Lightweight [GKC22, PZL^{+}19]. Like [NLBB23]. Limit [KEKK16, LS98]. Limited [JMSG02, uHKAMFM16a, uHKAMFM16b, GT86]. Limits [SS99]. Line [SR90, TFMP97, ZC09]. Linear [CCG^{+}14, CBR17, CJS21, DWS16, FLMR02, HKJ^{+}18, JLMW15, KS90, KFC08, KTRZ^{+}17, LDHL05, MP04, SMM94, Gao86]. Link [STB^{+}18]. Linked [HGT^{+}12, HTmG^{+}12, vdSGBW08]. Links [NIK00]. List [AF15, DS97, EM14, LBT17, SL14, vdSGBW08]. List-based [SL14]. Literature [IR19, dMMKdLN22]. Live [DST21, WHC^{+}17, ZXY^{+}15]. LLVM [RMG^{+}13]. Load [ASW^{+}15, BG96, EWHS11, JK03, MMD21, RLH14, RSJ^{+}14, YHGW16]. Load-Balance-Aware [YHGW16]. Load-Store [BG96]. Loads [AZK^{+}18]. Loads/Stores [AZK^{+}18]. Local [LLSS03, LYG^{+}18]. Locality [AMP01, AAB^{+}16, BE14, CAK17, JG97, KP01, LVJ22, LS98, LM00, PMHC03, Won02, XH98]. Locality-Aware [AAB^{+}16]. Localization [GWYQ18, OB13]. Locally [DCX^{+}17, SNB04, TV15]. Location [YFC21]. Location-based [YFC21]. Lock [AR16, MARC24, ZLD15]. Lock-Free [AR16, MARC24]. Locking [YLB19]. Log [Mar09]. Logic [AR16, AVPG00, KBD03, Lin91a, SAB11, BH87, Con88, Kas86, SRV88, Tin88]. Logic-Based [KBD03]. Logical [GZJ18, LWF^{+}19]. Look [MP04, NLBB23]. Loop [AMP01, CL96, DH00, GVB^{+}06, GMB95, GL95, HC17, IKN00, KDV22, LSL94, LCL17, NG92, OBB^{+}24, RAP95, WdSAM^{+}17, WMC98, YAI95, LP94]. Loops [Col95, GL95, MS11, MJ02, OGP^{+}16, QRW00, Sar01, TFNG09, WLL17, Wol86, YKM03, LAV98]. Loosely [LLM16]. Loss [AAN^{+}20, HZL16]. Lossless [HNC^{+}16]. Lossy [SAI^{+}20]. Low [Bos12, FVvL^{+}16, HZL16, NBN^{+}15, PO07, RSP20, Roy10, SAI^{+}20, SWF^{+}17, YZZ^{+}19]. Low-Latency [Bos12]. Low-Power [NBN^{+}15, PO07]. Low-Radix [SWF^{+}17]. LSA [UWF^{+}20]. LSH [RBR22, TRB^{+}24]. LSM [PYX17]. LSM-Tree [PYX17]. LTE [LF15].

M [FKD^{+}97, KHT21]. M-Machine

[FKD^{+}97]. Machine

[CHPC96, CZ12, CDDM18, DS16, FKD^{+}97, FKM^{+}11, GmWHR98, HHW10, HBC23, JQWG15, LVM16, MPR^{+}05, ÖA21, SHZ^{+}14, XLWX19, ZD19, Ali86, GS90, Ken94, PW92]. machine-independent [Ken94]. Machines [ABASS12, BJM20b, Den94, EGJS15, KTBP18, MGL^{+}17, NG92, ZWJK05, SDJS98]. Macro [GG14]. MAI [GN20]. Main [SZH18]. Mainstream [DMK21]. Maintained [SNB04]. maintaining [DPL86]. Malicious [CLJH16]. Managed [RMH21]. Management [ANS20, AGPGF14, CSF^{+}20, GJR09, HRH08, Joh94, MARC24, PHS19, SL14, SMH21, VFIN12, YJY16, ZLJ^{+}17, JK86]. Manager [BEA^{+}19]. Managers [Dem11]. Managing [ANS^{+}12, RNJ^{+}12, TFMP97]. Manipulator [BUMS02]. MANNA [HMT^{+}96]. Manual [NAP02]. Many [CTK^{+}11, CFC^{+}19, HG18, uHKAMFM16a, uHKAMFM16b, LHP^{+}17, LZ17, MFU21, NdMCdMMW16, OBB^{+}24, PMM^{+}18, PHS19, QZP15, SASH12, SA10, XWH21, vNR11]. Many-Core [CFC^{+}19, uHKAMFM16b, LHP^{+}17, LZ17, MFU21, OBB^{+}24, PMM^{+}18, SASH12, SA10, vNR11, NdMCdMMW16, XWH21]. Many-Cores [CTK^{+}11, HG18]. Many-Field [QZP15]. Many-Task [PHS19]. Manycore [HMF^{+}13, RSJ^{+}19, SMH21]. Map [FBV21, LFD17]. Map-Reduce [FBV21]. Mapping [CKC22, HtBK^{+}10, MEP07, RGB^{+}08, SDJS98, LRG^{+}91, NK88, PW87]. MapReduce [IR19, LSYG15, LHL^{+}16, LXL17, MM16, Mat17, RBR22, SHC15, VCP^{+}13, WW17, ZC17]. Massey [Moh19]. Massive [EKU22]. Massively [CYS16, HP13, KGK20, LTF^{+}12, WQZ^{+}24]. Matching [OOR13, RWMF24, Sca11]. MATLAB [MGW99, SM09]. Matrices [LPB13, LTSD15, SDL17, LP94]. Matrix [BBR11a, CFC^{+}19, CFX^{+}20, DZW10, JLV21, uHKAMFM16a, uHKAMFM16b, KJPN10, LHLT19, LKS^{+}20, MMN15, MGW99, SMM94]. Maximal [BCC00]. Maximizing [BG17]. Maximum [Gao86, KSBN22]. MC [GN20]. MC-DS-CDMA [GN20]. MCP [PIP18].

Mean [AK96]. Means [LKS^{+}20]. Measure [KKSP18]. Measurement [WSC20]. Measurements [JJIL15]. Measures [Cza17]. Mechanism [ANS20, CHYP96, EM14, FFS18, GMB06, Sek09, SHC15, WTZ^{+}19, WTQ21]. Mechanisms [GBPK07, Gen16, NAS23, MO90]. Media [LJ09]. Mediate [BBB^{+}17]. Medical [HZZS20]. Medium [DV97, NRB94]. Medium-Grain [NRB94]. Meld [AKD98]. Membership [KJHB14]. Memetic [NB15, ÖO07]. MemJam [MWES19]. Memories [AM04, LPB13]. Memory [ABSSS19, AF15, ANS^{+}12, BS03, BdS07, CCG^{+}14, CHCL14, Cra88, CRM17, DS97, DZW10, FBGEL19, GVB^{+}06, GRC^{+}14, GV99, GG13, HML^{+}20, ID08, JG97, Joh94, JMSG02, KTF23, KEKK16, LVJ22, LS20, LSL94, LMHW18, Lub90, MMG04, MCWK01, MBE03, MS99, MKAP05, NIK00, NAP02, OVA04, PZL^{+}19, PMM^{+}18, PO07, RC16, RWMF24, RRH03, SNB04, SMC94, SZH18, SD11, SAL16, SW16, SHC15, SWL05, SSMO96, SH15, SY08, SASH12, TTMD23, TMHT96, TA99, VSH^{+}11, WS14, WQJY17, WHC^{+}17, qWlJzKhC17, WTZ^{+}19, WTQ21, YZZ20, YBRM14, ZK07, ZLD15, ZLJ^{+}17, ZSH^{+}12, Con88, EO88, FcF87, GHLN86, GS90, GT86, Hem89]. Memory-Divergent [LVJ22]. Memory-Level [SASH12]. Memory-Optimized [LS20]. Merge [JK03, JLV21]. Mesh [DMC91, HAA^{+}11, SMSH13, SKAT91]. Mesh-Connected [DMC91]. MeshCleaner [MCT^{+}18]. Meshes [MCT^{+}18, qWlJzKhC17]. Message [BB90, CB01, EWHS11, GS05, GCD^{+}03, JGZ^{+}20, MFU21, GZ87, Hua89]. Message-Passing [CB01, GCD^{+}03, GZ87]. Meta [KPS14]. Meta-Heuristics [KPS14]. Metacomputing [ES06]. Metadata [AGPGF14]. Metagenomics [LSM^{+}18].

Method [BP17, DMMP18, Ger10, GRAG00, GHC^{+}17, IS03, LNP91, LEA15, NdMM09, PCJ20, RAP95, SMN09, WQZ^{+}24, ZYOY13, Wol86]. Methodology [KDV22, MOL05, RSJ^{+}14, UWF^{+}20]. Methods [BCC^{+}05, CCL12, CAK17, CJS21, MT96, MWHS24, RLEJ19]. Metropolis [CHB06]. Metrowerks [PB04]. MIC [FFS18]. Micro [JS06b]. Micro-grids [JS06b]. Microarchitectural [API03, DKB^{+}09]. Microarchitecture [PJS^{+}05]. Microbenchmarks [IPR^{+}05]. Microcode [BABW14]. Microfluidic [ZC09]. Microgrids [SS10]. Microphone [RLK20]. Microprocessor [LJE05]. microprogramming [CB86]. Microthread [BHJ06]. Migration [CML04, DST21, DLX^{+}17, JG97, NLRH07, PTdSF^{+}12, WHC^{+}17, XLWX19]. MILC [SKG09]. Milepost [FKM^{+}11]. MIMD [GL92, SDJS98]. Mini [ZXY^{+}15]. Mini-intrusive [ZXY^{+}15]. Miniature [NBN^{+}15]. Minimal [BTB^{+}13, DWS16, YAI95, Zha89]. minimax [NPT86]. Minimization [GLLH17, Mon97, PB04]. Minimizing [CH95, EDA96]. Mining [ASG20, CPP^{+}12, FJA^{+}18, HP13, OB13, PCJ18, WSS18, YWW^{+}19]. Mining-Based [OB13]. Mirroring [SDL17]. Mispredicted [JSHP97]. Mispredicted-Path [JSHP97]. Misprediction [NBD98]. Missing [DMC^{+}20, STB^{+}18]. Mitigating [JDF20]. Mixed [BEG^{+}10, SDJS98]. Mixed-Mode [BEG^{+}10, SDJS98]. Mixing [MRLR16]. ML [AGT17]. MLFQ [CLL21]. Mobile [ES06, JM20, YFC21]. Mobility [MMD21]. Mode [BEG^{+}10, OP12, YYYX20, SDJS98]. Model [AG06, AATD20, AK96, BEJD21, BAF94, BdS07, CND95, DMMS91, DTLW16, DFA^{+}09, FCZ16, FPCD14, FBGEL19, HBC23, HLP11, HKJ^{+}18, JM20, JF21, LLM16, LHL^{+}16, LCL19, Liv91, NAS23, OGP^{+}16, OATGEL15b, RSV^{+}05, RK13, fSxWC18, TAY^{+}12, TESK06, WSC20, YS22, ZJL22, JK86]. Model-Based [BEJD21, RK13]. Modeling [AA15, Ano18a, AMP^{+}05, BS07, HYBA18, KMjC02, LEA15, Mar17, MCE13, MGJS15, MOL05, PCP^{+}13, PVG17, Pra86, PS23, RJO22, SDH22, SSM21, TLSG05, WTL^{+}23]. Modelling [BKK20, BKK23, VNU19]. Models [BFS05, CAT18, Den94, FLMR17b, HHC^{+}15, ID08, KP05, Mat17, NAP02, RNJ^{+}12, SMSH13, SS01, Ski91, SDL17, VMS15, VCP^{+}13, AD86, DM87, FLMR17a]. Modern [HYBA18, KPS14, LG10, LQWP10, ME15]. Modifications [Hue97]. Modular [NdMM09]. Module [AAN^{+}20]. Modules [DJR16, SQH92]. Modulo [AG98, EDA96, GRAG00, LJ08, Rau96]. Modulo-Scheduled [GRAG00]. Molecular [ACC^{+}02, BS07]. Molecule [KLK16]. Moment [SSB21]. Monitor [LTL15]. Monitored [LJE05]. Monitoring [GAK20, NBN^{+}15, ZXY^{+}15]. Monoparametric [IAR21]. Monte [BJM20b, PES^{+}18]. Monte-Carlo [BJM20b, PES^{+}18]. more^{T M} [Ano87d]. MORPHEUS [GMB^{+}11]. Mosaic [MPAG18]. Motion [MVD^{+}14, TSS99]. Motivation [HmWHR97]. Movement [CFB94]. Moving [HAA^{+}11, ZQT20]. MPI [AJF16, BS07, BEG^{+}10, ES11, FPY08b, GJR09, GSY^{+}13, HMK09, LSM^{+}18, LWP04, MOL05, MANR09, NAS23, NSS12, RA09, SS01]. MPI/PVM [ES11]. MPJ [JQJ^{+}16]. MPSoC [ID08, OPLS17, RGB^{+}08, SWZ^{+}15]. MPSoCs [GHR20]. Much [MT96]. Multi [AOAM21, AH08, AKHD13, ABvK^{+}13, AML^{+}10, ABB^{+}10, BEJD21, BM09, CSF^{+}20, CZ12, CB19, CTB14, DS97, DS16, DTLW16, DJR16, FLD15, GM20, Ged13, GMB06, GGV17, GS06, HML^{+}20, HtBK^{+}10, JCH^{+}08, JDF20, KBG^{+}08, LYG^{+}18, MXP14, MV17, MG15, MHCF98, MFGEL19, NdMCdMMW16, OATGEL15b, PCJ20, QZP15, RPF18, RC16, RG18, RTD20, RD08, RK13, SSP^{+}00, SSEA14, SAI^{+}20, fSxWC18, SSB^{+}17, SFAG14, STB^{+}18, Sun11, VSDK09, WQJY17, WLL17, WSC20, WK20, XOdFV^{+}09, YWW^{+}19, Zha10, ZGH^{+}15, Ali86, AGT17, QGT^{+}19]. Multi-agent [STB^{+}18]. Multi-app [DJR16]. Multi-attitude [WSC20]. Multi-BSP [GM20, AGT17]. Multi-Component [fSxWC18]. Multi-Core [ABvK^{+}13, AML^{+}10, ABB^{+}10, GGV17, RPF18, SSEA14, Zha10, BEJD21, CZ12, Ged13, HML^{+}20, MXP14, NdMCdMMW16, QZP15, RC16]. Multi-cores [RTD20]. Multi-device [MFGEL19]. Multi-dimensional [RG18, WLL17]. Multi-domain [RK13]. Multi-external-storage [CSF^{+}20]. Multi-Fault [AKHD13]. Multi-GPU [CTB14, SFAG14, WK20]. Multi-GPUs [QGT^{+}19]. Multi-layer [OATGEL15b]. Multi-Level [MHCF98, SSP^{+}00, XOdFV^{+}09, YWW^{+}19]. Multi-ML [AGT17]. Multi-Orientation [LYG^{+}18]. Multi-path [JDF20]. Multi-Prefetcher [GMB06]. Multi-process [PCJ20]. Multi-process/Multi-thread [PCJ20]. Multi-Processor [HtBK^{+}10, BM09, KBG^{+}08, ZGH^{+}15]. Multi-processors [AH08, DS97]. Multi-queue [CSF^{+}20]. multi-sequential [Ali86]. Multi-sink [SAI^{+}20]. Multi-socket [RC16]. Multi-Stencil [CB19]. Multi-tenanted [WQJY17]. Multi-thread [PCJ20]. Multi-Threaded [MG15, VSDK09, DS16, GS06, RD08]. Multi-threading [DTLW16]. Multi-variable [MV17]. Multi-versioned [SSB^{+}17]. Multi-Word [FLD15, Sun11]. Multi-Zone [JCH^{+}08]. Multicluster [FCJV99]. Multicomputer [FKD^{+}97, Fos89]. Multicomputers [LNP91, SKAT91]. Multicore [AER^{+}17, Ano16d, CHCL14, HHW10, HMF^{+}13, KJHB14, LLM^{+}12, LLM16, RSJ^{+}19, SDH22, SS17, TKN^{+}08, WLL17, ZC17]. MulticoreBSP [YBRM14]. Multicores [TFNG09]. Multidimensional [Fea92b, LLM^{+}12]. Multigrid [MT96]. Multilevel [APR^{+}18, ADC^{+}17]. Multilisp [Hal86]. Multimedia [BG03, KL00, SG00, ZK07]. Multiplayer [CYS16]. Multiple [AZK^{+}18, ANS^{+}12, BDD^{+}18, CND95, CKC22, Gsc07, LEA15, PIP18, SQH92, TF94]. Multiple-Register-File [CND95]. Multiplication [Bos12, uHKAMFM16a, uHKAMFM16b, KJPN10, LHLT19]. Multiplications [CFC^{+}19, CFX^{+}20]. Multiply [BBR11a]. Multiprocessing [HML^{+}20, Bro86]. Multiprocessor [AK96, DeB87, Gol88, Gsc07, MB12b, Pan08, PPEP08, SEP08, SR04, BH87, GHLN86, GZ87, GTK^{+}88, Hua89, PD89]. Multiprocessor-based [Pan08]. Multiprocessors [AO19, BBGM95, GRV^{+}17, GV99, IPR^{+}05, KSEG14, KT01, LS07, LSL94, MVB^{+}06, NP01, OP12, SNB04, SMC94, SS01, TGT18, TESK06, ZLD15, Con88]. Multiscalar [LZ17]. Multisplitting [CCL12]. Multisplitting-Newton [CCL12]. Multitemporal [LFHAM19]. Multithreaded [FSS06, HTZ^{+}97, HMT^{+}96, KMjC02, LS07, MB99, OB13, WS08]. Multithreading [HTDL18, LEL^{+}99, TESK06]. MUSE [AK92, AK90a, AK90b]. Muzzle [KSA^{+}18]. MXNet [LWL^{+}19]. My [MFU21].

Nano [Mis09]. Nano/Bio [Mis09].

Nano/Bio-Inspired [Mis09]. Nanotube [CDC09]. Nanotube-Based [CDC09]. NaraView [SJKA99]. Native [JQJ^{+}16]. Nature [KPS14, MHCF98].

Nature-Inspired [KPS14]. Navigational [PLN^{+}04]. NC [PS92]. Near [BB90, SdLC21]. Near-Data [SdLC21]. Near-Optimal [BB90]. Nearest [LTF^{+}12, VVCA23]. Nebelung [MFG^{+}08]. Need [KT01, Kuc94]. Negative [DKB^{+}09, WS15, LKS^{+}20]. Neighbor [LTF^{+}12, PK20, VVCA23]. Nested [AMP01, EW96, MMS07, QRW00, Sar01, aMST07]. Nests [AMP01, GL95]. Net [LWDL17, GG14, GSS10]. Nets [KMjC02, LWF^{+}19, QGT^{+}19, RA94]. Netuno [SCB^{+}14]. Network [AOAM21, Ano18b, CPT14, DM20, DFZ21, FCZ16, FPCD14, GCD^{+}03, HZZ^{+}19, HLS15, HS16, HL21, JACK20, JDF20, KKZN12, LSHK09, LYL14, LSYG15, LXL17, Liv91, ML15, MANR09, MSPR18, NSU22, NRGB17, PG07, SAI^{+}20, SZ17, SWF^{+}17, SBN03, TG21, YMW^{+}17, ZZS^{+}19, AD86]. Network-Aware [FPCD14]. Network-Failure-Tolerant [GCD^{+}03]. Network-on-Chip [JDF20]. Networking [CSCL20]. Networks [AAN^{+}20, AKA^{+}20, AATD20, AK17, BS15, CLJH16, GWHY19, GKC22, HSM^{+}24, IBA11, JLDF19, KAI20, Li03, LCL19, LS05, LWGZ18, MVB^{+}06, MMD21, PMV17, RY20, RY22, SAI^{+}20, WZG^{+}17, YYYX20, YMW^{+}17, YZZ^{+}19, AD89]. Networks-on-Chip [JLDF19]. NetWorkSpace [BCS^{+}09]. Neural [AMAH01, AOAM21, FCZ16, GKC22, HZZ^{+}19, HSM^{+}24, LYL14, LXL17, LJ08, LWGZ18, PMV17, WZG^{+}17, YZZ^{+}19]. Neuromimetic [RNJ^{+}12]. Neuronal [CPP^{+}12]. Neutron [Zey05, SDJS98]. New-Age [DKB^{+}09]. News [FCZ16]. Newton [CCL12]. Next [Dar05]. Nighttime [fS18]. NMRU [Roy10]. no [Swa88]. NoC [LMHW18]. NoC-Side [LMHW18]. NoCs [MEP07, TOM^{+}11]. Node [GAR^{+}16, HZZS20, JF21, LJ09, PK20].

Node-Level [JF21]. Node-to-Node

[HZZS20]. Nodes

[BEA^{+}19, NBN^{+}15, TB23]. Non

[BG17, CSTGL03, EKU22, LKS^{+}20, Spr92, Con88, LP94]. Non-Blocking [EKU22, BG17]. Non-negative [LKS^{+}20]. Non-overlapping [Spr92]. non-shared [Con88]. non-singular [LP94]. Non-Strict [CSTGL03]. Noncoherent [BBGM95]. noncyclic [JB98]. Nonnegative [DZW10]. Nonsingular [OK99]. Normal [TG05]. Normalization [QGT^{+}19]. Note [Ano14, Ano16a, Ano16b, Ano16d, Ano16c, Ano18b, Ano18a, Ano19, Ano20, Ano21a, Ano21b, BKK23, Kes20, NL23, RY22]. Novel [AATD20, CSCL20, DMMS91, LKS^{+}20, OXL^{+}17, QFRA19, WWG^{+}19]. NUMA [BFG^{+}10]. Number [ALTT17, EVK22, HR11]. Numerical [EFED05, PES^{+}18, YKLD17, Zey05]. NVM [GZJ18].

O [AKT^{+}14, CSF^{+}20, MG15]. O2000 [CML04]. Obfuscator [FDY^{+}19]. Object [BBC07, DJR16, FMSG17, GS11, GS13, JM20]. Object-Oriented [GS11, GS13]. Objects [GK94]. Obtain [NRR99]. Obtaining [XZT20]. OCaml [SCS23]. occam [Cam89]. ODE [MLdlP02]. Off [ZK07]. Off-Chip [ZK07]. Offloading [JM20]. OFScheduler [LSYG15]. OLPCA [DMMP18]. oM [CLL21]. oM-DRL [CLL21]. OMP [SGJ^{+}03]. OMP2001 [TSB03]. On-Chip [GG13, KKZN12, MVB^{+}06, OBB^{+}24, AH08]. On-Line [ZC09]. On-Site [GWPV21]. On-the-Fly [JDF20, KSJ14]. One [Fea92a, SKG09, WW17]. One-dimensional [Fea92a]. Online [CLJH16, CYS16, HZL16, RC16, SMSH13]. onto [SDJS98]. Ontology [AFM^{+}06]. Open [AML^{+}10, SJW22, Cie91]. Open-Source [SJW22]. OpenCL [JSS^{+}15, RG18, SSB^{+}17]. OpenHMPP [AAB^{+}16]. OpenMP [AM07b, ABB^{+}10, BdS07, BGdS09, BFG^{+}10, BS07, BEG^{+}10, CF19, DFC^{+}07, DFA^{+}09, FMSG17, FM09, GSA08, HMK09, HAA^{+}11, JCH^{+}08, KaM10, KSJ14, MG15, MFG^{+}08, MBE03, MMS07, NIO^{+}03, OOS^{+}08, OP10, SSB21, WPC07, YKLD17, aMST07]. OpenMP/MPI [BEG^{+}10, HMK09]. OpenUH [CEH13]. Operating [CYS16, JGZ^{+}20, NP01]. Operation [FLD15, NB15]. Operational [Cam89]. operationally [DM87]. Operations [ABASS12, BG17, FPY08b, HDK24, IBA11, ML15, SZH18]. Operator [LCF21]. Operators [DM17]. Opportunistic [YMW^{+}17]. OPS5 [GTK^{+}88]. Optical [DMC91]. Optimal [AG98, BB90, CS20, DV97, DPS90, DPL86, GAR^{+}16, MA87, Mer86, NG92, SMM94, YKM03, ZLJ^{+}17, EG86, RB86]. optimality [Gai89]. Optimisation [GL18, PPEP08]. Optimised [Zha10]. Optimising [VNU19]. Optimization [AZK^{+}18, ALPS19, CFB94, CSCL20, CPMC96, CS97, CRM17, DLX^{+}17, GLLH17, GmWHR98, HBC23, HZZS20, HTmG^{+}12, JGZ^{+}20, KAI20, LDHL05, LM00, MO91, dMMHdLN21, NIO^{+}03, NdMCdMMW16, ÖO07, PCP^{+}13, RY20, RY22, RLH14, SRS06, SSEA14, Sca11, SA19, SHZ^{+}14, YHGW16]. Optimization-Based [SHZ^{+}14]. Optimizations [BKT08, BG96, ID08, KSEG14, LHLT19, LEL^{+}99, MV17, MS11, SB90, SLZB13]. Optimize [ZLAV04]. Optimized [EKU22, LF15, LS20, MGW99, Sar01]. Optimizer [LSYG15]. Optimizing [BBR11b, CGN^{+}09, CFC^{+}19, HDK24, uHKAMFM16b, MBE03, ZSH^{+}12, MO90, uHKAMFM16a]. Optimum [EDA96]. Option [Ger10]. OR- [SH96]. OR-Parallel [AK90b, Lin91a, Ali86, Cie91, Tin88]. OR-Parallelism [AK90a]. Orchestration [TMXS24]. Order [BS15, Boh23, BP17, CSD21, JDF20, MSJ01, NPD89]. Order-Aware [JDF20]. Ordering [IS03, DM87]. orders [Pra86]. OREGAMI [LRG^{+}91]. Organization [AM04]. Organizations [GWPV21]. Orientation [LYG^{+}18]. Oriented [ADC^{+}17, FMSG17, GS11, GS13, KKSP18, KK20, LVM16, RGB^{+}08, SRS06, SSM21, AKT^{+}14, CZ12]. Origin [IPR^{+}05]. Orthogonal [SSB21]. OS-Based [FC11]. OSD [AGPGF14]. Osmotic [FBV21]. Other [OP10, SS89]. Out-of-Core [SHLJ17, SMDJ19]. Out-of-Order [BS15, CSD21, MSJ01]. Output [CDRV98]. Output-Dependences [CDRV98]. Overhead [CTB14, KCW^{+}05, OPLS17, SJBV06]. Overheads [BGdS09, LJ08]. Overlap [BG17]. Overlapping [IKN00, Spr92]. Overview [BML^{+}13].

P [Zha10]. P-SURF [Zha10]. P2P [GJR09]. P2P-MPI [GJR09]. PAB [GMB06]. PAB-Based [GMB06]. Package [KKSP18]. Packet [DJR16, JDF20, QZP15]. Packing [JLMW15]. PAD [AG15]. Page [CML04, ZLJ^{+}17]. Page-Level [ZLJ^{+}17]. PageRank [EKU22, LEA15]. Panda [SBC17]. Para [Hud86]. Para-Functional [Hud86]. Paradigm [EW96]. Paradigms [DX14, Gen16]. ParaGraph [BCL90]. Parallel [AKBPV19, APR^{+}18, AMAH01, AM04, AK17, ACD^{+}16, ABvK^{+}13, AA15, Ano16a, Ano18b, Ano21a, AVPG00, AJF16, BR14a, Bel94, BAF94, BARSW95, BGMR11, BS03, BNWL90, BR14b, BETK24, BUMS02, BDD^{+}18, BDH^{+}14, Bro15, Bro19, BJM20b, CGN^{+}09, CPP^{+}12, CY14, CSD21, CB86, Cra88, CSTGL03, CDDM18, CAP88, Cza17, CPL^{+}10, Dam07, DPT17, DDD^{+}19, DMK21, DMMS91, DE00, DM17, DS97, DS16, Den94, DX14, DZW10, DGMP09, DSR17, ECSS88, EHKT07, EK14, EK17, ELK18, EVK22, EGK23, ES11, FFS18, FCRC16, GGE19, GBLG10, Ger10, GS11, GS13, GP17, GF14, GK18, GYL92, Gre16, GB20, GTK^{+}88, GKC22, HSCI^{+}16, HK14, HMF^{+}13, HP13, HPVRPF15, HLS15, HS16, Hun91, HAA^{+}11, IH04, Jan15, JW16, JLMW15, JK03, JLV21, Joh94, KS90, KK11, KS97, Kes20]. Parallel [KJHB14, KFC08, KGK20, KBG^{+}08, Kuc94, KR87, LMP98, LTF^{+}12, LYL14, LHL^{+}16, LT17, LLL^{+}15, LY95, LSL94, LWLG11, LHLT19, LBT17, Low00, LCL17, LYG^{+}18, Lub90, Lys08, MXP14, MMN15, MLdlP02, Mar09, MAJD16, MFC21, MM16, MG15, MCA98, dMMHdLN21, Mer86, Mil88, Moh19, MVD^{+}14, MFGEL19, NB15, NRGB17, NdMM09, NdMCdMMW16, NdMMW16, NSS12, NST89, NL23, OOR13, OP10, OGP^{+}16, OBB^{+}24, ÖA21, ÖO07, OG11, PW92, PGLC^{+}18, PLN^{+}04, PTD^{+}06, PVAE98, PMV17, PR99, PCJ18, QFRA19, RK92, RK87, Ric90, RTD20, RSV^{+}05, RMG^{+}13, RGB^{+}08, SGK12, SH87, SI11, SS92, SMN09, SMSH13, SQH92, SSM21, Sek09, SF20, SM09, SAS18, SO89, SKAT91, Ski91, SR90, SSB21, Spr92, SS89, SZ17, SC88, SHZ^{+}91, Swa88, TSS99, TG21, TRL09, VK88, VRGC19, WCC16, WL16, qWlJzKhC17]. Parallel [WR18, WS15, WZB^{+}92, WE18, YWY^{+}18, YH18, YS22, YBRM14, Zey05, Zha89, Zha10, ZZS^{+}19, ZQT20, ZWJK05, uRHH14, ACD^{+}14, BCL90, BCK98, Con88, DPL86, EG86, EO88, GN89, GZ87, GKMB87, Hua89, JGA^{+}88, JB98, Ken94, KMV87, KM86, LRG^{+}91, LS92, Par86a, Par86b, Par86c, TSS86, Wai87, WB87, AK90b, Lin91a, Ali86, Cie91, SRV88, Tin88]. Parallel-Access [Joh94]. parallel for [NCR^{+}19]. Parallelisation [KH18]. Parallelising [GS13]. Parallelism [AER^{+}17, ADC^{+}17, ACC^{+}01, BS03, BJM20a, DV97, EW96, GVB^{+}06, GGV18, GHDF19, Gsc07, GL92, HPY01, JBB21, JF21, KP04, LFL^{+}17, LS20, LHP^{+}22, MT96, MMS07, RSK09, SSEA14, SSNS16, SH96, SASH12, Tou05, WTL^{+}23, WS08, WW17, XOdFV^{+}09, BS89, CG94, Sch92, VR88, AK90a]. Parallelization [AAB^{+}16, BG17, BS07, CZ12, Col95, CAZ02, CF19, ELGE16, FLMR17b, FCRC16, FJO^{+}16, GK94, GYP22, GMS00, HML^{+}20, Hue97, IS03, JCD^{+}14, LQWP10, LXL17, dMMKdLN22, MVD^{+}14, NN95, PPQV16, RAP95, RLEJ19, SSP^{+}00, SHK13, SJKA99, SKA96, SR15, SNS21, TFNG09, TH17, WNMW16, WdSAM^{+}17, WP00, aMST07, FLMR17a]. Parallelize [MRLR16]. Parallelized [CR19, ELGE17, HTK98, KSTF24, TMHT96]. Parallelizing [CHCL14, GS11, KTT^{+}99, ME15, WZG^{+}17]. Parameter [BR14a]. Parameterized [LW97]. pareil [Lin91b, Lin86, Lin87, Lin89, Lin90, Lin88b]. Parlog [FT87, Hun91]. Parsers [BNWL90]. Parses [IP90]. Parsing [IP90, Lan90, PW92]. Part [JS06a, Fea92b, KR87, RK87]. Partial [AmWHM99, CKC22, DM87, GM20, RSP20, RWMF24, Pra86, SZH18]. Partial-Duplication [CKC22]. Partial-PreSET [SZH18]. Partially [SY08]. Particle [NdMCdMMW16, RLH14]. Partition [WLL17]. Partitioned [AT91]. Partitioning [CPG01, EW96, FCJV99, GAR^{+}16, Iqb91, KEKK16, LGY16, Lys08, MRLR16, NS97b, OPLS17, PDN21, PS23, SMN09, SWZ^{+}15, SHC15, TG05, VVCA23, GZ87, KMV87, NK88, PD89]. Partitioning-Aware [PS23]. Partitioning-Independent [EW96]. Partitions [DM20]. ParTriCluster [AFO^{+}08]. Pass [NS97b]. Passenger [RLK20]. Passing [CB01, EWHS11, GCD^{+}03, JGZ^{+}20, MFU21, GZ87, Hua89]. Path [AT91, CSC^{+}00, JAW17, JSHP97, LPF16, LJ08, OATGEL15a, SK97, SHZ^{+}91, JDF20]. Path-based [LJ08]. Pathfinder [JAW17]. Pattern [ACD^{+}16, BBR11a, CEP97, CPL^{+}10, DDD^{+}19, EVK22, GHC^{+}17, LYG^{+}18, QA11, TRB^{+}24, WSS18].

Pattern-Based [EVK22, BBR11a]. Patterns [ALG^{+}95, BETK24, BDD^{+}18, DMK21, DM17, DS16, FPY08b, JBB21, KSTF24, LLL^{+}15, RTD20, SHK13, YLB19, ACD^{+}14]. Patterns-Based [BDD^{+}18]. PBX/VoIP [AML^{+}10]. PCIe [OXL^{+}17]. PCM [SZH18, ZLJ^{+}17]. PCM-Based [SZH18]. PEMPIs [MOL05]. Per-Core [SA10]. percolating [ACD^{+}14]. perfect [GE89]. Performance [AM95, APR^{+}18, ASW^{+}15, AK92, AD86, Ano19, AKT^{+}14, BE14, BS07, BEG^{+}10, Car09, CHYP96, CHPC96, Cza17, DFH17, DB08, DCX^{+}17, DMC^{+}18, EGK23, GWYQ18, GGE19, Gha19, GJK^{+}05, GSY^{+}13, GKMB87, HRH08, HF14a, HF14b, HTmG^{+}12, JSS^{+}15, JLJ^{+}18, JCH^{+}08, KaM10, KTRZ^{+}17, KJPN10, LPB13, LPF16, Li03, LWL^{+}19, LY95, LWP04, LLSS03, LCL17, LWGZ18, MB12a, MCWK01, MS11, MOL05, MSPR18, MMS07, ME15, NFC^{+}09, NdMM09, NP01, PJS^{+}05, PGLC^{+}18, PVAE98, PS23, RTD20, RSJ^{+}14, SGJ^{+}03, SSEA14, Sca11, SAI^{+}20, SAL16, SCB^{+}14, SA10, TSB03, TFEK16, TKN^{+}08, Tin88, VCP^{+}13, WCC16, WGW04, WK20, YZ13, YBRM14, ZLA21, ZWJK05, ZJG17, dMP^{+}03, BCK98, DST21, OXL^{+}17]. Performance-Efficient [LWGZ18]. Performance-Portable [JSS^{+}15]. Performances [DFZ21]. Persistent [GW19]. Personal [HOZ06]. Personalized [LCT^{+}20]. Perspective [KBG^{+}08, WEJS94]. Perspectives [Ano16c]. Pessimistic [VSH^{+}11]. Petaflops [ACC^{+}02]. Petascale [TAY^{+}12]. PETRA [ME15]. Petri [KMjC02, LWDL17, RA94]. PGAS [JF21]. Phase [JHLM01, LGY16]. Phi [BP17, Cza17, ELGE17, LLGC17, PES^{+}18]. philosophers [RB86]. Phrase [LKS^{+}20]. Physical [WLW^{+}17]. Phytium [CFX^{+}20]. PIMP [MFU21]. Pin [JK12]. Pin-Based [JK12]. Pinning [CR19]. Pipeline [DF98, GG13, GRAG00, JBB21, LJ08, MFU21, SR04, Gai89]. Pipeline-Integrated [MFU21]. Pipelined [AD89, Low00, MJ02, NdMMW16, SWG^{+}18, LAV98]. Pipelining [BTB^{+}13, GRAG00, OGP^{+}16, RA94, YKM03, Gao86, WEJS94]. Piranha [CGJK95]. Pitfalls [HML^{+}20]. Placement [ANS^{+}12, DCX^{+}17, JQWG15, SHZ^{+}14, TMXS24]. Plain [SS23]. Plane [Mer86]. Planes [LYG^{+}18]. Planning [KCW^{+}05, LCT^{+}20, SI11]. PLASMA [YKLD17]. Platform [DTLW16, DZW10, ELGE17, FSS06, GMB^{+}11, LLW^{+}17, SSEA14, ZJG17]. Platform-Independent [FSS06]. Platforms [BC15, FRT^{+}18, Gha19, HMF^{+}13, MXP14, MMN15, MVD^{+}14, PGLC^{+}18, PVF21, RGB^{+}08, VFIN12]. pocl [JSS^{+}15]. Point [KSA^{+}18, LTF^{+}12, NST89, Ano86a, EG86]. Points [Mer86, SS92]. Polaris [FWH^{+}94]. Policies [BEP13, CML04]. Policy [Roy10]. Polka [Dav87]. Pollination [MSJ20]. Polling [Lin91a]. Pollutant [RSV^{+}05]. Pollution [MKAP05]. Polygons [SS92]. Polyhedra [LW97, QRW00]. Polyhedral [DV97, IAR21, JCD^{+}14, PCP^{+}13, SA19]. PolyJIT [SA19]. Polymorphic [CGPS18]. Polynomial [SWL05, ZYOY13]. Polynomial-Time [SWL05]. Pool [ACD^{+}16]. Pools [HR11]. Port [CND95, IBA11]. Portability [EGK23, KaM10]. Portable [EAK21, EVK22, JSS^{+}15, JF21, LS91, NLBB23]. Porting [YKLD17]. positive [GHLN86]. Post [NS97b]. Post-Pass [NS97b]. Potential [HML^{+}20]. Potentials [PDN21]. Potentials-Based [PDN21]. Power [AOAM21, AVLV03, GHR20, JS10, NBN^{+}15, OBB^{+}24, PO07, RSP20, SDH22, SWZ^{+}15, SAI^{+}20, WMN^{+}17, ZLJ^{+}17, ZJG17]. Power-Aware [AOAM21, AVLV03]. PR [KAI20]. Practical [CAP88, EKU22, GH96, KKKS24, Ski91, WZTH13, Hun87]. PRAM [GRR98]. Pre [DJS12]. Pre-Execution [DJS12]. Precision [KJPN10, ML15]. Predicated [CSC^{+}00, CHPC96, TF96]. predicates [FK87]. Predication [AmWHM99]. Predictability [SS99]. Predicting [ÖA21]. Prediction [AOAM21, CEP97, JSHP97, LEG11, MOL05, RWMF24, SK14, TF96, ZWJK05, ZJG17]. Predictive [PCP^{+}13]. Predictor [CHYP96]. Predictors [KMG01, LJ08]. Preface [CY14, WNMW16]. Prefetch [FDY^{+}19, HGT^{+}12, WLL^{+}08]. Prefetch-Based [WLL^{+}08]. Prefetch-Obfuscator [FDY^{+}19]. Prefetcher [GMB06]. Prefetching [CTK^{+}11, DJS12, GRV^{+}17, GV99, HGT^{+}12, HTmG^{+}12, ZGH^{+}15]. prefix [MA87, SS89]. Pregel [TH17]. Presence [JSHP97]. Preserving [DC20]. PreSET [SZH18]. pressure [LAV98]. Prevent [GMB95]. Price [Ger10]. Pricing [WWG^{+}19]. Primitive [JLV21, JHLM01]. Primitives [DeB87, JK86]. Priority [BEP13, LLM16, NYHA14, SS17, CRM92]. Priority-Based [NYHA14]. Privacy [DC20]. Private [JJIL15]. Probabilistic [LY95, LC11]. Problem [AT91, AVPG00, BR14b, DE00, Fea92b, MB12b, OATGEL15a, WS15, Fea92a, LS92, RB86]. Problems [HAA^{+}11, Iqb91, LSM^{+}18, LHP^{+}17, MFC21, Cie91]. procedural [JB98]. Procedure [KKMS99, SMM94]. Procedures [CK02, GMS00]. Process [FPY08b, LCL17, QA11, SSEA14, KMV87]. Process-Based [LCL17]. process/Multi [PCJ20]. Processes [EAT14, Mai87, PW87, RS90, Tho87]. Processing [AM95, APR^{+}18, CPP^{+}12, CY14, DJR16, GG14, GL18, GSS10, GB20, HZL16, HSXH19, JGP^{+}18, JGM15, KIT^{+}20, LT17, LAD15, Lys08, Mil88, NS97b, PTD^{+}06, PS23, RSK09, RSA^{+}18, RG15, SAB11, SHLJ17, SMDJ19, SBN03, TSS99, TRD21, TA99, VVCA23, WZB^{+}92, WW17, ZLA21, ZHF^{+}19, Ano87e, ECSS88, WB87]. Processor [BGMR11, BKT08, FCJV99, FVvL^{+}16, GWPV21, HtBK^{+}10, JHLM01, KBD03, KTT^{+}99, SMM94, TKN^{+}08, WSS18, WS08, WSO^{+}07, XWH21, BM09, KBG^{+}08, ZGH^{+}15, Sca11]. Processors [AOAM21, AK96, AMKE18, BG03, Cra88, Giv07, Giv08, GE90, HYBA18, KLG08, KL00, LG10, LZ17, MSJ01, OBB^{+}24, PG07, QZP15, RPF18, SDH22, SKA96, SA10, WLL17, Zha10, AH08, DS97, Hem89, MA87, PW87]. PrODACT [FDY^{+}19]. Product [MWHS24]. Productive [GHDF19]. Productivity [BCS^{+}09, BS07, Car09, KaM10]. Profile [CMW^{+}94, CPMC96, GAG22]. Profile-assisted [CMW^{+}94]. Profile-Based [GAG22]. Profile-Driven [CPMC96]. Profiling [CPMC96, LPF16, WTL^{+}23, ZD19, ZSH^{+}12]. Profiling-Free [WTL^{+}23]. Program [Dar05, KH18, KKMS99, MCFM12, SNB04, SLZB13, VNU19, CRM92]. Programmable [CDC09, Dam07]. Programming [AGT17, Ano16a, Ano21a, AVPG00, BBC07, BARSW95, BETK24, BCL14, BCL17, Bro19, BJM20b, CBR17, CAT18, DPT17, DMK21, DK16, DeB87, DX14, EK14, ELK18, EAK21, EVK22, EGK23, FBGEL19, GGE19, GMP89, GK18, GJK^{+}05, Gre16, GW19, GRR98, HSCI^{+}16, HG18, HK14, HKJ^{+}18, Hud86, KS97, Kes20, KBG^{+}08, LHL^{+}16, Lin91a, Lub90, MRLR16, MFGEL19, NL23, NAP02, PLN^{+}04, PVAE98, QFRA19, RPF18, SQH92, SS01, SS23, SFAG14, Swa88, UKT00, VRGC19, YS22, YBRM14, ACD^{+}14, BCL90, BCK98, Ken94, Par86a, Par86c, Tin88]. Programs [AR16, AJF16, BAF94, BS03, BDH^{+}14, CB01, CZ12, Dab21, DJR16, DSR17, EHKT07, FCRC16, FJO^{+}16, IAR21, Jak19, Jan15, JW16, JLMW15, KSJ14, LMP98, LWDL17, LBT17, Low00, MGW99, MOL05, MBE03, NS97b, OB13, SHK13, SJKA99, SK97, SO89, SCS23, VNU19, WP00, BS89, Con88, Gai89, Gol88, JB98, Kas86, SRV88]. Progressive [QGT^{+}19]. Project [BCC^{+}05, MAB^{+}11]. Projection [fSxWC18]. PROLOG [Ali86, AK90a, AK90b, Cie91, MARC24, SB90, SH96, TSS86]. PROMIS [SSP^{+}00]. Promoting [WLW^{+}17]. proof [FcF87]. Propagation [LMP98, LXL17, MXP14, SF20]. Properties [MAJD16]. Property [LWDL17]. Proposal [DFC^{+}07, DFA^{+}09]. Protein [FJZ^{+}15, KLK16]. Protocol [BAP01, DeB87, GSY^{+}13, MPAG18, PK20, RA09]. Protocol-Based [DeB87]. Protocols [BHL21, SB91, BCK98]. Provenance [AKA^{+}20]. Provide [SS17]. Proximity [LTL15]. Pruning [HSM^{+}24, WHC^{+}17]. Pseudo [EVK22, WQZ^{+}24]. Pseudo-Random [EVK22]. Pseudo-Synchronous [WQZ^{+}24]. Pseudosimulation [GT86]. PTAS [JLMW15]. Pthreads [JBB21]. pull [Par86c]. Pure [SNS21]. Purge [SAL16]. Purpose [WP00]. Push [RKG04, Par86c]. PVM [ES11]. PyACTS [DGMP09]. Python [DGMP09].

Q&A [GM20]. QCD [SKG09]. QoE [RY20, RY22]. QoS [AH08, BDD^{+}18, RY20, RY22, SS17, uRHH14]. QoS-supported [AH08]. Quadrant [PK20]. Quality [KK20, RLK20]. Quantifying [MHCF98]. Quantitative [LVJ22, LAV98, Sca11]. Quantum [PG16]. Query [STM15, VVCA23, WWG^{+}19]. Queue [BBB^{+}17, NSS12, WZTH13, ZLD15, CSF^{+}20, CRM92]. Queue-Based [ZLD15]. Queueing [RKG04, TB23, AD86]. Queues [GL92, LLM16, RMH21]. Queuing [WZTH13].

R [TRL09]. Race [KSJ14, MTT15, YLB19]. Radial [DMC^{+}18]. Radiation [LG10, Zey05]. Radiation-Induced [LG10]. Radio [vNR11]. Radios [KWA^{+}10]. Radix [SWF^{+}17]. Railway [FLMR02]. Random [AK17, EVK22, GAR^{+}16, YWY^{+}18]. Randomized [DS97, Li03, JGA^{+}88]. Ranking [DS97, uRHH14]. RANSAC [HPVRPF15]. Rapid [TCUV14]. Rate [HCEP98]. Ray [STF^{+}12]. Ray-Traversal [STF^{+}12]. RDMA [GSY^{+}13, JLJ^{+}18, LWL^{+}19, LWP04, RA09, ZLA21]. RDMA-Based [LWP04, ZLA21]. RDMA-Enabled [GSY^{+}13, RA09]. Reachability [WZB^{+}92]. Reaction [HF14a, HF14b]. Reactive [BHL21]. Read [DCX^{+}17, MV17]. Real [BEJD21, EWHS11, FJA^{+}18, FJO^{+}16, GAK20, LCT^{+}20, RSA^{+}18, YLB19]. Real-Time [BEJD21, FJA^{+}18, FJO^{+}16, LCT^{+}20, RSA^{+}18, EWHS11]. Real-World [YLB19]. Realistic [GZJ18]. Really [Kuc94]. Realm [PCJ18]. Rearrangement [SJBV06]. Reasoning [HVF18]. Recognition [KSA^{+}18, PR99, SS92, SHK13, WZG^{+}17]. Recognizing [PS92]. Recommendation [YFC21]. Reconfigurable [GMB^{+}11, GBC^{+}08, KBD03, NdMMW16, NBN^{+}15, PJS^{+}05, TKN^{+}08, ZC09, CB86]. Reconfiguration [RSP20, SA10]. Recovery [CJS21, JSHP97, LJ09, NBD98]. Rectangles [Spr92]. RECU [YBDJ17]. Recurrence [LM00, Gao86]. Recurrences [SKA96]. Recursions [uHKAMFM16a, uHKAMFM16b]. Recursive [GMS00]. Red [IS03]. Red-Black [IS03]. RedThreads [HTDL18]. Reduce [FBV21, MKAP05]. Reduced [ALTT17, DV97, MB12b, OOR13]. Reducing [CEP97, CK02, CTB14, FCJV99, TB23, ZK07]. Reduction [ABASS12, AVLV03, JS10, KCW^{+}05, LHF^{+}15, LJ08, ML15, PO07, SK97, SWL05, JK86]. Redundant [CH95, EAT14, GV95, HTDL18, KTT^{+}99].

Refactoring [AKBPV19, BDH^{+}14, BJM20b, BJM20a]. Referees [Lin92, Lin88b, Lin91b, Lin86, Lin87, Lin89, Lin90]. Reference [ALG^{+}95, RRH03, WGW04]. Reference-Set [WGW04]. References [Fea91, MKAP05]. Referencing [TMHT96]. Refinement [SMSH13]. Regeneration [MJ02]. Region [CI96, HmWHR97]. Region-Based [HmWHR97]. Register [ALTT17, AVLV03, BE14, BGGT02, CND95, EDA96, LkCH94, LHLT19, Tou05, ZLAV04, LAV98, NP98]. Register-Aware [LHLT19]. Register-File [ALTT17]. Registers [CGPS18]. Regular [HLP11, Sca11]. Regularization [ADC^{+}17]. Rehab [SAL16]. Reinforcement [CR19, CLL21]. Related [Lub90]. Relation [BKK20, BKK23, VK88]. Relational [MAWD^{+}16]. Relationship [LkCH94, fS18]. Relative [SB91]. Relaxation [CCL12]. Relaxed [HVF18]. Relaxing [AKD98]. Relevant [NAS23]. Reliability [BBR11b]. Remaining [XLWX19, YYYX20]. Remote [JG97]. Remoting [MGL^{+}17]. Removal [CDRV98, WS14]. Renaming [CSC^{+}00, TA99]. Rendering [BGMR11, SC88]. Rendezvous [CMW90, RA09]. Reordering [JDF20, KP95]. Reorderings [MCWK01]. Repair [SDL17]. RePhrase [DDD^{+}19]. Replacement [BEP13, Roy10, TFMP97]. Representation [ASS24, CFB94, FWH^{+}94, GP94, GBC^{+}08, KSA^{+}18, LYG^{+}18, WGW04]. Represented [SSB21]. Reproduction [Li03]. Request [DCX^{+}17, HZL16]. Request-Aware [DCX^{+}17]. Requirement [MSJ01]. Requirements [CMW90, EDA96, JSHP97]. Rescheduling [BCC^{+}05, CS97]. Reservations [SL14]. Resizing [KD15]. Resolution [Hue97, YWY^{+}18]. Resource [BEA^{+}19, BS91, CTK^{+}11, DS20, GB20, RY20, RY22, SHS21, ZLD15, JK86].

Resource-Aware [CTK^{+}11, GB20].

Resource-Constrained [SHS21]. Restart [CRM17]. Restarted [LEA15]. Restoration [JBB21]. Restoring [EGJS15]. Restructuring [KAMAMA17, MP04, PMHC03]. Results [AK92, AW98, GTK^{+}88, Hun87]. Retraction [BKK23, RY22]. Reusability [KKSP18]. Reuse [CSD21, SBN03, XH98]. Reuse-Driven [XH98]. Reverse [AmWHM99, BE14]. Reversible [PG16]. Review [IR19, Mar09, dMMKdLN22]. Reviewers [Nic14]. Revisited [PM07, Wol86]. Revisiting [KD15]. Right [MP04]. Rinda [Sek09]. RISC [OBB^{+}24]. RISC-V [OBB^{+}24]. RMI [SW16]. Road [ANS20, GGE19]. Robot [JGZ^{+}20]. Robotic [BUMS02]. Robust [LLSS03, Zha10]. Rochester [YBDJ17]. Rock [WSC20]. Rollback [DDJ^{+}18]. RollSec [DDJ^{+}18]. Route [SI11]. Router [JDF20]. Routing [BB90, IBA11, JLDF19, KAI20, LNP91, PK20, TOM^{+}11, YMW^{+}17]. RT [KAMAMA17]. RT-CUDA [KAMAMA17]. RTOS [KHT21]. RTOS-Aware [KHT21]. Ruby [Sek09]. Rule [KGK20, MAWD^{+}16, KM86]. Rule-Based [KGK20, KM86]. Ruleset [Sca11]. Run [BG17, ELGE17, HtBK^{+}10, RSP20, RAP95, TTF^{+}08]. Run-Time [RAP95, BG17, HtBK^{+}10, TTF^{+}08, vdSGBW08]. Runge [BP17]. Runtime [CEH13, ELGE16, FSS06, FRT^{+}18, FR95, FCRC16, GWYQ18, GRV^{+}17, GG14, HRC17, HL21, Hue97, KTF23, Low00, NAP02, PHS19, TAY^{+}12, TCUV14, XLWX19]. Runtime-Assisted [GRV^{+}17].

S [GG14, GSS10]. S-Net [GG14, GSS10]. SAC [GS06]. Safe [ELK18, LSL94]. Sage [Lin88a]. Samal [Swa88]. SAMOS [MCE13, PVG17, RJO22, ORJ24]. Sampling [SHC15]. Sans [Lin91b, Lin86, Lin87, Lin89, Lin90, Lin88b].

SARP [HZL16]. SAT [VSDK09]. Satisfaction [BBB^{+}17, LVM16]. Satisfaction-Oriented [LVM16]. Satisfiability [CPL^{+}10]. Saturation [Tou05]. SB [GRR98]. SB-PRAM [GRR98]. SBAC [AG15]. Scalability [ACC^{+}02, BS07, CFX^{+}20, Lys08, RSK09, RNJ^{+}12, SH15]. Scalable [AGPGF14, Bel94, DLRS13, EK14, Fea06, GWYQ18, GG14, GH89, HC17, IBA11, JQJ^{+}16, MPAG18, MARC24, QZP15, RC16, RAP95, RBR22, TAY^{+}12, Won02, XZX^{+}15, ZLD15]. Scalar [Bos12, Fea91, SK97, SW95, AK96]. Scale [HC17, HKJ^{+}18, KKZN12, LLGC17, LSA^{+}07, LWGZ18, MAT23, SWF^{+}17, WW17]. Scaled [CSD21, Mar17]. Scaling [KSBN22, MAJD16]. Scan [Mat17, KBD03]. Scenario [UKT00]. Scenario-Based [UKT00]. Schedule [EDA96]. Scheduled [CHPC96, GRAG00, PPEP08]. Scheduler [CLL21]. Schedulers [SY08]. Schedules [DF98, NST89]. Scheduling [ABASS12, AKD98, AO19, AK90a, AK92, AG98, AHKR01, AF15, BMA02, BR97, BBB^{+}17, BHJ06, BCC^{+}05, BM09, CGJK95, CSC^{+}00, CAK17, Cie91, CML04, EDA96, Fea92b, Fea06, FPCD14, GAG22, GRAG00, LLW^{+}17, LSL94, LMHW18, MSJ20, MP95, MB12b, NIO^{+}03, NBA13, PDN21, PIP18, Rau96, RD08, SMM94, TF94, WLL17, WMC98, YH18, YHGW16, CMW^{+}94, Fea92a, GN89, HC17, NP98]. Schema [WTZ^{+}19, WWG^{+}19]. Scheme [AKA^{+}20, FC11, GN20, GL18, GV99, GL95, KAI20, LGY16, Lin91a, LSL94, LJE05, NYHA14, SHZ^{+}14, WLL^{+}08, YHGW16, AD86, Mil88]. Scheme-Based [Mil88]. Schemes [ASW^{+}15, AMAH01, CKC22, HC17]. School [WMK19]. Science [HLK^{+}09]. Scientific [CAK17, DGMP09, HML^{+}20, IPR^{+}05, MV17, SSB^{+}17, TTF^{+}08, WSO^{+}07]. SCnC [SSNS16]. Scratchpad [CHCL14]. SDN [AAN^{+}20, FBV21, SAI^{+}20, UWF^{+}20].

SDRAM [LPB13]. Search

[BJM20b, DS20, GAR^{+}16, Ged13, Hun91, KS90, LY95, MAT23, MB12b, MVD^{+}14, WMK19, AD89, DPL86, KR87, RK87]. Searchable [AAI^{+}20a, AAI^{+}20b]. Searches [LTF^{+}12]. Second [SS10]. Section [Ano16d, Ano16c]. Secure [AKA^{+}20, DDJ^{+}18]. Security [AATD20, MSPR18]. Segmentation [LF15]. Segmented [JLV21]. Seismic [PTdSF^{+}12, Wai87]. Selected [KPS14]. Selecting [Low00]. Selection [CS20, DE00, GAR^{+}16, KDV22, SAS18, WTZ^{+}19, WTQ21, uRHH14]. Selective [KMG01, TFMP97]. Self [DWS16, EFED05, FKM^{+}11, HHW10, HC17, KFC08, LSL94, LJE05, NSS12]. Self-Adapting [EFED05]. Self-Monitored [LJE05]. Self-Scheduling [LSL94, HC17]. Self-stabilizing [DWS16]. Self-Submitting [NSS12]. Self-tuning [FKM^{+}11]. Self-verified [KFC08]. Semantic [HHC^{+}15, KSF^{+}18, LQWP10]. Semantic-Aware [LQWP10]. Semantics [ACC^{+}01, Cam89, Hud86, Ric90]. Semi [GVB^{+}06, KMV87]. Semi-Automatic [GVB^{+}06, KMV87]. Sensitivity [SLZB13]. Sensor [CPT14, DM20, NBN^{+}15, RY20, RY22]. Separation [SS92]. Sequence [LHP^{+}17, SO89, ECSS88, Hua89]. Sequences [AK17, FJZ^{+}15]. Sequential [FCRC16, LNG12, TFNG09, WNMW16, Ali86]. Serial [NIK00]. Series [DMC^{+}20]. Server [AFM^{+}06, CYS16, LJ09]. Servers [EAT14, NYHA14, RC16, WLW^{+}17]. Service [DWQ17, GAK20, LJ09, YFC21, uRHH14]. Services [HZL16, HHC^{+}15]. Set [API03, CZTM03, DDD^{+}19, GFL00, HCEP98, Mer86, SRS06, WGW04, XZT20, SZH18]. Sets [DWS16, FR95, LHF^{+}15, NRR99, SS92, EG86]. several [Hem89]. SFLA [DS20]. SGI [CML04, IPR^{+}05]. Shape [CAZ02]. Share [TV15]. Shared [BS03, BS91, CCG^{+}14, Cra88, FBGEL19, GV99, GG13, HML^{+}20, HR11, LSL94, Lub90, MMG04, MBE03, NIK00, NAP02, SNB04, SR15, SMC94, SS01, SS17, SSMO96, SY08, WQJY17, YBRM14, ZLD15, Con88, FcF87, GHLN86, Hem89]. Shared-address [HR11]. Shared-Memory [BS03, CCG^{+}14, FBGEL19, GV99, HML^{+}20, LSL94, NIK00, NAP02, SMC94, YBRM14, GHLN86]. Sharing [CML04, GMB95, SNB04, YBDJ17]. Shifting [DH00]. SHMEM [SS01]. Shortest [AT91, OATGEL15a]. Shortest-Path [AT91]. shuffle [GE89]. SIC [GN20]. Side [Gha19, LMHW18]. Side-Channel [Gha19]. Signal [FVvL^{+}16, NS97b]. Signals [vNR11]. Signed [GWHY19]. Significance [VCP^{+}16]. SIMD [GS90, KJHB14, Moh19, PES^{+}18, SJBV06, SDJS98, TB23]. Similarity [Cza17, Ged13, RBR22]. SimilarityJoin [TRB^{+}24]. Simple [CL96, WS08, LS91]. simplicial [EG86]. Simplify [MFGEL19]. Simplifying [MCA98]. Simulating [BH87]. Simulation [ABvK^{+}13, AA15, Ano18a, CSD21, Dem11, GHR20, KWA^{+}10, KP05, LJE05, MCE13, MGJS15, MANR09, PVG17, PPQV16, RJO22, SAB11, SJW22, TGT18, Zey05, ZWJK05, GT86]. Simulation-Based [ZWJK05, KWA^{+}10]. Simulations [ASW^{+}15, CGN^{+}09, GZJ18, HLP11, HF14a, HF14b, LLGC17, PES^{+}18]. Simulator [WPC07]. Simulators [MPR^{+}05, PC13, TCUV14]. Simultaneous [LEL^{+}99, PIP18, WS08, WE18]. Single [CB01, Dab21, EGK23, Fos89, HF14a, HF14b, PM07]. Single-Source [EGK23]. Single-Valuedness [Dab21]. Singular [BUMS02, LP94]. Sink [PK20, SAI^{+}20]. SISAL [AM95]. Site [GWPV21]. Size [KDV22, Low00]. SKA1 [FVvL^{+}16]. SKA1-Low [FVvL^{+}16]. Skeleton [DK16, DM17, EK14, ELK18, GRC^{+}14, GGV17, IH04, MFC21, RPF18, SFAG14, STB^{+}18]. Skeleton-Based [GGV17]. Skeleton-Driven [GRC^{+}14]. Skeletons [CPT14, DMK21, EM14, EK17, GK18, HdMMK22, HK23, JCD^{+}14, KH18, dMMHdLN21, SM16, WE18, WMK19, WK20]. SkePU [ELK18, EAK21]. SkeTo [EM14]. Skew [HHW20]. Skewing [Won02, Wol86]. Skewness [IR19]. Sliding [NdMM09, SF20]. Sliding-Window [NdMM09, SF20]. Slope [WSC20]. Slots [BMA02]. SLR [BNWL90]. SMA [LJE05]. Small [HZL16, HLP11, Sca11, SNS21]. Small-Ruleset [Sca11]. Small-World [HLP11]. Smart [DK16, KIT^{+}20, MMD21, UWF^{+}20, SJT13]. Smith [FJZ^{+}15, HMF^{+}13, RSJ^{+}19, TG05, ZTY^{+}19]. SMPs [BS03]. SMSG [WTL^{+}23]. SMT [KLG08]. Snow [TRL09]. SOC [LVM16, AML^{+}10, KHH08, KBG^{+}08]. Social [CLJH16, GWHY19, LCL19]. socket [RC16]. Soft [ÖA21]. Software [AVM^{+}16, BTB^{+}13, BKK20, BKK23, CSCL20, CFF^{+}06, DDJ^{+}18, Dar05, DJS12, EFED05, GV99, GRAG00, HYBA18, HTK98, JACK20, KWA^{+}10, KAMAMA17, KIT^{+}20, KVG18, KCW^{+}05, Lys08, MMG04, MJ02, MBE03, MSA^{+}07, NSU22, OXL^{+}17, OPLS17, RA94, RLPN^{+}02, RMH21, SWZ^{+}15, SAL16, STM15, YZZ20, YKM03, ZLAV04, dMP^{+}03, LAV98, WEJS94]. Software-Defined [CSCL20, RMH21]. Solid [YJY16]. Solution [GM20, GHLN86, KS90, RB86]. Solutions [BBGM95, BFRPVR^{+}15, Fea92b, RK92, SdLC21, Fea92a]. Solve [LSM^{+}18]. Solver [CF19, CTB14, KFC08, LDHL05, MLdlP02, MRLR16]. Solvers [CJS21, GLLH17, NLRH07]. Solving [AVPG00, VSDK09]. Some [Fea92a, Fea92b, VRGC19]. Sort [JK03, WR18]. Sorting [DMMS91, SKAT91, SJC18]. Sound [RLK20]. Soundness [LWF^{+}19]. Source [AML^{+}10, BML^{+}13, EGK23, GK94, HML^{+}20, PK20, SJW22]. Source-Code [GK94]. Source-to-Source [BML^{+}13, HML^{+}20]. Space [CML04, DWQ17, FVvL^{+}16, HPVRPF15, KS90, KWA^{+}10, LCU92, LWLG11, MB12b, MWHS24, PG07, PCP^{+}13, SS01, SWL05, TAY^{+}12, XZX^{+}15, EG86, Hua89]. Space-and-Time [LWLG11]. Space-Efficient [XZX^{+}15]. Space-Sharing [CML04]. Spaces [HR11]. spanning [Zha89]. Spark [GSP^{+}17, HHW20, LXL17, RSA^{+}18]. Sparse [BBR11a, CFC^{+}19, CFX^{+}20, HP13, JLV21, KTRZ^{+}17, KSA^{+}18, LTSD15, LDHL05, LHLT19, LWGZ18, GN89, GHLN86]. SparseNN [LWGZ18]. Spatial [CS20, DMC^{+}18, FLMR17a, FLMR17b, HtBK^{+}10, fS18, VVCA23]. SPAWN [PDN21]. SPEC [SGJ^{+}03, TSB03]. Special [Ano16b, Ano16d, Ano16c, Ano18b, Ano18a, Ano19, Ano21a, Ano21b, AM07b, Bro19, Car09, GSA08, Gha19, Giv07, Giv08, HSXH19, JACK20, JS06b, MCE13, MGJS15, MB12a, Mis09, NS97a, ORJ24, Pan08, PP10, PVG17, RJO22, SS10, SZ17, TG21, TFPF18, WNMW16, ZZS^{+}19, BmH98, DB08]. Specialization [FRT^{+}18, GW19]. Species [FJZ^{+}15]. Species-Based [FJZ^{+}15]. Specific [API03, CZTM03, CB19, TFEK16, TOM^{+}11, WL16, WK20]. Specification [BdS07, BS91, PC13, RA94]. specifications [Wai87]. Spectral [CS20]. Speculation [BS15, KVG18, WS08]. Speculative [AK92, CHPC96, Col95, ELGE16, JCD^{+}14, KLG08, KJHB14, KT01, LEG11, MS99, MKAP05, PPQV16, RKG04, RA09, TFNG09]. Speculatively [ELGE17]. Speculatively-Parallelized [ELGE17]. Speech [PR99]. Speed [GE90, MSPR18, NSU22, PMV17, TGT18, EG86]. speed-up [EG86]. Speeded [Zha10]. Speeded-Up [Zha10]. Speeding [SAB11]. Speedup [Gai89]. Speedups [KS90, GS90]. SPICE3 [WPC07]. Spike [CPP^{+}12]. Spill [PB04]. Spin [HLP11]. SpiNNaker [RNJ^{+}12]. Spline [AP86]. Split [WR18]. Splitting [GFL00]. SPMD [Dab21]. SPP [SSMO96]. SPP-1000 [SSMO96]. Spread [LEA15]. SQL [HHW20]. SR8000 [TSB03]. SSD [OXL^{+}17]. stabilizing [DWS16]. Stack [BEP13]. Stacked [LHP^{+}17]. Stage [EDA96, PYC16]. Stand [DJR16]. Stand-Alone [DJR16]. Standard [FSS06, SÜCV17, YKLD17, NdMMW16]. Standard-Library [SÜCV17]. StarCore [PB04]. State [BR97, KS90, KPS14, LHL^{+}16, OOR13, YJY16]. State-of-the-Art [LHL^{+}16]. State-Space [KS90]. Stateful [ACC^{+}01, DM17]. States [DDJ^{+}18]. Static [BCC00, CB01, CSD21, HYBA18, Li03, MRLR16, NIO^{+}03, RRH03, Gao86]. Statically [BCL17]. Statistical [AAI^{+}20a, AAI^{+}20b, NSU22, PYC16]. Status [Ano16c]. Steal [TV15]. Stealing [HHW20, YH18]. Steiner [BR14b, MNN22]. Stencil [CB19, HdMMK22, HDK24, MS11, SBC17]. Stencil-Operations [HDK24]. Stiff [MLdlP02]. STL [HG18]. Stochastic [ASW^{+}15, RSV^{+}05]. Storage [AMAH01, CM06, JSHP97, LT17, NG92, WTZ^{+}19, WTQ21, AH86, CSF^{+}20]. Storage-Centric [CM06]. Store [BG96]. Stores [AZK^{+}18]. Storm [ZLA21]. Story [MSA^{+}07]. Straightforward [MCT^{+}18]. Strassen [uHKAMFM16a, uHKAMFM16b]. Strategies [CGJK95, CF19, FLMR17a, FLMR17b, LJ09, PK20, SAS18]. Strategy [GSP^{+}17, IS03, JM20, RBES00, WLL^{+}08, WR18, ZLJ^{+}17]. Stream [GSS10, GHDF19, GHR20, LHP^{+}22, RSA^{+}18, RGB^{+}08, TF94, ZK07, ZLA21, SRV88]. Stream-Conscious [ZK07]. Stream-Oriented [RGB^{+}08].

Streaming [ASS24, BBR11b, CHCL14, HtBK^{+}10, LJ09, MAB^{+}11, SSNS16, TB23, VNU19]. Streams [CPP^{+}12, DM17, Tic90]. Strict [CSTGL03]. Structural [AMP^{+}05]. Structure [EFED05, LWDL17, MGW99]. Structured [BABW14, Fea06, GGV18, HCEP98, MV17, MP95, NLRH07, SASH12]. Structures [BCL17, CL96, ELGE16, GL18, HGT^{+}12, HTmG^{+}12, JSHP97, RG15, SL14, SH15, vdSGBW08]. Student [FJA^{+}18]. studies [CG94]. Study [BKT08, DE00, FPY08b, HPVRPF15, HMT^{+}96, LVJ22, LDHL05, MS11, PMV17, Sca11, SPS14, KM86]. Styles [PC13]. Sub [LS05]. Sub-Networks [LS05]. Subdivision [BGMR11]. Subgroup [FG16]. Submission [LLL^{+}15]. Submitting [NSS12]. Subroutines [CCG^{+}14]. subscribers [Ano92]. Suitable [MVB^{+}06]. Suites [SGJ^{+}03]. Summarization [KSF^{+}18]. Summation [ML15]. sums [MA87]. Super [AK96, JLDS16, YWY^{+}18]. Super-Resolution [YWY^{+}18]. Super-scalar [AK96]. Supercomputer [MSA^{+}07]. Supercomputers [SBC17, qWlJzKhC17]. Supernode [SPS14]. Superscalar [MSJ01, VMS15]. Superthreaded [TJY99]. Support [EK17, GRR98, KTBP18, PB01, WGF^{+}16, BCL90]. Supported [SD11, AH08]. Supporting [BHJ06, CYS16, FMSG17, MMS07, OOS^{+}08, SQH92]. SURF [Zha10]. Surfaces [AP86, BGMR11, DMC^{+}18, PGLC^{+}18, SC88]. Surrogate [JM20]. Surveillance [fS18, WZG^{+}17]. Survey [BR14b, KPS14, LHL^{+}16, LMPS05, SJC18, YH18, YS22, Cie91]. Survive [ABB^{+}10]. SVM [CSCL20]. SW [KBG^{+}08]. Swap [FLD15, Sun11]. Swarm [dMMKdLN22, NdMCdMMW16, RLH14]. SWIMM [RSJ^{+}19]. Switched [FPY08a]. Switches [SWF^{+}17]. Symbol [MARC24]. Symbolic [ACD^{+}16, CFF^{+}06, KP05, MP04, GKMB87]. Symmetric [GMP89]. Symposium [Ano21b, DB08]. Synchronisation [BHJ06, FG16]. Synchronization [DSR17, GH89, GE90, HTK98, Jan15, JHLM01, KKZN12, Liv91, Lub90, NP01, AD86, HFM88, MO90]. Synchronizations [CH95]. Synchronous [BS15, WQZ^{+}24]. Synopsis [HZL16]. Synthesis [PG16]. Synthesizing [AMP01, AMAH01, AJF16]. System [AG06, AA15, BBB^{+}17, BCS^{+}09, BC10, CYS16, Cza17, DLX^{+}17, EFED05, GG14, GZJ18, GCD^{+}03, HSCI^{+}16, HMT^{+}96, JGZ^{+}20, KTF23, KFC08, Mil88, MMS07, NIK00, NSU22, NP01, RSP20, RLH14, RNJ^{+}12, SGJ^{+}03, SJKA99, SSMO96, SH15, TLSG05, TTF^{+}08, TRL09, WGF^{+}16, XZX^{+}15, ZLC^{+}19, ZXY^{+}15, KM86, Tin88, KK11]. System-level [BC10]. Systematic [IR19, TH17]. SystemC [BFS05, CSD21]. Systems [Ano16c, Ano18a, Ano21a, Ano21b, AF15, AMP^{+}05, ANS^{+}12, BAP01, BETK24, Bro15, Bro19, CHB06, CS97, CAT18, DK16, DLRS13, EWHS11, ELK18, EAK21, EVK22, EGK23, FLMR02, FBV21, FPCD14, FBGEL19, FJO^{+}16, GWYQ18, HC17, HRH08, HtBK^{+}10, HSXH19, HLK^{+}09, KKSP18, KK20, KTRZ^{+}17, Kuc94, LLM^{+}12, LFL^{+}17, LSA^{+}07, LMPS05, MP91, Mar17, MCE13, MGJS15, MBE03, Pan08, PP10, PB01, PM07, PVG17, PO07, PPEP08, RK92, RWMF24, RJO22, SGK12, SSM21, SEP08, SS17, SFAG14, STB^{+}18, TSS99, TKN^{+}08, US05, WS14, WLL^{+}08, ZC17, AH86, Cie91, Dav87, GHLN86, Par86b, PD89, PW87]. Systolic [AP86, Ano87e, IP90, Lan90].

T [Swa88]. Table

[CEP97, CHSC18, MARC24, OOR13].

Tabled [AR16]. Tackling

[DFH17, SLZB13]. Tag [PO07, VFIN12]. Tampering [ABSSS19]. Target [fSxWC18]. Targets [ZQT20]. Task [BM09, CKC22, FPCD14, FCRC16, GN89, GS13, GP94, HHW20, HR11, JM20, KSTF24, LPF16, LLW^{+}17, MB12b, NIO^{+}03, OP10, OGP^{+}16, PHS19, RLH14, SSNS16, TFEK16, WTQ21]. Task-Aware [WTQ21]. Task-Based [FCRC16, RLH14, TFEK16]. Tasking [DFA^{+}09, KaM10]. Tasks [BC10, DFA^{+}09, HR11, PS23]. TAU [MMS07]. TCP [LSHK09]. TCP/IP [LSHK09]. Technique [AKD98, AOAM21, CPMC96, Hue97, HAA^{+}11, KTT^{+}99, PB04, RGB^{+}08, SR04, TOM^{+}11, WLWZ15]. Technique-Application [PB04]. Techniques [AK96, CAZ02, DS16, GBLG10, KL00, KP04, LY95, NP19, SRS06, STF^{+}12, SK97, TAY^{+}12, TJY99, ZLAV04]. Technologies [MAB^{+}11]. Technology [GKC22, JACK20, Ken94]. Telegraphic [ES11]. Telescoping [CK02]. Temperature [DKB^{+}09]. Template [GF14, LCF21, LHP^{+}22, NCR^{+}19]. Temporal [PMHC03, YYYX20]. tenanted [WQJY17]. Tensor [KKKS24, Boh23]. TensorFlow [JLJ^{+}18]. Teradevice [WGF^{+}16]. Terascale [GCD^{+}03]. termination [Tho87]. Test [CPL^{+}10, KJHB14, SRS06, BS89]. Testing [TCUV14, ZC09, Mai87]. Tests [JW16]. Text [FCZ16, KSF^{+}18, LYL14]. Textually [Dab21]. TFlux [DTLW16]. Their [CGJK95, LW97, RG18, ACC^{+}01]. Theory [GRAG00, RSJ^{+}14, CP88]. Thermal [OBB^{+}24]. Things [HZZS20, JACK20, KIT^{+}20, KAI20, MMD21, PCJ18, SAI^{+}20]. Thread [AO19, AMKE18, CPL^{+}10, DSR17, JG97, KGK20, ZGH^{+}15, PCJ20, WS08]. Thread-level [WS08]. Thread-Parallel [CPL^{+}10]. Threaded [HGT^{+}12, HTmG^{+}12, MG15, VSDK09, DS16, GS06, RD08]. threading [DTLW16]. Three [ABASS12, HDK24]. Three-Argument [ABASS12]. Three-Dimensional [HDK24]. Threshold [CLL21]. Throughput [AKT^{+}14, BBR11b].

Throughput-oriented [AKT^{+}14]. Thrown [AHKR01]. TIDeFlow [OGP^{+}16]. Tightly [RWMF24, SS01]. Tightly-Coupled [SS01]. Tikhonov [ADC^{+}17]. Tile [KDV22, RMH21]. Tile-Based [RMH21]. Tiled [FC11, OOR13, SMH21]. Tiling [IAR21, MHCF98, XH98, ZK07]. Time [BBB^{+}17, BEJD21, DWS16, DMC^{+}20, FJA^{+}18, FCJV99, Fea92b, FJO^{+}16, GAK20, KCW^{+}05, LCU92, LLL^{+}15, LWLG11, LCT^{+}20, MWES19, PTdSF^{+}12, RSP20, RSA^{+}18, RAP95, RK13, SWZ^{+}15, SA19, SWL05, Won02, YFC21, YKM03, BG17, EWHS11, Fea92a, HtBK^{+}10, TTF^{+}08, vdSGBW08]. Time-aware [YFC21]. Timed [GHR20]. Timed-Value [GHR20]. Times [SB91]. Timing [FDY^{+}19, GHR20, MP91, WQJY17, WMN^{+}17, YDV19]. TINPAR [KTT^{+}99]. Tissue [LLGC17]. Tissue-Scale [LLGC17]. TLB [JS10, VFIN12]. TM [SÜCV17]. TM-Based [SÜCV17]. TMT [VFIN12]. Tokenization [Sca11]. Tolerance [AKHD13, NRR99, WGF^{+}16, ZLJA12]. Tolerant [DFZ21, EAT14, GCD^{+}03]. Tolerating [AK96, JG97, LG10]. Too [CHSC18, MT96]. Tool [DST21, FG16, KAMAMA17, KSJ14, ME15, PDN21, PVAE98, WMN^{+}17]. Tools [ALG^{+}95, ARB^{+}05, DGMP09, LRG^{+}91, Lub90, CB86]. Top [Sca11]. Top-Performance [Sca11]. Topological [GE89]. Topologies [MVB^{+}06]. Torus [IBA11]. Trace [Mai87, RLPN^{+}02, RLEJ19, RD08]. Trace-based [RD08]. Traces [MANR09]. Tracing [BEJD21, ZD19]. Traffic [ANS20, CLL21, GAK20, PYC16]. Train [KKKS24]. Training [JCW^{+}18, LYL14, QGT^{+}19, WTL^{+}23]. Transaction [AA15, NBA13]. Transaction-Based [AA15]. Transactional [CRM17, GRC^{+}14, MFG^{+}08, PMM^{+}18, SAL16, SW16, SH15, VSH^{+}11, WS14, YZZ20, ZSH^{+}12]. Transactions [CHSC18, DTLW16, SD11]. Transfer [SR04]. Transfers [ALPS19]. Transform [BC15, DLRS13]. Transformation [HSCI^{+}16, IKN00, KH18, fSxWC18, SASH12, vdSGBW08, LP94]. Transformations [AG06, AMP01, GVB^{+}06, GMB95, HRC17, JS10, KP95, KP01, MO90, OK99, SPS14, TH17, VNU19, WMC98, YAI95]. transformed [Ano86b]. Transforming [BS89, JBB21]. Transient [LG10]. Transition [OOR13]. Transitive [CAP88, KPRS96, VK88]. Translator [ABvK^{+}13]. Translators [KCW^{+}05]. Transparent [PSM97, PPQV16, YZZ20]. Transport [CJA00, VHK^{+}18, Zey05]. Transpose [LPB13]. Transpositions [JGM15]. Travel [LCT^{+}20]. Traversal [STF^{+}12]. Traversing [ZHF^{+}19]. Traversing-Based [ZHF^{+}19]. Tree [BR14b, BJM20b, GH89, KF99, MM16, PS92, PW92, SM16, SMC94, SWF^{+}17, WR18, YJY16, DPL86, MA87, STF^{+}12, PYX17]. Tree-Based [KF99, SWF^{+}17]. Trees [Li03, MNN22, Zha89]. Triangular [MMN15]. Triangulating [Mer86, EG86]. Trie [AR16]. Triggered [CJA00, VHK^{+}18]. Trin [JK12]. Triple [DC20]. True [BAF94]. TrustZone [KHT21]. TuCCompi [OATGEL15b]. Tuned [LAD15]. Tuning [BG17, CCG^{+}14, LEL^{+}99, OATGEL15b, FKM^{+}11, Ged13]. Tunnels [KLK16]. Two [BARSW95, EHKT07, FJO^{+}16, HFM88, JHLM01, LPB13, LLW^{+}17, LS05, SS92]. Two-Dimensional [BARSW95, EHKT07, LPB13]. Two-Level [LLW^{+}17]. Two-Phase [JHLM01]. Type [CP88, ELK18, HZZ^{+}19, VNU19]. Type-Driven [VNU19]. Type-Safe [ELK18]. Typed [BBC07, BCL17]. types [Win89]. TZmCFI [KHT21].

UAV [SI11]. Uintah [PHS19, dMP^{+}03].

Ultra [NSU22]. Ultra-High-Speed [NSU22]. Unbalanced [MFC21, OP10]. Uncover [WS08]. Understanding [STF^{+}12]. Unequal [YBDJ17]. Unicast [DFZ21]. UniCNN [SWG^{+}18]. Unification [SSNS16, CRM92]. Unified [DLRS13, HPY01, RK13]. Uniformed [SWG^{+}18]. Uninterpreted [CFF^{+}06]. Union [CAP88]. Union-Find [CAP88]. Unit [JW16]. UniTi [RK13]. Units [CPP^{+}12, JGM15, RG15, SAB11]. Universal [GP94]. Unroll [BTB^{+}13]. Unrolling [Sar01]. Unstructured [qWlJzKhC17]. Update [SMDJ19, WQZ^{+}24]. Updates [ASS24]. updating [Hun87]. Upon [GL92]. Urban [fS18]. URLs [CLJH16]. URSA [PVAE98]. Use [GmWHR98]. Useful [YYYX20]. User [LLL^{+}15, MTT15]. User-Guided [MTT15]. Users [AKA^{+}20, BBB^{+}17, Kuc94]. Using [AKBPV19, APR^{+}18, AAN^{+}20, ASG20, BR97, BKK20, BKK23, BEJD21, BAF94, BABW14, BJM20b, CHPC96, CPT14, CS20, Col95, CFF^{+}06, DeB87, Dem11, DS16, DTLW16, DMC^{+}18, DMC^{+}20, DJR16, DS20, ELGE17, FFS18, GAR^{+}16, GG14, GK94, GN20, GG13, GRAG00, GAK20, GH89, GE90, HG18, HLP11, HHW20, HP13, ID08, JM20, JG97, JCD^{+}14, Joh94, KH18, KKSP18, KMjC02, KHT21, KP95, uHKAMFM16a, uHKAMFM16b, KAI20, KP05, KGK20, LPB13, LLGC17, LQWP10, LS05, LNG12, LEA15, LM00, LBT17, LKS^{+}20, MSJ20, MCWK01, MANR09, MKAP05, NIK00, NIO^{+}03, NRR99, NBA13, ÖA21, PLN^{+}04, PPQV16, PMV17, RA94, RY20, RY22, RLH14, RSA^{+}18, RSJ^{+}14, RLK20, SDH22, SSEA14, SAB11, fS18, SSB21, Sun11, TSB03, TCUV14, TFMP97, VVCA23, ZC09, AD86, HAA^{+}11, IPR^{+}05, KIT^{+}20]. Utility [ASG20, YBDJ17]. Utilization [JHLM01, MGW99, ZLAV04]. Utilizations [GHC^{+}17]. Utilizing [CPL^{+}10].

V [IPR^{+}05, OBB^{+}24]. V-Class [IPR^{+}05]. Valedictory [Lin92]. Validation [AML^{+}10, DSR17]. Validity [OK99]. Valuations [Ger10]. Value [AK96, BUMS02, GHR20, LS98, LEG11, PZL^{+}19, SS99, SW95, SK14]. Value-Based [SW95]. Valuedness [Dab21]. Values [DMC^{+}20]. Variable [AW98, EM14, VHK^{+}18, MV17]. Variable-Length [EM14]. Variables [JW16, PPQV16]. Vector [AZK^{+}18, BBR11a, CFC^{+}19, CFX^{+}20, KTBP18, RSJ^{+}19, TSS99]. Vectorization [BGGT02, HRC17, CRM92]. Vectorizing [CK02, SG00]. Vectors [Cza17]. Vehicle [GAK20, fS18]. Vehicles [GKC22]. Vehicular [ANS20]. Verification [AG06, BFS05, CHB06, CFF^{+}06, FG16, LMPS05, SRS06, US05]. Verification-Oriented [SRS06]. verified [KFC08]. Verifying [AKBPV19, Win89]. Versatile [KSJ14]. Version [YAI95]. versioned [SSB^{+}17]. Vertices [LW97]. via [EDA96, FRT^{+}18, HCEP98, SSP^{+}96, ZK07]. ViBe [ZQT20]. Video [DLRS13, KBD03, RY20, RY22, SSEA14, TSS99]. Videos [WZG^{+}17]. Virtual [AATD20, EGJS15, HHW10, JQWG15, LCU92, LVM16, PO07, RLK20, SHZ^{+}14, XLWX19]. Virtualization [MGL^{+}17, WLW^{+}17, ZXY^{+}15]. Virtualized [VFIN12]. Visibility [DPS90]. Vision [NFC^{+}09]. Visual [CPT14, QFRA19]. Visualization [SJKA99]. VLIW [ABASS12, CND95, CS97, GBPK07, ZLAV04]. VLSI [PP10]. VM [WHC^{+}17]. VOD [LJ09]. VoIP [AML^{+}10]. Volatile [CCL12]. Volume [LYG^{+}18]. VORD [KSJ14]. VPPE [QFRA19]. vs [NAP02]. Vshadow [WLW^{+}17]. Vulnerability [ÖA21].

Wait [FLD15, LFD17, Sun11]. Wait-Free [FLD15, LFD17, Sun11]. Warm [LJE05]. Warm-Up [LJE05]. Warp [Lys08].

Waterman

[FJZ^{+}15, HMF^{+}13, RSJ^{+}19, ZTY^{+}19]. Watermarking [GP17]. Wave [LS07]. Waveform [CCL12]. Wavefront [LS20, Wol86]. Wavelet [BC15]. way [DPL86]. Weak [BAP01]. Weakly [DWS16]. Web [HHC^{+}15, NYHA14]. Weight [CM06]. Weighted [Ken01]. Which [Gen16]. while [GL95, Col95]. while-Loops [Col95]. Who [JK12]. Wide [TB23]. Wide-SIMD [TB23]. Window [DM17, NdMM09, SF20]. Window-Based [DM17]. Winograd [uHKAMFM16a, uHKAMFM16b]. Wireless [DM20, RY20, RY22]. within [LLL^{+}15]. Without [LPB13]. WolfPath [ZHF^{+}19]. Word [FLD15, Sun11]. Work [AK92, YH18]. WorkCrews [VR88]. Workflow [CAK17, DST21, LWF^{+}19, SDL17]. Workflows [BEA^{+}19, TTF^{+}08]. Working [FR95]. Worklist [GRC^{+}14]. Workload [OP12]. Workloads [LVJ22, VCP^{+}13]. Workshop [SS10]. Workstation [NIK00]. Workstations [LS05]. World [GHM14, HLP11, WLW^{+}17, YLB19]. Wormhole [LNP91]. Write [MV17]. Writes [WHC^{+}24]. Written [KaM10]. WSN [PK20].

X10 [ASS21]. x86 [MGL^{+}17]. XDP

[CFB94]. Xeon

[BP17, Cza17, LLGC17, ELGE17, PES^{+}18]. Xeon/Xeon [Cza17]. XI [MCE13]. XV [PVG17].

Y-Invalidate [BAP01]. YAKL [NLBB23]. Yield [SS17]. YuruBackup [XZX^{+}15].

Zone [JCH^{+}08, MS11].

References

Anane:2015:TBE

[AA15] Amine Anane and El Mostapha Aboulhamid. A transactionbased environment for system modeling and parallel simulation. International Journal of Parallel Programming, 43(1): 24–58, February 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0303-4.

Andion:2016:LAA

[AAB^{+}16] José M. Andión, Manuel Arenaz, François Bodin, Gabriel Rodríguez, and Juan Touriño. Localityaware automatic parallelization for GPGPU with OpenHMPP directives. International Journal of Parallel Programming, 44 (3):620–643, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0362-9.

Ahsan:2020:CCS

[AAI^{+}20a] M. A. Manazir Ahsan, Ihsan Ali, Mohd Yamani Idna Bin Idris, Muhammad Imran, and Muhammad Shoaib. Correction to: Countering Statistical Attacks in Cloud-Based Searchable Encryption. International Journal of Parallel Programming, 48(3):580, June 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0599-1.pdf.

See [AAI^{+}20b].

Ahsan:2020:CSA

[AAI^{+}20b] M. A. Manazir Ahsan, Ihsan Ali, Mohd Yamani Idna Bin Idris, Muhammad Imran, and Muhammad Shoaib. Countering statistical attacks in cloudbased searchable encryption. International Journal of Parallel Programming, 48(3):470–495, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). See correction [AAI^{+}20a].

Ahmad:2020:LBC

[AAN^{+}20] Mudassar Ahmad, Usman Ahmad, Md Asri Ngadi, Muhammad Asif Habib, Shehzad Khalid, and Rehan Ashraf. Loss based congestion control module for health centers deployed by using advanced IoT based SDN communication networks. International Journal of Parallel Programming, 48(2):213– 243, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Alabady:2020:NSM

[AATD20] Salah A. Alabady, Fadi Al-Turjman, and Sadia Din. A novel security model for cooperative virtual networks in the IoT era. International Journal of Parallel Programming, 48 (2):280–295, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Abboud:2012:CHR

[ABASS12] Fadi Abboud, Yosi Ben-Asher, Yousef Shajrawi, and Esti Stein. Combining height reduction and scheduling for VLIW machines enhanced with three-argument arithmetic operations. International Journal of Parallel Programming, 40(5):488–513, October 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=5&spage=488.

Ayguade:2010:EOS

[ABB^{+}10] Eduard Ayguadé, Rosa M.

Badia, Pieter Bellens, Daniel Cabrera, Alejandro Duran Roger Ferrer, Marc González, Francisco Igual, Daniel Jiménez-González, Jesús Labarta, Luis Martinell, Xavier Martorell, Rafael Mayo, Josep M. Pérez, Judit Planas, and Enrique S. Quintana-Ortí. Extending OpenMP to survive the heterogeneous multicore era. International Journal of Parallel Programming, 38(5–6):440– 459, October 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=440.

Aldaya:2019:MTA

[ABSSS19] Alejandro Cabrera Aldaya, Billy Bob Brumley, Alejandro J. Cabrera Sarmiento, and Santiago Sánchez-Solano. Memory tampering attack on binary GCD based inversion algorithms. International Journal of Parallel Programming, 47(4):621–640, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Amme:2000:DDA

[ABTZ00] Wolfram Amme, Peter Braun, François Thomasset, and Eberhard Zehendner. Data dependence analysis of assembly code. International Journal of Parallel Programming, 28(5):431–467, October 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=5&spage=431.

Almer:2013:PDB

[ABvK^{+}13] Oscar Almer, Igor Böhm, Tobias Edler von Koch, Björn Franke, Stephen Kyle, Volker Seeker, Christopher Thompson, and Nigel Topham. A parallel dynamic binary translator for efficient multi-core simulation. International Journal of Parallel Programming, 41(2):212–235, April 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0222-9.

Attali:2001:EFI

[ACC^{+}01] Isabelle Attali, Denis Caromel, Yung-Syau Chen, Jean-Luc Gaudiot, and Andrew L. Wendelborn. Enhancing functional and irregular parallelism: Stateful functions and their semantics. International Journal of Parallel Programming, 29(4):433–460, August 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/22/3/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/22/3/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=4&spage=
                                        433.

Almasi:2002:DSM

[ACC^{+}02] George S. Almasi, Călin Caşcaval, José G. Castaños, Monty Denneau, Wilm Donath, Maria Eleftheriou, Mark Giampapa, Howard Ho, Derek Lieber, José E. Moreira, Dennis Newns, Marc Snir, and Henry S. Warren, Jr. Demonstrating the scalability of a molecular dynamics application on a petaflops computer. International Journal of Parallel Programming, 30(4):317–351, August 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/5/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/5/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&

volume=30&issue=4&spage=
317.

Aldinucci:2014:DPP

[ACD^{+}14] M. Aldinucci, S. Campa, M. Danelutto, P. Kilpatrick, and M. Torquati. Design patterns percolating to parallel programming framework implementation. International Journal of Parallel Programming, 42(6):1012– 1031, December 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0273-6.

Aldinucci:2016:PEP

[ACD^{+}16] Marco Aldinucci, Sonia Campa, Marco Danelutto, Peter Kilpatrick, and Massimo Torquati. Pool evolution: a parallel pattern for evolutionary and symbolic computing. International Journal of Parallel Programming, 44(3): 531–551, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0358-5.

Almeida:1986:PAS

[AD86] Virgilio A. F. Almeida and Lawrence W. Dowdy. Performance analysis of a scheme for concurrency/synchronization using queueing network models. International Journal of Parallel Programming, 15(6):529–550, December 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=6&spage=529.

Akl:1989:PSC

[AD89] Selim G. Akl and Frank Dehne. Pipelined search on coarse grained networks. International Journal of Parallel Programming, 18(5):359–364, October 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=5&spage=359.

Arcucci:2017:DTR

[ADC^{+}17] Rossella Arcucci, Luisa D’Amore, Luisa Carracciuolo, Giuseppe Scotti, and Giuliano Laccetti. A decomposition of the Tikhonov regularization functional oriented to exploit hybrid multilevel parallelism. International Journal of Parallel Programming, 45(5):1214– 1235, October 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Aguilar:2017:TPE

[AER^{+}17] Miguel Angel Aguilar, Juan Fernando Eusse, Projjol Ray, Rainer Leupers, Gerd Ascheid, Weihua Sheng, and Prashant Sharma. Towards parallelism extraction for heterogeneous multicore Android devices. International Journal of Parallel Programming, 45(6):1592– 1624, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Arras:2015:LSE

[AF15] Paul-Antoine Arras and Didier Fuin. List scheduling in embedded systems under memory constraints. International Journal of Parallel Programming, 43(6):1103– 1128, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0338-1.

Aiello:2006:EOS

[AFM^{+}06] A. Aiello, M. Mango Furnari, A. Massarotti, S. Brandi, V. Caputo, and V. Barone. An experimental ontology server for an information Grid environment. International Journal of Parallel Programming, 34(6):489–508, December 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=6&spage=489.

Araujo:2008:PAG

[AFO^{+}08] Renata Braga Araújo, Guilherme Henrique Trielli Ferreira, Gustavo Henrique Orair, Wagner Meira, Renato Antônio Celso Ferreira, Dorgival Olavo Guedes Neto, and Mohammed Javeed Zaki. The ParTriCluster algorithm for gene expression analysis. International Journal of Parallel Programming, 36(2): 226–249, April 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=2&spage=226.

Altman:1998:OMS

[AG98] Erik R. Altman and Guang R. Gao. Optimal modulo scheduling through enumeration. International Journal of Parallel Programming, 26 (3):313–344, June 1998.

CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=26&issue=3&spage=
313.

Abdi:2006:VSL

[AG06] Samar Abdi and Daniel Gajski. Verification of system level model transformations. International Journal of Parallel Programming, 34(1):29–59, March 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=1&spage=29.

Araujo:2015:GES

[AG15] Guido Araujo and Jean-Luc Gaudiot. Guest editorial: SBAC–PAD 2013. International Journal of Parallel Programming, 43(6):961–964, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0377-2;
http://link.springer.
com/content/pdf/10.1007/
s10766-015-0377-2.pdf.

Aviles-Gonzalez:2014:SMM

[AGPGF14] Ana Avilés-González, Juan Piernas, and Pilar González-Férez. Scalable metadata management through OSD+ devices. International Journal of Parallel Programming, 42(1):4– 29, February 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0207-8.

Allombert:2017:MMP

[AGT17] V. Allombert, F. Gava, and J. Tesson. Multi-ML: Programming Multi-BSP algorithms in ML. International Journal of Parallel Programming, 45 (2):340–361, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-016-0417-6.

Ali:1986:GGC

[AH86] Khayri A. M. Ali and Seif Haridi. Global garbage collection for distributed heap storage systems. International Journal of Parallel Programming, 15(5):339–387, October 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=5&spage=339.

AlFaruque:2008:QSC

[AH08] Mohammad Abdullah Al Faruque and Jörg Henkel. QoS-supported on-chip communication for multiprocessors. International Journal of Parallel Programming, 36(1):114–139, February 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=1&spage=114.

Arnold:2001:EIB

[AHKR01] Matthew Arnold, Michael Hsiao, Ulrich Kremer, and Barbara G. Ryder. Exploring the interaction between Java’s implicitly thrown exceptions and instruction scheduling. International Journal of Parallel Programming, 29(2):111–137, April 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/20/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/20/1/fulltext.pdf;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=2&spage=
                                        111.

Aubrey-Jones:2016:SMI

[AJF16] Tristan Aubrey-Jones and Bernd Fischer. Synthesizing MPI implementations from functional dataparallel programs. International Journal of Parallel Programming, 44(3): 552–573, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0359-4.

Ali:1990:FPS

[AK90a] Khayri A. M. Ali and Roland Karlsson. Full Prolog and scheduling OR-Parallelism in Muse. International Journal of Parallel Programming, 19(6): 445–475, December 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=19&issue=6&spage=
                                        445.

Ali:1990:MAP

[AK90b] Khayri A. M. Ali and Roland Karlsson. The Muse approach to OR-Parallel Prolog. International Journal of Parallel Programming, 19(2): 129–162 (or 129–160??), April 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=2&spage=129.

Ali:1992:SSW

[AK92] Khayri A. M. Ali and Roland Karlsson. Scheduling speculative work in MUSE and performance results. International Journal of Parallel Programming, 21(6):449–476, December 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=6&spage=449.

Albonesi:1996:MVA

[AK96] David H. Albonesi and Israel Koren. A mean value analysis multiprocessor model incorporating super-scalar processors and latency tolerating techniques. International Journal of Parallel Programming, 24(3):235– 263, June 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Alam:2017:PAG

[AK17] Maksudul Alam and Maleq Khan. Parallel algorithms for generating random networks with given degree sequences. International Journal of Parallel Programming, 45(1): 109–127, February 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0389-y.

Ahmed:2020:SPS

[AKA^{+}20] Idrees Ahmed, Abid Khan, Adeel Anjum, Mansoor Ahmed, and Muhammad Asif Habib. A secure provenance scheme for detecting consecutive colluding users in distributed networks. International Journal of Parallel Programming, 48(2): 344–366, April 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Abadi:2019:VPC

[AKBPV19] Moria Abadi, Sharon Keidar-Barner, Dmitry Pidan, and Tatyana Veksler. Verifying parallel code after refactoring using equivalence checking. International Journal of Parallel Programming, 47(1): 59–73, February 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Abraham:1998:MST

[AKD98] Santosh G. Abraham, Vinod Kathail, and Brian L. Deitrich. Meld scheduling: a technique for relaxing scheduling constraints. International Journal of Parallel Programming, 26(4):349–381, August 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=4&spage=349.

Ali:2013:MFT

[AKHD13] Nawab Ali, Sriram Krishnamoorthy, Mahantesh Halappanavar, and Jeff Daily. Multi-fault tolerance for Cartesian data distributions. International Journal of Parallel Programming, 41(3): 469–493, June 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0218-5.

Arunagiri:2014:FTO

[AKT^{+}14] Sarala Arunagiri, Yipkei Kwok, Patricia J. Teller, Ricardo A. Portillo, and Seetharami R.

Seelam. FAIRIO: a throughput-oriented algorithm for differentiated I/O performance. International Journal of Parallel Programming, 42(1): 165–197, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0217-6.

Ayguade:1995:ARP

[ALG^{+}95] Eduard Ayguade, Jesus Labarta, Jordi Garcia, Merce Girones, and Mateo Valero. Analyzing reference patterns in automatic data distribution tools. International Journal of Parallel Programming, 23(6):515–535, December 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ali:1986:PEP

[Ali86] Khayri A. M. Ali. ORparallel execution of PRO-LOG on a multi-sequential machine. International Journal of Parallel Programming, 15(3):189–214, June 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=3&spage=189.

Part of the BC-machine project, SICS, Sweden.

Ashcraft:2019:COA

[ALPS19] Matthew B. Ashcraft, Alexander Lemon, David A. Penry, and Quinn Snell. Compiler optimization of accelerator data transfers. International Journal of Parallel Programming, 47 (1):39–58, February 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Asher:2017:GAR

[ALTT17] Yosi Ben Asher, Irina Lipov, Vladislav Tartakovsky, and Dror Tiv. Generating ASIPs with reduced number of connections to the register-file. International Journal of Parallel Programming, 45 (6):1461–1487, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Abramson:1995:EPS

[AM95] David Abramson and A. McKay. Evaluating the performance of a SISAL implementation of the Abingdon Cross Image Processing Benchmark. International Journal of Parallel Programming, 23(2):105–134, April 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Al-Mouhamed:2004:AOP

[AM04] Mayez Al-Mouhamed. Array organization in parallel memories. International Journal of Parallel Programming, 32(2):123– 163, April 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=2&spage=123.

Ayguade:2007:I

[AM07a] Eduard Ayguadé and Matthias S. Mueller. Introduction. International Journal of Parallel Programming, 35(5):437–439, October 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=5&spage=437.

Ayguade:2007:SIO

[AM07b] Eduard Ayguadé and Matthias S. Mueller. Special issue on OpenMP

— Guest Editors’ introduction. International Journal of Parallel Programming, 35(4):331–333, August 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=4&spage=331.

Al-Mouhamed:2001:ENG

[AMAH01] Mayez Al-Mouhamed and Hussam Abu-Haimed. Evaluation of neural and genetic algorithms for synthesizing parallel storage schemes. International Journal of Parallel Programming, 29(4):365–399, August 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/22/1/abstract.htm;
http://ipsapp009.lwwonline.
com/content/getfile/
4773/22/1/fulltext.pdf;
http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=4&spage=
365.

Arandi:2018:DDT

[AMKE18] Samer Arandi, George Matheou, Costas Kyriacou, and Paraskevas Evripidou. Data-driven thread execution on heterogeneous processors. International Journal of Parallel Programming, 46 (2):198–224, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Apostolakos:2010:DIV

[AML^{+}10] Spyros Apostolakos, Apostolos Meliones, George Lykakis, Emmanuel Touloupis, and Vassilis Vlagoulis. Design, implementation and validation of an open source IP-PBX/VoIP gateway multi-core SoC. International Journal of Parallel Programming, 38(3–4): 288–302, June 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=3&spage=288.

Ahmed:2001:STL

[AMP01] Nawaaz Ahmed, Nikolay Mateev, and Keshav Pingali. Synthesizing transformations for locality enhancement of imperfectlynested loop nests. International Journal of Parallel Programming, 29(5):493– 544, October 2001. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/23/3/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/23/3/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&

volume=29&issue=5&spage=
493.

August:2005:ASC

[AMP^{+}05] David I. August, Sharad Malik, Li-Shiuan Peh, Vijay Pai, Manish Vachharajani, and Paul Willmann. Achieving structural and composable modeling of complex systems. International Journal of Parallel Programming, 33(2–3): 81–101, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=2&spage=81.

anMey:2007:NPO

[aMST07] Dieter an Mey, Samuel Sarholz, and Christian Terboven. Nested parallelization with OpenMP. International Journal of Parallel Programming, 35 (5):459–476, October 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=35&issue=5&spage=
459.

August:1999:PRI

[AmWHM99] David I. August, Wen mei W. Hwu, and Scott A. Mahlke. The partial reverse if-conversion framework for balancing control flow and predication. International Journal of Parallel Programming, 27(5):381–423, October 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=5&spage=381.

Anonymous:1986:CLF

[Ano86a] Anonymous. The church of the least fixed point. International Journal of Parallel Programming, 15(5): 457, October 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=5&spage=457.

Anonymous:1986:EJT

[Ano86b]  Anonymous.  Editorial:

a journal transformed. International Journal of Parallel Programming, 15 (1):3–4, February 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=15&issue=1&spage=

Anonymous:1986:H

[Ano86c] Anonymous. Hotspotting.

International Journal of Parallel Programming, 15 (4):337, August 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=4&spage=337.

Anonymous:1986:IA

[Ano86d] Anonymous. Important announcement. International Journal of Parallel Programming, 15(1):1, February 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=1&spage=1.

Anonymous:1987:AL

[Ano87a]  Anonymous.  Amdahl’s

law. International Journal of Parallel Programming, 16(1):85, February 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=1&spage=85.

Anonymous:1987:C

[Ano87b]  Anonymous.  Connec-

tionism. International Journal of Parallel Programming, 16(4):339, August 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=4&spage=339.

Anonymous:1987:FDL

[Ano87c]  Anonymous.  Fixpoints

in Daily Life. International Journal of Parallel Programming, 16(5):425, October 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=5&spage=425.

Anonymous:1987:ICI

[Ano87d] Anonymous. Isomorphic computers inc.: With isomorphic computers, more is more^{T M}. International Journal of Parallel Programming, 16(2):179–182, April 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=2&spage=179.

Anonymous:1987:SP

[Ano87e]  Anonymous.  Systolic

processing. International Journal of Parallel Programming, 16(3): 261, June 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=3&spage=261.

Anonymous:1992:IAS

[Ano92] Anonymous. Important announcement to subscribers. International Journal of Parallel Programming, 21(6):387, December 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=6&spage=387.

Anonymous:2000:GEI

[Ano00a] Anonymous. Guest Editor’s introduction. International Journal of Parallel Programming, 28(2): 133–134, April 2000. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=2&spage=133.

Anonymous:2000:Ia

[Ano00b]  Anonymous.  Introduc-

tion. International Journal of Parallel Programming, 28(4):321–323, August 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=4&spage=321.

Anonymous:2000:Ib

[Ano00c]  Anonymous.  Introduc-

tion. International Journal of Parallel Programming, 28(5):429–430, October 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=5&spage=429.

Anonymous:2001:I

[Ano01]  Anonymous.  Introduc-

tion. International Journal of Parallel Programming, 29(1):1–2, February 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/13/1/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/13/1/fulltext.pdf;

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=1&spage=
1.

Anonymous:2003:E

[Ano03]  Anonymous.  Erratum.

International Journal of Parallel Programming, 31 (3):179, June 2003. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/
4773/33/6/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/33/6/fulltext.pdf;
http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=3&spage=
179.

Anonymous:2014:EN

[Ano14] Anonymous. Editor’s note.

International Journal of Parallel Programming, 42 (2):383, April 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-013-0298-x.pdf.

Anonymous:2016:ENH

[Ano16a] Anonymous. Editor’s note:

High-level parallel programming and applications (HLPP). International Journal of Parallel Programming, 44(3): 381–382, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0402-0.

Anonymous:2016:ENS

[Ano16b] Anonymous. Editor’s note: Special issue on computing frontiers. International Journal of Parallel Programming, 44(5): 923, October 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0446-1.pdf.

Anonymous:2016:ENSb

[Ano16c] Anonymous. Editor’s note: Special section on concurrent systems: Status and perspectives. International Journal of Parallel Programming, 44(2): 308, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0405-x.pdf.

Anonymous:2016:ENSa

[Ano16d] Anonymous. Editor’s note: Special section on dataflow for multicore. International Journal of Parallel Programming, 44(2):

207, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0404-y.pdf.

Anonymous:2018:ENSb

[Ano18a] Anonymous. Editor’s note:

Special issue on embedded computer systems: Architectures, modeling and simulation. International Journal of Parallel Programming, 46(6):1184, December 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0596-4.pdf.

Anonymous:2018:ENSa

[Ano18b] Anonymous. Editor’s note:

Special issue on network and parallel computing for new architectures and applications. International Journal of Parallel Programming, 46(4): 647, August 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0588-4.pdf.

Anonymous:2019:ENS

[Ano19] Anonymous. Editor’s note:

Special issue on high-level languages and frameworks for high-performance computing. International Journal of Parallel Programming, 47(5–6):1045, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00644-z;
https://link.springer.
com/content/pdf/10.1007/
s10766-019-00644-z.pdf.

Anonymous:2020:EN

[Ano20] Anonymous. Editor’s note. International Journal of Parallel Programming, 48 (4):729, August 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00666-y.

Anonymous:2021:ENSa

[Ano21a]  Anonymous.  Editor’s

note: Special issue on High-level Programming for Heterogeneous Parallel Systems (2019). International Journal of Parallel Programming, 49(2): 135, April 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00705-2.

Anonymous:2021:ENSb

[Ano21b] Anonymous. Editor’s note:

Special Issue on International Embedded Systems Symposium (2019). International Journal of Parallel Programming, 49(2): 199, April 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00706-1.

Awasthi:2012:MDP

[ANS^{+}12] M. Awasthi, D. Nellans, K. Sudan, R. Balasubramonian, and A. Davis. Managing data placement in memory systems with multiple memory controllers. International Journal of Parallel Programming, 40(1):57–83, February 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=1&spage=57.

Ahmad:2020:CHV

[ANS20] Iftikhar Ahmad, Rafidah Md Noor, and Muhammad Shoaib. A cooperative heterogeneous vehicular clustering mechanism for road traffic management. International Journal of Parallel Programming, 48(5): 870–889, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-019-00629-y.

Akturk:2019:ATS

[AO19] Ismail Akturk and Ozcan Ozturk. Adaptive thread scheduling in chip multiprocessors. International Journal of Parallel Programming, 47(5–6):1014– 1044, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-019-00637-y.

Al-Obaidy:2021:PAH

[AOAM21] Furat Al-Obaidy, Arghavan Asad, and Farah A. Mohammadi. A poweraware hybrid cache for chip-multi processors based on neural network prediction technique. International Journal of Parallel Programming, 49(3): 326–346, June 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00691-5.

Ajjanagadde:1986:SAB

[AP86] Venkatramana G. Ajjanagadde and L. M. Patnaik. Systolic architecture for Bspline surfaces. International Journal of Parallel Programming, 15(6):551– 565, December 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=6&spage=551.

Atasu:2003:AAS

[API03] Kubilay Atasu, Laura Pozzi, and Paolo Ienne. Automatic applicationspecific instruction-set extensions under microarchitectural constraints. International Journal of Parallel Programming, 31 (6):411–428, December 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/2/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=6&spage=
411.

Ahmad:2018:MDP

[APR^{+}18] Awais Ahmad, Anand Paul, Sadia Din M. Mazhar Rathore, Gyu Sang Choi, and Gwanggil Jeon. Multilevel data processing using parallel algorithms for analyzing big data in highperformance computing. International Journal of Parallel Programming, 46 (3):508–527, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Areias:2016:LFH

[AR16] Miguel Areias and Ricardo Rocha. A lock-free hash trie design for concurrent tabled logic programs. International Journal of Parallel Programming, 44 (3):386–406, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0346-1.

Azevedo:2005:AAD

[ARB^{+}05] Rodolfo Azevedo, Sandro Rigo, Marcus Bartholomeu, Guido Araujo, Cristiano Araujo, and Edna Barros. The ArchC architecture description language and tools. International Journal of Parallel Programming, 33(5):453–484, October 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=5&spage=453.

Arunkumar:2020:HUI

[ASG20] M. S. Arunkumar, P. Suresh, and C. Gunavathi. High utility infrequent itemset mining using a customized ant colony algorithm. International Journal of Parallel Programming, 48(5):833–849, October 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0621-7.

Anand:2021:FAX

[ASS21] Anshu S. Anand, Karthik Sayani, and R. K. Shyamasundar. Fortress abstractions in X10 framework. International Journal of Parallel Programming, 49(6):911–933, December 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00719-w.

Ahmed:2024:GHR

[ASS24] Alif Ahmed, Farzana Ahmed Siddique, and Kevin Skadron. GraphTango: a hybrid representation format for efficient streaming graph updates and analysis. International Journal of Parallel Programming, 52 (3):147–170, June 2024.

CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00768-x.

Ahn:2015:FAP

[ASW^{+}15] Tae-Hyuk Ahn, Adrian Sandu, Layne T. Watson, Clifford A. Shaffer, Yang Cao, and William T. Baumann. A framework to analyze the performance of load balancing schemes for ensembles of stochastic simulations. International Journal of Parallel Programming, 43(4): 597–630, August 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0309-6.

Adamson:1991:GPA

[AT91] P. Adamson and E. Tick. Greedy partitioned algorithms for the shortestpath problem. International Journal of Parallel Programming, 20(4):271– 298, August 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=4&spage=271.

Ayala:2003:PAC

[AVLV03] José L. Ayala, Alexander Veidenbaum, and Marisa López-Vallejo. Poweraware compilation for register file energy reduction. International Journal of Parallel Programming, 31(6):451–467, December 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/4/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=6&spage=
451.

Anghel:2016:IAH

[AVM^{+}16] Andreea Anghel, Laura Mihaela Vasilescu, Giovanni Mariani, Rik Jongerius, and Gero Dittmann. An instrumentation approach for hardware-agnostic software characterization. International Journal of Parallel Programming, 44 (5):924–948, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0410-0.

Arias:2000:PLP

[AVPG00] Ramiro Varela Arias, Camino Rodríguez Vela, Jorge Puente Peinador, and Cesar Alonso Gonzalez. Parallel logic programming for problem solving. International Journal of Parallel Programming, 28(3):275–319, June 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=3&spage=275.

Autrey:1998:IRG

[AW98] Tito Autrey and Michael Wolfe. Initial results for glacial variable analysis. International Journal of Parallel Programming, 26(1):43–64, February 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=1&spage=43.

Ayguade:2003:GEI

[Ayg03] Eduard Ayguade. Guest Editor’s introduction. International Journal of Parallel Programming, 31 (3):181–183, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://ipsapp007.kluweronline.
com/content/getfile/
4773/33/1/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/33/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=3&spage=
181.

Aleen:2018:ACO

[AZK^{+}18] Farhana Aleen, Vyacheslav P. Zakharin, Rakesh Krishnaiyer, Garima Gupta, David Kreitzer, and Chang-Sun Lin, Jr. Automated compiler optimization of multiple vector loads/ stores. International Journal of Parallel Programming, 46(2):471–503, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Borin:2014:MCU

[BABW14] Edson Borin, Guido Araujo, Mauricio Breternitz, Jr., and Youfeng Wu. Microcode compression using structured–constrained clustering. International Journal of Parallel Programming, 42(1):140–164, February 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

                                        com/article/10.1007/
                                        s10766-012-0206-9.

Ben-Asher:1994:UTC

[BAF94] Yosi Ben-Asher and Eitan Farchi. Using true concurrency to model execution of parallel programs. International Journal of Parallel Programming, 22 (4):375–407, August 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Banerjee:1994:EI

[Ban94] Utpal Banerjee. Editor’s introduction. International Journal of Parallel Programming, 22(5):483– ??, October 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Banerjee:2004:GEIa

[Ban04a] Utpal Banerjee. Guest Editor’s introduction. International Journal of Parallel Programming, 32(3): 165–166, June 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        32&issue=3&spage=165.

Banerjee:2004:GEIb

[Ban04b] Utpal Banerjee. Guest Editor’s introduction. International Journal of Parallel Programming, 32(4):

259–261, August 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=32&issue=4&spage=
259.

Ben-Asher:2001:INP

[BAP01] Yosi Ben-Asher and Dimitry Podvolny. Y-invalidate: a new protocol for implementing weak consistency in DSM systems. International Journal of Parallel Programming, 29(6):583– 606, December 2001. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/24/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/24/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=6&spage=
583.

Ben-Asher:1995:FPF

[BARSW95] Yosi Ben-Asher, Gudula Runger, Assaf Schuster, and Reinhard Wilhelm. 2DT-FP: a parallel functional programming language on two-dimensional data. International Journal of Parallel Programming, 23(5):389–422, October 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Blough:1990:NOM

[BB90] Douglas M. Blough and Nader Bagherzadeh. Nearoptimal message routing and broadcasting in faulty hypercubes. International Journal of Parallel Programming, 19(5):405–423, October 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        19&issue=5&spage=405.

Barone:2017:AFQ

[BBB^{+}17] G. B. Barone, V. Boccia, D. Bottalico, R. Campagna, L. Carracciuolo, G. Laccetti, and M. Lapegna. An approach to forecast queue time in adaptive scheduling: How to mediate system efficiency and users satisfaction. International Journal of Parallel Programming, 45(5): 1164–1193, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Baduel:2007:ATO

[BBC07] Laurent Baduel, Françoise Baude, and Denis Caromel. Asynchronous typed object groups for Grid programming. International Journal of Parallel Programming, 35(6):573–614, December 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=6&spage=573.

Bernstein:1995:SDD

[BBGM95] David Bernstein, Mauricio Breternitz, Jr., Ahmed M. Gheith, and Bilha Mendelson. Solutions and debugging for data consistency in multiprocessors with noncoherent caches. International Journal of Parallel Programming, 23 (1):83–103, February 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Belgin:2011:LPB

[BBR11a] Mehmet Belgin, Godmar Back, and Calvin J. Ribbens. A library for pattern-based sparse matrix vector multiply. International Journal of Parallel Programming, 39(1): 62–87, February 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        39&issue=1&spage=62.

Benoit:2011:ORS

[BBR11b] Anne Benoit, Hinde Lilia Bouziane, and Yves Robert. Optimizing the reliability of streaming applications under throughput constraints. International Journal of Parallel Programming, 39(5):584–614, October 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=5&spage=584.

Bordoloi:2010:GBA

[BC10] Unmesh D. Bordoloi and Samarjit Chakraborty. GPU-based acceleration of system-level design tasks. International Journal of Parallel Programming, 38(3–4):225–253, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=3&spage=225.

Bernabe:2015:AEF

[BC15] Gregorio Bernabé and Javier Cuenca. An autotuning engine for the 3D fast wavelet transform on clusters with hybrid CPU + GPU platforms. International Journal of Parallel Programming, 43(6):1160–1191, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0328-3.

Barthou:2000:MSE

[BCC00] Denis Barthou, Albert Cohen, and Jean-François Collard. Maximal static expansion. International Journal of Parallel Programming, 28(3):213–243, June 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=3&spage=213.

Berman:2005:NGS

[BCC^{+}05] F. Berman, H. Casanova, A. Chien, K. Cooper, H. Dail, A. Dasgupta, W. Deng, J. Dongarra, L. Johnsson, K. Kennedy, C. Koelbel, B. Liu, X. Liu, A. Mandal, G. Marin, M. Mazina, J. Mellor-Crummey, C. Mendes, A. Olugbile, M. Patel, D. Reed, Z. Shi, O. Sievert, H. Xia, and A. YarKhan. New Grid scheduling and rescheduling methods in the GrADS Project. International Journal of Parallel Programming, 33(2–3): 209–229, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=2&spage=209.

Bianchini:1998:EEC

[BCK98] Ricardo Bianchini, Enrique V. Carrera, and Leonidas Kontothanassis. Evaluating the effect of coherence protocols on the performance of parallel programming constructs. International Journal of Parallel Programming, 26(2):143–181, April 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=2&spage=143.

Bailey:1990:PGE

[BCL90] Duane A. Bailey, Janice E. Cuny, and Craig P. Loomis. ParaGraph: Graph editor support for parallel programming environments. International Journal of Parallel Programming, 19(2):75–110, April 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=2&spage=75.

Bourgoin:2014:EAG

[BCL14] Mathias Bourgoin, Emmanuel Chailloux, and Jean-Luc Lamotte. Efficient abstractions for GPGPU programming. International Journal of Parallel Programming, 42 (4):583–600, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0261-x.

Bourgoin:2017:HLD

[BCL17] Mathias Bourgoin, Emmanuel Chailloux, and Jean-Luc Lamotte. High level data structures for GPGPU programming in a statically typed language. International Journal of Parallel Programming, 45 (2):242–261, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0424-7.

Bjornson:2009:NCS

[BCS^{+}09] Robert D. Bjornson, Nicholas J.

Carriero, Martin H. Schultz, Patrick M. Shields, and Stephen B. Weston. Net-WorkSpace: a coordination system for highproductivity environments. International Journal of Parallel Programming, 37 (1):106–125, February 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=37&issue=1&spage=
106.

Brogi:2018:AMQ

[BDD^{+}18] Antonio Brogi, Marco Danelutto, Daniele De Sensi, Ahmad Ibrahim, Jacopo Soldani, and Massimo Torquati. Analysing multiple QoS attributes in parallel design patternsbased applications. International Journal of Parallel Programming, 46(1): 81–100, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Brown:2014:CDR

[BDH^{+}14] Christopher Brown, Marco Danelutto, Kevin Hammond, Peter Kilpatrick, and Archibald Elliott. Cost-directed refactoring for parallel Erlang programs. International Journal of Parallel Programming, 42(4):564–582, August 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0266-5.

Bronevetsky:2007:CFS

[BdS07] Greg Bronevetsky and Bronis R. de Supinski. Complete formal specification of the OpenMP memory model. International Journal of Parallel Programming, 35(4):335–392, August 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=4&spage=335.

Bahi:2014:IRC

[BE14] Mouad Bahi and Christine Eisenbeis. Impact of reverse computing on information locality in register allocation for high performance computing. International Journal of Parallel Programming, 42 (1):49–76, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0212-y.

Badosa:2019:HBR

[BEA^{+}19] Ferran Badosa, Antonio Espinosa, Cesar Acevedo, Gonzalo Vera, and Ana Ripoll. A historybased resource manager for genome analysis workflows applications on clusters with heterogeneous nodes. International Journal of Parallel Programming, 47(2):317–342, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0600-z.pdf.

Bull:2010:PEM

[BEG^{+}10] J. Mark Bull, James Enright, Xu Guo, Chris Maynard, and Fiona Reid. Performance evaluation of mixed-mode OpenMP/MPI implementations. International Journal of Parallel Programming, 38(5–6):396– 417, October 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=396.

Beamonte:2021:AGM

[BEJD21] Raphael Beamonte, Naser Ezzati-Jivan, and Michel R. Dagenais. Automated generation of model-based constraints for common multi-core and real-time applications using execution tracing. International Journal of Parallel Programming, 49(1): 104–134, February 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00689-5.

Bell:1994:SPC

[Bel94] Gordon Bell. Scalable, parallel computers: Alternatives, issues, and challenges. International Journal of Parallel Programming, 22(1):3–46, February 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Bilardi:2013:ESD

[BEP13] Gianfranco Bilardi, Kattamuri Ekanadham, and Pratap Pattnaik. Efficient stack distance computation for a class of priority replacement policies. International Journal of Parallel Programming, 41 (3):430–468, June 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0200-2.

Birath:2024:HLP

[BETK24] Björn Birath, August Ernstsson, John Tinnerholm, and Christoph Kessler. High-level programming of FPGA-accelerated systems with parallel patterns. International Journal of Parallel Programming, 52(4):253–273, August 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00770-3.

Broquedis:2010:FEO

[BFG^{+}10] François Broquedis, Nathalie Furmento, Brice Goglin, Pierre-André Wacrenier, and Raymond Namyst. ForestGOMP: an efficient OpenMP environment for NUMA architectures. International Journal of Parallel Programming, 38(5–6):418– 439, October 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=418.

Boton-Fernandez:2015:CAA

[BFRPVR^{+}15] María Botón-Fernández, Manuel Rodríguez-Pascual, Miguel A. Vega-Rodríguez, Francisco Prieto-Castrillo, and Rafael Mayo-García. A comparative analysis of adaptive solutions for grid environments. International Journal of Parallel Programming, 43(5): 786–811, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0342-5.

Bruschi:2005:FFV

[BFS05] Francesco Bruschi, Fabrizio Ferrandi, and Donatella Sciuto. A framework for the functional verification of SystemC models. International Journal of Parallel Programming, 33(6):667–695, December 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=6&spage=667.

Bodik:1996:ADF

[BG96] Rastislav Bodik and Rajiv Gupta. Array data flow analysis for load-store optimizations in fine-grain architectures. International Journal of Parallel Programming, 24(6): 481–512, December 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Bulic:2003:EAC

[BG03] Patricio Bulić and Veselko Guštin. An extended ANSI C for processors with a multimedia extension. International Journal of Parallel Programming, 31(2):107–136, April 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/32/2/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/32/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=2&spage=
107.

Barigou:2017:MCC

[BG17] Youcef Barigou and Edgar Gabriel. Maximizing communication–computation overlap through automatic parallelization and run-time tuning of nonblocking collective operations. International Journal of Parallel Programming, 45(6):1390–1416, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Bronevetsky:2009:CAC

[BGdS09] Greg Bronevetsky, John Gyllenhaal, and Bronis R. de Supinski. CLOMP: Accurately characterizing OpenMP application overheads. International Journal of Parallel Programming, 37(3):250–265, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=3&spage=250.

Bik:2002:AIR

[BGGT02] Aart J. C. Bik, Milind Girkar, Paul M. Grey, and Xinmin Tian. Automatic intra-register vectorization for the Intel(R) architecture. International Journal of Parallel Programming, 30(2): 65–98, April 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/26/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/26/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=2&spage=
65.

Benini:2011:PRA

[BGMR11] L. Benini, R. Grottesi, S. Morigi, and M. Ruggiero. Parallel rendering and animation of subdivision surfaces on the Cell BE processor. International Journal of Parallel Programming, 39(4):494– 521, August 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=4&spage=494.

Bradley:1987:SLC

[BH87] E. Bradley and R. H. Halstead, Jr. Simulating logic circuits: a multiprocessor application. International Journal of Parallel Programming, 16(4):305– 338, August 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=4&spage=305.

Bell:2006:SMS

[BHJ06] Ian Bell, Nabil Hasasneh, and Chris Jesshope. Supporting microthread scheduling and synchronisation in CMPs. International Journal of Parallel Programming, 34(4):343– 381, August 2006. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        34&issue=4&spage=343.

Bai:2021:CCC

[BHL21] Yang Bai, Dinghuang Hu, and Xiangke Liao. CCRP: Converging credit-based and reactive protocols in datacenters. International Journal of Parallel Programming, 49(5): 685–699, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00698-y.

Brown:2020:RGG

[BJM20a] Christopher Brown, Vladimir Janjic, and Kenneth MacKenzie. Refactoring GrPPI: Generic refactoring for generic parallelism in C++. International Journal of Parallel Programming, 48(4):603–625, August 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00667-x.

Brown:2020:PHP

[BJM20b] Christopher Brown, Vladimir Janjic, and J. McCall. Programming heterogeneous parallel machines using refactoring and Monte-Carlo tree search. International Journal of Parallel Programming, 48(4): 583–602, August 2020.

CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-020-00665-z.

BalaAnand:2020:DFC

[BKK20] M. BalaAnand, N. Karthikeyan, and S. Karthik. Designing a framework for communal software: Based on the assessment using relation modelling. International Journal of Parallel Programming, 48(2):329– 343, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). See retraction notice [BKK23].

BalaAnand:2023:RND

[BKK23] M. BalaAnand, N. Karthikeyan, and S. Karthik. Retraction note: Designing a framework for communal software: Based on the assessment using relation modelling. International Journal of Parallel Programming, 51(1): 107, February 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00751-4. See

[BKK20].

Bik:2008:CSC

[BKT08] Aart J. C. Bik, David L. Kreitzer, and Xinmin Tian. A case study on compiler optimizations for the Intel^{(}R) Core^{T M} 2 Duo processor. International Journal of Parallel Programming, 36(6):571– 591, December 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=6&spage=571.

Bonyadi:2009:BGA

[BM09] Mohammad Reza Bonyadi and Mohsen Ebrahimi Moghaddam. A bipartite genetic algorithm for multi-processor task scheduling. International Journal of Parallel Programming, 37(5):462–487, October 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=5&spage=462.

Baev:2002:BBI

[BMA02] Ivan D. Baev, Waleed M.

Meleis, and Santosh G. Abraham. Backtrackingbased instruction scheduling to fill branch delay slots. International Journal of Parallel Programming, 30(6):397–418, December 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/30/1/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/30/1/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=30&issue=6&spage=
                                        397.

Beaty:1998:FSI

[BmH98] Steve Beaty and Wen mei Hwu. Foreword to the Special Issue. International Journal of Parallel Programming, 26(4):345– 347, August 1998. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=4&spage=345.

Bae:2013:CSS

[BML^{+}13] Hansang Bae, Dheya Mustafa, Jae-Woo Lee, Aurangzeb, Hao Lin, Chirag Dave, Rudolf Eigenmann, and Samuel P. Midkiff. The Cetus sourceto-source compiler infrastructure: Overview and evaluation. International Journal of Parallel Programming, 41(6):753–767, December 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0211-z.

Bermudez:1990:PCS

[BNWL90] Manuel E. Bermudez, Richard Newman-Wolfe, and George Logothetis. Parallel construction of SLR(1) and LALR(1) parsers. International Journal of Parallel Programming, 19(3):163–184, June 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=3&spage=163.

Bohaienko:2023:CDO

[Boh23] Vsevolod Bohaienko. Calculation of distributedorder fractional derivative on Tensor coresenabled GPU. International Journal of Parallel Programming, 51(4–5): 256–270, October 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-023-00754-9.

Bos:2012:LLE

[Bos12] Joppe W. Bos. Lowlatency elliptic curve scalar multiplication. International Journal of Parallel Programming, 40(5):532– 550, October 2012. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=5&spage=532.

Bylina:2017:EFO

[BP17] Beata Bylina and Joanna Potiopa. Explicit fourthorder Runge–Kutta method on Intel Xeon Phi coprocessor. International Journal of Parallel Programming, 45(5):1073–1090, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-016-0458-x.pdf.

Bala:1997:EIS

[BR97] Vasanth Bala and Norman Rubin. Efficient instruction scheduling using finite state automata. International Journal of Parallel Programming, 25(2): 53–82, April 1997. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Baumann:2014:PPI

[BR14a] Thomas Baumann and Michael Resch. Parallel parameter identification in industrial biotechnology. International Journal of Parallel Programming, 42 (3):490–504, June 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0244-y.

Bezensek:2014:SPD

[BR14b] Mitja Bezensek and Borut Robic. A survey of parallel and distributed algorithms for the Steiner tree problem. International Journal of Parallel Programming, 42(2):287–319, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0243-z.

Brooks:1986:BBM

[Bro86] Eugene D. Brooks, II. The butterfly barrier (multiprocessing). International Journal of Parallel Programming, 15(4):295–307, August 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=4&spage=295.

Brown:2015:HLH

[Bro15] Christopher Brown. Highlevel heterogeneous and hierarchical parallel systems (HLPGPU 2014). International Journal of Parallel Programming, 43(5):892– 893, October 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-015-0367-4.pdf.

Brown:2019:GES

[Bro19] Christopher Brown. Guest editorial special issue: High-level programming for heterogeneous parallel systems. International Journal of Parallel Programming, 47(1):1– 2, February 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0576-8.pdf.

Bansal:1989:TGT

[BS89] Arvind K. Bansal and Leon S. Sterling. Transforming generate-and-test programs to execute under committed-choice ANDparallelism. International Journal of Parallel Programming, 18(5):401–446, October 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=5&spage=401.

Broy:1991:SDS

[BS91] Manfred Broy and Thomas Streicher. Specification and design of shared resource arbitration. International Journal of Parallel Programming, 20(1): 1–22, February 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=1&spage=1.

Benkner:2003:EDM

[BS03] Siegfried Benkner and Viera Sipkova. Exploiting distributed-memory and shared-memory parallelism on clusters of SMPs with data parallel programs. International Journal of Parallel Programming, 31(1):3–19, February 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/
4773/31/2/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/31/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&

                                        volume=31&issue=1&spage=
                                        3.

Brown:2007:HSP

[BS07] Russell Brown and Ilya Sharapov. High-scalability parallelization of a molecular modeling application: Performance and productivity comparison between OpenMP and MPI implementations. International Journal of Parallel Programming, 35(5):441–458, October 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=5&spage=441.

Baudisch:2015:ESO

[BS15] Daniel Baudisch and Klaus Schneider. Evaluation of speculation in out-of-order execution of synchronous dataflow networks. International Journal of Parallel Programming, 43(1): 86–129, February 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0277-2.

Bachir:2013:MUF

[BTB^{+}13] Mounira Bachir, Sid-Ahmed-Ali Touati, Frederic Brault, David Gregg, and Albert Cohen. Minimal unroll factor for code generation of software pipelining. International Journal of Parallel Programming, 41(1):1– 58, February 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0203-z.

Braun:2002:PAS

[BUMS02] Tracy D. Braun, Renard Ulrey, Anthony A. Maciejewski, and Howard Jay Siegel. Parallel approaches for singular value decomposition as applied to robotic manipulator Jacobians. International Journal of Parallel Programming, 30(1):1–35, February 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/25/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/25/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=1&spage=
1.

Choi:2017:DLA

[CAK17] Jieun Choi, Theodora Adufu, and Yoonhee Kim.

Data-locality aware scientific workflow scheduling methods in HPC cloud environments. International Journal of Parallel Programming, 45(5):1128– 1141, October 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Camilleri:1989:OSO

[Cam89] Juanito Camilleri. An operational semantics for occam. International Journal of Parallel Programming, 18(5):365–400 (or 149–167??), October 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=5&spage=365.

Cybenko:1988:PPU

[CAP88] G. Cybenko, T. G. Allen, and J. E. Polito. Practical parallel union-find algorithms for transitive closure and clustering. International Journal of Parallel Programming, 17(5):403–423, October 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&

issn=0885-7458&volume=
17&issue=5&spage=403.

Carriero:2009:GEI

[Car09] Nicholas Carriero. Guest Editor introduction: Special issue on high performance computing for high productivity environments. International Journal of Parallel Programming, 37(1):1–2, February 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=1.

Cuomo:2018:GEP

[CAT18] Salvatore Cuomo, Marco Aldinucci, and Massimo Torquati. Guest editorial for programming models and algorithms for data analysis in HPC systems. International Journal of Parallel Programming, 46(3):505–507, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0531-0.pdf.

Corbera:2002:NSA

[CAZ02] Francisco Corbera, Rafael Asenjo, and Emilio Zapata. New shape analysis and interprocedural techniques for automatic parallelization of C codes. International Journal of Parallel Programming, 30(1): 37–63, February 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/25/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/25/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=1&spage=
37.

Chiarulli:1986:PMT

[CB86] Donald M. Chiarulli and Duncan A. Buell. Parallel microprogramming tools for a horizontally reconfigurable architecture. International Journal of Parallel Programming, 15(2): 151–162, April 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=2&spage=151.

Chakrabarti:2001:SSA

[CB01] Dhruva R. Chakrabarti and Prithviraj Banerjee. Static single assignment form for message-passing programs. International Journal of Parallel Programming, 29(2):139–184, April 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/20/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/20/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=2&spage=
139.

Coullon:2019:ECM

[CB19] Hélène Coullon and Julien Bigot. Extensibility and composability of a multistencil domain specific framework. International Journal of Parallel Programming, 47(5–6):1046– 1085, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-017-0539-5.

Che:2017:PGG

[CBR17] Shuai Che, Bradford M.

Beckmann, and Steven K. Reinhardt. Programming GPGPU graph applications with linear algebra building blocks. International Journal of Parallel Programming, 45(3):657– 679, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Camara:2014:EIL

[CCG^{+}14] Jesús Cámara, Javier Cuenca, Domingo Giménez, Luis Pedro García, and Antonio M. Vidal. Empirical installation of linear algebra shared-memory subroutines for auto-tuning. International Journal of Parallel Programming, 42 (3):408–434, June 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0249-6.

Charr:2012:AEM

[CCL12] Jean-Claude Charr, Raphaël Couturier, and David Laiymani. Adaptation and evaluation of the multisplitting-Newton and waveform relaxation methods over distributed volatile environments. International Journal of Parallel Programming, 40(2):164– 183, April 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=2&spage=164.

Chilstedt:2009:DEC

[CDC09] Scott Chilstedt, Chen Dong, and Deming Chen. Design and evaluation of a carbon nanotubebased programmable architecture. International Journal of Parallel Programming, 37(4):389–416, August 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=4&spage=389.

Cuomo:2018:PIM

[CDDM18] Salvatore Cuomo, Pasquale De Michele, Emanuel Di Nardo, and Livia Marcellino. Parallel implementation of a machine learning algorithm on GPU. International Journal of Parallel Programming, 46(5):923–942, October 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Calland:1998:RAO

[CDRV98] Pierre-Yves Calland, Alain Darte, Yves Robert, and Frederic Vivien. On the removal of anti- and outputdependences. International Journal of Parallel Programming, 26(3):285– 312, June 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=3&spage=285.

Chapman:2013:EDO

[CEH13] Barbara Chapman, Deepak Eachempati, and Oscar Hernandez. Experiences developing the OpenUH compiler and runtime infrastructure. International Journal of Parallel Programming, 41(6): 825–854, December 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0230-9.

Chang:1997:IBP

[CEP97] Po-Yung Chang, Marius Evers, and Yale N. Patt. Improving branch prediction accuracy by reducing pattern history table interference. International Journal of Parallel Programming, 25(5):339–362, October 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Crivellini:2019:OPS

[CF19] Andrea Crivellini and Matteo Franciolini. OpenMP parallelization strategies for a discontinuous Galerkin solver. International Journal of Parallel Programming, 47(5–6):838–873, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0589-3.

Carter:1994:XCI

[CFB94] Larry Carter, Jeanne Ferrante, and Vasanth Bala. XDP: a compiler intermediate language extension for the representation and optimization of data movement. International Journal of Parallel Programming, 22(5):485–518, October 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Chen:2019:OSM

[CFC^{+}19] Donglin Chen, Jianbin Fang, Shizhao Chen, Chuanfu Xu, and Zheng Wang. Optimizing sparse matrix–vector multiplications on an ARMv8-based many-core architecture. International Journal of Parallel Programming, 47 (3):418–432, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Currie:2006:ESV

[CFF^{+}06] David Currie, Xiushan Feng, Masahiro Fujita, Alan J. Hu, Mark Kwan, and Sreeranga Rajan. Embedded software verification using symbolic execution and uninterpreted functions. International Journal of Parallel Programming, 34(1):61–91, March 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=1&spage=61.

Chen:2020:CSS

[CFX^{+}20] Donglin Chen, Jianbin Fang, Chuanfu Xu, Shizhao Chen, and Zheng Wang. Characterizing scalability of sparse matrix– vector multiplications on Phytium FT-2000+. International Journal of Parallel Programming, 48(1): 80–97, February 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Carriero:1994:CSA

[CG94] Nicholas Carriero and David Gelernter. Case studies in asynchronous data parallelism. International Journal of Parallel Programming, 22(2):129– 149, April 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Carriero:1995:PSS

[CGJK95] Nicholas Carriero, David Gelernter, Marc Jourdenais, and David Kaminsky. Piranha scheduling: Strategies and their implementation. International Journal of Parallel Programming, 23(1):5–33, February 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Cao:2009:OCP

[CGN^{+}09] Jun Cao, Ayush Goyal, Krista A. Novstrup, Samuel P. Midkiff, and James M. Caruthers. An optimizing compiler for parallel chemistry simulations. International Journal of Parallel Programming, 37(2): 127–152, April 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=2&spage=127.

Ciobanu:2018:CPR

[CGPS18] Catalin Bogdan Ciobanu, Georgi Gaydadjiev, Christian Pilato, and Donatella Sciuto. The case for polymorphic registers in dataflow computing. International Journal of Parallel Programming, 46(6):1185–1219, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-017-0494-1.pdf.

Chao:1995:MRD

[CH95] Heng-Yi Chao and Mary P. Harper. Minimizing redundant dependencies and interprocessor synchronizations. International Journal of Parallel Programming, 23(3):245–262, June 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Chen:2006:VAM

[CHB06] Xi Chen, Harry Hsieh, and Felice Balarin. Verification approach of Metropolis design framework for embedded systems. International Journal of Parallel Programming, 34(1): 3–27, March 2006. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        34&issue=1&spage=3.

Chen:2014:PCS

[CHCL14] Shin-Kai Chen, Cheng-Yu Hung, Ching-Chih Chen, and Chih-Wei Liu. Parallelizing complex streaming applications on distributed scratchpad memory multicore architecture. International Journal of Parallel Programming, 42(6): 875–899, December 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0256-7.

Chang:1996:UPE

[CHPC96] Po-Yung Chang, Eric Hao, Yale N. Patt, and Pohua P. Chang. Using predicated execution to improve the performance of a dynamically scheduled machine with speculative execution. International Journal of Parallel Programming, 24(3):209–234, June 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Conte:1999:EIa

[CHS99] Thomas Conte, Wen-Mei Hwu, and Mark Smotherman. Editors’ introduction. International Journal of Parallel Programming, 27(5):325–326, October 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=5&spage=325.

Chen:2018:YCE

[CHSC18] Zhiwen Chen, Xin He, Jianhua Sun, and Hao Chen. Have your cake and eat it (too): a concurrent hash table with hardware transactions. International Journal of Parallel Programming, 46(4):699–709, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Chang:1996:BCN

[CHYP96] Po-Yung Chang, Eric Hao, Tse-Yu Yeh, and Yale Patt. Branch classification: a new mechanism for improving branch predictor performance. International Journal of Parallel Programming, 24(2):133– 158, April 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Creusillet:1996:IAR

[CI96] Beatrice Creusillet and François Irigoin. Interprocedural array region analyses. International Journal of Parallel Programming, 24(6):513–546, December 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ciepielewski:1991:SPP

[Cie91] Andrzej Ciepielewski. Scheduling in OR-parallel Prolog systems: survey and open problems. International Journal of Parallel Programming, 20(6):421– 451, December 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=6&spage=421.

Corporaal:2000:CCT

[CJA00] Henk Corporaal, Johan Janssen, and Marnix Arnold. Computation in the context of transport triggered architectures. International Journal of Parallel Programming, 28(4):401–427, August 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=4&spage=401.

Coleman:2021:FRM

[CJS21] Evan Coleman, Erik J.

Jensen, and Masha Sosonkina. Fault recovery methods for asynchronous linear solvers. International Journal of Parallel Programming, 49(1): 51–80, February 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

                                        com/article/10.1007/
                                        s10766-020-00676-w.

Chauhan:2002:RVP

[CK02] Arun Chauhan and Ken Kennedy. Reducing and vectorizing procedures for telescoping languages. International Journal of Parallel Programming, 30(4):291–315, August 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/4/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/4/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=30&issue=4&spage=
                                        291.

Cui:2022:EEP

[CKC22] Minyu Cui, Angeliki Kritikakou, and Emmanuel Casseau. Energy-efficient partial-duplication task mapping under multiple DVFS schemes. International Journal of Parallel Programming, 50(2): 267–294, April 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00724-7.

Cosnard:1996:SAG

[CL96] M. Cosnard and M. Loi.

A simple algorithm for the generation of efficient loop structures. International Journal of Parallel Programming, 24(3):265– 289, June 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Cao:2016:DFB

[CLJH16] Jian Cao, Qiang Li, Yuede Ji, and Yukun He. Detection of forwarding-based malicious URLs in online social networks. International Journal of Parallel Programming, 44(1): 163–180, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0330-9.

Chen:2021:MDD

[CLL21] Tianba Chen, Wei Li, and Yunchun Li. oM-DRL: Deep reinforcement learning based coflow traffic scheduler with MLFQ threshold adaption. International Journal of Parallel Programming, 49(5): 646–657, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

com/article/10.1007/
s10766-021-00711-4.

Chockler:2006:LWL

[CM06] Gregory Chockler and Dahlia Malkhi. Lightweight leases for storagecentric coordination. International Journal of Parallel Programming, 34 (2):143–170, April 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=34&issue=2&spage=
143.

Conte:1999:EIb

[CmHS99] Thomas Conte, Wen mei Hwu, and Mark Smotherman. Editors’ introduction. International Journal of Parallel Programming, 27(6):425–426, December 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=6&spage=425.

Corbalan:2004:PMD

[CML04] Julita Corbalan, Xavier Martorell, and Jesus Labarta. Page migration with dynamic spacesharing scheduling policies: The case of the SGI O2000. International Journal of Parallel Programming, 32(4):263–288, August 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=4&spage=263.

Clapp:1990:CCR

[CMW90] Russell M. Clapp, Trevor N.

Mudge, and Donald C. Winsor. Cache coherence requirements for interprocess rendezvous. International Journal of Parallel Programming, 19(1): 31–51, February 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=1&spage=31.

Chen:1994:PAI

[CMW^{+}94] William Y. Chen, Scott A.

Mahlke, Nancy J. Warter, Sadun Anik, and Wen-Mei W. Hwu. Profileassisted instruction scheduling. International Journal of Parallel Programming, 22(2):151–181, April 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Capitanio:1995:HBM

[CND95] Andrea Capitanio, Alexandru Nicolau, and Nikil Dutt. A hypergraph-based model for port allocation on multiple-register-file VLIW architectures. International Journal of Parallel Programming, 23 (6):499–513, December 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Collard:1995:APW

[Col95]  Jean-François  Collard.

Automatic parallelization of while-loops using speculative execution. International Journal of Parallel Programming, 23(2):191– 219, April 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Conery:1988:BEP

[Con88] John S. Conery. Binding environments for parallel logic programs in non-shared memory multiprocessors. International Journal of Parallel Programming, 17(2):125–152, April 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        17&issue=2&spage=125.

Cleaveland:1988:TTC

[CP88] Rance Cleaveland and Prakash Panangaden. Type theory and concurrency. International Journal of Parallel Programming, 17 (2):153–206, April 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=17&issue=2&spage=
153.

Carroll:2004:FIE

[CP04] Steven Carroll and Constantine Polychronopoulos. A framework for incremental extensible compiler construction. International Journal of Parallel Programming, 32(4):289– 316, August 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=4&spage=289.

Canal:2001:DCP

[CPG01] Ramon Canal, Joan-Manuel Parcerisa, and Antonio González. Dynamic code partitioning for clustered architectures. International Journal of Parallel Programming, 29(1): 59–79, February 2001. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.

                                        com/content/getfile/
                                        4773/13/4/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/13/4/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=1&spage=
                                        59.

Czutro:2010:TPI

[CPL^{+}10] Alexander Czutro, Ilia Polian, Matthew Lewis, Piet Engelke, Sudhakar M. Reddy, et al. Threadparallel integrated test pattern generator utilizing satisfiability analysis. International Journal of Parallel Programming, 38(3–4):185–202, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=3&spage=185.

Conte:1996:HBP

[CPMC96] Thomas M. Conte, Burzin A. Patel, Kishore N. Menezes, and J. Stan Cox. Hardwarebased profiling: an effective technique for profiledriven optimization. International Journal of Parallel Programming, 24(2): 187–206, April 1996. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Cao:2012:PMN

[CPP^{+}12] Yong Cao, Debprakash Patnaik, Sean Ponce, Jeremy Archuleta, Patrick Butler, et al. Parallel mining of neuronal spike streams on graphics processing units. International Journal of Parallel Programming, 40(6):605– 632, December 2012. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=6&spage=605.

Chessa:2014:EEE

[CPT14] Stefano Chessa, Susanna Pelagatti, and Nicoletta Triolo. Engineering energy efficient visual sensor network applications using skeletons. International Journal of Parallel Programming, 42(4): 663–680, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0260-y.

Chasparis:2019:EDP

[CR19] Georgios C. Chasparis and Michael Rossbory. Efficient dynamic pinning of parallelized applications by distributed reinforcement learning. International Journal of Parallel Programming, 47(1): 24–38, February 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Crammond:1988:GCA

[Cra88]  Jim Crammond.  A

garbage collection algorithm for shared memory parallel processors. International Journal of Parallel Programming, 17(6): 497–522, December 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=17&issue=6&spage=
                                        497.

Chuang:1992:APU

[CRM92] Ling-Yu Chuang, Vernon Rego, and Aditya Mathur. An application of program unification to priority queue vectorization. International Journal of Parallel Programming, 21(3): 193–224, June 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=3&spage=193.

Cvetanovic:2017:ROT

[CRM17] Milos Cvetanović, Zaharije Radivojević, and Veljko Milutinović. Restart optimization for transactional memory with lazy conflict detection. International Journal of Parallel Programming, 45(3):482– 507, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Conte:1997:OVC

[CS97] Thomas M. Conte and Sumedh W. Sathaye. Optimization of VLIW compatibility systems employing dynamic rescheduling. International Journal of Parallel Programming, 25 (2):83–112, April 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Chabkinian:2016:FL

[CS16] Juan Chabkinian and Thomas J. E. Schwarz SJ. Fast LH∗. International Journal of Parallel Programming, 44(4): 709–734, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0371-8.

Chidambaram:2020:OFS

[CS20] S. Chidambaram and A. Sumathi. Optimal feature selection for the classification of hyperspectral imagery using adaptive spectral–spatial clustering. International Journal of Parallel Programming, 48(5):813–832, October 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0607-5.

Carter:2000:PAR

[CSC^{+}00] Lori Carter, Beth Simon, Brad Calder, Larry Carter, and Jeanne Ferrante. Path analysis and renaming for predicated instruction scheduling. International Journal of Parallel Programming, 28(6):563–588, December 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=6&spage=563.

Chiang:2020:NAB

[CSCL20] Hsiu-Sen Chiang, Arun Kumar Sangaiah, Mu-Yen Chen, and Jia-Yu Liu. A novel artificial bee colony optimization algorithm with SVM for bio-inspired software-defined networking. International Journal of Parallel Programming, 48(2):310–328, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Cheng:2021:SSA

[CSD21] Zhongqi Cheng, Tim Schmidt, and Rainer Dömer. Scaled static analysis and IP reuse for out-of-order parallel SystemC simulation. International Journal of Parallel Programming, 49(2): 200–215, April 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00668-w.

Chen:2020:CCG

[CSF^{+}20] Shuo Chen, Zhan Shi, Dan Feng, Shang Liu, Fang Wang, Lei Yang, and Ruili Yu. CSMq-Graph: Coarse-grained and multi-external-storage multi-queue I/O management for graph computing. International Journal of Parallel Programming, 48 (1):98–118, February 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Chen:1989:HEH

[CSG89] Woei-Kae Chen, Matthias F. M. Stallmann, and Edward F. Gehringer. Hypercube embedding heuristics: an evaluation. International Journal of Parallel Programming, 18(6):505–549, December 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=6&spage=505.

Cristobal-Salas:2003:NSE

[CSTGL03] Alfredo Cristobal-Salas, Andrei Tchernykh, Jean-Luc Gaudiot, and Wen-Yen Lin. Non-strict execution in parallel and distributed computing. International Journal of Parallel Programming, 31(2): 77–105, April 2003. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/
4773/32/1/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/32/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=2&spage=
77.

Czapinski:2014:RCO

[CTB14] Michal Czapiński, Chris Thompson, and Stuart Barnes. Reducing communication overhead in multi-GPU hybrid solver for 2D Laplace’s equation. International Journal of Parallel Programming, 42(6):1032–1047, December 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0293-2.

Caragea:2011:RAC

[CTK^{+}11] George C. Caragea, Alexandros Tzannes, Fuat Keceli, Rajeev Barua, and Uzi Vishkin. Resourceaware compiler prefetching for fine-grained manycores. International Journal of Parallel Programming, 39(5):615–638, October 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=5&spage=615.

Cascaval:2013:GEC

[CTP13] Calin Cascaval, Pedro Trancoso, and Viktor Prasanna. Guest editorial: Computing frontiers. International Journal of Parallel Programming, 41 (3):355–356, June 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0240-2;
http://link.springer.
com/content/pdf/10.1007/
s10766-013-0240-2.pdf.

Carretero:2014:PDP

[CY14] Jesus Carretero and Laurence T. Yang. Parallel and distributed processing with applications: Preface. International Journal of Parallel Programming, 42(3):405–407, June 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-013-0254-9.pdf.

Chiang:2016:OSE

[CYS16] Mei-Ling Chiang, Bo-Wen Yu, and Chi-Shian Shia. Operating system enhancement for supporting massively multiplayer online games in a server cluster. International Journal of Parallel Programming, 44(1): 46–67, February 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0341-6.

Ching:2012:APA

[CZ12] Wai-Mee Ching and Da Zheng. Automatic parallelization of array-oriented programs for a multi-core machine. International Journal of Parallel Programming, 40(5):514–531, October 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=5&spage=514.

Czarnul:2017:BPH

[Cza17] Pawel Czarnul. Benchmarking performance of a hybrid Intel Xeon/Xeon Phi system for parallel computation of similarity measures between large vectors. International Journal of Parallel Programming, 45(5):1091– 1107, October 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0455-0.pdf.

Clark:2003:ADA

[CZTM03] Nathan Clark, Hongtao Zhong, Wilkin Tang, and Scott Mahlke. Automatic design of application specific instruction set extensions through dataflow graph exploration. International Journal of Parallel Programming, 31 (6):429–449, December 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/3/abstract.

                                        htm; http://ipsapp007.
                                        kluweronline.com/content/
                                        getfile/4773/37/3/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=6&spage=
                                        429.

Dabrowski:2021:SVT

[Dab21] Frédéric Dabrowski. On single-valuedness in textually aligned SPMD programs. International Journal of Parallel Programming, 49(6):802–819, December 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00710-5.

Damaj:2007:PAD

[Dam07] Issam W. Damaj. Parallel algorithms development for programmable devices with application from cryptography. International Journal of Parallel Programming, 35(6):529– 572, December 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=6&spage=529.

Darema:2005:NGS

[Dar05] Frederica Darema. The next generation software program. International Journal of Parallel Programming, 33(2–3):73–79, June 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=2&spage=73.

Davison:1987:BSP

[Dav87] A. Davison. Blackboard systems in Polka. International Journal of Parallel Programming, 16(5):401– 424, October 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=5&spage=401.

DeSouza:2008:ISI

[DB08] Alberto F. De Souza and Rajkumar Buyya. Introduction to the Special Issue on the 18th International Symposium on Computer Architecture and High Performance Computing. International Journal of Parallel Programming, 36(2): 163–165, April 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        36&issue=2&spage=163.

Devi:2020:TPP

[DC20] R. Ramya Devi and V. Vijaya Chamundeeswari. Triple DES: Privacy preserving in big data healthcare. International Journal of Parallel Programming, 48(3):515–533, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Deng:2017:GDG

[DCX^{+}17] Mingzhu Deng, Wei Chen, Nong Xiao, Songping Yu, and Yupeng Hu. GLE-Dedup: a globally-locally even deduplication by request-aware placement for better read performance. International Journal of Parallel Programming, 45(4):946–964, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Danelutto:2019:REP

[DDD^{+}19] Marco Danelutto, Tiziano De Matteis, Daniele De Sensi, Gabriele Mencagli, Massimo Torquati, Marco Aldinucci, and Peter Kilpatrick. The RePhrase extended pattern set for data intensive parallel computing. International Journal of Parallel Programming, 47(1):74–93, February 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Dai:2018:RAS

[DDJ^{+}18] Weiqi Dai, Yukun Du, Hai Jin, Weizhong Qiang, Deqing Zou, Shouhuai Xu, and Zhongze Liu. RollSec: Automatically secure software states against general rollback. International Journal of Parallel Programming, 46(4):788–805, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Daumas:2000:PIS

[DE00] Marc Daumas and Paraskevas Evripidou. Parallel implementations of the selection problem: a case study. International Journal of Parallel Programming, 28(1):103–131, February 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=1&spage=103.

DeBenedictis:1987:MUP

[DeB87] E. P. DeBenedictis. A multiprocessor using protocolbased programming primitives. International Journal of Parallel Programming, 16(1):53–84, February 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=1&spage=53.

Demsky:2011:UDE

[Dem11] Brian Demsky. Using discrete event simulation to analyze contention managers. International Journal of Parallel Programming, 39(6):783–808, December 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=6&spage=783.

Dennis:1994:MMP

[Den94] Jack B. Dennis. Machines and models for parallel computing. International Journal of Parallel Programming, 22(1): 47–77, February 1994. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Donaldson:1998:AAP

[DF98] Val Donaldson and Jeanne Ferrante. Analyzing asynchronous pipeline schedules. International Journal of Parallel Programming, 26(1):5–42, February 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=1&spage=5.

Duran:2009:PEO

[DFA^{+}09] Alejandro Duran, Roger Ferrer, Eduard Ayguadé, Rosa M. Badia, and Jesus Labarta. A proposal to extend the OpenMP tasking model with dependent tasks. International Journal of Parallel Programming, 37(3):292– 305, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=3&spage=292.

Duran:2007:PEH

[DFC^{+}07] Alejandro Duran, Roger Ferrer, Juan José Costa, Marc Gonzàlez, Xavier Martorell, Eduard Ayguadé, and Jesús Labarta. A proposal for error handling in OpenMP. International Journal of Parallel Programming, 35(4):393–416, August 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=4&spage=393.

Darlington:2017:TCH

[DFH17] J. Darlington, A. J. Field, and L. Hakim. Tackling complexity in high performance computing applications. International Journal of Parallel Programming, 45(2):402–420, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-016-0422-9.pdf.

Dong:2021:FTU

[DFZ21] Hui Dong, Jianxi Fan, and Jingya Zhou. Faulttolerant and unicast performances of the data center network HSDC. International Journal of Parallel Programming, 49(5): 700–714, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00699-x.

Drummond:2009:PPB

[DGMP09] L. Anthony Drummond, Vicente Galiano, Violeta Migallón, and Jose Penadés. PyACTS: a Python based interface to ACTS tools and parallel scientific applications. International Journal of Parallel Programming, 37(1): 58–77, February 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=58.

Darte:2000:LSL

[DH00] Alain Darte and Guillaume Huard. Loop shifting for loop compaction. International Journal of Parallel Programming, 28(5):499–534, October 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=5&spage=499.

Duncan:2016:UPP

[DJR16] Ralph Duncan, Peder Jungck, and Kenneth Ross. Using packet processing object modules interchangeably as standalone programs or “multiapp” components. International Journal of Parallel Programming, 44(1): 26–45, February 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0332-7.

Dudas:2012:SCA

[DJS12] Ákos Dudás, Sándor Juhász, and Tamás Schrádi. Software controlled adaptive pre-execution for data prefetching. International Journal of Parallel Programming, 40(4):381–396, August 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=4&spage=381.

Dastgeer:2016:SCS

[DK16] Usman Dastgeer and Christoph Kessler. Smart containers and skeleton programming for GPUbased systems. International Journal of Parallel Programming, 44(3): 506–530, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0357-6.

DeBole:2009:NAN

[DKB^{+}09] Michael DeBole, Ramakrishnan Krishnan, Varsha Balakrishnan, Wenping Wang, Hong Luo, et al. New-Age: a negative bias temperature instabilityestimation framework for microarchitectural components. International Journal of Parallel Programming, 37(4):417–431, August 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=4&spage=417.

Dias:2013:SUT

[DLRS13] Tiago Dias, Sebastián López, Nuno Roma, and Leonel Sousa. Scalable unified transform architecture for advanced video coding embedded systems. International Journal of Parallel Programming, 41 (2):236–260, April 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0221-x.

Du:2017:ODA

[DLX^{+}17] Jiayi Du, Renfa Li, Zheng Xiao, Zhao Tong, and Li Zhang. Optimization of data allocation on CMP embedded system with data migration. International Journal of Parallel Programming, 45 (4):965–981, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Degano:1987:POM

[DM87] Pierpaolo Degano and Sergio Marchetti. Partial ordering models for concurrency can be defined operationally. International Journal of Parallel Programming, 16(6):451–478, December 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=6&spage=451.

DeMatteis:2017:PPW

[DM17] Tiziano De Matteis and Gabriele Mencagli. Parallel patterns for windowbased stateful operators on data streams: an algorithmic skeleton approach. International Journal of Parallel Programming, 45 (2):382–401, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-016-0413-x.

Devi:2020:IPW

[DM20] E. Anna Devi and J. Martin Leo Manickam. Identifying partitions in wireless sensor network. International Journal of Parallel Programming, 48(2):296– 309, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Dehne:1991:OCM

[DMC91] Frank Dehne, Russ Miller, and Andrew Rau Chaplin. Optical clustering on a mesh-connected computer. International Journal of Parallel Programming, 20(6):475–486, December 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=6&spage=475.

Ding:2018:PEG

[DMC^{+}18] Zengyu Ding, Gang Mei, Salvatore Cuomo, Nengxiong Xu, and Hong Tian. Performance evaluation of GPU-accelerated spatial interpolation using radial basis functions for building explicit surfaces. International Journal of Parallel Programming, 46(5):963– 991, October 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ding:2020:CEM

[DMC^{+}20] Zengyu Ding, Gang Mei, Salvatore Cuomo, Yixuan Li, and Nengxiong Xu.

Comparison of estimating missing values in IoT time series data using different interpolation algorithms. International Journal of Parallel Programming, 48 (3):534–548, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Danelutto:2021:ASP

[DMK21] Marco Danelutto, Gabriele Mencagli, and Peter Kilpatrick. Algorithmic skeletons and parallel design patterns in mainstream parallel programming. International Journal of Parallel Programming, 49 (2):177–198, April 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00684-w.

Menezes:2021:HLP

[dMMHdLN21] Breno A. de Melo Menezes, Nina Herrmann, and Fernando Buarque de Lima Neto. High-level parallel ant colony optimization with algorithmic skeletons. International Journal of Parallel Programming, 49 (6):776–801, December 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL https://link.springer.
com/article/10.1007/
s10766-021-00714-1.

Menezes:2022:PSI

[dMMKdLN22] Breno Augusto de Melo Menezes, Herbert Kuchen, and Fernando Buarque de Lima Neto. Parallelization of swarm intelligence algorithms: Literature review. International Journal of Parallel Programming, 50(5–6): 486–514, December 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00736-3.

DeMichele:2018:GIO

[DMMP18] Pasquale De Michele, Francesco Maiorano, Livia Marcellino, and Francesco Piccialli. A GPU implementation of OLPCA method in hybrid environment. International Journal of Parallel Programming, 46(3):528–542, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Das:1991:PSA

[DMMS91] Amitabha Das, Louise E.

Moser, and P. M. Melliar-Smith. A parallel sorting algorithm for a novel model of computation. International Journal of Parallel Programming, 20 (5):403–419, October 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=20&issue=5&spage=
403.

deStGermain:2003:PAI

[dMP^{+}03] J. Davison de St.Germain, Alan Morris, Steven G. Parker, Allen D. Malony, and Sameer Shende. Performance analysis integration in the Uintah software development cycle. International Journal of Parallel Programming, 31(1):35–53, February 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/31/4/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/31/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=1&spage=
35.

Dekel:1986:OPA

[DPL86] Eliezer Dekel, Shietung Peng, and S. Sitharma Lyengar. Optimal parallel algorithms for constructing and maintaining a balanced m-way search tree. International Journal of Parallel Programming, 15(6):503–528, December 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=6&spage=503.

Dehne:1990:OVA

[DPS90] Frank Dehne, Quoc T.

Pham, and Ivan Stojmenović. Optimal visibility algorithms for binary images on the hypercube. International Journal of Parallel Programming, 19(3): 213–224, June 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=3&spage=213.

Danelutto:2017:GEH

[DPT17] Marco Danelutto, Susanna Pelagatti, and Massimo Torquati. Guest editorial: High-level parallel programming and applications. International Journal of Parallel Programming, 45(2):199–202, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0442-5.pdf.

Dehne:1997:RPL

[DS97] Frank Dehne and Siang W.

Song. Randomized parallel list ranking for distributed memory multi-processors. International Journal of Parallel Programming, 25 (1):1–16, February 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Deniz:2016:UML

[DS16] Etem Deniz and Alper Sen. Using machine learning techniques to detect parallel patterns of multithreaded applications. International Journal of Parallel Programming, 44 (4):867–900, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0396-z.

Durgadevi:2020:RAC

[DS20] P. Durgadevi and S. Srinivasan. Resource allocation in cloud computing using SFLA and cuckoo search hybridization. International Journal of Parallel Programming, 48(3):549– 565, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Dutta:2017:SVC

[DSR17] Sudakshina Dutta, Dipankar Sarkar, and Arvind Rawat. Synchronization validation for crossthread dependences in parallel programs. International Journal of Parallel Programming, 45(6):1326– 1365, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Di:2021:HPM

[DST21] Zhanyuan Di, En Shao, and Guangming Tan. High-performance migration tool for live container in a workflow. International Journal of Parallel Programming, 49(5): 658–670, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00697-z.

Diavastos:2016:ITD

[DTLW16] Andreas Diavastos, Pedro Trancoso, Mikel Luján, and Ian Watson. Integrating transactions into the data-driven multithreading model using the TFlux platform. International Journal of Parallel Programming, 44(2): 257–277, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0369-2.

Darte:1997:OFM

[DV97] Alain Darte and Frédéric Vivien. Optimal fine and medium grain parallelism detection in polyhedral reduced dependence graphs. International Journal of Parallel Programming, 25(6):447–496, December 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        25&issue=6&spage=447.

Du:2017:CSC

[DWQ17] Yuyue Du, Lu Wang, and Man Qi. Constructing service clusters based on service space. International Journal of Parallel Programming, 45(4):982– 1000, August 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ding:2016:LTS

[DWS16] Yihua Ding, James Z. Wang, and Pradip K. Srimani. A linear time selfstabilizing algorithm for minimal weakly connected dominating sets. International Journal of Parallel Programming, 44(1): 151–162, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0335-4.

Dobre:2014:PPP

[DX14] Ciprian Dobre and Fatos Xhafa. Parallel programming paradigms and frameworks in big data era. International Journal of Parallel Programming, 42(5):710–738, October 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
013-0272-7.

Dong:2010:PNM

[DZW10] Chao Dong, Huijie Zhao, and Wei Wang. Parallel nonnegative matrix factorization algorithm on the distributed memory platform. International Journal of Parallel Programming, 38(2):117–137, April 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=2&spage=117.

Eigenmann:2009:GEI

[EA09] Rudolf Eigenmann and Eduard Ayguadé. Guest editors’ introduction. International Journal of Parallel Programming, 37 (3):247–249, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=37&issue=3&spage=
247.

Ernstsson:2021:SPH

[EAK21] August Ernstsson, Johan Ahlqvist, and Christoph Kessler. SkePU 3: Portable high-level programming of heterogeneous systems and HPC clusters. International Journal of Parallel Programming, 49(6):846–866, December 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00704-3.

Enokido:2014:EER

[EAT14] Tomoya Enokido, Ailixier Aikebaier, and Makoto Takizawa. Energy-efficient redundant execution of processes in a faulttolerant cluster of servers. International Journal of Parallel Programming, 42 (5):798–819, October 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0270-9.

Edmiston:1988:PPB

[ECSS88] Elizabeth W. Edmiston, Nolan G. Core, Joel H. Saltz, and Roger M. Smith. Parallel processing of biological sequence comparison algorithms. International Journal of Parallel Programming, 17(3): 259–275, June 1988. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Eichenberger:1996:MRR

[EDA96] Alexandre E. Eichenberger, Edward S. Davidson, and Santosh G. Abraham. Minimizing register requirements of a modulo schedule via optimum stage scheduling. International Journal of Parallel Programming, 24(2):103– 132, April 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Eijkhout:2005:CSS

[EFED05] Victor Eijkhout, Erika Fuentes, Thomas Eidson, and Jack Dongarra. The component structure of a self-adapting numerical software system. International Journal of Parallel Programming, 33(2–3): 137–143, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=2&spage=137.

El-Gindy:1986:OSP

[EG86] Hossam El-Gindy. An optimal speed-up parallel algorithm for triangulating simplicial point sets in space. International Journal of Parallel Programming, 15(5):389–398, October 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        15&issue=5&spage=389.

Egger:2015:ERV

[EGJS15] Bernhard Egger, Erik Gustafsson, Changyeon Jo, and Jeongseok Son. Efficiently restoring virtual machines. International Journal of Parallel Programming, 43(3): 421–439, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0295-0.

Ernstsson:2023:AAE

[EGK23] August Ernstsson, Dalvan Griebler, and Christoph Kessler. Assessing application efficiency and performance portability in single-source programming for heterogeneous parallel systems. International Journal of Parallel Programming, 51(1): 61–82, February 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00746-1.

Emoto:2007:CFD

[EHKT07] Kento Emoto, Zhenjiang Hu, Kazuhiko Kakehi, and Masato Takeichi. A compositional framework for developing parallel programs on two-dimensional arrays. International Journal of Parallel Programming, 35(6):615–658, December 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=6&spage=615.

Ernsting:2014:SFS

[EK14] Steffen Ernsting and Herbert Kuchen. A scalable farm skeleton for hybrid parallel and distributed programming. International Journal of Parallel Programming, 42(6): 968–987, December 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0269-2.

Ernsting:2017:DPA

[EK17] Steffen Ernsting and Herbert Kuchen. Data parallel algorithmic skeletons with accelerator support. International Journal of Parallel Programming, 45 (2):283–299, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-016-0416-7.

Eedi:2022:IOP

[EKU22] Hemalatha Eedi, Sahith Karra, and Rahul Utkoor. An improved/optimized practical non-blocking PageRank algorithm for massive graphs*. International Journal of Parallel Programming, 50(3– 4):381–404, August 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00725-6.

Estebanez:2016:NDS

[ELGE16] Alvaro Estebanez, Diego R. Llanos, and Arturo Gonzalez-Escribano. New data structures to handle speculative parallelization at runtime. International Journal of Parallel Programming, 44(3):407–426, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
014-0347-0.

Estebanez:2017:UXP

[ELGE17] Alvaro Estebanez, Diego R.

Llanos, and Arturo Gonzalez-Escribano. Using the Xeon Phi platform to run speculatively-parallelized codes. International Journal of Parallel Programming, 45(2):225–241, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
016-0421-x.

Ernstsson:2018:SFT

[ELK18] August Ernstsson, Lu Li, and Christoph Kessler. SkePU 2: Flexible and type-safe skeleton programming for heterogeneous parallel systems. International Journal of Parallel Programming, 46 (1):62–80, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/content/pdf/10.1007/
s10766-017-0490-5.pdf.

Eigenmann:2013:CI

[EM13] Rudi Eigenmann and Sam Midkiff. Compiler infrastructure. International Journal of Parallel Programming, 41(6): 751–752, December 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0250-0;
                                        http://link.springer.
                                        com/content/pdf/10.1007/
                                        s10766-013-0250-0.pdf.

Emoto:2014:AFM

[EM14] Kento Emoto and Kiminori Matsuzaki. An automatic fusion mechanism for variable-length list skeletons in SkeTo. International Journal of Parallel Programming, 42 (4):546–563, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0263-8.

Ebcioglu:1997:GEI

[EmH97] Kemal Ebcioğlu and Wen mei Hwu. Guest Editors’ introduction. International Journal of Parallel Programming, 25(2): 51–??, April 1997. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ellis:1988:APM

[EO88] Carla Schlatter Ellis and Thomas J. Olson. Algorithms for parallel memory allocation. International Journal of Parallel Programming, 17(4):303– 345, August 1988. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=4&spage=303.

Evripidou:2006:MMA

[ES06] Paraskevas Evripidou and George Samaras. Metacomputing with mobile agents. International Journal of Parallel Programming, 34(5):429–458, October 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=5&spage=429.

Ewedafe:2011:PID

[ES11] Simon Uzezi Ewedafe and Rio Hirowati Shariffudin. Parallel implementation of 2-D telegraphic equation on MPI/ PVM cluster. International Journal of Parallel Programming, 39(2):202– 231, April 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=2&spage=202.

Ernstsson:2022:DPP

[EVK22] August Ernstsson, Nicolas Vandenbergen, and Christoph Kessler. A deterministic portable parallel pseudo-random number generator for patternbased programming of heterogeneous parallel systems. International Journal of Parallel Programming, 50(3–4):319–340, August 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00726-5.

Evripidou:2000:I

[Evr00] Paraskevas Evripidou. Introduction. International Journal of Parallel Programming, 28(6):535–536, December 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=6&spage=535.

Engelhardt:1996:PIP

[EW96] Dean Engelhardt and Andrew Wendelborn. A partitioning-independent paradigm for nested data parallelism. International Journal of Parallel Programming, 24(4):291–317, August 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

ElKabbany:2011:DLB

[EWHS11] Ghada F. El Kabbany, Nayer M. Wanas, Nadia H. Hegazi, and Samir I. Shaheen. A dynamic load balancing framework for real-time applications in message passing systems. International Journal of Parallel Programming, 39(2):143– 182, April 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=2&spage=143.

Fresno:2019:HDP

[FBGEL19] Javier Fresno, Daniel Barba, Arturo Gonzalez-Escribano, and Diego R. Llanos. HitFlow: a dataflow programming model for hybrid distributedand shared-memory systems. International Journal of Parallel Programming, 47(1):3–23, February 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Fazio:2021:MRA

[FBV21] Maria Fazio, Alina Buzachis, and Massimo Villari. A map-reduce approach for the Dijkstra algorithm in SDN over osmotic computing systems. International Journal of Parallel Programming, 49(3): 347–375, June 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00693-3.

Fensch:2011:EBC

[FC11] Christian Fensch and Marcelo Cintra. An evaluation of an OS-based coherence scheme for tiled CMPs. International Journal of Parallel Programming, 39(3):271–295, June 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=3&spage=271.

Fujimoto:1987:SMA

[FcF87] Richard M. Fujimoto and Hwa chung Feng. A shared memory algorithm and proof for the generalized alternative construct in CSP. International Journal of Parallel Programming, 16(3):215– 241, June 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=3&spage=215.

Farkas:1999:MAR

[FCJV99] Keith I. Farkas, Paul Chow, Norman P. Jouppi, and Zvonko Vranesic. The multicluster architecture: Reducing processor cycle time through partitioning. International Journal of Parallel Programming, 27(5):327–356, October 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=5&spage=327.

Fonseca:2016:APE

[FCRC16] Alcides Fonseca, Bruno Cabral, João Rafael, and Ivo Correia. Automatic parallelization: Executing sequential programs on a task-based parallel runtime. International Journal of Parallel Programming, 44(6):1337– 1358, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-016-0426-5.

Fan:2016:TCA

[FCZ16] Zhaoxin Fan, Shuoying Chen, and Li Zha. A text clustering approach of Chinese news based on neural network language model. International Journal of Parallel Programming, 44(1): 198–206, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0329-2.

Fang:2019:PPO

[FDY^{+}19] Hongyu Fang, Sai Santosh Dayapule, Fan Yao, Miloš Doroslovački, and Guru Venkataramani. PrO-DACT: Prefetch-obfuscator to defend against cache timing channels. International Journal of Parallel Programming, 47(4):571– 594, August 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Feautrier:1991:DAA

[Fea91] Paul Feautrier. Dataflow analysis of array and scalar references. International Journal of Parallel Programming, 20 (1):23–53 (or 23–52??), February 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=1&spage=23.

Feautrier:1992:SESa

[Fea92a] Paul Feautrier. Some efficient solutions to the affine scheduling problem. I. onedimensional time. International Journal of Parallel Programming, 21(5):313– 347, October 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=5&spage=313.

Feautrier:1992:SESb

[Fea92b] Paul Feautrier. Some efficient solutions to the affine scheduling problem. Part II. multidimensional time. International Journal of Parallel Programming, 21(6):389–420, December 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=6&spage=389.

Feautrier:2006:SSS

[Fea06] Paul Feautrier. Scalable and structured scheduling. International Journal of Parallel Programming, 34(5):459–487, October 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        34&issue=5&spage=459.

Fan:2018:ADL

[FFS18] Sijiang Fan, Jiawei Fei, and Li Shen. Accelerating deep learning with a parallel mechanism using CPU + MIC. International Journal of Parallel Programming, 46(4):660–673, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Fortin:2016:BWT

[FG16] Jean Fortin and Frédéric Gava. BSP-Why: a tool for deductive verification of BSP algorithms with subgroup synchronisation. International Journal of Parallel Programming, 44 (3):574–597, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0360-y.

Fummi:2005:E

[FH05] Franco Fummi and Ian G.

Harris. Editorial. International Journal of Parallel Programming, 33(6):583– 584, December 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=6&spage=583.

Farhan:2018:RTD

[FJA^{+}18] Muhammad Farhan, Sohail Jabbar, Muhammad Aslam, Awais Ahmad, Muhammad Munwar Iqbal, Murad Khan, and Martinez-Enriquez Ana Maria. A real-time data mining approach for interaction analytics assessment: IoT based student interaction framework. International Journal of Parallel Programming, 46(5):886–903, October 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Frieb:2016:PAH

[FJO^{+}16] Martin Frieb, Ralf Jahr, Haluk Ozaktas, Andreas Hugl, Hans Regler, and Theo Ungerer. A parallelization approach for hard real-time systems and its application on two industrial programs.

International Journal of Parallel Programming, 44 (6):1296–1336, December 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
016-0432-7.

Feng:2015:ASW

[FJZ^{+}15] Xiaowen Feng, Hai Jin, Ran Zheng, Lei Zhu, and Weiqi Dai. Accelerating Smith–Waterman alignment of species-based protein sequences on GPU. International Journal of Parallel Programming, 43 (3):359–380, June 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0284-3.

Francez:1987:FAC

[FK87] Nissim Francez and Shmuel Katz. Fairness and the axioms of control predicates. International Journal of Parallel Programming, 16(4):263–278, August 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=4&spage=263.

Fillo:1997:MMM

[FKD^{+}97] Marco Fillo, Stephen W.

Keckler, William J. Dally, Nicholas P. Carter, Andrew Chang, Yevgeny Gurevich, and Whay S. Lee. The M-machine multicomputer. International Journal of Parallel Programming, 25(3):183– 212, June 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Fursin:2011:MGM

[FKM^{+}11] Grigori Fursin, Yuriy Kashnikov, Abdul Wahid Memon, Zbigniew Chamski, Olivier Temam, et al. Milepost GCC: Machine learning enabled selftuning compiler. International Journal of Parallel Programming, 39(3): 296–327, June 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=3&spage=296.

Franke:2012:GEC

[FKT12] Hubertus Franke, Paul H. J. Kelly, and Pedro Trancoso. Guest editorial: Computing frontiers. International Journal of Parallel Programming, 40(6):551–552, December 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=6&spage=551.

Feldman:2015:WFM

[FLD15] Steven Feldman, Pierre LaBorde, and Damian Dechev. A wait-free multi-word compare-and-swap operation. International Journal of Parallel Programming, 43(4): 572–596, August 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0308-7.

Favati:2002:RCI

[FLMR02] Paola Favati, Grazia Lotti, Ornella Menchi, and Francesco Romani. Railway computation for infinite linear systems. International Journal of Parallel Programming, 30 (6):419–439, December 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp009.
lwwonline.com/content/
getfile/4773/30/2/abstract.
htm;   http://ipsapp009.
lwwonline.com/content/
getfile/4773/30/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=

article&issn=0885-7458&
volume=30&issue=6&spage=
419.

Fachada:2017:EPS

[FLMR17a] Nuno Fachada, Vitor V.

Lopes, Rui C. Martins, and Agostinho C. Rosa. Erratum to: Parallelization Strategies for Spatial Agent-Based Models. International Journal of Parallel Programming, 45(6):1625–1626, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0504-3.pdf.

See [FLMR17b].

Fachada:2017:PSS

[FLMR17b] Nuno Fachada, Vitor V.

Lopes, Rui C. Martins, and Agostinho C. Rosa. Parallelization strategies for spatial agent-based models. International Journal of Parallel Programming, 45(3):449–481, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). See erratum [FLMR17a].

Furlinger:2009:CAE

[FM09] Karl Fürlinger and Shirley Moore. Capturing and analyzing the execution control flow of OpenMP applications. International Journal of Parallel Programming, 37(3):266– 276, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=3&spage=266.

Farrens:1996:GEI

[FmH96] Matthew Farrens and Wen mei Hwu. Guest Editors’ introduction. International Journal of Parallel Programming, 24(1):1– ??, February 1996. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Fan:2017:SEE

[FMSG17] Xing Fan, Mostafa Mehrabi, Oliver Sinnen, and Nasser Giacaman. Supporting enhanced exception handling with OpenMP in object–oriented languages. International Journal of Parallel Programming, 45 (6):1366–1389, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Foster:1989:MGC

[Fos89]  Ian Foster.  A multi-

computer garbage collector for a single assignment language. International Journal of Parallel Programming, 18(3):181– 203, June 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=3&spage=181.

Fiore:2014:CBD

[FPCD14] Ugo Fiore, Francesco Palmieri, Aniello Castiglione, and Alfredo De Santis. A clusterbased data-centric model for network-aware task scheduling in distributed systems. International Journal of Parallel Programming, 42(5):755–775, October 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0289-y.

Faraj:2008:BEA

[FPY08a] Ahmad Faraj, Pitch Patarasuk, and Xin Yuan. Bandwidth efficient all-to-all broadcast on switched clusters. International Journal of Parallel Programming, 36(4):426–453, August 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        36&issue=4&spage=426.

Faraj:2008:SPA

[FPY08b] Ahmad Faraj, Pitch Patarasuk, and Xin Yuan. A study of process arrival patterns for MPI collective operations. International Journal of Parallel Programming, 36(6):543– 570, December 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=6&spage=543.

Feitelson:1995:CBR

[FR95] Dror G. Feitelson and Larry Rudolph. Coscheduling based on runtime identification of activity working sets. International Journal of Parallel Programming, 23(2):135–160, April 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Farooqui:2018:ADA

[FRT^{+}18] Naila Farooqui, Indrajit Roy, Yuan Chen Vanish Talwar, Rajkishore Barik, Brian Lewis, Tatiana Shpeisman, and Karsten Schwan. Accelerating data analytics on integrated GPU platforms via runtime specialization. International Journal of Parallel Programming, 46(2): 336–375, April 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Song:2018:VDU

[fS18] Jun fang Song. Vehicle detection using spatial relationship GMM for complex urban surveillance in daytime and nighttime. International Journal of Parallel Programming, 46(5): 859–872, October 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Factor:2006:PID

[FSS06] Michael Factor, Assaf Schuster, and Konstantin Shagin. A platformindependent distributed runtime for standard multithreaded Java. International Journal of Parallel Programming, 34(2): 113–142, April 2006. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=2&spage=113.

Song:2018:TDB

[fSxWC18] Jun fang Song, Wei xing Wang, and Feng Chen. Target detection based on 3D multi-component model and inverse projection transformation. International Journal of Parallel Programming, 46(5): 873–885, October 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Foster:1987:FPB

[FT87] Ian Foster and Stephen Taylor. Flat Parlog: a basis for comparison. International Journal of Parallel Programming, 16(2): 87–125, April 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=2&spage=87.

Furnari:1995:GEI

[Fur95] Mario Mango Furnari. Guest Editor’s introduction. International Journal of Parallel Programming, 23(6):497–??, December 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Fiorin:2016:EDS

[FVvL^{+}16] Leandro Fiorin, Erik Vermij, Jan van Lunteren, Rik Jongerius, and Christoph Hagleitner. Exploring the design space of an energyefficient accelerator for the SKA1-low central signal processor. International Journal of Parallel Programming, 44(5): 1003–1027, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0420-y.

Faigin:1994:PIR

[FWH^{+}94] Keith A. Faigin, Stephen A.

Weatherford, Jay P. Hoeflinger, David A. Padua, and Paul M. Petersen. The Polaris internal representation. International Journal of Parallel Programming, 22(5):553–586, October 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Geng:2022:PBA

[GAG22] Tongsheng Geng, Marcos Amaris, and Jean-Luc Gaudiot. A profilebased AI-assisted dynamic scheduling approach for heterogeneous architectures. International Journal of Parallel Programming, 50(1):115–151, February 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00721-2.

Gait:1989:SOP

[Gai89]  Jason Gait.  Speedup

and optimality in pipeline programs. International Journal of Parallel Programming, 18(4):277–290, August 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=4&spage=277.

Gowtham:2020:EMR

[GAK20] P. Gowtham, V. P. Arunachalam, and S. Karthik. An efficient monitoring of real time traffic clearance for an emergency service vehicle using IOT. International Journal of Parallel Programming, 48(5): 786–812, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-018-0603-9.

Gao:1986:MPL

[Gao86] Guang R. Gao. Maximum pipelining linear recurrence on static data flow computers. International Journal of Parallel Programming, 15(2):127– 149, April 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=2&spage=127.

Gadde:2016:AOI

[GAR^{+}16] Srimanth Gadde, William Acosta, Jordan Ringenberg, Robert Green, and Vijay Devabhaktuni. Achieving optimal inter-node communication in graph partitioning using random selection and breadthfirst search. International Journal of Parallel Programming, 44(4): 772–800, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0374-5.

Gaudiot:1996:GEI

[Gau96] Jean-Luc Gaudiot. Guest Editor’s introduction. International Journal of Parallel Programming, 24 (3):207–??, June 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Grelck:2020:RAD

[GB20] Clemens Grelck and Cédric Blom. Resource-aware data parallel array processing. International Journal of Parallel Programming, 48(4):652–674, August 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-020-00664-0.

Guo:2008:CIR

[GBC^{+}08] Zhi Guo, Betul Buyukkurt, John Cortes, Abhishek Mitra, and Walild Najjar. A compiler intermediate representation for reconfigurable fabrics. International Journal of Parallel Programming, 36(5):493– 520, October 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=5&spage=493.

Gaster:2010:CTH

[GBLG10] Benedict R. Gaster, Tim Bainbridge, David Lacey, and David Gardner. Compilation techniques for high level parallel code. International Journal of Parallel Programming, 38 (1):4–18, February 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=38&issue=1&spage=
                                        4.

Gangwar:2007:EBB

[GBPK07] Anup Gangwar, M. Balakrishnan, Preeti Ranjan Panda, and Anshul Kumar. Evaluation of bus based interconnect mechanisms in clustered VLIW architectures. International Journal of Parallel Programming, 35(6):507– 527, December 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=6&spage=507.

Graham:2003:NFT

[GCD^{+}03] Richard L. Graham, Sung-Eun Choi, David J. Daniel, Nehal N. Desai, Ronald G. Minnich, Craig E. Rasmussen, L. Dean Risinger, and Mitchel W. Sukalski. A network-failure-tolerant message-passing system for terascale clusters. International Journal of Parallel Programming, 31 (4):285–303, August 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://ipsapp007.kluweronline.
com/content/getfile/
4773/34/2/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/34/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=4&spage=

                                        285.

Ginosar:1989:TCP

[GE89] Ran Ginosar and David Egozi. Topological comparison of perfect shuffle and hypercube. International Journal of Parallel Programming, 18(1): 37–68, February 1989. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=1&spage=37.

Gupta:1990:HSS

[GE90] Rajiv Gupta and Michael Epstein. High speed synchronization of processors using fuzzy barriers. International Journal of Parallel Programming, 19(1): 53–73, February 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        19&issue=1&spage=53.

Gedik:2013:ATS

[Ged13]  Bugra Gedik.  Auto-

tuning similarity search algorithms on multi-core architectures. International Journal of Parallel Programming, 41(5): 595–620, October 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0239-8.

Gentleman:2016:CPC

[Gen16] W. Morven Gentleman.

Concurrency paradigms: Competitive, coordinated, and collaborative: Which control mechanisms are appropriate? International Journal of Parallel Programming, 44(2):325– 336, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0370-9.

Gerbessiotis:2010:POP

[Ger10] Alexandros V. Gerbessiotis. Parallel option price valuations with the explicit finite difference method. International Journal of Parallel Programming, 38(2):159–182, April 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=2&spage=159.

Gonzalez:2014:ATD

[GF14] Carlos H. González and Basilio B. Fraguela. An algorithm template for domain-based parallel irregular algorithms. International Journal of Parallel Programming, 42(6): 948–967, December 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0268-3.

Griebl:2000:ISS

[GFL00] Martin Griebl, Paul Feautrier, and Christian Lengauer. Index set splitting. International Journal of Parallel Programming, 28(6): 607–631, December 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=28&issue=6&spage=
607.

Gou:2013:AGC

[GG13] Chunyang Gou and Georgi N. Gaydadjiev. Addressing GPU on-chip shared memory bank conflicts using elastic pipeline. International Journal of Parallel Programming, 41(3): 400–429, June 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0201-1;
http://link.springer.

com/content/pdf/10.1007/
s10766-012-0201-1.pdf.

Gijsbers:2014:ESR

[GG14] Bert Gijsbers and Clemens Grelck. An efficient scalable runtime system for macro data flow processing using S-Net. International Journal of Parallel Programming, 42(6): 988–1011, December 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0271-8.

Garcia:2019:GEH

[GGE19] J. Daniel García and Arturo Gonzalez-Escribano. Guest editorial: Highlevel parallel programming and the road to high performance. International Journal of Parallel Programming, 47(2):161– 163, April 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0606-6.pdf.

Goli:2017:ACS

[GGV17] Mehdi Goli and Horacio González-Vélez. Autonomic coordination of skeleton-based applications over CPU/GPU multi-core architectures.

International Journal of Parallel Programming, 45 (2):203–224, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0419-4.

Goli:2018:FCI

[GGV18] Mehdi Goli and Horacio González-Vélez. Formalised composition and interaction for heterogeneous structured parallelism. International Journal of Parallel Programming, 46(1):120–151, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gupta:1989:SIB

[GH89] Rajiv Gupta and Charles R. Hill. A scalable implementation of barrier synchronization using an adaptive combining tree. International Journal of Parallel Programming, 18(3): 161–180, June 1989. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=3&spage=161.

Ghiya:1996:CAP

[GH96] Rakesh Ghiya and Laurie J. Hendren. Connection analysis: a practical interprocedural heap analysis for C. International Journal of Parallel Programming, 24(6):547–578, December 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ghalaty:2019:ESI

[Gha19] Nahid Farhady Ghalaty.

Editorial: Special issue on side-channel and fault analysis of highperformance computing platforms. International Journal of Parallel Programming, 47(4):535–537, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-019-00636-z.pdf.

Gu:2017:DEP

[GHC^{+}17] Zhuoer Gu, Ligang He, Cheng Chang, Jianhua Sun, Hao Chen, and Chenlin Huang. Developing an efficient pattern discovery method for CPU utilizations of computers. International Journal of Parallel Programming, 45(4):853–878, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/

                                        content/pdf/10.1007/
                                        s10766-016-0439-0.pdf.

Griebler:2019:HLP

[GHDF19] Dalvan Griebler, Renato B. Hoffmann, Marco Danelutto, and Luiz G. Fernandes. High-level and productive stream parallelism for Dedup, Ferret, and Bzip2. International Journal of Parallel Programming, 47(2):253– 271, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

George:1986:SSP

[GHLN86] Alan George, Michael T. Heath, Joseph Liu, and Esmond Ng. Solution of sparse positive definite systems on a shared-memory multiprocessor. International Journal of Parallel Programming, 15(4):309–325, August 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        15&issue=4&spage=309.

Gaspar:2014:BCW

[GHM14] Nuno Gaspar, Ludovic Henrio, and Eric Madelaine. Bringing Coq into the world of GCM distributed applications.

International Journal of Parallel Programming, 42 (4):643–662, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0264-7.

Gruttner:2020:TVS

[GHR20] Kim Grüttner, Philipp A.

Hartmann, and Wolfgang Rosenstiel. A timedvalue stream based ESL timing and power estimation and simulation framework for heterogeneous MPSoCs. International Journal of Parallel Programming, 48(6): 957–1007, December 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00656-0.

Givargis:2007:SIE

[Giv07] Tony Givargis. Special issue on embedded processors — Guest Editor introduction. International Journal of Parallel Programming, 35(2): 99–100, April 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=2&spage=99.

Givargis:2008:GEI

[Giv08] Tony Givargis. Guest Editor introduction: Special issue on embedded processors. International Journal of Parallel Programming, 36(5):455–456, October 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=5&spage=455.

Gregor:2005:GPH

[GJK^{+}05] Douglas Gregor, Jaakko Järvi, Mayuresh Kulkarni, Andrew Lumsdaine, David Musser, and Sibylle Schupp. Generic programming and high-performance libraries. International Journal of Parallel Programming, 33(2–3):145– 164, June 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=2&spage=145.

Genaud:2009:FMP

[GJR09] Stéphane Genaud, Emmanuel Jeannot, and Choopan Rattanapoka. Fault-management in P2P-MPI. International Journal of Parallel Programming, 37(5):433–461, October 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=5&spage=433.

Gilder:1994:ASC

[GK94] Mark R. Gilder and Mukkai S. Krishnamoorthy. Automatic sourcecode parallelization using HICOR objects. International Journal of Parallel Programming, 22(3):303– 350, June 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gorlatch:2018:GEH

[GK18] Sergei Gorlatch and Herbert Kuchen. Guest editorial: High-level parallel programming with algorithmic skeletons. International Journal of Parallel Programming, 46(1): 1–3, February 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0512-3.pdf.

Gupta:2022:LAI

[GKC22] Neeraj Gupta, Mahdi Khosravy, and Rubén González Crespo. Lightweight artificial intelligence technology for health diagnosis of agriculture vehicles: Parallel evolving artificial neural networks by genetic algorithm. International Journal of Parallel Programming, 50(1):1– 26, February 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00671-1.

Guzman:1987:PSA

[GKMB87] Adolfo Guzman, Edward J. Krall, Patrick F. McGehearty, and Nader Bagherzadeh. Performance of symbolic applications on a parallel architecture. International Journal of Parallel Programming, 16(3):183– 214, June 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=3&spage=183.

Gupta:1992:EPF

[GL92] Rajiv Gupta and Sunah Lee. Exploiting parallelism on a fine-grained MIMD architecture based upon channel queues. International Journal of Parallel Programming, 21(3):169– 192, June 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=3&spage=169.

Griebl:1995:CSD

[GL95] Martin Griebl and Christian Lengauer. A communication scheme for the distributed execution of loop nests with while loops. International Journal of Parallel Programming, 23(5):471–496, October 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gorawski:2018:EPL

[GL18] Marcin Gorawski and Michal Lorek. Efficient processing of large data structures on GPUs: Enumeration scheme based optimisation. International Journal of Parallel Programming, 46(6):1063– 1093, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0515-0.pdf.

Gao:2017:AOM

[GLLH17] Jiaquan Gao, Zejie Li, Ronghua Liang, and Guixia He. Adaptive optimization l_{1}-minimization solvers on GPU. International Journal of Parallel Programming, 45(3):508–529, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gava:2020:AIC

[GM20] F. Gava and Y. Marquer. Axiomatization and imperative characterization of multi-BSP algorithms: A Q&A on a partial solution. International Journal of Parallel Programming, 48(4): 626–651, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-020-00669-9.

Granston:1995:LTP

[GMB95] Elana D. Granston, Thierry Montaut, and François Bodin. Loop transformations to prevent false sharing. International Journal of Parallel Programming, 23(4):263–301, August 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gendler:2006:PBM

[GMB06] Alexander Gendler, Avi Mendelson, and Yitzhak Birk. A PAB-based multi-prefetcher mechanism. International Journal of Parallel Programming, 34(2):171–188, April 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=2&spage=171.

Grasset:2011:MHD

[GMB^{+}11] Arnaud Grasset, Philippe Millet, Philippe Bonnot, Sami Yehia, Wolfram Putzke-Roeming, et al. The MOR-PHEUS heterogeneous dynamically reconfigurable platform. International Journal of Parallel Programming, 39(3):328–356, June 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=3&spage=328.

Giacalone:1989:FSI

[GMP89] Alessandro Giacalone, Prateek Mishra, and Sanjiva Prasad. FACILE: a symmetric integration of concurrent and functional programming. International Journal of Parallel Programming, 18(2):121– 160, April 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=2&spage=121.

Also in TAPSOFT ’89, ed. J. Diaz and F. Orejas, pp. 184-209, Springer-Verlag, Lecture Notes in Computer Science 352 (1989).

Gupta:2000:APR

[GMS00] Manish Gupta, Sayak Mukhopadhyay, and Navin Sinha. Automatic parallelization of recursive procedures. International Journal of Parallel Programming, 28(6):537–562, December 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=6&spage=537.

Gyllenhaal:1998:OMD

[GmWHR98] John C. Gyllenhaal, Wen mei W. Hwu, and B. Ramakrishna Rau. Optimization of machine descriptions for efficient use. International Journal of Parallel Programming, 26(4):417–447, August 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&

issn=0885-7458&volume=
26&issue=4&spage=417.

Geist:1989:TSP

[GN89] G. A. Geist and E. Ng.

Task scheduling for parallel sparse Cholesky factorization. International Journal of Parallel Programming, 18(4):291–314, August 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=4&spage=291.

Gnanasekar:2020:EMC

[GN20] A. K. Gnanasekar and V. Nagarajan. Efficient MAI cancellation scheme in MC-DS-CDMA using SIC. International Journal of Parallel Programming, 48(3):416–430, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Goldberg:1988:MEF

[Gol88] Benjamin Goldberg. Multiprocessor execution of functional programs. International Journal of Parallel Programming, 17 (5):425–473, October 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=

                                        article&issn=0885-7458&
                                        volume=17&issue=5&spage=
                                        425.

Girkar:1994:HTG

[GP94] Milind Girkar and Constantine D. Polychronopoulos. The hierarchical task graph as a universal intermediate representation. International Journal of Parallel Programming, 22(5):519–551, October 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Glowacz:2017:IDW

[GP17] Andrzej Glowacz and Marcin Pietroń. Implementation of digital watermarking algorithms in parallel hardware accelerators. International Journal of Parallel Programming, 45(5):1108–1127, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Govindarajan:2000:ECS

[GRAG00] R. Govindarajan, N. S. S. Narasimha Rao, E. R. Altman, and Guang R. Gao. Enhanced coscheduling: a software pipelining method using modulo-scheduled pipeline theory. International Journal of Parallel Programming, 28(1):1–46, February 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=1&spage=1.

Goes:2014:ASD

[GRC^{+}14] Luís Fabrício Wanderley Góes, Christiane Pousa Ribeiro, Márcio Castro, Jean-François Méhaut, Murray Cole, and Marcelo Cintra. Automatic skeletondriven memory affinity for transactional worklist applications. International Journal of Parallel Programming, 42(2): 365–382, April 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0253-x.

Grelck:2016:GEH

[Gre16] Clemens Grelck. Guest editorial for high-level parallel programming and applications. International Journal of Parallel Programming, 44(3): 383–385, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0423-8.pdf.

Grun:1998:SEP

[GRR98] Thomas Grün, Thomas Rauber, and Jochen Röhrig. Support for efficient programming on the SB-PRAM. International Journal of Parallel Programming, 26(3):209–240, June 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=3&spage=209.

Garcia:2017:ARA

[GRV^{+}17] Victor Garcia, Alejandro Rico, Carlos Villavieja, Paul Carpenter, Nacho Navarro, and Alex Ramirez. Adaptive runtime-assisted block prefetching on chipmultiprocessors. International Journal of Parallel Programming, 45(3):530– 550, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Greenlaw:1990:ASA

[GS90] Raymond Greenlaw and Lawrence Snyder. Achieving speedups for APL on an SIMD distributed memory machine. International Journal of Parallel Programming, 19(2):111– 127, April 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=2&spage=111.

Gaudiot:2005:MGE

[GS05] Jean-Luc Gaudiot and Siang Wun Song. Message from the Guest Editors. International Journal of Parallel Programming, 33(5):451–452, October 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=5&spage=451.

Grelck:2006:SFA

[GS06] Clemens Grelck and Sven-Bodo Scholz. SAC — a functional array language for efficient multi-threaded execution. International Journal of Parallel Programming, 34(4):383–427, August 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=4&spage=383.

Giacaman:2011:PIP

[GS11] Nasser Giacaman and Oliver Sinnen. Parallel iterator for parallelizing object-oriented applications. International Journal of Parallel Programming, 39(2):232–269, April 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=2&spage=232.

Giacaman:2013:PTP

[GS13] Nasser Giacaman and Oliver Sinnen. Parallel task for parallelising object-oriented desktop applications. International Journal of Parallel Programming, 41(5): 621–681, October 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0238-9.

Gao:2008:GEI

[GSA08] Guang R. Gao, Mitsuhisa Sato, and Eduard Ayguadé. Guest Editors introduction: Special issue on OpenMP. International Journal of Parallel Programming, 36(3): 287–288, June 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

issn=0885-7458&volume=
36&issue=3&spage=287.

Gschwind:2007:CBE

[Gsc07] Michael Gschwind. The Cell Broadband Engine: Exploiting multiple levels of parallelism in a chip multiprocessor. International Journal of Parallel Programming, 35(3):233– 262, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=3&spage=233.

Geng:2017:LED

[GSP^{+}17] Yuanzhen Geng, Xuanhua Shi, Cheng Pei, Hai Jin, and Wenbin Jiang. LCS: an efficient data eviction strategy for Spark. International Journal of Parallel Programming, 45 (6):1285–1297, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Grelck:2010:ASP

[GSS10] Clemens Grelck, Sven-Bodo Scholz, and Alex Shafarenko. Asynchronous stream processing with S-Net. International Journal of Parallel Programming, 38(1):38–67, February 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=1&spage=38.

Gu:2013:PCI

[GSY^{+}13] Zheng Gu, Matthew Small, Xin Yuan, Aniruddha Marathe, and David K. Lowenthal. Protocol customization for improving MPI performance on RDMA-enabled clusters. International Journal of Parallel Programming, 41(5):682–703, October 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
013-0242-0.

Groselj:1986:PAD

[GT86] B. Grošelj and C. Tropper. Pseudosimulation: an algorithm for distributed simulation with limited memory. International Journal of Parallel Programming, 15(5):413–456, October 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=5&spage=413.

Gupta:1988:PIO

[GTK^{+}88] Anoop Gupta, Milind Tambe, Dirk Kalp, Charles Forgy, and Allen Newell. Parallel implementation of OPS5 on the Encore multiprocessor: results and analysis. International Journal of Parallel Programming, 17(2):95–124, April 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=2&spage=95.

Granston:1995:CFD

[GV95] Elana D. Granston and Alexander V. Veidenbaum. Combining flow and dependence analyses to expose redundant array accesses. International Journal of Parallel Programming, 23(5):423–470, October 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gornish:1999:IHS

[GV99] Edward H. Gornish and Alexander Veidenbaum. An integrated hardware/ software data prefetching scheme for shared-memory multiprocessors. International Journal of Parallel Programming, 27(1):

35–70, February 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=1&spage=35.

Girbal:2006:SAC

[GVB^{+}06] Sylvain Girbal, Nicolas Vasilache, Cédric Bastoul, Albert Cohen, David Parello, Marc Sigler, and Olivier Temam. Semiautomatic composition of loop transformations for deep parallelism and memory hierarchies. International Journal of Parallel Programming, 34(3):261– 317, June 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=3&spage=261.

Grelck:2019:PAA

[GW19] Clemens Grelck and Heinrich Wiesinger. Persistent asynchronous adaptive specialization for generic array programming. International Journal of Parallel Programming, 47(2): 164–183, April 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/

content/pdf/10.1007/
s10766-018-0567-9.pdf.

Gao:2019:ASG

[GWHY19] Jianliang Gao, Jianxin Wang, Jianbiao He, and Fengxia Yan. Against signed graph deanonymization attacks on social networks. International Journal of Parallel Programming, 47(4):725–739, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Gesper:2021:EDP

[GWPV21] Sven Gesper, Moritz Weißbrich, and Guillermo Payá-Vayá. Evaluation of different processor architecture organizations for on-site electronics in harsh environments. International Journal of Parallel Programming, 49(4): 541–569, August 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00686-8.

Gao:2018:SRF

[GWYQ18] Jian Gao, Hongmei Wei, Kang Yu, and Peng Qing. A scalable runtime fault localization framework for high-performance computing systems. International Journal of Parallel Programming, 46(4):749–761, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Govindarajan:1992:AGP

[GYL92] R. Govindarajan, S. Yu, and V. S. Lakshmanan. Attempting guards in parallel: a data flow approach to execute generalized guarded commands. International Journal of Parallel Programming, 21(4):225–268, August 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=4&spage=225.

Guo:2022:EEI

[GYP22] Jichi Guo, Qing Yi, and Kleanthis Psarris. Enhancing the effectiveness of inlining in automatic parallelization. International Journal of Parallel Programming, 50(1): 65–88, February 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00722-1.

Gilbert:1987:PGP

[GZ87] John R. Gilbert and Earl Zmijewski. A parallel graph partitioning algorithm for a messagepassing multiprocessor. International Journal of Parallel Programming, 16 (6):427–449, December 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=16&issue=6&spage=
427.

Gouk:2018:ERL

[GZJ18] Donghyun Gouk, Jie Zhang, and Myoungsoo Jung. Enabling realistic logical device interface and driver for NVM express enabled full system simulations. International Journal of Parallel Programming, 46(4):710–721, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hussain:2011:PIA

[HAA^{+}11] Masroor Hussain, Muhammad Abid, Mushtaq Ahmad, Ashfaq Khokhar, and Arif Masud. A parallel implementation of ALE moving mesh technique for FSI problems using OpenMP. International Journal of Parallel Programming, 39(6):717–745, December 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=6&spage=717.

Halstead:1986:AML

[Hal86] Robert H. Halstead, Jr. An assessment of Multilisp — lessons from experience. International Journal of Parallel Programming, 15(6):459–501, December 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        15&issue=6&spage=459.

(Due to publishing delays, this issue did not appear until late 1987.).

Hakimi:2023:HML

[HBC23] Yacine Hakimi, Riyadh Baghdadi, and Yacine Challal. A hybrid machine learning model for code optimization. International Journal of Parallel Programming, 51(6): 309–331, December 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-023-00758-5.

Han:2017:SLS

[HC17] Yiming Han and Anthony T. Chronopoulos. Scalable loop selfscheduling schemes for large-scale clusters and cloud systems. International Journal of Parallel Programming, 45(3):595– 611, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hao:1998:IIF

[HCEP98] Eric Hao, Po-Yung Chang, Marius Evers, and Yale N. Patt. Increasing the instruction fetch rate via block-structured instruction set architectures. International Journal of Parallel Programming, 26(4):449–478, August 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=4&spage=449.

Herrmann:2024:OTD

[HDK24] Nina Herrmann, Justus Dieckmann, and Herbert Kuchen. Optimizing three-dimensional stencil-operations on heterogeneous computing environments. International Journal of Parallel Programming, 52(4): 274–297, August 2024. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-024-00769-w.

Herrmann:2022:SCA

[HdMMK22] Nina Herrmann, Breno A. de Melo Menezes, and Herbert Kuchen. Stencil calculations with algorithmic skeletons for heterogeneous computing environments. International Journal of Parallel Programming, 50(5–6): 433–453, December 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00735-4.

Hemmendinger:1989:IMS

[Hem89] David Hemmendinger. Initializing memory shared by several processors. International Journal of Parallel Programming, 18 (4):241–253, August 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=18&issue=4&spage=
                                        241.

Henriksen:2021:BCG

[Hen21] Troels Henriksen. Bounds checking on GPU. International Journal of Parallel Programming, 49(6):

761–775, December 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00703-4.

Harris:2006:GEI

[HF06] Ian G. Harris and Franco Fummi. Guest editorś introduction. International Journal of Parallel Programming, 34(1):1– 2, March 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=1&spage=1.

Holmen:2014:ASI

[HF14a] John K. Holmen and David L. Foster. Accelerating single iteration performance of CUDA–based 3D reaction–diffusion simulations. International Journal of Parallel Programming, 42(2):343–363, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-

013-0251-z. See erratum [HF14b].

Holmen:2014:EAS

[HF14b] John K. Holmen and David L. Foster. Erratum to: Accelerating single iteration performance of CUDA–based 3D reaction–diffusion simulations. International Journal of Parallel Programming, 42(2):364, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://link.springer.
com/content/pdf/10.1007/
s10766-014-0305-x.pdf.

See [HF14a].

Hensgen:1988:TAB

[HFM88] Debra Hensgen, Raphael Finkel, and Udi Manber. Two algorithms for barrier synchronization. International Journal of Parallel Programming, 17(1):1– 17, February 1988. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=1&spage=1.

Haidl:2018:HLP

[HG18] Michael Haidl and Sergei Gorlatch. High-level programming for manycores using C++14 and the STL. International Journal of Parallel Programming, 46(1):23–41, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Huang:2012:EEP

[HGT^{+}12] Yan Huang, Zhi-Min Gu, Jie Tang, Min Cai, Jianxun Zhang, et al. Estimating effective prefetch distance in threaded prefetching for linked data structures. International Journal of Parallel Programming, 40(5):465–487, October 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=5&spage=465.

Huang:2015:ACH

[HHC^{+}15] Hui Huang, Ligang He, Xueguang Chen, Minghui Yu, and Zhiwu Wang. Automatic composition of heterogeneous models based on semantic Web services. International Journal of Parallel Programming, 43(3): 339–358, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0279-0.

Haase:2010:SDV

[HHW10] Jan Haase, Andreas Hofmann, and Klaus Waldschmidt. A self distributing virtual machine for adaptive multicore environments. International Journal of Parallel Programming, 38(1):19–37, February 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=1&spage=19.

He:2020:HDS

[HHW20] Zeyu He, Qiuli Huang, and Chuliang Weng. Handling data skew for aggregation in Spark SQL using task stealing. International Journal of Parallel Programming, 48(6): 941–956, December 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00657-z.

Hains:2014:GEH

[HK14] Gaetan Hains and Youry Khmelevsky. Guest editorial for high-level parallel programming and applications. International Journal of Parallel Programming, 42(4):525–528, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-013-0297-y.pdf.

Herrmann:2023:DCA

[HK23] Nina Herrmann and Herbert Kuchen. Distributed calculations with algorithmic skeletons for heterogeneous computing environments. International Journal of Parallel Programming, 51(2–3):172– 185, June 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00742-5.

Horn:2018:GPI

[HKJ^{+}18] William Horn, Manoj Kumar, Joefon Jann, José Moreira, Pratap Pattnaik, Mauricio Serrano, Gabriel Tanase, and Hao Yu. Graph Programming Interface (GPI): a linear algebra programming model for large scale graph computations. International Journal of Parallel Programming, 46(2):412–440, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hu:2021:CHA

[HL21] Xiao Hu and Zhonghai Lu. A configurable hardware architecture for runtime application of network calculus. International Journal of Parallel Programming, 49(5):

745–760, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00700-7.

Hudak:2009:CSI

[HLK^{+}09] David E. Hudak, Neil Ludban, Ashok Krishnamurthy, Vijay Gadepally, Siddharth Samsi, et al. A computational science IDE for HPC systems: Design and applications. International Journal of Parallel Programming, 37(1):91– 105, February 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=91.

Hawick:2011:RLS

[HLP11] K. A. Hawick, A. Leist, and D. P. Playne. Regular lattice and smallworld spin model simulations using CUDA and GPUs. International Journal of Parallel Programming, 39(2):183–201, April 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=2&spage=183.

Hsu:2015:NPC

[HLS15] Ching-Hsien Hsu, Xiaoming Li, and Xuanhua Shi. Network and parallel computing. International Journal of Parallel Programming, 43(3): 311–315, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-014-0311-z.pdf.

Hamidouche:2013:PSW

[HMF^{+}13] Khaled Hamidouche, Fernando Machado Mendonca, Joel Falcou, Alba Cristina Magalhaes Alves de Melo, and Daniel Etiemble. Parallel Smith– Waterman comparison on multicore and manycore computing platforms with BSP++. International Journal of Parallel Programming, 41(1):111–136, February 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0209-6.

Hilbrich:2009:MCC

[HMK09] Tobias Hilbrich, Matthias S.

Müller, and Bettina Krammer. MPI correctness checking for OpenMP/ MPI applications. International Journal of Parallel Programming, 37(3): 277–291, June 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=3&spage=277.

Harel:2020:SSP

[HML^{+}20] Re’em Harel, Idan Mosseri, Harel Levin, Lee or Alon, Matan Rusanovsky, and Gal Oren. Source-to-source parallelization compilers for scientific sharedmemory multi-core and accelerated multiprocessing: Analysis, pitfalls, enhancement and potential. International Journal of Parallel Programming, 48(1):1– 31, February 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hum:1996:SEM

[HMT^{+}96] Herbert H. J. Hum, Olivier Maquelin, Kevin B. Theobald, Xinmin Tian, Guang R. Gao, and Laurie J. Hendren. A study of the EARTH-MANNA multithreaded system. International Journal of Parallel Programming, 24 (4):319–348, August 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hank:1997:RBC

[HmWHR97] Richard E. Hank, Wen mei W. Hwu, and B. Ramakrishna Rau. Regionbased compilation: Introduction, motivation, and initial experience. International Journal of Parallel Programming, 25(2):113– 146, April 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hwu:1994:GE

[HN94] Wen-Mei Hwu and Alex Nicolau. From the Guest Editors. International Journal of Parallel Programming, 22(3):207–??, June 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Huang:2016:CFL

[HNC^{+}16] Xiaomeng Huang, Yufang Ni, Dexun Chen, Songbin Liu, Haohuan Fu, and Guangwen Yang. Czip: a fast lossless compression algorithm for climate data. International Journal of Parallel Programming, 44(6):1248–1267, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0403-z.

Holobar:2006:DJJ

[HOZ06] Ales Holobar, Milan Ojstersek, and Damjan Zazula. Distributed Jacobi joint diagonalization on clusters of personal computers. International Journal of Parallel Programming, 34(6):509–530, December 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=6&spage=509.

Heinecke:2013:EAE

[HP13] Alexander Heinecke and Dirk Pflüger. Emerging architectures enable to boost massively parallel data mining using adaptive sparse grids. International Journal of Parallel Programming, 41(3): 357–399, June 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0202-0.

Hidalgo-Paniagua:2015:CSP

[HPVRPF15] Alejandro Hidalgo-Paniagua, Miguel A. Vega-Rodríguez, Nieves Pavón, and Joaquín Ferruz. A comparative study of parallel RANSAC implementations in 3D space. International Journal of Parallel Programming, 43(5): 703–720, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0316-7.

Hoeflinger:2001:UIP

[HPY01] Jay P. Hoeflinger, Yunheung Paek, and Kwang Yi. Unified interprocedural parallelism detection. International Journal of Parallel Programming, 29(2):185–215, April 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/20/3/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/20/3/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=2&spage=
185.

Hoffmann:2011:ATP

[HR11] Ralf Hoffmann and Thomas Rauber. Adaptive task pools: Efficiently balancing large number of tasks on shared-address spaces. International Journal of Parallel Programming, 39(5):553–581, October 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=5&spage=553.

Hallou:2017:RVT

[HRC17] Nabil Hallou, Erven Rohou, and Philippe Clauss. Runtime vectorization transformations of binary code. International Journal of Parallel Programming, 45 (6):1536–1565, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hassanein:2008:AEH

[HRH08] Wessam M. Hassanein, Layali K. Rashid, and Moustafa A. Hammad. Analyzing the effects of hyperthreading on the performance of data management systems. International Journal of Parallel Programming, 36(2):206– 225, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=2&spage=206.

Hsu:2016:NPC

[HS16] Ching-Hsien Hsu and Valentina Salapura. Network and parallel computing. International Journal of Parallel Programming, 44(1):1–4, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0345-2;
http://link.springer.
com/content/pdf/10.1007/
s10766-014-0345-2.pdf.

Habel:2016:CDC

[HSCI^{+}16] Rachid Habel, Frédérique Silber-Chaussumier, François Irigoin, Elisabeth Brunet, and François Trahay. Combining data and computation distribution directives for hybrid parallel programming: a transformation system. International Journal of Parallel Programming, 44(6):1268– 1295, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0428-3.

Heidorn:2024:HAE

[HSM^{+}24] Christian Heidorn, Muhammad Sabih, Nicolai Meyerhöfer, Christian Schinabeck, Jürgen Teich, and Frank Hannig. Hardwareaware evolutionary explainable filter pruning for convolutional neural networks. International Journal of Parallel Programming, 52(1–2): 40–58, April 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-024-00760-5.

Hua:2019:GES

[HSXH19] Qiang-Sheng Hua, Xuanhua Shi, Yinglong Xia, and Howie Huang. Guest editorial: Special issue on algorithms and systems on big graph processing. International Journal of Parallel Programming, 47(4):641– 643, August 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-019-00635-0.pdf.

Holzenspies:2010:RTS

[HtBK^{+}10] Philip K. F. Hölzenspies, Timon D. ter Braak, Jan Kuper, Gerard J. M. Smit, and Johann M. Hurink. Run-time spatial mapping of streaming applications to heterogeneous multiprocessor systems. International Journal of Parallel Programming, 38(1): 68–83, February 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&

issn=0885-7458&volume=
38&issue=1&spage=68.

Hukerikar:2018:RIA

[HTDL18] Saurabh Hukerikar, Keita Teranishi, Pedro C. Diniz, and Robert F. Lucas. RedThreads: an interface for application-level fault detection/correction through adaptive redundant multithreading. International Journal of Parallel Programming, 46 (2):225–251, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Han:1998:EBS

[HTK98] Hwansoo Han, Chau-Wen Tseng, and Pete Keleher. Eliminating barrier synchronization for compilerparallelized codes on software DSMs. International Journal of Parallel Programming, 26(5):591–612, October 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=5&spage=591.

Huang:2012:POT

[HTmG^{+}12] Yan Huang, Jie Tang, Zhi min Gu, Min Cai, Jianxun Zhang, and Ninghan Zheng. The performance optimization of threaded prefetching for linked data structures. International Journal of Parallel Programming, 40(2):141–163, April 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=2&spage=141.

Hendren:1997:CCE

[HTZ^{+}97] Laurie J. Hendren, Xinan Tang, Yingchun Zhu, Shereen Ghobrial, Guang R. Gao, Xun Xue, Haiying Cai, and Pierre Ouellet. Compiling C for the EARTH multithreaded architecture. International Journal of Parallel Programming, 25(4):305–338, August 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Huang:1989:SEP

[Hua89] Xiaoqiu Huang. A space– efficient parallel sequence comparison algorithm for a message–passing multiprocessor. International Journal of Parallel Programming, 18(3):223–239, June 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

issn=0885-7458&volume=
18&issue=3&spage=223.

Hudak:1986:DSP

[Hud86] Paul R. Hudak. The denotational semantics of a para-functional programming language. International Journal of Parallel Programming, 15(2):103– 125, April 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=2&spage=103.

Also Research Report YALEU/DCS/TR-484, Department of Computer Science, Yale University, New Haven, CT (1986).

Huelsbergen:1997:DRR

[Hue97] Lorenz Huelsbergen. Dynamic resolution: a runtime technique for the parallelization of modifications to directed acyclic graphs. International Journal of Parallel Programming, 25(5):385–417, October 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Hunt:1987:EAU

[Hun87] Frances E. Hunt. Experiments with applicative updating: practical results. International Journal of Parallel Programming, 16(4):279–303, August 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=4&spage=279.

Huntbach:1991:PBB

[Hun91] Matthew Huntbach. Parallel branch-and-bound search in Parlog. International Journal of Parallel Programming, 20(4):299– 314, August 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        20&issue=4&spage=299.

He:2018:GRA

[HVF18] Mengda He, Viktor Vafeiadis, and Shengchao Qin João F. Ferreira. GPS+: Reasoning about fences and relaxed atomics. International Journal of Parallel Programming, 46(6):1157– 1183, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Haj-Yihia:2018:SSE

[HYBA18] Jawad Haj-Yihia and Yosi Ben-Asher. Software static energy modeling for modern processors. International Journal of Parallel Programming, 46(2):284– 312, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Han:2016:SSB

[HZL16] Rui Han, Jianfeng Zhan, and Jose Vazquez-Poletti Luis. SARP: Synopsis– based approximate request processing for low latency and small correctness loss in cloud online services. International Journal of Parallel Programming, 44(5): 1054–1077, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0406-9.

Han:2019:FFE

[HZZ^{+}19] Dong Han, Shengyuan Zhou, Tian Zhi, Yibo Wang, and Shaoli Liu. Float-Fix: an efficient and hardware-friendly data type for deep neural network. International Journal of Parallel Programming, 47(3):345–359, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Han:2020:CAD

[HZZS20] Tao Han, Miaowang Zeng, Lijuan Zhang, and Arun Kumar Sangaiah. A channelaware duty cycle optimization for node-to-node communications in the Internet of medical things. International Journal of Parallel Programming, 48 (2):264–279, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Iooss:2021:MTP

[IAR21] Guillaume Iooss, Christophe Alias, and Sanjay Rajopadhye. Monoparametric tiling of polyhedral programs. International Journal of Parallel Programming, 49(3): 376–409, June 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00694-2.

Imre:2011:ESR

[IBA11] Kayhan M. Imre, Cesur Baransel, and Harun Artuner. Efficient and scalable routing algorithms for collective communication operations on 2D all-port torus networks. International Journal of Parallel Programming, 39(6):746– 782, December 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

issn=0885-7458&volume=
39&issue=6&spage=746.

Issenin:2008:UFM

[ID08] Ilya Issenin and Nikil Dutt. Using FORAY models to enable MP-SoC memory optimizations. International Journal of Parallel Programming, 36(1):93–113, February 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=1&spage=93.

Iwasaki:2004:NPS

[IH04] Hideya Iwasaki and Zhenjiang Hu. A new parallel skeleton for general accumulative computations. International Journal of Parallel Programming, 32(5):389–414, October 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=5&spage=389.

Ishizaki:2000:LTA

[IKN00] Kazuaki Ishizaki, Hideaki Komatsu, and Toshio Nakatani. A loop transformation algorithm for communication overlapping.

International Journal of Parallel Programming, 28 (2):135–154, April 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=28&issue=2&spage=
135.

Introduction:1998:EA

[Int98] Editorial Introduction. Editor’s announcement. International Journal of Parallel Programming, 26 (1):1–2, February 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=26&issue=1&spage=
1.

Ibarra:1990:EAP

[IP90] Oscar H. Ibarra and Michael A. Palis. An efficient all-parses systolic algorithm for general context-free parsing. International Journal of Parallel Programming, 19(4):295–331, August 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=4&spage=295.

Iyer:2005:EEH

[IPR^{+}05] Ravi Iyer, Jack Perdue, Lawrence Rauchwerger, Nancy M. Amato, and Laxmi Bhuyan. An experimental evaluation of the HP V-Class and SGI Origin 2000 multiprocessors using microbenchmarks and scientific applications. International Journal of Parallel Programming, 33(4):307–350, August 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=4&spage=307.

Iqbal:1991:AAP

[Iqb91] Mohammad Ashraf Iqbal.

Approximate algorithms for partitioning problems. International Journal of Parallel Programming, 20(5):341–361, October 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=5&spage=341.

Irandoost:2019:MDS

[IR19] Mohammad Amin Irandoost and Amir Masoud Rahmani. MapReduce data skewness handling: a systematic literature review. International Journal of Parallel Programming, 47(5–6): 907–950, December 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00627-0.

Iwashita:2003:BRB

[IS03] Takeshi Iwashita and Masaaki Shimasaki. Block red-black ordering: a new ordering strategy for parallelization of ICCG method. International Journal of Parallel Programming, 31(1):55–75, February 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/
4773/31/5/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/31/5/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=1&spage=
55.

Jeon:2020:GES

[JACK20] Gwanggil Jeon, Awais Ahmad, Salvatore Cuomo, and Burak Kantarci. Guest editorial: Special issue on emerging technology for software define network enabled Internet of Things. International Journal of Parallel Programming, 48(2):157–161, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jakobsson:2019:ACA

[Jak19] Arvid Jakobsson. Automatic cost analysis for imperative BSP programs. International Journal of Parallel Programming, 47 (2):184–212, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jannesari:2015:DHL

[Jan15] Ali Jannesari. Detection of high-level synchronization anomalies in parallel programs. International Journal of Parallel Programming, 43(4): 656–678, August 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0313-x.

Jin:2017:PAA

[JAW17] Hai Jin, Aaqif Afzaal Abbasi, and Song Wu. Pathfinder: Applicationaware distributed path computation in clouds. International Journal of Parallel Programming, 45 (6):1273–1284, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

John:1998:CCP

[JB98] Ajita John and James C. Browne. Compilation of constraint programs with noncyclic and cyclic dependencies to procedural parallel programs. International Journal of Parallel Programming, 26(1): 65–119, February 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=26&issue=1&spage=
65.

Janjic:2021:RLP

[JBB21] Vladimir Janjic, Christopher Brown, and Adam D. Barwell. Restoration of legacy parallelism: Transforming Pthreads into farm and pipeline patterns. International Journal of Parallel Programming, 49(6):886–910, December 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00716-z.

Jimborean:2014:DSP

[JCD^{+}14] Alexandra Jimborean, Philippe Clauss, Jean-François Dollinger, Vincent Loechner, and Juan Manuel Martinez Caamaño. Dynamic and speculative polyhedral parallelization using compiler-generated skeletons. International Journal of Parallel Programming, 42(4):529–545, August 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0259-4.

Jin:2008:PEM

[JCH^{+}08] Haoqiang Jin, Barbara Chapman, Lei Huang, Dieter an Mey, and Thomas Reichstein. Performance evaluation of a multi-zone application in different OpenMP approaches. International Journal of Parallel Programming, 36(3):312–325, June 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=3&spage=312.

Ju:2018:DDB

[JCW^{+}18] Xiangyu Ju, Quan Chen, Zhenning Wang, Minyi Guo, and Guang R. Gao. DCF: a dataflowbased collaborative filtering training algorithm. International Journal of Parallel Programming, 46 (4):686–698, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jin:2020:DOA

[JDF20] Kang Jin, Dezun Dong, and Binzhang Fu. Dancer-Fly: An order-aware network-on-chip router onthe-fly mitigating multipath packet reordering. International Journal of Parallel Programming, 48 (4):730–749, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-019-00648-9.

Jungblut:2021:PNL

[JF21] Pascal Jungblut and Karl Fürlinger. Portable nodelevel parallelism for the PGAS model. International Journal of Parallel Programming, 49(6): 867–885, December 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00718-x.

Jenks:1997:ELT

[JG97] Stephen Jenks and Jean-Luc Gaudiot. Exploiting locality and tolerating remote memory access latency using thread migration. International Journal of Parallel Programming, 25(4):281–304, August 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Janakiram:1988:RPB

[JGA^{+}88] V. K. Janakiram, E. F.

Gehringer, D. P. Agrawal, Mehrotra, and R. A randomized parallel branchand-bound algorithm. International Journal of Parallel Programming, 17 (3):277–301, June 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jodra:2015:ETG

[JGM15] Jose L. Jodra, Ibai Gurrutxaga, and Javier Muguerza. Efficient 3D transpositions in graphics processing units. International Journal of Parallel Programming, 43(5): 876–891, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0366-5.

Jaeger:2018:FAP

[JGP^{+}18] David Jaeger, Hendrik Graupner, Chris Pelchen, Feng Cheng, and Christoph Meinel. Fast automated processing and evaluation of identity leaks. International Journal of Parallel Programming, 46(2):441– 470, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jiang:2020:MPO

[JGZ^{+}20] Ziyue Jiang, Yifan Gong, Jidong Zhai, Yu-Ping Wang, Wei Liu, Hao Wu, and Jiangming Jin. Message passing optimization in robot operating system. International Journal of Parallel Programming, 48(1):119–136, February 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jung:2001:TPB

[JHLM01] Inbum Jung, Jongwoong Hyun, Joonwon Lee, and Joongsoo Ma. Two-phase barrier: a synchronization primitive for improving the processor utilization. International Journal of Parallel Programming, 29(6):607–627, December 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/24/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/

4773/24/2/fulltext.pdf;
http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=6&spage=
607.

Jin:2015:IPB

[JJIL15] Hai Jin, Honglei Jiang, Shadi Ibrahim, and Xiaofei Liao. Inaccuracy in private BitTorrent measurements. International Journal of Parallel Programming, 43 (3):528–547, June 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0282-5.

Jayaraman:1986:PRM

[JK86] Bharat Jayaraman and Robert M. Keller. Primitives for resource management in a demand-driven reduction model. International Journal of Parallel Programming, 15(3): 215–244, June 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=3&spage=215.

Jeon:2003:PMS

[JK03] Minsoo Jeon and Dongseung Kim. Parallel merge sort with load balancing. International Journal of Parallel Programming, 31(1):21–33, February 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/31/3/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/31/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=1&spage=
21.

Jalan:2012:TTW

[JK12] Rohit Jalan and Arun Kejariwal. Trin–Trin: Who’s calling? A pin-based dynamic call graph extraction framework. International Journal of Parallel Programming, 40(4):410– 442, August 2012. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=4&spage=410.

Jin:2019:HHA

[JLDF19] Kang Jin, Cunlu Li, Dezun Dong, and Binzhang Fu. HARE: History-aware adaptive routing algorithm for endpoint congestion in networks-on-chip. International Journal of Parallel Programming, 47(3): 433–450, June 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0614-6.pdf.

Jindal:2016:EGS

[JLDS16] Nakul Jindal, Victor Lotrich, Erik Deumens, and Beverly A. Sanders. Exploiting GPUs with the super instruction architecture. International Journal of Parallel Programming, 44(2):309– 324, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0319-4.

Jia:2018:IPD

[JLJ^{+}18] Chengfan Jia, Junnan Liu, Xu Jin, Han Lin, Hong An, Wenting Han, Zheng Wu, and Mengxian Chi. Improving the performance of distributed TensorFlow with RDMA. International Journal of Parallel Programming, 46(4):674– 685, August 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jelic:2015:FPI

[JLMW15] Slobodan Jelić, Sören Laue, Domagoj Matijević, and Patrick Wijerama. A fast parallel implementation of a PTAS for fractional packing and covering linear programs. International Journal of Parallel Programming, 43(5): 840–875, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0352-y.

Ji:2021:SMN

[JLV21] Haonan Ji, Shibo Lu, and Brian Vinter. Segmented merge: A new primitive for parallel sparse matrix computations. International Journal of Parallel Programming, 49(5): 732–744, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00695-1.

Jeevan:2020:DDT

[JM20] A. N. Gnana Jeevan and M. A. Maluk Mohamed. DyTO: Dynamic task offloading strategy for mobile cloud computing using surrogate object model. International Journal of Parallel Programming, 48 (3):399–415, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Joisha:2002:EAJ

[JMSG02] Pramod G. Joisha, Samuel P.

Midkiff, Mauricio J. Serrano, and Manish Gupta. Efficiently adapting Java binaries in limited memory contexts. International Journal of Parallel Programming, 30(4):257– 289, August 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/28/3/abstract.htm;
http://ipsapp009.lwwonline.
com/content/getfile/
4773/28/3/fulltext.pdf;
http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=4&spage=
257.

Joe:1999:GEI

[Joe99] Kazuki Joe. Guest Editor’s introduction. International Journal of Parallel Programming, 27(2): 71–72, April 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=2&spage=71.

Joe:2003:GEI

[Joe03] Kazuki Joe. Guest Editor’s introduction. International Journal of Parallel Programming, 31(1):1– 2, February 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/
4773/31/1/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/
4773/31/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=1&spage=
1.

Johnson:1994:PAM

[Joh94] Theodore Johnson. Parallelaccess memory management using fast-fits. International Journal of Parallel Programming, 22(6): 617–648, December 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Javed:2016:TSJ

[JQJ^{+}16] Ansar Javed, Bibrak Qamar, Mohsan Jameel, Aamir Shafi, and Bryan Carpenter. Towards scalable Java HPC with hybrid and native communication devices in MPJ express. International Journal of Parallel Programming, 44(6):1142–1172, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

com/article/10.1007/
s10766-015-0375-4.

Jin:2015:CCC

[JQWG15] Hai Jin, Hanfeng Qin, Song Wu, and Xuerong Guo. CCAP: a cache contention-aware virtual machine placement approach for HPC cloud. International Journal of Parallel Programming, 43 (3):403–420, June 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0286-1.

Jesshope:2006:GEI

[JS06a] Chris Jesshope and Alex Shafarenko. Guest Editor’s introduction á <part 2>. International Journal of Parallel Programming, 34(4):319–322, August 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=4&spage=319.

Jesshope:2006:SIM

[JS06b] Chris Jesshope and Alex Shafarenko. Special issue on micro-grids — Guest Editor introduction. International Journal of Parallel Programming, 34(3):

189–192, June 2006. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        34&issue=3&spage=189.

Jeyapaul:2010:CTT

[JS10] Reiley Jeyapaul and Aviral Shrivastava. Code transformations for TLB power reduction. International Journal of Parallel Programming, 38(3–4): 254–276, June 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=3&spage=254.

Jourdan:1997:RRB

[JSHP97] Stephan Jourdan, Jared Stark, Tse-Hao Hsing, and Yale N. Patt. Recovery requirements of branch prediction storage structures in the presence of mispredicted-path execution. International Journal of Parallel Programming, 25(5):363–383, October 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Jaaskelainen:2015:PPP

[JSS^{+}15] Pekka Jääskeläinen, Carlos Sánchez de La Lama, Erik Schnetter, Kalle Raiskila, Jarmo Takala, and Heikki Berg. pocl: a performance-portable OpenCL implementation. International Journal of Parallel Programming, 43 (5):752–785, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0320-y.

Jannesari:2016:AGU

[JW16] Ali Jannesari and Felix Wolf. Automatic generation of unit tests for correlated variables in parallel programs. International Journal of Parallel Programming, 44(3): 644–662, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0363-8.

Khan:2020:EOP

[KAI20] Fakhri Alam Khan, Awais Ahmad, and Muhammad Imran. Energy optimization of PR–LEACH routing scheme using distance awareness in Internet of Things networks. International Journal of Parallel Programming, 48(2):244– 263, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kapinos:2010:PPP

[KaM10] Paul Kapinos and Dieter an Mey. Productivity and performance portability of the OpenMP 3.0 tasking concept when applied to an engineering code written in Fortran 95. International Journal of Parallel Programming, 38(5–6):379– 395, October 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=5&spage=379.

Khan:2017:RCS

[KAMAMA17] Ayaz H. Khan, Mayez Al-Mouhamed, Muhammed Al-Mulhem, and Adel F. Ahmed. RT-CUDA: a software tool for CUDA code restructuring. International Journal of Parallel Programming, 45(3):551– 594, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kasif:1986:CDD

[Kas86] S. Kasif. Control and data driven execution of logic programs: A comparison. International Journal of Parallel Programming, 15(1):73–99, February 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=1&spage=73.

Kachris:2003:RLB

[KBD03] C. Kachris, N. Bourbakis, and A. Dollas. A reconfigurable logic-based processor for the SCAN image and video encryption algorithm. International Journal of Parallel Programming, 31 (6):489–506, December 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/6/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/6/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=6&spage=
489.

Kriaa:2008:PPM

[KBG^{+}08] Lobna Kriaa, Aimen Bouchhima, Marius Gligor, Anne-Marie Fouillart, Fréderic Pétrot, and Ahmed-Amine Jerraya. Parallel programming of multi-processor SoC: a HW–SW interface perspective. International Journal of Parallel Programming, 36(1): 68–92, February 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=1&spage=68.

Kumar:2005:CTP

[KCW^{+}05] Naveen Kumar, Bruce R. Childers, Daniel Williams, Jack W. Davidson, and Mary Lou Soffa. Compiletime planning for overhead reduction in software dynamic translators. International Journal of Parallel Programming, 33(2–3): 103–114, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=2&spage=103.

Keramidas:2015:RCR

[KD15] Georgios Keramidas and Chrysovalantis Datsios. Revisiting cache resizing. International Journal of Parallel Programming, 43 (1):59–85, February 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0291-4.

Kelefouras:2022:MET

[KDV22] Vasilios Kelefouras, Karim Djemame, and Nikolaos Voros. A methodology for efficient tile size selection for affine loop kernels. International Journal of Parallel Programming, 50(3–4):405– 432, August 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00734-5.

Kultursay:2016:MPL

[KEKK16] Emre Kültürsay, Kemal Ebcioglu, Gürhan Küçük, and Mahmut T. Kandemir. Memory partitioning in the limit. International Journal of Parallel Programming, 44(2): 337–380, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0380-7.

Kennedy:1994:CTM

[Ken94] Ken Kennedy. Compiler technology for machineindependent parallel programming. International Journal of Parallel Programming, 22(1):79–98, February 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kennedy:2001:FGW

[Ken01] Ken Kennedy. Fast greedy weighted fusion. International Journal of Parallel Programming, 29(5):463– 491, October 2001. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/23/2/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/23/2/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=5&spage=
                                        463.

Kessler:2020:GEN

[Kes20] Christoph Kessler. Guest Editor’s note: High-Level Parallel Programming 2019. International Journal of Parallel Programming, 48(4): 581–582, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-020-00672-0.

Kistler:1999:TBA

[KF99] Thomas Kistler and Michael Franz. A tree-based alternative to Java bytecodes. International Journal of Parallel Programming, 27(1):21–33, February 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=1&spage=21.

Kolberg:2008:DLS

[KFC08] Mariana Luderitz Kolberg, Luiz Gustavo Fernandes, and Dalcidio Moraes Claudio. Dense linear system: a parallel self-verified solver. International Journal of Parallel Programming, 36(4):412–425, August 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=4&spage=412.

Koster:2020:MPR

[KGK20] M. Köster, J. Groß, and A. Krüger. Massively parallel rule-based interpreter execution on GPUs using thread compaction. International Journal of Parallel Programming, 48 (4):675–691, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

                                        com/article/10.1007/
                                        s10766-020-00670-2.

Kannan:2018:FPT

[KH18] Venkatesh Kannan and G. W. Hamilton. Functional program transformation for parallelisation using skeletons. International Journal of Parallel Programming, 46(1): 152–172, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kalla:2008:FFC

[KHH08] Praveen Kalla, X. Sharon Hu, and Jörg Henkel. A flexible framework for communication evaluation in SoC design. International Journal of Parallel Programming, 36(5):457– 477, October 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=5&spage=457.

Kawada:2021:TRA

[KHT21] Tomoaki Kawada, Shinya Honda, and Hiroaki Takada. TZmCFI: RTOS-aware control-flow integrity using TrustZone for Armv8-M. International Journal of Parallel Programming, 49(2):216–236, April 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00673-z.

Khan:2020:BDP

[KIT^{+}20] Murad Khan, Javed Iqbal, Muhammad Talha, Muhammad Arshad, Muhammad Diyan, and Kijun Han. Big data processing using Internet of Software Defined Things in smart cities. International Journal of Parallel Programming, 48(2):178–191, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ko:2014:SPD

[KJHB14] Yousun Ko, Minyoung Jung, Yo-Sub Han, and Bernd Burgstaller. A speculative parallel DFA membership test for multicore, SIMD and cloud computing environments. International Journal of Parallel Programming, 42 (3):456–489, June 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0258-5.

Kumar:2010:FBH

[KJPN10] Vinay B. Y. Kumar, Siddharth Joshi, Sachin B. Patkar, and H. Narayanan.

FPGA based high performance double-precision matrix multiplication. International Journal of Parallel Programming, 38 (3–4):322–338, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=38&issue=3&spage=
322.

Kella:2011:AAP

[KK11] Kush K. Kella and Aasia Khanum. APCFS: Autonomous and Parallel Compressed File System. International Journal of Parallel Programming, 39(4):522–532, August 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=4&spage=522.

Kaur:2020:FAE

[KK20] Puneet Jai Kaur and Sakshi Kaushal. A fuzzy approach for estimating quality of aspect oriented systems. International Journal of Parallel Programming, 48(5): 850–869, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0618-2.

Kokhazadeh:2024:PAE

[KKKS24] Milad Kokhazadeh, Georgios Keramidas, Vasilios Kelefouras, and Iakovos Stamoulis. A practical approach for employing tensor train decomposition in edge devices. International Journal of Parallel Programming, 52 (1–2):20–39, April 2024. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00762-3.

Kirovski:1999:PBP

[KKMS99] Darko Kirovski, Johnson Kin, and William H. Mangione-Smith. Procedure based program compression. International Journal of Parallel Programming, 27(6):457–475, December 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=6&spage=457.

Kaur:2018:FAR

[KKSP18] Puneet Jai Kaur, Sakshi Kaushal, Arun Kumar Sangaiah, and Francesco Piccialli. A framework for assessing reusability using package cohesion measure in aspect oriented systems. International Journal of Parallel Programming, 46(3):543–564, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kavadias:2012:CIN

[KKZN12] Stamatis Kavadias, Manolis Katevenis, Michail Zampetakis, and Dimitrios S. Nikolopoulos. Cache-integrated network interfaces: Flexible onchip communication and synchronization for largescale CMPs. International Journal of Parallel Programming, 40(6):583–604, December 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=6&spage=583.

Krall:2000:CTM

[KL00] Andreas Krall and Sylvain Lelait. Compilation techniques for multimedia processors. International Journal of Parallel Programming, 28(4):347–361, August 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=4&spage=347.

Kang:2008:ISE

[KLG08] Dongsoo Kang, Chen Liu, and Jean-Luc Gaudiot. The impact of speculative execution on SMT processors. International Journal of Parallel Programming, 36(4):361–385, August 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=4&spage=361.

Kim:2016:GAF

[KLK16] Byungjoo Kim, Jung Eun Lee, and Young J. Kim. GPU accelerated finding of channels and tunnels for a protein molecule. International Journal of Parallel Programming, 44(1): 87–108, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0331-8.

Krall:1986:CSP

[KM86] Edward J. Krall and Patrick F. McGehearty. A case study of parallel execution of a rule-based expert system. International Journal of Parallel Programming, 15(1):5– 32, February 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=1&spage=5.

Klauser:2001:SBI

[KMG01] Artur Klauser, Srilatha Manne, and Dirk Grunwald. Selective branch inversion: Confidence estimation for branch predictors. International Journal of Parallel Programming, 29(1):81–110, February 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/13/5/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/13/5/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=1&spage=
81.

Kavi:2002:MMA

[KMjC02] Krishna M. Kavi, Alireza Moshtaghi, and Deng jyi Chen. Modeling multithreaded applications using Petri nets. International Journal of Parallel Programming, 30(5):353– 371, October 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/29/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/29/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=5&spage=
353.

Koelbel:1987:SAP

[KMV87] Charles Koelbel, Piyush Mehrotra, and John Van Rosendale. Semi-automatic process partitioning for parallel computation. International Journal of Parallel Programming, 16 (5):365–382, October 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=16&issue=5&spage=
365.

Kelly:1995:UAC

[KP95] Wayne Kelly and William Pugh. Using affine closure to find legal reordering transformations. International Journal of Parallel Programming, 23(4):303– 325, August 1995. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kodukula:2001:DCT

[KP01] Induprakas Kodukula and Keshav Pingali. Datacentric transformations for locality enhancement. International Journal of Parallel Programming, 29 (3):319–364, June 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/21/4/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/21/4/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=3&spage=
                                        319.

Kyriakopoulos:2004:DDA

[KP04] Konstantinos Kyriakopoulos and Kleanthis Psarris. Data dependence analysis techniques for increased accuracy and extracted parallelism. International Journal of Parallel Programming, 32(4):317–359, August 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        32&issue=4&spage=317.

Koelbl:2005:CEF

[KP05] Alfred Koelbl and Carl Pixley. Constructing efficient formal models from high-level descriptions using symbolic simulation. International Journal of Parallel Programming, 33(6):645–666, December 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=6&spage=645.

Kelly:1996:TCI

[KPRS96] Wayne Kelly, William Pugh, Evan Rosser, and Tatiana Shpeisman. Transitive closure of infinite graphs and its applications. International Journal of Parallel Programming, 24(6):579–598, December 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kromer:2014:NIM

[KPS14] Pavel Krömer, Jan Platos, and Václav Snásel. Natureinspired meta-heuristics on modern GPUs: State of the art and brief survey of selected algorithms. International Journal of Parallel Programming, 42(5): 681–709, October 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0292-3.

Kumar:1987:PDF

[KR87] Vipin Kumar and V. Nageshwara Rao. Parallel depth first search. Part II. Analysis. International Journal of Parallel Programming, 16(6):501–519, December 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=6&spage=501.

Kale:1990:PSS

[KS90] L. V. Kalé and Vikram A. Saletore. Parallel statespace search for a first solution with consistent linear speedups. International Journal of Parallel Programming, 19(4):251– 293, August 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        19&issue=4&spage=251.

Kessler:1997:FPP

[KS97] Christoph W. Kessler and Helmut Seidl. The Fork95 parallel programming language: Design, implementation, application. International Journal of Parallel Programming, 25(1): 17–50, February 1997. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kumar:2018:GSR

[KSA^{+}18] Santosh Kumar, Sanjay Kumar Singh, Ali Imam Abidi, Deepanwita Datta, and Arun Kumar Sangaiah. Group sparse representation approach for recognition of cattle on muzzle point images. International Journal of Parallel Programming, 46(5):812–837, October 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Khatri:2022:SMF

[KSBN22] Jash Khatri, Arihant Samar, Bikash Behera, and Rupesh Nasre. Scaling the maximum flow computation on GPUs. International Journal of Parallel Programming, 50(5–6): 515–561, December 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00740-7.

Kayi:2014:BAC

[KSEG14] Abdullah Kayi, Olivier Serres, and Tarek El-Ghazawi. Bandwidth adaptive cache coherence optimizations for chip multiprocessors. International Journal of Parallel Programming, 42(3): 435–455, June 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0247-8.

Khan:2018:ATS

[KSF^{+}18] Atif Khan, Naomie Salim, Haleem Farman, Murad Khan, Bilal Jan, Awais Ahmad, Imran Ahmed, and Anand Paul. Abstractive text summarization based on improved semantic graph approach. International Journal of Parallel Programming, 46(5): 992–1016, October 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kim:2014:VVF

[KSJ14] Young-Joo Kim, Sejun Song, and Yong-Kee Jun. VORD: a versatile on-the-fly race detection tool in OpenMP programs. International Journal of Parallel Programming, 42(6): 900–930, December 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

com/article/10.1007/
s10766-013-0257-6.

Knorr:2024:ADC

[KSTF24] Fabian Knorr, Philip Salzmann, Peter Thoman, and Thomas Fahringer. Automatic discovery of collective communication patterns in parallelized task graphs. International Journal of Parallel Programming, 52(3): 171–186, June 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00767-y.

Krishnan:2001:NFC

[KT01] Venkata Krishnan and Josep Torrellas. The need for fast communication in hardware-based speculative chip multiprocessors. International Journal of Parallel Programming, 29(1):3–33, February 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/13/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/13/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&

                                        volume=29&issue=1&spage=
                                        3.

Kyrkou:2018:BHE

[KTBP18] Christos Kyrkou, Theocharis Theocharides, Christos-Savvas Bouganis, and Marios Polycarpou. Boosting the hardware-efficiency of cascade support vector machines for embedded classification applications. International Journal of Parallel Programming, 46 (6):1220–1246, December 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Knorr:2023:DDF

[KTF23] Fabian Knorr, Peter Thoman, and Thomas Fahringer. Declarative data flow in a graph-based distributed memory runtime system. International Journal of Parallel Programming, 51 (2–3):150–171, June 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00743-4.

Kreutzer:2017:GBB

[KTRZ^{+}17] Moritz Kreutzer, Jonas Thies, Melven Röhrig-Zöllner, Andreas Pieper, Faisal Shahzad, Martin Galgon, Achim Basermann, Holger Fehske, Georg Hager, and Gerhard Wellein. GHOST:

Building blocks for high performance sparse linear algebra on heterogeneous systems. International Journal of Parallel Programming, 45(5):1046– 1072, October 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kubota:1999:TER

[KTT^{+}99] Atsushi Kubota, Shogo Tatsumi, Toshihiko Tanaka, Masahiro Goshima, Shin ichiro Mori, Hiroshi Nakashima, and Shinji Tomita. A technique to eliminate redundant inter-processor communication on parallelizing compiler TIN-PAR. International Journal of Parallel Programming, 27(2):97–109, April 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=2&spage=97.

Kuck:1994:WDU

[Kuc94] David J. Kuck. What do users of parallel computer systems really need? International Journal of Parallel Programming, 22 (1):99–127, February 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Koduru:2018:SSC

[KVG18] Sai Charan Koduru, Keval Vora, and Rajiv Gupta. Software speculation on caching DSMs. International Journal of Parallel Programming, 46(2):313– 332, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kempf:2010:ASB

[KWA^{+}10] T. Kempf, S. Wallentowitz, G. Ascheid, R. Leupers, and H. Meyr. Analytical and simulationbased design space exploration of software defined radios. International Journal of Parallel Programming, 38(3–4):303– 321, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=3&spage=303.

Lobeiras:2015:BTB

[LAD15] J. Lobeiras, M. Amor, and R. Doallo. BPLG: a tuned butterfly processing library for GPU architectures. International Journal of Parallel Programming, 43(6):1078– 1102, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0323-8.

Langlois:1990:SPC

[Lan90] Laurent Langlois. Systolic parsing of contextfree languages. International Journal of Parallel Programming, 19(4):333– 355, August 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=4&spage=333.

Llosa:1998:QER

[LAV98] Josep Llosa, Eduard Ayguadé, and Mateo Valero. Quantitative evaluation of register pressure on software pipelined loops. International Journal of Parallel Programming, 26(2):121– 142, April 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=2&spage=121.

Loulergue:2017:CPP

[LBT17] Frédéric Loulergue, Wadoud Bousdira, and Julien Tesson. Calculating parallel programs in Coq using list homomorphisms. International Journal of Parallel Programming, 45 (2):300–319, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0415-8.

Lu:2011:PAA

[LC11] Yu-Min Lu and Peng-Sheng Chen. Probabilistic alias analysis of executable code. International Journal of Parallel Programming, 39(6):663– 693, December 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=6&spage=663.

Li:2021:CAO

[LCF21] Jiansong Li, Wei Cao, and Xiaobing Feng. Compilerassisted operator template library for DNN accelerators. International Journal of Parallel Programming, 49(5): 628–645, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00701-6.

Lu:2017:PEE

[LCL17] Xingjing Lu, Long Chen, and Zhiyuan Li. Performance evaluation and enhancement of processbased parallel loop execution. International Journal of Parallel Programming, 45(1):185–198, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0394-1.

Li:2019:ABE

[LCL19] Deng Li, Zhujun Chen, and Jiaqi Liu. Analysis for behavioral economics in social networks: An altruism-based dynamic cooperation model. International Journal of Parallel Programming, 47(4): 686–708, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Liu:2020:PAA

[LCT^{+}20] Zelin Liu, Jian Cao, Yudong Tan, Quanwu Xiao, and Mukesh Prasad. Planning above the API clouds before flying above the clouds: a real-time personalized air travel planning approach. International Journal of Parallel Programming, 48(1): 137–156, February 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Li:1992:VTV

[LCU92] Xining Li, John Cleary, and Brian Unger. Virtual time and virtual space. International Journal of Parallel Programming, 21(2):123–150, April 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=2&spage=123.

Lee:2005:EOS

[LDHL05] Yoon-Ju Lee, Pedro C. Diniz, Mary W. Hall, and Robert Lucas. Empirical optimization for a sparse linear solver: a case study. International Journal of Parallel Programming, 33(2–3):165– 181, June 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=2&spage=165.

Liu:2015:PCU

[LEA15] Zifan Liu, Nahid Emad, and Soufian Ben Amor. PageRank computation using a multiple implicitly restarted Arnoldi method for modeling epidemic spread. International Journal of Parallel Programming, 43(6):1028– 1053, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0344-3.

Liu:2011:VPS

[LEG11] Shaoshan Liu, Christine Eisenbeis, and Jean-Luc Gaudiot. Value prediction and speculative execution on GPU. International Journal of Parallel Programming, 39(5):533– 552, October 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=5&spage=533.

Lo:1999:TCO

[LEL^{+}99] Jack L. Lo, Susan J.

Eggers, Henry M. Levy, Sujay S. Parekh, and Dean M. Tullsen. Tuning compiler optimizations for simultaneous multithreading. International Journal of Parallel Programming, 27(6):477–503, December 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=6&spage=477.

Lenzi:2015:FOC

[LF15] Karlo G. Lenzi and Felipe A. P. Figueiredo. Fully optimized code block segmentation algorithm for LTE–Advanced. International Journal of Parallel Programming, 43(6): 988–1003, December 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0324-7.

Laborde:2017:WFH

[LFD17] Pierre Laborde, Steven Feldman, and Damian Dechev. A wait-free hash map. International Journal of Parallel Programming, 45(3):421–448, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Lopez-Fandino:2019:GFC

[LFHAM19] Javier López-Fandiño, Dora B. Heras, Francisco Argüello, and Mauro Dalla Mura. GPU framework for change detection in multitemporal hyperspectral images. International Journal of Parallel Programming, 47 (2):272–292, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Laccetti:2017:GEH

[LFL^{+}17] Giuliano Laccetti, Ian Foster, Marco Lapegna, Paul Messina, Raffaele Montella, and Almerico Murli. Guest editorial for hybrid parallelism in new HPC systems. International Journal of Parallel Programming, 45(5):1021– 1025, October 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0465-y.pdf.

Li:2010:TRI

[LG10] Xiaobin Li and Jean-Luc Gaudiot. Tolerating radiation-induced transient faults in modern processors. International Journal of Parallel Programming, 38(2): 85–116, April 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=2&spage=85.

Liao:2016:PBA

[LGY16] Xiaofei Liao, Rentong Guo, and Danping Yu. A phase behavior aware dynamic cache partitioning scheme for CMPs. International Journal of Parallel Programming, 44 (1):68–86, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0334-5.

Liu:2015:DRA

[LHF^{+}15] Songbin Liu, Xiaomeng Huang, Haohuan Fu, Guangwen Yang, and Zhenya Song. Data reduction analysis for climate data sets. International Journal of Parallel Programming, 43(3): 508–527, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0287-0.

Li:2016:MPP

[LHL^{+}16] Ren Li, Haibo Hu, Heng Li, Yunsong Wu, and Jianxi Yang. MapReduce parallel programming model: A state-of-the-art survey. International Journal of Parallel Programming, 44(4): 832–866, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0395-0.

Liu:2019:RAO

[LHLT19] Junhong Liu, Xin He, Weifeng Liu, and Guangming Tan. Register-aware optimizations for parallel sparse matrix–matrix multiplication. International Journal of Parallel Programming, 47(3):403– 417, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Liu:2017:SMC

[LHP^{+}17] Pei Liu, Ahmed Hemani, Kolin Paul, Christian Weis, Matthias Jung, and Norbert Wehn. 3D-stacked many-core architecture for biological sequence analysis problems. International Journal of Parallel Programming, 45(6):1420– 1460, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0495-0.pdf.

Loff:2022:DCT

[LHP^{+}22] Júnior Löff, Renato B.

Hoffmann, Ricardo Pieper, Dalvan Griebler, and Luiz G. Fernandes. DSPar-Lib: a C++ template library for distributed stream parallelism. International Journal of Parallel Programming, 50(5–6): 454–485, December 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

                                        com/article/10.1007/
                                        s10766-022-00737-2.

Li:2003:PRE

[Li03] Keqin Li. On the performance of randomized embedding of reproduction trees in static networks. International Journal of Parallel Programming, 31(5):393–406, October 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        ipsapp007.kluweronline.
                                        com/content/getfile/
                                        4773/36/2/abstract.htm;
                                        http://ipsapp007.kluweronline.
                                        com/content/getfile/
                                        4773/36/2/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=5&spage=
                                        393.

Lindstrom:1986:SPR

[Lin86] Gary Lindstrom. Sans pareil: Referees. International Journal of Parallel Programming, 15(6):567– 568, December 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        15&issue=6&spage=567.

Lindstrom:1987:SPR

[Lin87] Gary Lindstrom. Sans pareil: Referees. International Journal of Parallel Programming, 16(6):521– 522, December 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=6&spage=521.

Lindstrom:1988:SC

[Lin88a] Gary Lindstrom. Sage commentary. International Journal of Parallel Programming, 17(1):93, February 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=1&spage=93.

Linstrom:1988:SPR

[Lin88b]  Gary Linstrom.  Sans

pareil: referees. International Journal of Parallel Programming, 17(6):529– 530, December 1988. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=6&spage=529.

Lindstrom:1989:SPR

[Lin89] Gary Lindstrom. Sans pareil: Referees. International Journal of Parallel Programming, 18(6):551– 552, December 1989. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=6&spage=551.

Lindstrom:1990:SPR

[Lin90] Gary Lindstrom. Sans pareil: Referees. International Journal of Parallel Programming, 19(6):511– 512, December 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        19&issue=6&spage=511.

Lin:1991:DFP

[Lin91a] Zheng Lin. A distributed fair polling scheme applied to OR-Parallel logic programming. International Journal of Parallel Programming, 20(4):315–339, August 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        20&issue=4&spage=315.

Lindstorm:1991:SPR

[Lin91b] Gary Lindstorm. Sans pareil: Referees. International Journal of Parallel Programming, 20(6):487– 488, December 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=6&spage=487.

Lindstrom:1992:RV

[Lin92] Gary Lindstrom. Referees and valedictory. International Journal of Parallel Programming, 21(6):477– 479, December 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=6&spage=477.

Livesey:1991:NMB

[Liv91] Mike Livesey. A network model of barrier synchronization algorithms. International Journal of Parallel Programming, 20(1): 55–74, February 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=1&spage=55.

Loh:2008:MPH

[LJ08] Gabriel H. Loh and Daniel A. Jiménez. Modulo path history for the reduction of pipeline overheads in path-based neural branch predictors. International Journal of Parallel Programming, 36(2): 267–286, April 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=2&spage=267.

Lee:2009:RSS

[LJ09] Joahyoung Lee and Inbum Jung. Recovery strategies for streaming media service in a clusterbased VOD server with a fault node. International Journal of Parallel Programming, 37(2):175– 194, April 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=2&spage=175.

Luo:2005:SSM

[LJE05] Yue Luo, Lizy K. John, and Lieven Eeckhout. SMA: a self-monitored adaptive cache warm-up scheme for microprocessor simulation. International Journal of Parallel Programming, 33(5):561– 581, October 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=5&spage=561.

Lai:1994:CRI

[LkCH94] Feipei Lai, Yung kuang Chao, and Chia-Jung Hsieh. The complementary relationship of interprocedural register allocation and inlining. International Journal of Parallel Programming, 22(4):409– 434, August 1994. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Lydia:2020:CDC

[LKS^{+}20] E. Laxmi Lydia, P. Krishna Kumar, K. Shankar, S. K. Lakshmanaprabu, R. M. Vidhyavathi, and Andino Maseleno. Charismatic document clustering through novel K-means non-negative matrix factorization (KNMF) algorithm using key phrase extraction. International Journal of Parallel Programming, 48(3):496–514, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Langguth:2017:ADT

[LLGC17] Johannes Langguth, Qiang Lan, Namit Gaur, and Xing Cai. Accelerating detailed tissue-scale 3D cardiac simulations using heterogeneous CPU–Xeon Phi computing. International Journal of Parallel Programming, 45(5): 1236–1258, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Liang:2015:PJE

[LLL^{+}15] Feng Liang, Yunzhen Liu, Hai Liu, Shilong Ma, and Bettina Schnor. A parallel job execution time estimation approach based on user submission patterns within computational grids. International Journal of Parallel Programming, 43(3): 440–454, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0294-1.

Laccetti:2012:DAA

[LLM^{+}12] Giuliano Laccetti, Marco Lapegna, Valeria Mele, Diego Romano, and Almerico Murli. A double adaptive algorithm for multidimensional integration on multicore based HPC systems. International Journal of Parallel Programming, 40(4):397–409, August 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=4&spage=397.

Laccetti:2016:LCM

[LLM16] Giuliano Laccetti, Marco Lapegna, and Valeria Mele. A loosely coordinated model for heapbased priority queues in multicore environments. International Journal of Parallel Programming, 44 (4):901–921, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0398-x.

Lu:2003:ABH

[LLSS03] Zhijian Lu, John Lach, Mircea R. Stan, and Kevin Skadron. Alloyed branch history: Combining global and local branch history for robust performance. International Journal of Parallel Programming, 31(2):137–177, April 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp007.kluweronline.
com/content/getfile/
4773/32/3/abstract.htm;
http://ipsapp007.kluweronline.
com/content/getfile/

                                        4773/32/3/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=2&spage=
                                        137.

Li:2017:TLT

[LLW^{+}17] Jing Li, Lei Liu, Yuan Wu, Xiaobing Feng, and Chengyong Wu. Twolevel task scheduling for irregular applications on GPU platform. International Journal of Parallel Programming, 45(1): 79–93, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0387-0.

Loechner:2000:COA

[LM00] Vincent Loechner and Catherine Mongenet. Communication optimization for affine recurrence equations using broadcast and locality. International Journal of Parallel Programming, 28(1):47–102, February 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=1&spage=47.

Liu:2018:DNS

[LMHW18] Wenjie Liu, Sheng Ma, Libo Huang, and Zhiying Wang. The design of NoC-side memory access scheduling for energyefficient GPGPUs. International Journal of Parallel Programming, 46(4): 722–735, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Lee:1998:CPA

[LMP98] Jaejin Lee, Samuel P. Midkiff, and David A. Padua. A constant propagation algorithm for explicitly parallel programs. International Journal of Parallel Programming, 26(5):563– 589, October 1998. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=5&spage=563.

Loghi:2005:DFV

[LMPS05] Mirko Loghi, Tiziana Margaria, Graziano Pravadelli, and Bernhard Steffen. Dynamic and formal verification of embedded systems: A comparative survey. International Journal of Parallel Programming, 33(6):585–611, December 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=6&spage=585.

Lin:2012:ESC

[LNG12] Changhui Lin, Vijay Nagarajan, and Rajiv Gupta. Efficient sequential consistency using conditional fences. International Journal of Parallel Programming, 40(1):84–117, February 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=1&spage=84.

Li:1991:ECM

[LNP91] Kai Li, Jeffrey F. Naughton, and James S. Plank. An efficient checkpointing method for multicomputers with wormhole routing. International Journal of Parallel Programming, 20(3):159– 180, June 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        20&issue=3&spage=159.

Lowenthal:2000:ASB

[Low00] David K. Lowenthal. Accurately selecting block size at runtime in pipelined parallel programs. International Journal of Parallel Programming, 28(3): 245–274, June 2000. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=3&spage=245.

Li:1994:SLT

[LP94] Wei Li and Keshav Pingali. A singular loop transformation framework based on non-singular matrices. International Journal of Parallel Programming, 22(2):183–205, April 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Langemeyer:2013:USM

[LPB13] Stefan Langemeyer, Peter Pirsch, and Holger Blume. Using SDRAM memories for high-performance accesses to two-dimensional matrices without transpose. International Journal of Parallel Programming, 41(2):331–354, April 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://link.springer.
com/article/10.1007/
s10766-012-0225-6.

Lattuada:2016:PET

[LPF16] Marco Lattuada, Christian Pilato, and Fabrizio Ferrandi. Performance estimation of task graphs based on path profiling. International Journal of Parallel Programming, 44 (4):735–771, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0372-7.

Liao:2010:SAA

[LQWP10] Chunhua Liao, Daniel J. Quinlan, Jeremiah J. Willcock, and Thomas Panas. Semantic-aware automatic parallelization of modern applications using high-level abstractions. International Journal of Parallel Programming, 38(5–6):361–378, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=5&spage=361.

Lo:1991:OTM

[LRG^{+}91] Virginia M. Lo, Sanjay Rajopadhye, Samik Gupta, David Keldsen, Moataz A. Mohamed, Bill Nitzberg, Jan Arne Tell, and Xiaoxiong Zhong. OREGAMI: Tools for mapping parallel computations to parallel architectures. International Journal of Parallel Programming, 20(3): 237–270, June 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=3&spage=237.

Lee:2014:BCA

[LRG14] Changmin Lee, Won Woo Ro, and Jean-Luc Gaudiot. Boosting CUDA applications with CPU– GPU hybrid computing. International Journal of Parallel Programming, 42 (2):384–404, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0252-y.

Lin:1991:PIS

[LS91] Calvin Lin and Lawrence Snyder. A portable implementation of SIM-PLE. International Journal of Parallel Programming, 20(5):363–401, October 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=5&spage=363.

Loots:1992:PAK

[LS92] W. Loots and T. H. C. Smith. A parallel algorithm for the 0-1 knapsack problem. International Journal of Parallel Programming, 21(5):349– 362, October 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=5&spage=349.

Lipasti:1998:EVL

[LS98] Mikko H. Lipasti and John Paul Shen. Exploiting value locality to exceed the dataflow limit. International Journal of Parallel Programming, 26(4):505–538, August 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=4&spage=505.

Lin:2005:EBH

[LS05] Chao Lin and Jang-Ping Sheu. Efficient broadcast in heterogeneous networks of workstations using two sub-networks. International Journal of Parallel Programming, 33(4):351– 391, August 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=4&spage=351.

Laudon:2007:CWM

[LS07] James Laudon and Lawrence Spracklen. The coming wave of multithreaded chip multiprocessors. International Journal of Parallel Programming, 35(3):299– 330, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=3&spage=299.

Li:2020:MOW

[LS20] Yuanzhe Li and Loren Schwiebert. Memoryoptimized wavefront parallelism on GPUs. International Journal of Parallel Programming, 48(6):1008– 1031, December 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00658-y.

Lee:2007:DBI

[LSA^{+}07] Gregory L. Lee, Martin Schulz, Dong H. Ahn, Andrew Bernat, Bronis R. de Supinski, Steven Y.

Ko, and Barry Rountree. Dynamic binary instrumentation and data aggregation on large scale systems. International Journal of Parallel Programming, 35(3):207–232, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=3&spage=207.

Larsen:2009:ABE

[LSHK09] Steen Larsen, Parthasarathy Sarangam, Ram Huggahalli, and Siddharth Kulkarni. Architectural breakdown of end-to-end latency in a TCP/IP network. International Journal of Parallel Programming, 37(6):556–571, December 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=6&spage=556.

Liu:1994:SSS

[LSL94] Jie Liu, Vikram A. Saletore, and Ted G. Lewis. Safe self-scheduling: a parallel loop scheduling scheme for shared-memory multiprocessors. International Journal of Parallel Programming, 22(6): 589–616, December 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Lin:2018:CHM

[LSM^{+}18] Han Lin, Zhichao Su, Xiandong Meng, Xu Jin, Zhong Wang, Wenting Han, Hong An, Mengxian Chi, and Zheng Wu. Combining Hadoop with MPI to solve metagenomics problems that are both data- and computeintensive. International Journal of Parallel Programming, 46(4):762–775, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Li:2015:ODN

[LSYG15] Zhao Li, Yao Shen, Bin Yao, and Minyi Guo. OF-Scheduler: a dynamic network optimizer for MapReduce in heterogeneous cluster. International Journal of Parallel Programming, 43(3): 472–488, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0281-6.

Li:2017:GEP

[LT17] Maozhen Li and Zhuo Tang. Guest editorial:

The parallel storage, processing and analysis for big data. International Journal of Parallel Programming, 45(4):731–733, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0475-9.pdf.

Leite:2012:NNS

[LTF^{+}12] Pedro Leite, João Marcelo Teixeira, Thiago Farias, Bernardo Reis, Veronica Teichrieb, and Judith Kelner. Nearest neighbor searches on the GPU: a massively parallel approach for dynamic point clouds. International Journal of Parallel Programming, 40(3):313– 330, June 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=3&spage=313.

Li:2015:CCM

[LTL15] Guohong Li, Olivier Temam, and Zhenyu Liu. Cluster cache monitor: Leveraging the proximity data in CMP. International Journal of Parallel Programming, 43(6):1054–1077, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0339-0;
http://link.springer.
com/content/pdf/10.1007/
s10766-014-0339-0.pdf.

Langr:2015:DAL

[LTSD15] Daniel Langr, Pavel Tvrdík, Ivan Simecek, and Tomás Dytrych. Downsampling algorithms for large sparse matrices. International Journal of Parallel Programming, 43(5): 679–702, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0315-8.

Lubachevsky:1990:SBR

[Lub90] Boris D. Lubachevsky.

Synchronization barrier and related tools for shared memory parallel programming. International Journal of Parallel Programming, 19(3):225– 250, June 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=3&spage=225.

Lal:2022:QSL

[LVJ22] Sohan Lal, Bogaraju Sharatchandra Varma, and Ben Juurlink. A quantitative study of locality in GPU caches for memorydivergent workloads. International Journal of Parallel Programming, 50 (2):189–216, April 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00729-2.

Li:2016:SSO

[LVM16] Xi Li, Anthony Ventresque, and John Murphy. SOC: Satisfactionoriented virtual machine consolidation in enterprise data centers. International Journal of Parallel Programming, 44(1): 130–150, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0333-6.

Loechner:1997:PPT

[LW97] Vincent Loechner and Doran K. Wilde. Parameterized polyhedra and their vertices. International Journal of Parallel Programming, 25(6):525–549, December 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
25&issue=6&spage=525.

Liu:2017:DPA

[LWDL17] Wei Liu, Lu Wang, Yuyue Du, and Maozhen Li. Deadlock property analysis of concurrent programs based on Petri net structure. International Journal of Parallel Programming, 45(4):879–898, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Liu:2019:SAC

[LWF^{+}19] Wei Liu, Lu Wang, Xin Feng, Man Qi, Chun Yan, and Maozhen Li. Soundness analytics of composed logical workflow nets. International Journal of Parallel Programming, 47 (4):709–724, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Lu:2018:SPE

[LWGZ18] Yuntao Lu, Chao Wang, Lei Gong, and Xuehai Zhou. SparseNN: a performance-efficient accelerator for large-scale sparse neural networks. International Journal of Parallel Programming, 46 (4):648–659, August 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Li:2019:IPD

[LWL^{+}19] Mingfan Li, Ke Wen, Han Lin, Xu Jin, Zheng Wu, Hong An, and Mengxian Chi. Improving the performance of distributed MXNet with RDMA. International Journal of Parallel Programming, 47 (3):467–480, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Liu:2011:STE

[LWLG11] Shaoshan Liu, Ligang Wang, Xiao-Feng Li, and Jean-Luc Gaudiot. Spaceand-time efficient parallel garbage collector for data-intensive applications. International Journal of Parallel Programming, 39(4):451–472, August 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=4&spage=451.

Liu:2004:HPR

[LWP04] Jiuxing Liu, Jiesheng Wu, and Dhabaleswar K. Panda. High performance RDMA-based MPI implementation over InfiniBand. International Journal of Parallel Programming, 32(3):167–198, June 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=3&spage=167.

Liu:2017:PBP

[LXL17] Yang Liu, Lixiong Xu, and Maozhen Li. The parallelization of back propagation neural network in MapReduce and Spark. International Journal of Parallel Programming, 45 (4):760–779, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Lin:1995:PPA

[LY95] Wei-Ming Lin and Bo Yang.

Probabilistic performance analysis for parallel search techniques. International Journal of Parallel Programming, 23(2):161–189, April 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Li:1998:Ia

[LY98a] Zhiyuan Li and Pen-Chung Yew. Introduction. International Journal of Parallel Programming, 26(5):539–540, October 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=5&spage=539.

Li:1998:Ib

[LY98b] Zhiyuan Li and Pen-Chung Yew. Introduction. International Journal of Parallel Programming, 26(6):639–640, December 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=6&spage=639.

Lu:2018:PHK

[LYG^{+}18] Wei Lu, Xiaomin Yang, Xu Gou, Lihua Jian, Wei Wu, and Gwanggil Jeon. Parallel heat kernel volume based local binary pattern on multiorientation planes for face representation. International Journal of Parallel Programming, 46(5):943– 962, October 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Li:2014:PTI

[LYL14] Cheng Hua Li, Laurence T. Yang, and Man Lin. Parallel training of an improved neural network for text categorization. International Journal of Parallel Programming, 42(3):505–523, June 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0245-x.

Lysecky:2008:SPE

[Lys08] Roman Lysecky. Scalability and parallel execution of Warp processing: Dynamic hardware/software partitioning. International Journal of Parallel Programming, 36(5):478–492, October 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=5&spage=478.

Lu:2017:IFM

[LZ17] Yaojie Lu and Sotirios G.

Ziavras. Instruction fusion for multiscalar and manycore processors. International Journal of Parallel Programming, 45(1): 67–78, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

                                        com/article/10.1007/
                                        s10766-015-0386-1.

Meijer:1987:OCP

[MA87] Henk Meijer and Selim G. Akl. Optimal computation of prefix sums on a binary tree of processors. International Journal of Parallel Programming, 16(2): 127–136, April 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=2&spage=127.

Muller:2010:GEI

[MA10] Matthias S. Müller and Eduard Ayguadé. Guest editors’ introduction. International Journal of Parallel Programming, 38 (5–6):339–340, October 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

                                        URL http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=38&issue=5&spage=
                                        339.

Munk:2011:APA

[MAB^{+}11] Harm Munk, Eduard Ayguadé, Cédric Bastoul, Paul Carpenter, Zbigniew Chamski, et al. ACOTES Project: Advanced compiler technologies for embedded streaming. International Journal of Parallel Programming, 39(3): 397–450, June 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=3&spage=397.

Main:1987:TFT

[Mai87] Michael G. Main. Trace, failure and testing equivalences for communicating processes. International Journal of Parallel Programming, 16(5):383– 400, October 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=5&spage=383.

Mariani:2016:SPP

[MAJD16] Giovanni Mariani, Andreea Anghel, Rik Jongerius, and Gero Dittmann. Scaling properties of parallel applications to exascale. International Journal of Parallel Programming, 44(5):975–1002, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
016-0412-y.

Miguel-Alonso:2009:INS

[MANR09] J. Miguel-Alonso, J. Navaridas, and F. J. Ridruejo. Interconnection network simulation using traces of MPI applications. International Journal of Parallel Programming, 37(2): 153–174, April 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=2&spage=153.

Margaris:2009:LFF

[Mar09] Athanasios I. Margaris. Log file formats for parallel applications: a review. International Journal of Parallel Programming, 37(2):195–222, April 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=2&spage=195.

Marowka:2017:EAM

[Mar17] Ami Marowka. Energyaware modeling of scaled heterogeneous systems. International Journal of Parallel Programming, 45 (5):1026–1045, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Moreno:2024:YAL

[MARC24] Pedro Moreno, Miguel Areias, Ricardo Rocha, and Vítor Santos Costa. Yet another lock-free atom table design for scalable symbol management in Prolog. International Journal of Parallel Programming, 52(3): 187–206, June 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00766-z.

Matsuzaki:2017:FMH

[Mat17] Kiminori Matsuzaki. Functional models of Hadoop MapReduce with application to scan. International Journal of Parallel Programming, 45(2):362– 381, April 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0414-9.

MacGregor:2023:GEC

[MAT23] Ruairidh MacGregor, Blair Archibald, and Phil Trinder. Generic exact combinatorial search at HPC scale. International Journal of Parallel Programming, 51 (1):83–106, February 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00744-3.

Martinez-Angeles:2016:RLG

[MAWD^{+}16] Carlos Alberto Martínez-Angeles, Haicheng Wu, Inês Dutra, Vítor Santos Costa, and Jorge Buenabad-Chávez. Relational learning with GPUs: Accelerating rule coverage. International Journal of Parallel Programming, 44 (3):663–685, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0364-7.

Mendelson:1999:DAM

[MB99] Avi Mendelson and Michael Bekerman. Design alternatives of multithreaded architecture. International Journal of Parallel Programming, 27(3):161– 193, June 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        27&issue=3&spage=161.

Meira:2012:SIC

[MB12a] Wagner Meira and Ricardo Bianchini. Special issue on computer architecture and high-performance computing. International Journal of Parallel Programming, 40(3):259–261, June 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=3&spage=259.

Moghaddam:2012:IBG

[MB12b] Mohsen Ebrahimi Moghaddam and Mohammad Reza Bonyadi. An immunebased genetic algorithm with reduced search space coding for multiprocessor task scheduling problem. International Journal of Parallel Programming, 40 (2):225–257, April 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=40&issue=2&spage=
225.

Min:2003:OOP

[MBE03] Seung-Jai Min, Ayon Basumallik, and Rudolf Eigenmann. Optimizing OpenMP programs on software distributed shared memory systems. International Journal of Parallel Programming, 31 (3):225–249, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL /

                                        ips/frames/Refs/referenceskapmain.
                                        asp?J=4773&I=33&A=5&
                                        LK=NM;   http://ipsapp007.
                                        kluweronline.com/content/
                                        getfile/4773/33/5/abstract.
                                        htm;   http://ipsapp007.
                                        kluweronline.com/content/
                                        getfile/4773/33/5/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=3&spage=
                                        225.

Mellor-Crummey:1998:SCF

[MCA98]  John  Mellor-Crummey

and Vikram Adve. Simplifying control flow in compiler-generated parallel code. International Journal of Parallel Programming, 26(5):613–638, October 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=5&spage=613.

McAllister:2013:GES

[MCE13] John McAllister, Luigi Carro, and Skevos Evripidou. Guest editorial: Special issue on 2011 International Conference on Embedded Computer Systems: Architectures, Modeling and Simulation (SAMOS XI). International Journal of Parallel Programming, 41(2):161– 162, April 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0233-6;
http://link.springer.
com/content/pdf/10.1007/
s10766-012-0233-6.pdf.

Menotti:2012:LLP

[MCFM12] Ricardo Menotti, João M. P. Cardoso, Marcio M. Fernandes, and Eduardo Marques. LALP: a language to program custom FPGA-based acceleration engines. International Journal of Parallel Programming, 40(3):262– 289, June 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=3&spage=262.

McKee:2007:GEI

[McK07] Sally A. McKee. Guest Editor’s introduction. International Journal of Parallel Programming, 35(3): 179–180, June 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=3&spage=179.

Mei:2018:MGS

[MCT^{+}18] Gang Mei, Salvatore Cuomo, Hong Tian, Nengxiong Xu, and Linjun Peng. MeshCleaner: a generic and straightforward algorithm for cleaning finite element meshes. International Journal of Parallel Programming, 46(3):565– 583, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Mellor-Crummey:2001:IMH

[MCWK01] John Mellor-Crummey, David Whalley, and Ken Kennedy. Improving memory hierarchy performance for irregular applications using data and computation reorderings. International Journal of Parallel Programming, 29(3): 217–247, June 2001. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/21/1/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/21/1/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=3&spage=
                                        217.

Mustafa:2015:PPE

[ME15] Dheya Mustafa and Rudolf Eigenmann. PETRA: Performance evaluation tool for modern parallelizing compilers. International Journal of Parallel Programming, 43(4): 549–571, August 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0307-8.

Manolache:2007:FAC

[MEP07] Sorin Manolache, Petru Eles, and Zebo Peng. Fault-aware communication mapping for NoCs with guaranteed latency. International Journal of Parallel Programming, 35 (2):125–156, April 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=35&issue=2&spage=
125.

Merks:1986:OPA

[Mer86] Ed Merks. An optimal parallel algorithm for triangulating a set of points in the plane. International Journal of Parallel Programming, 15(5):399–411, October 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        15&issue=5&spage=399.

Martinez:2021:PSD

[MFC21] Millán A. Martínez, Basilio B. Fraguela, and José C. Cabaleiro. A parallel skeleton for divide-and-conquer unbalanced and deep problems. International Journal of Parallel Programming, 49(6): 820–845, December 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00709-y.

Milovanovic:2008:NEE

[MFG^{+}08] Milos Milovanović, Roger Ferrer, Vladimir Gajinov, Osman S. Unsal, Adrian Cristal, Eduard Ayguadé, and Mateo Valero. Nebelung: Execution environment for transactional OpenMP. International Journal of Parallel Programming, 36(3): 326–346, June 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=3&spage=326.

Moreton-Fernandez:2019:MDC

[MFGEL19] Ana Moreton-Fernandez, Arturo Gonzalez-Escribano, and Diego R. Llanos. Multi-device controllers: a library to simplify parallel heterogeneous programming. International Journal of Parallel Programming, 47(1):94–113, February 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Mische:2021:PMM

[MFU21] Jörg Mische, Martin Frieb, and Theo Ungerer. PIMP my many-core: Pipelineintegrated message passing. International Journal of Parallel Programming, 49(4):487–505, August 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL https:

//link.springer.com/
article/10.1007/s10766-
020-00685-9.

Mehta:2015:MTP

[MG15] Kshitij Mehta and Edgar Gabriel. Multi-threaded parallel I/O for OpenMP applications. International Journal of Parallel Programming, 43(2):286– 309, April 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0306-9.

Melo:2014:GE

[MGD^{+}14] Alba Melo, Jean-Luc Gaudiot, Luiz DeRose, Kunle Olukotun, and Albert Zomaya. Guest editorial. International Journal of Parallel Programming, 42(1):1–3, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-013-0255-8.pdf.

McAllister:2015:GES

[MGJS15] John McAllister, David Guevorkian, Hartwig Jeschke, and Mihai Sima. Guest editorial: Special issue on embedded computer systems: Architectures, modeling and simulation. International Journal of Parallel Programming, 43 (1):1–2, February 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0321-x.

Montella:2017:VCB

[MGL^{+}17] Raffaele Montella, Giulio Giunta, Giuliano Laccetti, Marco Lapegna, Carlo Palmieri, Carmine Ferraro, Valentina Pelliccia, Cheol-Ho Hong, Ivor Spence, and Dimitrios S. Nikolopoulos. On the virtualization of CUDA based GPU remoting on ARM and x86 machines in the GVirtuS framework. International Journal of Parallel Programming, 45(5): 1142–1163, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Marsolf:1999:UMS

[MGW99] Bret A. Marsolf, Kyle A.

Gallivan, and Harry A. G. Wijshoff. The utilization of matrix structure to generate optimized code from MATLAB programs. International Journal of Parallel Programming, 27 (2):73–96, April 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=2&spage=73.

Mitchell:1998:QML

[MHCF98] Nicholas Mitchell, Karin Högstedt, Larry Carter, and Jeanne Ferrante. Quantifying the multilevel nature of tiling interactions. International Journal of Parallel Programming, 26(6):641–670, December 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=6&spage=641.

Maydan:1995:EDD

[MHL95] Dror E. Maydan, John L. Hennessy, and Monica S. Lam. Effectiveness of data dependence analysis. International Journal of Parallel Programming, 23 (1):63–81, February 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Miller:1988:ISB

[Mil88] James S. Miller. Implementing a Scheme-Based parallel processing system. International Journal of Parallel Programming, 17(5):367–402, October 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        17&issue=5&spage=367.

Mishra:2009:GEI

[Mis09] Prabhat Mishra. Guest Editor introduction: Special issue on nano/bioinspired applications and architectures. International Journal of Parallel Programming, 37(4):343– 344, August 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=4&spage=343.

Milicev:2002:CFR

[MJ02] Dragan Milicev and Zoran Jovanovic. Control flow regeneration for software pipelined loops with conditions. International Journal of Parallel Programming, 30(3): 149–179, June 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/27/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/27/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=3&spage=
149.

Mutlu:2005:UFL

[MKAP05] Onur Mutlu, Hyesoon Kim, David N. Armstrong, and Yale N. Patt. Using the first-level caches as filters to reduce the pollution caused by speculative memory references. International Journal of Parallel Programming, 33(5):529–559, October 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        33&issue=5&spage=529.

Michelogiannakis:2015:ESP

[ML15] George Michelogiannakis and Xiaoye S. Li. Extending summation precision for network reduction operations. International Journal of Parallel Programming, 43(6):1218– 1243, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0326-5.

MantasRuiz:2002:CBD

[MLdlP02] Jose M. Mantas Ruiz, Julio Ortega Lopera, and Jose A. Carrillo de la Plata. Component-based derivation of a parallel stiff ODE solver implemented in a cluster of computers. International Journal of Parallel Programming, 30(2):99–148, April 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/26/2/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/26/2/fulltext.pdf;
                                         http://www.springerlink.

com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=2&spage=
99.

Matsuzaki:2016:PTA

[MM16] Kiminori Matsuzaki and Reina Miyazaki. Parallel tree accumulations on MapReduce. International Journal of Parallel Programming, 44(3): 466–485, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0355-8.

Mukherjee:2021:IHM

[MMD21] Amartya Mukherjee, Prateeti Mukherjee, and Nilanjan Dey. iGridEdge-Drone: Hybrid mobility aware intelligent load forecasting by edge enabled Internet of Drone Things for smart grid networks. International Journal of Parallel Programming, 49 (3):285–325, June 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00675-x.

Manoj:2004:CDC

[MMG04] N. P. Manoj, K. V. Manjunath, and R. Govindarajan. CAS-DSM: a compiler assisted software distributed shared memory. International Journal of Parallel Programming, 32(2):77–122, April 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=2&spage=77.

Mahfoudhi:2015:PCA

[MMN15] Ryma Mahfoudhi, Zaher Mahjoub, and Wahid Nasri. Parallel communicationavoiding algorithm for triangular matrix inversion on homogeneous and heterogeneous platforms. International Journal of Parallel Programming, 43 (4):631–655, August 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0310-0.

Morris:2007:SNO

[MMS07] Alan Morris, Allen D. Malony, and Sameer S. Shende. Supporting nested OpenMP parallelism in the TAU performance system. International Journal of Parallel Programming, 35(4):417–436, August 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=4&spage=417.

Muniasamy:2022:ACS

[MNN22] Rajesh Pandian Muniasamy, Rupesh Nasre, and N. S. Narayanaswamy. Accelerating computation of Steiner trees on GPUs. International Journal of Parallel Programming, 50(1):152–185, February 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00723-0.

McNamee:1990:TOI

[MO90] Carole M. McNamee and Ronald A. Olsson. Transformations for optimizing interprocess communication and synchronization mechanisms. International Journal of Parallel Programming, 19(5):357–387, October 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=5&spage=357.

McNamee:1991:AGA

[MO91] Carole M. McNamee and Ronald A. Olsson. An attribute grammar approach to compiler optimization of IntraModule interprocess communication. International Journal of Parallel Programming, 20(3):181–202, June 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=3&spage=181.

Mohebbi:2019:PSC

[Moh19] Hamidreza Mohebbi. Parallel SIMD CPU and GPU implementations of Berlekamp–Massey algorithm and its error correction application. International Journal of Parallel Programming, 47(1): 137–160, February 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Midorikawa:2005:PNM

[MOL05] Edson Toshimi Midorikawa, Helio Marci Oliveira, and Jean Marcos Laine. PEM-PIs: a new methodology for modeling and prediction of MPI programs performance. International Journal of Parallel Programming, 33(5):499–527, October 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=5&spage=499.

Mongenet:1997:ADC

[Mon97] Catherine Mongenet. Affine dependence classification for communications minimization. International Journal of Parallel Programming, 25(6):497–524, December 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
25&issue=6&spage=497.

Mall:1991:FTA

[MP91] R. Mall and L. M. Patnaik.

Formal timing analysis of distributed systems. International Journal of Parallel Programming, 20(2): 75–94, April 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=2&spage=75.

Melvin:1995:EIS

[MP95] Stephen Melvin and Yale Patt. Enhancing instruction scheduling with a block-structured ISA. International Journal of Parallel Programming, 23 (3):221–243, June 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Menon:2004:LLL

[MP04] Vijay Menon and Keshav Pingali. Look left, look right, look left again: an application of fractal symbolic analysis to linear algebra code restructuring. International Journal of Parallel Programming, 32(6):501–523, December 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        32&issue=6&spage=501.

Menezo:2018:MSC

[MPAG18] Lucia G. Menezo, Valentin Puente, Pablo Abad, and Jose-Angel Gregorio. Mosaic: a scalable coherence protocol. International Journal of Parallel Programming, 46(6):1110– 1138, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Moss:2005:CCB

[MPR^{+}05] J. Eliot B. Moss, Trek Palmer, Timothy Richards, Edward K. Walters, and Charles C. Weems. CISL: a class-based machine description language for cogeneration of compilers and simulators. International Journal of Parallel Programming, 33(2–3): 231–246, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=2&spage=231.

Mendelson:2006:I

[MPZ06] Bilha Mendelson, Shlomit S.

Pinter, and Ayal Zaks. Introduction. International Journal of Parallel Programming, 34(2):111– 112, April 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=2&spage=111.

Menouer:2016:MSD

[MRLR16] Tarek Menouer, Mohamed Rezgui, Bertrand Le Cun, and Jean-Charles Régin. Mixing static and dynamic partitioning to parallelize a constraint programming solver. International Journal of Parallel Programming, 44(3): 486–505, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0356-7.

Moshovos:1999:SMC

[MS99] Andreas Moshovos and Gurindar S. Sohi. Speculative memory cloaking and bypassing. International Journal of Parallel Programming, 27(6):427– 456, December 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        27&issue=6&spage=427.

Meng:2011:PSI

[MS11] Jiayuan Meng and Kevin Skadron. A performance study for iterative stencil loops on GPUs with ghost zone optimizations. International Journal of Parallel Programming, 39(1):115–142, February 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=1&spage=115.

Moreira:2007:BGS

[MSA^{+}07] José E. Moreira, Valentina Salapura, George Almasi, Charles Archer, Ralph Bellofatto, Peter Bergner, Randy Bickford, Mathias Blumrich, José R. Brunheroto, Arthur A. Bright, Michael Brutman, José G. Castaños, Dong Chen, Paul Coteus, Paul Crumley, Sam Ellis, Thomas Engelsiepen, Alan Gara, Mark Giampapa, Tom Gooding, Shawn Hall, Ruud A. Haring, Roger Haskin, Philip Heidelberger, Dirk Hoenicke, Todd Inglett, Gerrard V. Kopcsay, Derek Lieber, David Limpert, Pat Mc-Carthy, Mark Megerian, Mike Mundy, Martin Ohmacht, Jeff Parker, Rick A. Rand, Don Reed, Ramendra Sahoo, Alda Sanomiya, Richard Shok, Brian Smith, Gordon G. Stewart, Todd Takken, Pavlos Vranas, Brian Wallenfelt, Michael Blocksome, and Joe Ratterman. The Blue Gene/L supercomputer: a hardware and software story. International Journal of Parallel Programming, 35(3):181– 206, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=3&spage=181.

Michaud:2001:EIF

[MSJ01] Pierre Michaud, André Seznec, and Stéphan Jourdan. An exploration of instruction fetch requirement in out-of-order superscalar processors. International Journal of Parallel Programming, 29(1):35–58, February 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/13/3/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/13/3/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=29&issue=1&spage=
                                        35.

Mahmud:2020:ALS

[MSJ20] Ayman A. Ataher Mahmud, Satakshi, and W. Jeberson. Aircraft landing scheduling using embedded flower pollination algorithm. International Journal of Parallel Programming, 48(5): 771–785, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-018-0602-x.

Mohanraj:2018:HPG

[MSPR18] Vanitha Mohanraj, R. Sakthivel, Anand Paul, and Seungmin Rho. High performance GCM architecture for the security of high speed network. International Journal of Parallel Programming, 46(5):904– 922, October 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Matheson:1996:PMM

[MT96] Lesley R. Matheson and Robert E. Tarjan. Parallelism in multigrid methods: How much is too much? International Journal of Parallel Programming, 24(5):397–432, October 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Metzger:2015:UGD

[MTT15] Markus Metzger, Xinmin Tian, and Walfred Tedeschi. User-guided dynamic data race detection. International Journal of Parallel Programming, 43 (2):159–179, April 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0296-z.

Malakar:2017:HRW

[MV17] Preeti Malakar and Venkatram Vishwanath. Hierarchical read–write optimizations for scientific applications with multi-variable structured datasets. International Journal of Parallel Programming, 45(1):94–108, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0388-z.

Martinez:2006:DGN

[MVB^{+}06] Carmen Martínez, Enrique Vallejo, Ramón Beivide, Cruz Izu, and Miquel Moretó. Dense Gaussian networks: Suitable topologies for on-chip multiprocessors. International Journal of Parallel Programming, 34(3):193– 211, June 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=3&spage=193.

Monteiro:2014:PFS

[MVD^{+}14] Eduarda Monteiro, Bruno Vizzotto, Cláudio Diniz, Marilena Maule, Bruno Zatt, and Sergio Bampi. Parallelization of full search motion estimation algorithm for parallel and distributed platforms. International Journal of Parallel Programming, 42 (2):239–264, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0216-7.

Moghimi:2019:MFD

[MWES19] Ahmad Moghimi, Jan Wichelmann, Thomas Eisenbarth, and Berk Sunar. MemJam: a false dependency attack against constant-time crypto implementations. International Journal of Parallel Programming, 47(4):538– 570, August 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Muller:2024:IMA

[MWHS24] Luise Müller, Philipp Wanko, Christian Haubelt, and Torsten Schaub. Investigating methods for ASPmT-based design space exploration in evolutionary product design. International Journal of Parallel Programming, 52(1–2): 59–92, April 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

                                        com/article/10.1007/
                                        s10766-024-00763-2.

Ma:2014:DPI

[MXP14] Nam Ma, Yinglong Xia, and Viktor K. Prasanna. Data parallel implementation of belief propagation in factor graphs on multicore platforms. International Journal of Parallel Programming, 42(1): 219–237, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0246-9.

Nikolopoulos:2002:RVM

[NAP02] Dimitrios S. Nikolopoulos, Eduard Ayguadé, and Constantine D. Polychronopoulos. Runtime vs. manual data distribution for architecture-agnostic shared-memory programming models. International Journal of Parallel Programming, 30(4):225– 255, August 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/2/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/2/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&

volume=30&issue=4&spage=
225.

Nansamba:2023:FMR

[NAS23] Grace Nansamba, Amani Altarawneh, and Anthony Skjellum. A fault-model-relevant classification of consensus mechanisms for MPI and HPC. International Journal of Parallel Programming, 51(2– 3):128–149, June 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00749-y.

Nalepa:2015:COP

[NB15] Jakub Nalepa and Miroslaw Blocho. Co-operation in the parallel memetic algorithm. International Journal of Parallel Programming, 43(5):812–839, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-014-0343-4.pdf.

Nicacio:2013:TSU

[NBA13] Daniel Nicácio, Alexandro Baldassin, and Guido Araújo. Transaction scheduling using dynamic conflict avoidance. International Journal of Parallel Programming, 41(1):

89–110, February 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0205-x.

Nanda:1998:MRC

[NBD98] Ashwini K. Nanda, James O. Bondi, and Simonjit Dutta. The misprediction recovery cache. International Journal of Parallel Programming, 26(4):383–415, August 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=4&spage=383.

Nylanden:2015:LPR

[NBN^{+}15] Teemu Nyländen, Jani Boutellier, Karri Nikunen, Jari Hannuksela, and Olli Silvén. Low-power reconfigurable miniature sensor nodes for condition monitoring. International Journal of Parallel Programming, 43(1):3– 23, February 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0302-5.

Navarro:2019:HPT

[NCR^{+}19] Angeles Navarro, Francisco Corbera, Andres Rodriguez, Antonio Vilches, and Rafael Asenjo. Het-

erogeneous parallel for

template for CPU–GPU chips. International Journal of Parallel Programming, 47(2):213–233, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Nedjah:2016:PIC

[NdMCdMMW16] Nadia Nedjah, Rogério de M. Calazan, Luiza de Macedo Mourelle, and Chao Wang. Parallel implementations of the cooperative particle swarm optimization on many-core and multi-core architectures. International Journal of Parallel Programming, 44(6):1173–1199, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0368-3.

Nedjah:2009:HPH

[NdMM09] Nadia Nedjah and Luiza de Macedo Mourelle. High-performance hardware of the sliding-window method for parallel computation of modular exponentiations. International Journal of Parallel Programming, 37(6):537–555, December 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=6&spage=537.

Nedjah:2016:PYP

[NdMMW16] Nadia Nedjah, Luiza de Macedo Mourelle, and Chao Wang. A parallel yet pipelined architecture for efficient implementation of the Advanced Encryption Standard algorithm on reconfigurable hardware. International Journal of Parallel Programming, 44(6):1102– 1117, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-016-0408-7.

Nageswaran:2009:BDV

[NFC^{+}09] Jayram Moorkanikara Nageswaran, Andrew Felch, Ashok Chandrasekhar, Nikil Dutt, Richard Granger, et al. Brain derived vision algorithm on high performance architectures. International Journal of Parallel Programming, 37(4):345– 369, August 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&

issn=0885-7458&volume=
37&issue=4&spage=345.

Ning:1992:OLS

[NG92] Qi Ning and Guang R.

Gao. Optimal loop storage allocation for argumentfetching dataflow machines. International Journal of Parallel Programming, 21(6):421–448, December 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=6&spage=421.

Nicolau:2014:AR

[Nic14]  Alex Nicolau.  Ac-

knowledgment to reviewers. International Journal of Parallel Programming, 42(6):873–874, December 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0314-9.

Nakajo:2000:DSM

[NIK00] Hironori Nakajo, Akihiro Ichikawa, and Yukio Kaneda. A distributed shared-memory system on a workstation cluster using fast serial links. International Journal of Parallel Programming, 28(2):

179–194, April 2000. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=2&spage=179.

Nakano:2003:SCG

[NIO^{+}03] Hirofumi Nakano, Kazuhisa Ishizaka, Motoki Obata, Keiji Kimura, and Hironori Kasahara. Static coarse grain task scheduling with cache optimization using OpenMP. International Journal of Parallel Programming, 31(3): 211–223, June 2003. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL /ips/

frames/Refs/referenceskapmain.
asp?J=4773&I=33&A=4&
LK=NM;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/4/abstract.
htm;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=3&spage=
211.

Ni:1988:PMH

[NK88] Lionel M. Ni and Chung-Ta King. On partitioning and mapping for hypercube computing. International Journal of Parallel Programming, 17(6):475– 495, December 1988. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=6&spage=475.

Niculescu:2023:GEN

[NL23] Virginia Niculescu and Frédéric Loulergue. Guest Editor’s note: High– level parallel programming 2021. International Journal of Parallel Programming, 51(4–5):271– 273, October 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-023-00752-x.

Norman:2023:PCC

[NLBB23] Matthew Norman, Isaac Lyngaas, Abhishek Bagusetty, and Mark Berrill. Portable C++ code that can look and feel like Fortran code with Yet Another Kernel Launcher (YAKL). International Journal of Parallel Programming, 51(4– 5):209–230, October 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00739-0.

Norden:2007:DDM

[NLRH07] Markus Nordén, Henrik Löf, Jarmo Rantakokko, and Sverker Holmgren. Dynamic data migration for structured AMR solvers. International Journal of Parallel Programming, 35(5):477–491, October 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=5&spage=477.

Novack:1995:HAI

[NN95] Steven Novack and Alexandru Nicolau. A hierarchical approach to instructionlevel parallelization. International Journal of Parallel Programming, 23(1): 35–62, February 1995. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Norris:1998:ECR

[NP98] Cindy Norris and Lori L. Pollock. Experiences with cooperating register allocation and instruction scheduling. International Journal of Parallel Programming, 26(3):241–283, June 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/

openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=3&spage=241.

Nikolopoulos:2001:AOS

[NP01] Dimitrios S. Nikolopoulos and Theodore S. Papatheodorou. The architectural and operating system implications on the performance of synchronization on ccNUMA multiprocessors. International Journal of Parallel Programming, 29(3): 249–282, June 2001. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/21/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/21/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=3&spage=
249.

Niedzielski:2019:AED

[NP19] David Niedzielski and Kleanthis Psarris. An analytical evaluation of data dependence analysis techniques. International Journal of Parallel Programming, 47(5–6): 781–804, December 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

                                        com/article/10.1007/
                                        s10766-018-0577-7.

Neirynck:1989:EAH

[NPD89] Anne Neirynck, Prakash Panangaden, and Alan J. Demers. Effect analysis in higher-order languages. International Journal of Parallel Programming, 18(1):1–36, February 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=1&spage=1.

Nau:1986:EAM

[NPT86] D. Nau, P. Purdom, and Chun-Hung Tzeng. Experiments on alternatives to minimax. International Journal of Parallel Programming, 15(2):163– 183, April 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        15&issue=2&spage=163.

Najjar:1994:EMG

[NRB94] Walid A. Najjar, Lucas Roh, and A. P. Wim Böhm. An evaluation of medium-grain dataflow code. International Journal of Parallel Programming, 22(3):209–242, June 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Natesan:2017:HBP

[NRGB17] P. Natesan, R. R. Rajalaxmi, G. Gowrison, and P. Balasubramanie. Hadoop based parallel binary bat algorithm for network intrusion detection. International Journal of Parallel Programming, 45(5):1194–1213, October 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Narasimhan:1999:UDF

[NRR99] Ragini Narasimhan, Daniel J.

Rosenkrantz, and S. S. Ravi. Using data flow information to obtain efficient check sets for algorithm-based fault tolerance. International Journal of Parallel Programming, 27(4):289–323, August 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=4&spage=289.

Najjar:1997:FSI

[NS97a] Walid A. Najjar and Gabriel M. Silberman. Foreword to the special issues. International Journal of Parallel Programming, 25(4):243–??, August 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Newburn:1997:PPP

[NS97b] Chris J. Newburn and John Paul Shen. Postpass partitioning of signal processing programs. International Journal of Parallel Programming, 25 (4):245–280, August 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Neuberger:2012:MIS

[NSS12] John M. Neuberger, Nándor Sieben, and James W. Swift. An MPI implementation of a selfsubmitting parallel job queue. International Journal of Parallel Programming, 40(4):443–464, August 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=4&spage=443.

Nicol:1989:DPS

[NST89] David M. Nicol, Joel H. Saltz, and James C. Townsend. Delay point schedules for irregular parallel computations. International Journal of Parallel Programming, 18(1): 69–90, February 1989. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=1&spage=69.

Naqash:2022:SAB

[NSU22] Talha Naqash, Sajjad Hussain Shah, and Muhammad Najam Ul Islam. Statistical analysis based intrusion detection system for ultra-high-speed software defined network. International Journal of Parallel Programming, 50 (1):89–114, February 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00715-0.

Nafea:2014:PBA

[NYHA14] Ibtehal Nafea, Muhammad Younas, Robert Holton, and Irfan Awan. A priority-based admission control scheme for commercial Web servers. International Journal of Parallel Programming, 42(5): 776–797, October 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0290-5.

Oz:2021:PSE

[ÖA21] Isil Öz and Sanem Arslan. Predicting the soft error vulnerability of parallel applications using machine learning. International Journal of Parallel Programming, 49(3): 410–439, June 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00707-0.

Ortega-Arranz:2015:CEN

[OATGEL15a]  Hector  Ortega-Arranz,

Yuri Torres, Arturo Gonzalez-Escribano, and Diego R. Llanos. Comprehensive evaluation of a new GPU-based approach to the shortest path problem. International Journal of Parallel Programming, 43(5):918–938, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        article/10.1007/s10766-
                                        015-0351-z.

Ortega-Arranz:2015:TML

[OATGEL15b]  Hector  Ortega-Arranz,

Yuri Torres, Arturo Gonzalez-Escribano, and Diego R. Llanos. TuCCompi: a multi-layer model for distributed heterogeneous computing with tuning capabilities. International Journal of Parallel Programming, 43(5): 939–960, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0349-6.

Ossner:2013:GMB

[OB13] Christopher Oßner and Klemens Böhm. Graphs for mining-based defect localization in multithreaded programs. International Journal of Parallel Programming, 41(4): 570–593, August 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0237-2.

Ottaviano:2024:CRV

[OBB^{+}24]  Alessandro  Ottaviano,

Robert Balas, Giovanni Bambini, Antonio Del Vecchio, Maicol Ciani, Davide Rossi, Luca Benini, and Andrea Bartolini. ControlPULP: a RISC-V onchip parallel power controller for many-core HPC processors with FPGAbased hardware-in-the-loop power and thermal emulation. International Journal of Parallel Programming, 52(1–2): 93–123, April 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-024-00761-4.

Ozturan:2011:GEP

[OG11] Can Ozturan and Dan Grigoras. Guest editorial: Parallel and distributed computing. International Journal of Parallel Programming, 39(5):582–583, October 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=5&spage=582.

Orozco:2016:DIT

[OGP^{+}16] Daniel Orozco, Elkin Garcia, Robert Pavel, Jaime Arteaga, and Guang Gao. The design and implementation of TIDeFlow: A dataflow-inspired execution model for parallel loops and task pipelining. International Journal of Parallel Programming, 44 (2):278–307, April 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0373-6.

OBoyle:1999:NDT

[OK99] Michael F. P. O’Boyle and Peter M. W. Knijnenburg. Nonsingular data transformations: Definition, validity, and applications. International Journal of Parallel Programming, 27(3):131–159, June 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=3&spage=131.

Ozcan:2007:MAP

[ÖO07] Ender Özcan and Esin Onbasioglu. Memetic algorithms for parallel code optimization. International Journal of Parallel Programming, 35(1): 33–61, February 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=1&spage=33.

Oh:2013:GFP

[OOR13] Yunho Oh, Doohwan Oh, and Won W. Ro. GPUfriendly parallel genome matching with tiled access and reduced state transition table. International Journal of Parallel Programming, 41(4): 526–551, August 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0234-5.

OBrien:2008:SOC

[OOS^{+}08] Kevin OB́rien, Kathryn OB́rien, Zehra Sura, Tong Chen, and Tao Zhang. Supporting OpenMP on Cell. International Journal of Parallel Programming, 36(3):289–311, June 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=36&issue=3&spage=
                                        289.

Olivier:2010:COO

[OP10] Stephen L. Olivier and Jan F. Prins. Comparison of OpenMP 3.0 and other task parallel frameworks on unbalanced task graphs. International Journal of Parallel Programming, 38(5–6):341– 360, October 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=5&spage=341.

Otoom:2012:WMI

[OP12] Mwaffaq Otoom and JoAnn M. Paul. Workload mode identification for chip heterogeneous multiprocessors. International Journal of Parallel Programming, 40(2):184–224, April 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=2&spage=184.

Ouyang:2017:HSP

[OPLS17] Aijia Ouyang, Xuyu Peng, Jing Liu, and Ahmed Sallam. Hardware/software partitioning for heterogeneous MPSoC considering communication overhead. International Journal of Parallel Programming, 45 (4):899–922, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Orailoglu:2003:GEI

[Ora03] Alex Orailoglu. Guest Editor’s introduction. International Journal of Parallel Programming, 31 (6):407–409, December 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/1/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/1/fulltext.

                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=6&spage=
                                        407.

Orailoglu:2024:SIS

[ORJ24] Alex Orailoglu, Marc Reichenbach, and Matthias Jung. Special issue on SAMOS 2022. International Journal of Parallel Programming, 52(1–2): 1–2, April 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-024-00765-0.

Ortega:2004:DMI

[OVA04] Daniel Ortega, Mateo Valero, and Eduard Ayguadé. Dynamic memory instruction bypassing. International Journal of Parallel Programming, 32(3):199– 224, June 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        32&issue=3&spage=199.

Ou:2017:GNH

[OXL^{+}17] Yang Ou, Nong Xiao, Fang Liu, Zhiguang Chen, Wei Chen, and Lizhou Wu. Gemini: a novel hardware and software implementation of high-performance PCIe SSD. International Journal of Parallel Programming, 45(4):923–945, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Panda:2008:GEI

[Pan08] Preeti Ranjan Panda.

Guest Editor introduction: Special issue on multiprocessor-based embedded systems. International Journal of Parallel Programming, 36(1): 1–2, February 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=1&spage=1.

Parallax:1986:BPP

[Par86a]  Parallax.  The bards

on parallel programming. International Journal of Parallel Programming, 15 (3):277, June 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=3&spage=277.

Parallax:1986:HPS

[Par86b] Parallax. How are parallel systems invented? International Journal of Parallel Programming, 15(1): 101–102, February 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=15&issue=1&spage=
101.

Parallax:1986:WPB

[Par86c] Parallax. When is pull better than push? (parallel programming). International Journal of Parallel Programming, 15(2): 185–188, April 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=2&spage=185.

Pandey:2001:SIE

[PB01] Raju Pandey and James C. Browne. Support for implementation of evolutionary concurrent systems. International Journal of Parallel Programming, 29(4):401–431, August 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/22/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/22/2/fulltext.pdf;

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=4&spage=
401.

Palanciuc:2004:SCM

[PB04] Virgil Palanciuc and Dragos Badea. A spill code minimization techniqueapplication in the Metrowerks StarCore C compiler. International Journal of Parallel Programming, 32(6):475–499, December 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=6&spage=475.

Penry:2013:ABS

[PC13] David A. Penry and Kurtis D. Cahill. ADL-based specification of implementation styles for functional simulators. International Journal of Parallel Programming, 41(2): 163–211, April 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0223-8.

Piccialli:2018:PAD

[PCJ18] Francesco Piccialli, Salvatore Cuomo, and Gwanggil Jeon. Parallel approaches for data mining in the Internet of Things realm. International Journal of Parallel Programming, 46(5):807– 811, October 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0565-y.pdf.

Park:2020:DAM

[PCJ20] Jihyun Park, Byoungju Choi, and Seungyeun Jang. Dynamic analysis method for concurrency bugs in multiprocess/multi-thread environments. International Journal of Parallel Programming, 48(6):1032– 1060, December 2020. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00661-3.

Park:2013:PMP

[PCP^{+}13] Eunjung Park, John Cavazos, Louis-Noël Pouchet, Cédric Bastoul, Albert Cohen, and P. Sadayappan. Predictive modeling in a polyhedral optimization space. International Journal of Parallel Programming, 41(5): 704–750, October 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0241-1.

Park:1989:DPM

[PD89] Kee-Hyun Park and Lawrence W.

Dowdy. Dynamic partitioning of multiprocessor systems. International Journal of Parallel Programming, 18(2): 91–120, April 1989. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=2&spage=91.

Papin:2021:SIP

[PDN21] Jean-Charles Papin, Christophe Denoual, and Raymond Namyst. SPAWN: An iterative, potentials-based, dynamic scheduling and partitioning tool. International Journal of Parallel Programming, 49(1): 81–103, February 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00677-9.

Plazolles:2018:SMC

[PES^{+}18] Bastien Plazolles, Didier El Baz, Martin Spel, Vincent Rivola, and Pascal Gegout. SIMD Monte-Carlo numerical simulations accelerated on GPU and Xeon Phi. International Journal of Parallel Programming, 46(3):584– 606, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Pai:2007:FFE

[PG07] Rajani Pai and R. Govindarajan. FEADS: a framework for exploring the application design space on network processors. International Journal of Parallel Programming, 35(1):1– 31, February 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=1&spage=1.

Pavlidis:2016:HSQ

[PG16] Archimedes Pavlidis and Dimitris Gizopoulos. Hierarchical synthesis of quantum and reversible architectures. International Journal of Parallel Programming, 44(5): 1028–1053, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-016-0407-8.

Palomar:2018:HPC

[PGLC^{+}18] Rafael Palomar, Juan Gómez-Luna, Faouzi A. Cheikh, Joaquín Olivares-Bueno, and Ole J. Elle. High-performance computation of Bézier surfaces on parallel and heterogeneous platforms. International Journal of Parallel Programming, 46(6):1035– 1062, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0506-1.pdf.

Peterson:2019:AHM

[PHS19] Brad Peterson, Alan Humphrey, and Dan Sunderland. Automatic halo management for the Uintah GPU–heterogeneous asynchronous many-task runtime. International Journal of Parallel Programming, 47(5–6):1086– 1116, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0619-1.

Pinter:1995:I

[Pin95] Shlomit S. Pinter. Introduction. International Journal of Parallel Programming, 23(1):3–??, February 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Pinter:1999:I

[Pin99] Shlomit S. Pinter. Introduction. International Journal of Parallel Programming, 27(4):227–228, August 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        27&issue=4&spage=227.

Popa:2018:AMH

[PIP18] Emilia Popa, Mauro Iacono, and Florin Pop. Adapting MCP and HLFET algorithms to multiple simultaneous scheduling. International Journal of Parallel Programming, 46 (3):607–629, June 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Padmanabhan:2005:EIM

[PJS^{+}05] Shobana Padmanabhan, Phillip Jones, David V. Schuehler, Scott J. Friedman, Praveen Krishnamurthy, Huakai Zhang, Roger Chamberlain, Ron K. Cytron, Jason Fritts, and John W. Lockwood. Extracting and improving microarchitecture performance on reconfigurable architectures. International Journal of Parallel Programming, 33(2–3):

115–136, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=2&spage=115.

Paulswamy:2020:QBN

[PK20] Sathees Lingam Paulswamy and Hariharan Kaluvan. Quadrant based neighbor to sink and neighbor to source routing protocol and alternate node deployment strategies for WSN. International Journal of Parallel Programming, 48 (3):447–469, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Pan:2004:DPC

[PLN^{+}04] Lei Pan, MingKin Lai, Koji Noguchi, Javid J. Huseynov, Lubomir F. Bic, and Michael B. Dillencourt. Distributed parallel computing using navigational programming. International Journal of Parallel Programming, 32(1):1–37, February 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=1&spage=1.

Paul:2007:ALR

[PM07] JoAnn M. Paul and Brett H. Meyer. Amdahl’s Law revisited for single chip systems. International Journal of Parallel Programming, 35(2): 101–123, April 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=2&spage=101.

Pingali:2003:RCT

[PMHC03] Venkata K. Pingali, Sally A. McKee, Wilson C. Hsieh, and John B. Carter. Restructuring computations for temporal data cache locality. International Journal of Parallel Programming, 31(4):305–338, August 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        ipsapp007.kluweronline.
                                        com/content/getfile/
                                        4773/34/3/abstract.htm;
                                        http://ipsapp007.kluweronline.
                                        com/content/getfile/
                                        4773/34/3/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=4&spage=
                                        305.

Papagiannopoulou:2018:HTM

[PMM^{+}18] Dimitra Papagiannopoulou, Andrea Marongiu, Tali Moreshet, Luca Benini, Maurice Herlihy, and R. Iris Bahar. Hardware transactional memory exploration in coherence-free many-core architectures. International Journal of Parallel Programming, 46 (6):1304–1328, December 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Phan:2017:ESI

[PMV17] Kien Tuong Phan, Tomas Henrique Maul, and Tuong Thuy Vu. An empirical study on improving the speed and generalization of neural networks using a parallel circuit approach. International Journal of Parallel Programming, 45(4):780– 796, August 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Petrov:2007:DTR

[PO07] Peter Petrov and Alex Orailoglu. Dynamic tag reduction for lowpower caches in embedded systems with virtual memory. International Journal of Parallel Programming, 35(2):157– 177, April 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=2&spage=157.

Panda:2010:GES

[PP10] Preeti Ranjan Panda and Rajendran Panda. Guest editorial: Special issue on VLSI design and embedded systems. International Journal of Parallel Programming, 38(3–4): 183–184, June 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=3&spage=183.

Pop:2008:AOH

[PPEP08] Traian Pop, Paul Pop, Petru Eles, and Zebo Peng. Analysis and optimisation of hierarchically scheduled multiprocessor embedded systems. International Journal of Parallel Programming, 36(1): 37–67, February 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=1&spage=37.

Pellegrini:2016:TSP

[PPQV16] Alessandro Pellegrini, Sebastiano Peluso, Francesco Quaglia, and Roberto Vitali. Transparent speculative parallelization of discrete event simulation applications using global variables. International Journal of Parallel Programming, 44(6):1200– 1247, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0429-2.

Phillips:1999:PSR

[PR99] Steven Phillips and Anne Rogers. Parallel speech recognition. International Journal of Parallel Programming, 27(4):257–288, August 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=4&spage=257.

Pratt:1986:MCP

[Pra86] Vaughan R. Pratt. Modeling concurrency with partial orders. International Journal of Parallel Programming, 15(1):33–71, February 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=1&spage=33.

Palis:1992:NAR

[PS92] Michael A. Palis and Sunil M. Shende. An NC algorithm for recognizing tree adjoining languages. International Journal of Parallel Programming, 21(2):151–167, April 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=2&spage=151.

Presser:2023:PAP

[PS23] Daniel Presser and Frank Siqueira. Partitioningaware performance modeling of distributed graph processing tasks. International Journal of Parallel Programming, 51(4– 5):231–255, October 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-023-00753-w.

Park:1997:AGT

[PSM97] Daeyeon Park, Rafael H. Saavedra, and Sungdo Moon. Adaptive granularity: Transparent integration of fine- and coarse-grain communication. International Journal of Parallel Programming, 25(5):419–446, October 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Panesar:2006:DPP

[PTD^{+}06] Gajinder Panesar, Daniel Towner, Andrew Duller, Alan Gray, and Will Robbins. Deterministic parallel processing. International Journal of Parallel Programming, 34(4):323– 341, August 2006. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=4&spage=323.

Panetta:2012:ATD

[PTdSF^{+}12] Jairo Panetta, Thiago Teixeira, Paulo R. P. de Souza Filho, Carlos A. da Cunha Filho, David Sotelo, Fernando M. Roxo da Motta, Silvio Sinedino Pinheiro, Andre L. Romanelli Rosa, Luiz R. Monnerat, Leandro T. Carneiro, and Carlos H. B. de Albrecht. Accelerating time and depth seismic migration by CPU and GPU cooperation. International Journal of Parallel Programming, 40(3): 290–312, June 2012. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=3&spage=290.

Park:1998:PPP

[PVAE98] Insung Park, Michael Voss, Brian Armstrong, and Rudolf Eigenmann. Parallel programming and performance evaluation with the URSA tool family. International Journal of Parallel Programming, 26(5):541–561, October 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        26&issue=5&spage=541.

Pothos:2021:DLI

[PVF21] V. Pothos, E. Vassalos, and N. Fragoulis. Deep learning inference with dynamic graphs on heterogeneous platforms. International Journal of Parallel Programming, 49(2): 158–176, April 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-020-00654-2.

Paya-Vaya:2017:GES

[PVG17] Guillermo Payá-Vayá and Andreas Gerstlauer. Guest editorial: Special issue on the 2015 International Conference on Embedded Computer Systems — Architectures, Modeling and Simulation (SAMOS XV). International Journal of Parallel Programming, 45(6):1417–1419, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0500-7.pdf.

Pinter:1987:MPP

[PW87] Shlomit S. Pinter and Yaron Wolfstahl. On mapping processes to processors in distributed systems. International Journal of Parallel Programming, 16(1):1–15, February 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=1&spage=1.

Palis:1992:PPT

[PW92] Michael A. Palis and David S. L. Wei. Parallel parsing of tree adjoining grammars on the Connection Machine. International Journal of Parallel Programming, 21(1): 1–38, February 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=1&spage=1.

Peng:2016:ESF

[PYC16] Lizhi Peng, Bo Yang, and Yuehui Chen. Effectiveness of statistical features for early stage Internet traffic identification. International Journal of Parallel Programming, 44(1): 181–197, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0337-2.

Pan:2017:DDC

[PYX17] Fengfeng Pan, Yinliang Yue, and Jin Xiong. dCompaction: Delayed compaction for the LSM-Tree. International Journal of Parallel Programming, 45(6):1310–1325, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Pan:2019:LAM

[PZL^{+}19] Cheng Pan, Lan Zhou, Yingwei Luo, Xiaolin Wang, and Zhenlin Wang. Lightweight and accurate memory allocation in key– value cache. International Journal of Parallel Programming, 47(3):451– 466, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Qian:2011:PAP

[QA11] Ying Qian and Ahmad Afsahi. Process arrival pattern aware Alltoall and Allgather on InfiniBand clusters. International Journal of Parallel Programming, 39(4):473–493, August 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=4&spage=473.

Quiroz-Fabian:2019:VNV

[QFRA19] José L. Quiroz-Fabián and Graciela Román-Alonso. VPPE: a novel visual parallel programming environment. International Journal of Parallel Programming, 47(5–6):1117– 1151, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00639-w.

Qin:2019:TDN

[QGT^{+}19] Lianke Qin, Yifan Gong, Tianqi Tang, Yutian Wang, and Jiangming Jin. Training deep nets with progressive batch normalization on Multi-GPUs. International Journal of Parallel Programming, 47 (3):373–387, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Quillere:2000:GEN

[QRW00] Fabien Quillere, Sanjay Rajopadhye, and Doran Wilde. Generation of efficient nested loops from polyhedra. International Journal of Parallel Programming, 28(5):469–498, October 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=5&spage=469.

Wang:2017:PAG

[qWlJzKhC17] Xiao qing Wang, Xian long Jin, Da zhi Kou, and Jia hui Chen. A parallel approach for the generation of unstructured meshes with billions of elements on distributedmemory supercomputers. International Journal of Parallel Programming, 45 (3):680–710, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Qu:2015:DBA

[QZP15] Yun R. Qu, Shijie Zhou, and Viktor K. Prasanna. A decomposition-based approach for scalable manyfield packet classification on multi-core processors. International Journal of Parallel Programming, 43(6):965–987, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0325-6.

Rajagopalan:1994:SSP

[RA94] M. Rajagopalan and V. H.

Allan. Specification of software pipelining using Petri nets. International Journal of Parallel Programming, 22(3):273– 301, June 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rashti:2009:SAM

[RA09] Mohammad J. Rashti and Ahmad Afsahi. A speculative and adaptive MPI rendezvous protocol over RDMA-enabled interconnects. International Journal of Parallel Programming, 37(2):223– 246, April 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=2&spage=223.

Rauchwerger:1995:SMR

[RAP95] Lawrence Rauchwerger, Nancy M. Amato, and David A. Padua. A scalable method for run-time loop parallelization. International Journal of Parallel Programming, 23(6): 537–576, December 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rau:1996:IMS

[Rau96] B. Ramakrishna Rau. Iterative modulo scheduling. International Journal of Parallel Programming, 24 (1):3–64, February 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rana:1986:ODS

[RB86] S. P. Rana and D. K. Banerji. An optimal distributed solution to the dining philosophers problem. International Journal of Parallel Programming, 15(4):327–335, August 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&

issn=0885-7458&volume=
15&issue=4&spage=327.

Rohou:2000:HGC

[RBES00] Erven Rohou, François Bodin, Christine Eisenbeis, and Andre Seznec. Handling global constraints in compiler strategy. International Journal of Parallel Programming, 28(4): 325–345, August 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=28&issue=4&spage=
325.

Rivault:2022:SSJ

[RBR22] Sébastien Rivault, Mostafa Bamha, and Sophie Robert. A scalable similarity join algorithm based on MapReduce and LSH. International Journal of Parallel Programming, 50(3– 4):360–380, August 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00733-6.

Rahman:2016:ASO

[RC16] Musfiq Rahman and Bruce R.

Childers. Asteroid: Scalable online memory diagnostics for multi-core, multi-socket servers. International Journal of Parallel Programming, 44(5): 949–974, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0400-2.

Rounce:2008:DIS

[RD08] Peter A. Rounce and Alberto F. De Souza. Dynamic instruction scheduling in a trace-based multithreaded architecture. International Journal of Parallel Programming, 36 (2):184–205, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=36&issue=2&spage=
184.

Reguly:2015:FEA

[RG15] I. Z. Reguly and M. B. Giles. Finite element algorithms and data structures on graphical processing units. International Journal of Parallel Programming, 43(2):203– 239, April 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0301-6.

Rasch:2018:MDH

[RG18] Ari Rasch and Sergei Gorlatch. Multi-dimensional homomorphisms and their implementation in OpenCL. International Journal of Parallel Programming, 46 (1):101–119, February 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ruggiero:2008:FAT

[RGB^{+}08] Martino Ruggiero, Alessio Guerri, Davide Bertozzi, Michela Milano, and Luca Benini. A fast and accurate technique for mapping parallel applications on stream-oriented MPSoC platforms with communication awareness. International Journal of Parallel Programming, 36(1): 3–36, February 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=1&spage=3.

Rice:1990:SDP

[Ric90] Michael D. Rice. Semantics for data parallel computation. International Journal of Parallel Programming, 19(6):477–509, December 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        19&issue=6&spage=477.

Reichenbach:2022:GES

[RJO22] Marc Reichenbach, Matthias Jung, and Alex Orailoglu. Guest editorial: Special issue on 2020 IEEE International Conference on Embedded Computer Systems: Architectures, Modeling and Simulation (SAMOS 2020). International Journal of Parallel Programming, 50(2): 187–188, April 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00732-7.

Rao:1987:PDF

[RK87] V. Nageshwara Rao and Vipin Kumar. Parallel depth first search. Part I. Implementation. International Journal of Parallel Programming, 16(6):479– 499, December 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        16&issue=6&spage=479.

Ramkumar:1992:JAC

[RK92]  Balkrishna  Ramkumar

and Laxmikant V. Kalé.

A join algorithm for combining AND parallel solutions in AND/OR parallel systems. International Journal of Parallel Programming, 21(1):67–107, February 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=1&spage=67.

Rovers:2013:UUC

[RK13] Kenneth C. Rovers and Jan Kuper. UniTi: Unified composition and time for multi-domain modelbased design. International Journal of Parallel Programming, 41(2):261– 304, April 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0226-5.

Rajwar:2004:IQS

[RKG04] Ravi Rajwar, Alain Kägi, and James R. Goodman. Inferential queueing and speculative push. International Journal of Parallel Programming, 32(3): 225–258, June 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        32&issue=3&spage=225.

Reumont-Locke:2019:EMT

[RLEJ19] Fabien Reumont-Locke and Naser Ezzati-Jivan. Efficient methods for trace analysis parallelization. International Journal of Parallel Programming, 47(5–6):951–972, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-019-00631-4.

Ramezani:2014:TBS

[RLH14] Fahimeh Ramezani, Jie Lu, and Farookh Khadeer Hussain. Task-based system load balancing in cloud computing using particle swarm optimization. International Journal of Parallel Programming, 42(5):739–754, October 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0275-4.

Ryu:2020:ACE

[RLK20] Seokhoon Ryu, Young-Sup Lee, and Seonghyun Kim. Active control of engine sound quality in a passenger car using a virtual error microphone. International Journal of Parallel Programming, 48(5): 909–927, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00633-2.

Ramirez:2002:STC

[RLPN^{+}02] Alex Ramirez, Josep Ll.

Larriba-Pey, Carlos Navarro, Mateo Valero, and Josep Torrellas. Software trace cache for commercial applications. International Journal of Parallel Programming, 30(5):373–395, October 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.
com/content/getfile/
4773/29/2/abstract.htm;
http://ipsapp009.lwwonline.
com/content/getfile/
4773/29/2/fulltext.pdf;
http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=5&spage=
373.

Rodriguez:2013:CAC

[RMG^{+}13] Gabriel Rodríguez, María J.

Martín, Patricia González, Juan Touriño, and Ramón Doallo. Compiler-assisted checkpointing of parallel codes: The Cetus and LLVM experience. International Journal of Parallel Programming, 41(6): 782–805, December 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0231-8.

Rheindt:2021:DDS

[RMH21] Sven Rheindt, Sebastian Maier, and Andreas Herkersdorf. DySHARQ: Dynamic software-defined hardware-managed queues for tile-based architectures. International Journal of Parallel Programming, 49(4):506–540, August 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00687-7.

Rast:2012:MBS

[RNJ^{+}12] Alexander D. Rast, Javier Navaridas, Xin Jin, Francesco Galluppi, Luis A. Plana, et al. Managing burstiness and scalability in eventdriven models on the SpiN-Naker neuromimetic system. International Journal of Parallel Programming, 40(6):553–582, December 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/

openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=6&spage=553.

Roy:2010:HNE

[Roy10] Sourav Roy. H-NMRU:

an efficient cache replacement policy with low area. International Journal of Parallel Programming, 38(3–4):277–287, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=3&spage=277.

Rad:2018:EPS

[RPF18] Mina Hosseini Rad, Ahmad Patooghy, and Mahdi Fazeli. An efficient programming skeleton for clusters of multi-core processors. International Journal of Parallel Programming, 46(6):1094– 1109, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rus:2003:HAS

[RRH03] Silvius Rus, Lawrence Rauchwerger, and Jay Hoeflinger. Hybrid analysis: Static & dynamic memory reference analysis. International Journal of Parallel Programming, 31(4):251–283, August 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/34/1/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/34/1/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=4&spage=
251.

Reif:1990:DFA

[RS90] John H. Reif and Scott A. Smolka. Data flow analysis of distributed communicating processes. International Journal of Parallel Programming, 19(1): 1–30, February 1990. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=1&spage=1.

Rathore:2018:RTB

[RSA^{+}18] M. Mazhar Rathore, Hojae Son, Awais Ahmad, Anand Paul, and Gwanggil Jeon. Real-time big data stream processing using GPU with Spark over Hadoop ecosystem. International Journal of Parallel Programming, 46(3): 630–646, June 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rosas:2014:IPD

[RSJ^{+}14] Claudia Rosas, Anna Sikora, Josep Jorba, Andreu Moreno, and Eduardo César. Improving performance on data– intensive applications using a load balancing methodology based on divisible load theory. International Journal of Parallel Programming, 42(1): 94–118, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0199-4.

Rucci:2019:SES

[RSJ^{+}19] Enzo Rucci, Carlos Garcia Sanchez, Guillermo Botella Juan, Armando De Giusti, Marcelo Naiouf, and Manuel Prieto-Matias. SWIMM 2.0: Enhanced Smith– Waterman on Intel’s multicore and manycore architectures based on AVX-512 vector extensions. International Journal of Parallel Programming, 47 (2):296–316, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Rasmussen:2009:PSI

[RSK09] Morten S. Rasmussen, Matthias B. Stuart, and Sven Karlsson. Parallelism and scalability in an image processing application. International Journal of Parallel Programming, 37(3):306–323, June 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=3&spage=306.

Ram:2020:AEL

[RSP20] R. Saravana Ram, A. Gopi Saminathan, and S. Arun Prakash. An area efficient and low power consumption of run time digital system based on dynamic partial reconfiguration. International Journal of Parallel Programming, 48(3):431–446, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Roberti:2005:PIL

[RSV^{+}05] Debora R. Roberti, Roberto P. Souto, Haroldo F. Campos Velho, Gervasio A. Degrazia, and Domenico Anfossi. Parallel implementation of a Lagrangian stochastic model for pollutant dispersion. International Journal of Parallel Programming, 33(5):485– 498, October 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=5&spage=485.

Rinaldi:2020:IPA

[RTD20] Luca Rinaldi, Massimo Torquati, and Marco Danelutto. Improving the performance of actors on multi-cores with parallel patterns. International Journal of Parallel Programming, 48(4): 692–712, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00663-1.

Razilov:2024:AIP

[RWMF24] Viktor Razilov, Robert Wittig, Emil Matúš, and Gerhard Fettweis. Access interval prediction by partial matching for tightly coupled memory systems. International Journal of Parallel Programming, 52 (1–2):3–19, April 2024. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00764-1.

Ramesh:2020:QQE

[RY20] S. Ramesh and C. Yaashuwanth.

QoS and QoE enhanced resource allocation for wireless video sensor networks using hybrid optimization algorithm. International Journal of Parallel Programming, 48(2):192– 212, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). See retraction notice [RY22].

Ramesh:2022:RNQ

[RY22] S. Ramesh and C. Yaashuwanth. Retraction note: QoS and QoE enhanced resource allocation for wireless video sensor networks using hybrid optimization algorithm. International Journal of Parallel Programming, 50(5– 6):562, December 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00738-1. See

[RY20].

Suri:2010:IAP

[SA10] Tameesh Suri and Aneesh Aggarwal. Improving adaptability and per-core performance of many-core processors through reconfiguration. International Journal of Parallel Programming, 38(3–4):203– 224, June 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=3&spage=203.

Simburger:2019:PPO

[SA19] Andreas Simbürger and Sven Apel. PolyJIT: Polyhedral optimization just in time. International Journal of Parallel Programming, 47(5–6): 874–906, December 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0597-3.

Sen:2011:SCB

[SAB11] Alper Sen, Baris Aksanli, and Murat Bozkurt. Speeding up cycle based logic simulation using graphics processing units. International Journal of Parallel Programming, 39 (5):639–661, October 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=39&issue=5&spage=
639.

Shabbir:2020:NPE

[SAI^{+}20] Ghulam Shabbir, Adeel Akram, Muhammad Munwar Iqbal, Sohail Jabbar, Mai Alfawair, and Junaid Chaudhry. Network performance enhancement of multi-sink enabled low power lossy networks in SDN based Internet of Things. International Journal of Parallel Programming, 48(2):367–398, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Siddique:2016:PRE

[SAL16] Abubakar Siddique, Mohammad Ansari, and Mikel Luján. Purge– Rehab: Eager software transactional memory with high performance under contention. International Journal of Parallel Programming, 44(6):1359– 1383, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-016-0427-4.pdf.

Sarkar:2001:OUN

[Sar01]  Vivek Sarkar.  Opti-

mized unrolling of nested loops. International Journal of Parallel Programming, 29(5):545–581, October 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/ipsapp009.lwwonline.

com/content/getfile/
4773/23/4/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/23/4/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=29&issue=5&spage=
545.

Silva:2018:PAS

[SAS18] Jorge Silva, Ana Aguiar, and Fernando Silva. Parallel asynchronous strategies for the execution of feature selection algorithms. International Journal of Parallel Programming, 46 (2):252–283, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Sung:2012:DLT

[SASH12] I-Jui Sung, Nasser Anssari, John A. Stratton, and Wen-Mei W. Hwu. Data layout transformation exploiting memory-level parallelism in structured grid many-core applications. International Journal of Parallel Programming, 40(1):4–24, February 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=1&spage=4.

Sengupta:1990:CCO

[SB90] Soumitra Sengupta and Arthur J. Bernstein. Concurrency control optimizations in a Prolog database. International Journal of Parallel Programming, 19(3):185–211, June 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        19&issue=3&spage=185.

Singh:1991:RET

[SB91] Gurdip Singh and Arthur J. Bernstein. On the relative execution times of distributed protocols. International Journal of Parallel Programming, 20(3): 203–235, June 1991. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        20&issue=3&spage=203.

Sourouri:2017:PCF

[SBC17] Mohammed Sourouri, Scott B. Baden, and Xing Cai. Panda: a compiler framework for concurrent CPU + GPU execution of 3D stencil computations on GPU-accelerated supercomputers. International Journal of Parallel Programming, 45(3):711–729, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Surendra:2003:EFA

[SBN03] G. Surendra, S. Banerjee, and S. K. Nandy. On the effectiveness of flow aggregation in improving instruction reuse in network processing applications. International Journal of Parallel Programming, 31 (6):469–487, December 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/5/abstract.
htm; http://ipsapp007.
kluweronline.com/content/
getfile/4773/37/5/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=6&spage=
469.

Stepoway:1988:PRF

[SC88] Stephen L. Stepoway and Michael Christiansen. Parallel rendering of fractal surfaces. International Journal of Parallel Programming, 17(1): 43–58, February 1988. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        17&issue=1&spage=43.

Scarpazza:2011:TPT

[Sca11] Daniele Paolo Scarpazza. Top-performance tokenization and small-ruleset regular expression matching: a quantitative performance analysis and optimization study on the Cell/B.E. Processor. International Journal of Parallel Programming, 39(1):3–32, February 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=1&spage=3.

Silva:2014:EDE

[SCB^{+}14] Gabriel P. Silva, Juliana Correa, Cristiana Bentes, Sergio Guedes, and Mariela Gabioux. The experience in designing and evaluating the high performance cluster Netuno. International Journal of Parallel Programming, 42(2):265– 286, April 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0224-7.

Schwab:1992:EPG

[Sch92] Stephen A. Schwab. Extended parallelism in the Gröbner basis algorithm. International Journal of Parallel Programming, 21(1):39–66, February 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=1&spage=39.

Sylvestre:2023:AOP

[SCS23] Loïc Sylvestre, Emmanuel Chailloux, and Jocelyn Sérot. Accelerating OCaml programs on FPGA. International Journal of Parallel Programming, 51(2– 3):186–207, June 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00748-z.

Shriraman:2011:ACH

[SD11] Arrvindh Shriraman and Sandhya Dwarkadas. Analyzing conflicts in hardwaresupported memory transactions. International Journal of Parallel Programming, 39(1):33–61, February 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=1&spage=33.

Sagi:2022:FGP

[SDH22] Mark Sagi, Nguyen Anh Vu Doan, and Andreas Herkersdorf. Fine-grained power modeling of multicore processors using FFNNs. International Journal of Parallel Programming, 50(2):243–266, April 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL https:

//link.springer.com/
article/10.1007/s10766-
022-00730-9.

So:1998:MCG

[SDJS98] John John E. So, Thomas J. Downar, Raghunandan Janardhan, and Howard Jay Siegel. Mapping conjugate gradient algorithms for neutron diffusion applications onto SIMD, MIMD, and mixed-mode machines. International Journal of Parallel Programming, 26(2):183–207, April 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=2&spage=183.

Sun:2017:RWM

[SDL17] Yanan Sun, Yuyue Du, and Maozhen Li. A repair of workflow models based on mirroring matrices. International Journal of Parallel Programming, 45(4):1001–1020, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Santos:2021:END

[SdLC21] Paulo C. Santos, João P. C. de Lima, and Luigi Carro. Enabling neardata accelerators adoption by through investigation of datapath solutions. International Journal of Parallel Programming, 49 (2):237–252, April 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00674-y.

Sehr:1998:GEI

[Seh98] David Sehr. Guest Editor’s introduction. International Journal of Parallel Programming, 26(1): 3–4, February 1998. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=1&spage=3.

Seki:2009:DRI

[Sek09] Masatoshi Seki. dRuby and Rinda: Implementation and application of Distributed Ruby and its parallel coordination mechanism. International Journal of Parallel Programming, 37(1):37–57, February 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=37.

Shee:2008:AEH

[SEP08] Seng Lin Shee, Andrea Erdos, and Sri Parameswaran. Architectural exploration of heterogeneous multiprocessor systems for JPEG. International Journal of Parallel Programming, 36(1):140–162, February 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=1&spage=140.

Shan:2020:GAP

[SF20] Bowei Shan and Yong Fang. GPU accelerated parallel algorithm of sliding-window belief propagation for LDPC codes. International Journal of Parallel Programming, 48(3):566–579, June 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Steuwer:2014:IIA

[SFAG14] Michel Steuwer, Malte Friese, Sebastian Albers, and Sergei Gorlatch. Introducing and implementing the allpairs skeleton for programming multi-GPU systems. International Journal of Parallel Programming, 42(4): 601–618, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0265-6.

Sreraman:2000:VCM

[SG00] N. Sreraman and R. Govindarajan. A vectorizing compiler for multimedia extensions. International Journal of Parallel Programming, 28(4):363–400, August 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=4&spage=363.

Saito:2003:LSP

[SGJ^{+}03] Hideki Saito, Greg Gaertner, Wesley Jones, Rudolf Eigenmann, Hidetoshi Iwashita, Ron Lieberman, Matthijs van Waveren, and Brian Whitney. Large system performance of SPEC OMP benchmark suites. International Journal of Parallel Programming, 31(3):197–209, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL /ips/frames/Refs/
referenceskapmain.asp?
J=4773&I=33&A=3&LK=NM;
 http://ipsapp007.kluweronline.
com/content/getfile/
4773/33/3/abstract.htm;
 http://ipsapp007.kluweronline.
com/content/getfile/
4773/33/3/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=3&spage=
197.

Salapura:2012:GEP

[SGK12] Valentina Salapura, Michael Gschwind, and Jens Knoop. Guest editorial: Parallel systems and compilers. International Journal of Parallel Programming, 40(1):1–3, February 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=1&spage=1.

Samal:1987:PCL

[SH87] Ashok Samal and Tom Henderson. Parallel consistent labeling algorithms. International Journal of Parallel Programming, 16(5):341–364, October 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=5&spage=341.

Shen:1996:HLC

[SH96] Kish Shen and Manuel V.

Hermenegildo. Highlevel characteristics of OR- and Independent AND-parallelism in Prolog. International Journal of Parallel Programming, 24(5):433–478, October 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Su:2015:SDD

[SH15] Gong Su and Stephen Heisig. The scalability of disjoint data structures on a new hardware transactional memory system. International Journal of Parallel Programming, 43(6):1192– 1217, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0322-9.

Slagter:2015:AME

[SHC15] Kenn Slagter, Ching-Hsien Hsu, and Yeh-Ching Chung. An adaptive and memory efficient sampling mechanism for partitioning in MapReduce. International Journal of Parallel Programming, 43(3): 489–507, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0288-z.

Sarvestani:2013:ERA

[SHK13] Amin Shafiee Sarvestani, Erik Hansson, and Christoph Kessler. Extensible recognition of algorithmic patterns in DSP programs for automatic parallelization. International Journal of Parallel Programming, 41(6): 806–824, December 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0229-2.

Shao:2017:FFC

[SHLJ17] Zhiyuan Shao, Jian He, Huiming Lv, and Hai Jin. FOG: a fast out-of-core graph processing framework. International Journal of Parallel Programming, 45(6):1259–1272, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Stahl:2021:DFD

[SHS21] Rafael Stahl, Alexander Hoffman, and Ulf Schlichtmann. DeeperThings: Fully distributed CNN inference on resourceconstrained edge devices. International Journal of Parallel Programming, 49 (4):600–624, August 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00712-3.

Susswein:1991:PPC

[SHZ^{+}91] Steven Y. Susswein, Thomas C.

Henderson, Joseph L. Zachary, Chuck Hansen, Paul Hinker, and Gary C. Marsden. Parallel path consistency. International Journal of Parallel Programming, 20(6):453–473, December 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
20&issue=6&spage=453.

Song:2014:OBS

[SHZ^{+}14] Fei Song, Daochao Huang, Huachun Zhou, Hongke Zhang, and Ilsun You. An optimization-based scheme for efficient virtual machine placement. International Journal of Parallel Programming, 42(5): 853–872, October 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0274-5.

Sanci:2011:PAU

[SI11] Seçkin Sanci and Veysi Isler. A parallel algorithm for UAV flight route planning on GPU. International Journal of Parallel Programming, 39(6):809– 837, December 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=6&spage=809.

Shahbahrami:2006:ACR

[SJBV06] Asadollah Shahbahrami, Ben Juurlink, Demid Borodin, and Stamatis Vassiliadis. Avoiding conversion and rearrangement overhead in SIMD architectures. International Journal of Parallel Programming, 34(3):237– 260, June 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=3&spage=237.

Singh:2018:SGB

[SJC18] Dhirendra Pratap Singh, Ishan Joshi, and Jaytrilok Choudhary. Survey of GPU based sorting algorithms. International Journal of Parallel Programming, 46(6):1017– 1034, December 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Sasakura:1999:NIV

[SJKA99] Mariko Sasakura, Kazuki Joe, Yoshitoshi Kunieda, and Keijiro Araki. NaraView: an interactive 3D visualization system for parallelization of programs. International Journal of Parallel Programming, 27(2):111–129, April 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=2&spage=111.

Sundararajan:2013:SCE

[SJT13] Karthik T. Sundararajan, Timothy M. Jones, and Nigel P. Topham. The Smart cache: an energyefficient cache architecture through dynamic adaptation. International Journal of Parallel Programming, 41(2):305–330, April 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0220-y.

Steiner:2022:DOS

[SJW22] Lukas Steiner, Matthias Jung, and Norbert Wehn. DRAMSys4.0: an opensource simulation framework for in-depth DRAM analyses. International Journal of Parallel Programming, 50(2):217–242, April 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL https:

//link.springer.com/
article/10.1007/s10766-
022-00727-4.

Schlansker:1997:TCP

[SK97] Michael Schlansker and Vinod Kathail. Techniques for critical path reduction of scalar programs. International Journal of Parallel Programming, 25(3): 147–181, June 1997. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Sun:2014:AVP

[SK14] Enqiang Sun and David Kaeli. Aggressive value prediction on a GPU. International Journal of Parallel Programming, 42 (1):30–48, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0232-7.

Schlansker:1996:PCR

[SKA96] Michael Schlansker, Vinod Kathail, and Sadun Anik. Parallelization of control recurrences for ILP processors. International Journal of Parallel Programming, 24(1):65–102, February 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Singh:1991:EAP

[SKAT91] V. Singh, V. Kumar, G. Agha, and C. Tomlinson. Efficient algorithms for parallel sorting on mesh multicomputers. International Journal of Parallel Programming, 20(2):95–131, April 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        20&issue=2&spage=95.

Shi:2009:BIO

[SKG09] Guochun Shi, Volodymyr Kindratenko, and Steven Gottlieb. The bottomup implementation of one MILC lattice QCD application on the Cell Blade. International Journal of Parallel Programming, 37(5):488–507, October 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=5&spage=488.

Skillicorn:1991:MPP

[Ski91] D. B. Skillicorn. Models for practical parallel computation. International Journal of Parallel Programming, 20(2):133– 158, April 1991. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        20&issue=2&spage=133.

Actually appeared in 1992.

Schneider:2014:LBD

[SL14] Joerg Schneider and Barry Linnert. List-based data structures for efficient management of advance reservations. International Journal of Parallel Programming, 42(1): 77–93, February 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0219-4.

Shen:2013:ITI

[SLZB13] Xipeng Shen, Yixun Liu, Eddy Z. Zhang, and Poornima Bhamidipati. An infrastructure for tackling input-sensitivity of GPU program optimizations. International Journal of Parallel Programming, 41(6):855–869, December 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0236-3.

Sharma:2009:MLP

[SM09] Gaurav Sharma and Jos Martin. MATLAB^{(}R): a language for parallel computing. International Journal of Parallel Programming, 37(1):3– 36, February 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=3.

Sato:2016:GIT

[SM16] Shigeyuki Sato and Kiminori Matsuzaki. A generic implementation of tree skeletons. International Journal of Parallel Programming, 44(3): 686–707, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0365-6.

Scott:1994:FCF

[SMC94] Michael L. Scott and John M. Mellor-Crummey. Fast, contention-free combining tree barriers for shared-memory multiprocessors. International Journal of Parallel Programming, 22(4):449–481, August 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Shao:2019:BEB

[SMDJ19] Zhiyuan Shao, Zhenjie Mei, Xiaofeng Ding, and Hai Jin. BlockGraphChi: Enabling block update in out-of-core graph processing. International Journal of Parallel Programming, 47(4):668–685, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Srivatsa:2021:DDC

[SMH21] Akshay Srivatsa, Mostafa Mansour, and Andreas Herkersdorf. DynaCo: Dynamic coherence management for tiled manycore architectures. International Journal of Parallel Programming, 49(4): 570–599, August 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00688-6.

Stojcev:1994:OSP

[SMM94] M. K. Stojčev, E. I. Milovanović, and I. Ž. Milovanović. An optimal scheduling procedure for matrix inversion on linear array at a processor level. International Journal of Parallel Programming, 22 (4):435–448, August 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Salapura:2011:GEI

[SMM11] Valentina Salapura, José E.

Moreira, and Sally A. McKee. Guest editors introduction. International Journal of Parallel Programming, 39(1):1– 2, February 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/

                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        39&issue=1&spage=1.

Schepke:2009:PLB

[SMN09] Claudio Schepke, Nicolas Maillard, and Philippe O. A. Navaux. Parallel lattice Boltzmann method with blocked partitioning. International Journal of Parallel Programming, 37(6):593–611, December 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=6&spage=593.

Schepke:2013:OMR

[SMSH13] Claudio Schepke, Nicolas Maillard, Joerg Schneider, and Hans-Ulrich Heiss. Online mesh refinement for parallel atmospheric models. International Journal of Parallel Programming, 41(4): 552–569, August 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0235-4.

Sarojadevi:2004:CPE

[SNB04] H. Sarojadevi, S. K. Nandy, and S. Balakrishnan. On the correctness of program execution when cache coherence is maintained locally at datasharing boundaries in distributed shared memory multiprocessors. International Journal of Parallel Programming, 32(5):415– 446, October 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=5&spage=415.

Suss:2021:PFC

[SNS21] Tim Süß, Lars Nagel, and Thomas Soddemann. Pure functions in C: A small keyword for automatic parallelization. International Journal of Parallel Programming, 49(1):1– 24, February 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00660-4.

Singh:1989:DEP

[SO89] Ambuj K. Singh and Ross Overbeek. Derivation of efficient parallel programs: an example from genetic sequence analysis. International Journal of Parallel Programming, 18(6):447–484, December 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=6&spage=447.

Sprague:1992:PAC

[Spr92] Alan P. Sprague. A parallel algorithm to construct a dominance graph on non-overlapping rectangles. International Journal of Parallel Programming, 21(4):303–312, August 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        21&issue=4&spage=303.

Steinbrecher:2014:CSI

[SPS14] Johann Steinbrecher, Cesar J. Philippidis, and Weijia Shang. A case study of implementing supernode transformations. International Journal of Parallel Programming, 42 (2):320–342, April 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0248-7.

Seevers:1992:PPE

[SQH92] Bradley K. Seevers, Michael J. Quinn, and Philip J. Hatcher. A parallel programming environment supporting multiple dataparallel modules. International Journal of Parallel Programming, 21(5):363– 386, October 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
21&issue=5&spage=363.

Sosic:1990:PAL

[SR90] Rok Sosic and Richard F.

Riesenfeld. Parallel algorithms for line generation. International Journal of Parallel Programming, 19(5):389–404, October 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=5&spage=389.

Souravlas:2004:PTD

[SR04] Stavros Souravlas and Manos Roumeliotis. A pipeline technique for dynamic data transfer on a multiprocessor Grid. International Journal of Parallel Programming, 32 (5):361–388, October 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=

                                        article&issn=0885-7458&
                                        volume=32&issue=5&spage=
                                        361.

Schreiber:2015:ICB

[SR15] Martin Schreiber and Christoph Riesinger. Invasive compute balancing for applications with shared and hybrid parallelization. International Journal of Parallel Programming, 43(6):1004– 1027, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-014-0336-3.

Sanchez:2006:ETA

[SRS06] Ernesto Sánchez, Matteo Sonza Reorda, and Giovanni Squillero. Efficient techniques for automatic verification-oriented test set optimization. International Journal of Parallel Programming, 34 (1):93–109, March 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=34&issue=1&spage=
                                        93.

Somogyi:1988:BAS

[SRV88] Z. Somogyi, K. Ramamohanarao, and J. Vaghani. A backtracking algorithm for the stream ANDparallel execution of logic programs. International Journal of Parallel Programming, 17(3):207–257, June 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Springsteel:1989:PGP

[SS89] Frederick Springsteel and Ivan Stojmenović. Parallel general prefix computations with geometric, algebraic, and other applications. International Journal of Parallel Programming, 18(6):485–503, December 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
18&issue=6&spage=485.

Sarkar:1992:PAS

[SS92] Dilip Sarkar and Ivan Stojmenović. Parallel algorithms for separation of two sets of points and recognition of digital convex polygons. International Journal of Parallel Programming, 21(2):109– 121, April 1992. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        21&issue=2&spage=109.

Sazeides:1999:LDV

[SS99] Yiannakis Sazeides and James E. Smith. Limits of data value predictability. International Journal of Parallel Programming, 27(4):229–256, August 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        27&issue=4&spage=229.

Shan:2001:CMS

[SS01] Hongzhang Shan and Jaswinder Pal Singh. A comparison of MPI, SHMEM and cache-coherent shared address space programming models on a tightly-coupled multiprocessors. International Journal of Parallel Programming, 29(3):283–318, June 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/21/3/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/21/3/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&

volume=29&issue=3&spage=
283.

Scholz:2010:GEE

[SS10] Sven-Bodo Scholz and Alex Shafarenko. Guest editors’ editorial: Special issue on the Second International Workshop on Microgrids. International Journal of Parallel Programming, 38(1): 1–3, February 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
38&issue=1&spage=1.

Sivakumaran:2017:PBY

[SS17] Krupa Sivakumaran and Arul Siromoney. Priority based yield of shared cache to provide cache QoS in multicore systems. International Journal of Parallel Programming, 45(3): 634–656, June 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Silva:2023:EHL

[SS23] Rui S. Silva and João L.

Sobral. Efficient highlevel programming in plain Java. International Journal of Parallel Programming, 51(1): 22–42, February 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00747-0.

Sotomayor:2017:ACG

[SSB^{+}17] Rafael Sotomayor, Luis Miguel Sanchez, Javier Garcia Blas, Javier Fernandez, and J. Daniel Garcia. Automatic CPU/GPU generation of multi-versioned OpenCL kernels for C++ scientific applications. International Journal of Parallel Programming, 45 (2):262–282, April 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0425-6.

Spiliotis:2021:PCD

[SSB21] Iraklis M. Spiliotis, Charalampos Sitaridis, and Michael P. Bekakos. Parallel computation of discrete orthogonal moment on block represented images using OpenMP. International Journal of Parallel Programming, 49(3): 440–462, June 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00713-2.

Sankaraiah:2014:POV

[SSEA14] S. Sankaraiah, Lam Hai Shuan, C. Eswaran, and Junaidi Abdullah. Performance optimization of video coding process on multi-core platform using GOP level parallelism. International Journal of Parallel Programming, 42(6):931–947, December 2014. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0267-4.

Seewald:2021:CGC

[SSM21] Adam Seewald, Ulrik Pagh Schultz, and Henrik Skov Midtiby. Coarse-grained computation-oriented energy modeling for heterogeneous parallel embedded systems. International Journal of Parallel Programming, 49(2):136– 157, April 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00645-y.

Sterling:1996:EEC

[SSMO96] Thomas Sterling, Daniel Savarese, Phillip Merkey, and Kevin Olson. An empirical evaluation of the convex SPP-1000 hierarchical shared memory system. International Journal of Parallel Programming, 24(4):377–396, August 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Sbirlea:2016:SEU

[SSNS16] Dragos Sbîrlea, Jun Shirako, Ryan Newton, and Vivek Sarkar. SCnC: Efficient unification of streaming with dynamic task parallelism. International Journal of Parallel Programming, 44(2): 233–256, April 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0353-x.

Sheffler:1996:EDA

[SSP^{+}96] Thomas J. Sheffler, Robert Schreiber, William Pugh, John R. Gilbert, and Siddhartha Chatterjee. Efficient distribution analysis via graph contraction. International Journal of Parallel Programming, 24(6):599–620, December 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Saito:2000:DPC

[SSP^{+}00] Hideki Saito, Nicholas J.

Stavrakos, Constantine D. Polychronopoulos, et al. The design of the PROMIS compiler-towards multilevel parallelization. International Journal of Parallel Programming, 28(2): 195–212, April 2000. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
28&issue=2&spage=195.

Stypka:2018:MLN

[STB^{+}18] Jan Stypka, Wojciech Turek, Aleksander Byrski, Marek Kisiel-Dorohinicki, Adam D. Barwell, Christopher Brown, Kevin Hammond, and Vladimir Janjic. The missing link! A new skeleton for evolutionary multi-agent systems in Erlang. International Journal of Parallel Programming, 46(1):4– 22, February 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0503-4.pdf.

Santos:2012:UEK

[STF^{+}12] Artur Santos, João Marcelo Teixeira, Thiago Farias, Veronica Teichrieb, and Judith Kelner. Understanding the efficiency of kD-tree ray-traversal techniques over a GPGPU architecture. International Journal of Parallel Programming, 40(3):331– 352, June 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=3&spage=331.

Sukhwani:2015:HSA

[STM15] Bharat Sukhwani, Mathew Thoennes, and Hong Min. A hardware/software approach for database query acceleration with FP-GAs. International Journal of Parallel Programming, 43(6):1129–1159, December 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0327-4.

Smiljkovic:2017:DSL

[SÜCV17] Vesna Smiljković, Osman Ünsal, Adrián Cristal, and Mateo Valero. Determinism at standardlibrary level in TM-based applications. International Journal of Parallel Programming, 45(1): 17–29, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0383-4.

Sundell:2011:WFM

[Sun11] Håkan Sundell. Wait-free multi-word compare-and-swap using greedy helping and grabbing. International Journal of Parallel Programming, 39(6):694– 716, December 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=6&spage=694.

Stoltz:1995:DVB

[SW95] Eric Stoltz and Michael Wolfe. Detecting valuebased scalar dependence. International Journal of Parallel Programming, 23 (4):327–358, August 1995. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Siek:2016:ARD

[SW16] Konrad Siek and Pawel T.

Wojciechowski. Atomic RMI: a distributed transactional memory framework. International Journal of Parallel Programming, 44(3):598–619, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-015-0361-x.pdf.

Swain:1988:CSH

[Swa88] Michael J. Swain. Comments on A. Samal and T. Henderson: “Parallel consistent labeling algorithms” [Internat. J. Parallel Programming 16 (1987), no. 5, 341– 364]. International Journal of Parallel Programming, 17(6):523–528, December 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=6&spage=523.

Su:2017:HLS

[SWF^{+}17] Yong Su, Zhan Wang, Zhiguo Fan, Zheng Cao, Xiaoli Liu, En Shao, Xuejun An, and Ninghui Sun. HyperFatTree: a large-scale tree-based network with low-radix switches. International Journal of Parallel Programming, 45(1):172–184, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0393-2.

Sun:2018:UPA

[SWG^{+}18] Fan Sun, Chao Wang, Lei Gong, Yiwei Zhang, Chongchong Xu, Yuntao Lu, Xi Li, and Xuehai Zhou. UniCNN: a pipelined accelerator towards uniformed computing for CNNs. International Journal of Parallel Programming, 46(4):776– 787, August 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Song:2005:PTA

[SWL05] Yonghong Song, Cheng Wang, and Zhiyuan Li. A polynomial-time algorithm for memory space reduction. International Journal of Parallel Programming, 33(1):1–33, February 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=1&spage=1.

Sha:2015:PEH

[SWZ^{+}15] Edwin Sha, Li Wang, Qingfeng Zhuge, Jun Zhang, and Jing Liu. Power efficiency for hardware/software partitioning with time and area constraints on MPSoC. International Journal of Parallel Programming, 43 (3):381–402, June 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0283-4.

Subramani:2008:DIS

[SY08] K. Subramani and Kiran Yellajyosula. On the design and implementation of a shared memory dispatcher for partially clairvoyant schedulers. International Journal of Parallel Programming, 36(4):386–411, August 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=4&spage=386.

Srinivasan:2017:SIN

[SZ17] Vijayalakshmi Srinivasan and Yunquan Zhang. Special issue on network and parallel computing. International Journal of Parallel Programming, 45(1): 1–3, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-015-0381-6.pdf.

Shi:2018:PPE

[SZH18] Yang Shi, Yanmin Zhu, and Linpeng Huang. Partial-PreSET: Enhancing lifetime of PCM-based main memory with fine-grained SET operations. International Journal of Parallel Programming, 46(4):736– 748, August 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Tyson:1999:MRF

[TA99] Gary S. Tyson and Todd M.

Austin. Memory renaming: Fast, early and accurate processing of memory communication. International Journal of Parallel Programming, 27(5):357– 380, October 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=5&spage=357.

Tipparaju:2012:RTE

[TAY^{+}12] Vinod Tipparaju, Edoardo Apra, Weikuan Yu, Xinyu Que, and Jeffrey S. Vetter. Runtime techniques to enable a highly-scalable global address space model for petascale computing. International Journal of Parallel Programming, 40(6):633–655, December 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=6&spage=633.

Timcheck:2023:INR

[TB23] Stephen Timcheck and Jeremy Buhler. Interruptible nodes: Reducing queueing costs in irregular streaming dataflow applications on wide-SIMD architectures. International Journal of Parallel Programming, 51(1): 43–60, February 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00745-2.

Tomic:2014:UDR

[TCUV14] Sasa Tomić, Adrián Cristal, Osman Unsal, and Mateo Valero. Using dynamic runtime testing for rapid development of architectural simulators. International Journal of Parallel Programming, 42(1): 119–139, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0208-7.

Trancoso:2006:CCM

[TESK06] Pedro Trancoso, Paraskevas Evripidou, Kyriakos Stavrou, and Costas Kyriacou. A case for chip multiprocessors based on the data-driven multithreading model. International Journal of Parallel Programming, 34(3):213–235, June 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
34&issue=3&spage=213.

Tyson:1994:CSM

[TF94] Gary Tyson and Matthew Farrens. Code scheduling for multiple instruction stream architectures. International Journal of Parallel Programming, 22 (3):243–272, June 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Tyson:1996:EEP

[TF96] Gary Tyson and Matthew Farrens. Evaluating the effects of predicated execution on branch prediction. International Journal of Parallel Programming, 24(2):159–186, April 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Tan:2016:ATB

[TFEK16] Antoine Tran Tan, Joel Falcou, Daniel Etiemble, and Hartmut Kaiser. Automatic task-based code generation for high performance domain specific embedded language. International Journal of Parallel Programming, 44(3): 449–465, June 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0354-9.

Tyson:1997:MDC

[TFMP97] Gary Tyson, Matthew Farrens, John Matthews, and Andrew R. Pleszkun. Managing data caches using selective cache line replacement. International Journal of Parallel Programming, 25(3):213– 242, June 1997. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Tian:2009:SPS

[TFNG09] Chen Tian, Min Feng, Vijay Nagarajan, and Rajiv Gupta. Speculative parallelization of sequential loops on multicores. International Journal of Parallel Programming, 37(5):508–535, October 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

issn=0885-7458&volume=
37&issue=5&spage=508.

Tumeo:2018:GES

[TFPF18] Antonino Tumeo, Hubertus Franke, Gianluca Palermo, and John Feo. Guest editorial: Special issue on computing frontiers. International Journal of Parallel Programming, 46(2):333– 335, April 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0556-z.pdf.

Tseng:2005:AAP

[TG05] Eric Hung-Yu Tseng and Jean-Luc Gaudiot. Automatic array partitioning based on the Smith Normal Form. International Journal of Parallel Programming, 33(1): 35–56, February 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=1&spage=35.

Tan:2021:GES

[TG21] Guangming Tan and Guang R.

Gao. Guest editorial: Special issue on network and parallel computing for emerging architectures and applications. International Journal of Parallel Programming, 49(5): 625–627, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00720-3.

Thompson:2018:HSC

[TGT18] Christopher Thompson, Miles Gould, and Nigel Topham. High speed cycle-approximate simulation of embedded cacheincoherent and coherent chip-multiprocessors. International Journal of Parallel Programming, 46 (6):1247–1282, December 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0566-x.pdf.

Tung:2017:TSP

[TH17] Le-Duc Tung and Zhenjiang Hu. Towards systematic parallelization of graph transformations over Pregel. International Journal of Parallel Programming, 45(2):320– 339, April 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

com/article/10.1007/
s10766-016-0418-5.

Thomsen:1987:IPE

[Tho87] Kristine Stougaard Thomsen. Inheritance on processes, exemplified on distributed termination detection. International Journal of Parallel Programming, 16(1):17–52, February 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=1&spage=17.

Tick:1990:ECL

[Tic90]  E. Tick.  Execution

characteristics of layered streams. International Journal of Parallel Programming, 19(6):425–443, December 1990. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
19&issue=6&spage=425.

Tinker:1988:PPL

[Tin88] P. A. Tinker. Performance of an OR-parallel logic programming system. International Journal of Parallel Programming, 17(1):59–92, February 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=1&spage=59.

Tsai:1999:CTS

[TJY99] Jenn-Yuan Tsai, Zhenzhen Jiang, and Pen-Chung Yew. Compiler techniques for the superthreaded architectures. International Journal of Parallel Programming, 27(1):1– 19, February 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
27&issue=1&spage=1.

Taubenfeld:1989:IFD

[TKM89] Gadi Taubenfeld, Shmuel Katz, and Shlomo Moran. Initial failures in distributed computations. International Journal of Parallel Programming, 18 (4):255–276, August 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=18&issue=4&spage=
255.

Tao:2008:PAR

[TKN^{+}08] Jie Tao, Marcel Kunze, Fabian Nowak, Rainer Buchty, and Wolfgang Karl. Performance advantage of reconfigurable cache design on multicore processor systems. International Journal of Parallel Programming, 36(3): 347–360, June 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=3&spage=347.

Talpin:2005:CBM

[TLSG05] Jean-Pierre Talpin, Paul Le Guernic, Sandeep Kumar Shukla, and Rajesh Gupta. A compositional behavioral modeling framework for embedded system design and conformance checking. International Journal of Parallel Programming, 33(6):613– 643, December 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=6&spage=613.

Torrie:1996:MRB

[TMHT96] Evan Torrie, Margaret Martonosi, Mary W. Hall, and Chau-Wen Tseng. Memory referencing behavior in compilerparallelized applications. International Journal of Parallel Programming, 24 (4):349–376, August 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Tzenetopoulos:2024:OEI

[TMXS24] Achilleas Tzenetopoulos, Dimosthenis Masouros, Sotirios Xydis, and Dimitrios Soudris. Orchestration extensions for interference- and heterogeneityaware placement for dataanalytics. International Journal of Parallel Programming, 52(4):298–323, August 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-024-00771-2.

Tornero:2011:CDR

[TOM^{+}11] R. Tornero, J. M. Orduña, A. Mejia, J. Flich, and J. Duato. A communicationdriven routing technique for application-specific NoCs. International Journal of Parallel Programming, 39(3):357–374, June 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=3&spage=357.

Touati:2005:RSI

[Tou05] Sid-Ahmed-Ali Touati.

Register saturation in instruction level parallelism. International Journal of Parallel Programming, 33(4):393–449, August 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=4&spage=393.

Tonci:2024:LSP

[TRB^{+}24] Nicolò Tonci, Sébastien Rivault, Mostafa Bamha, Sophie Robert, Sébastien Limet, and Massimo Torquati. LSH SimilarityJoin pattern in Fast-Flow. International Journal of Parallel Programming, 52(3):207–230, June 2024. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL https:

//link.springer.com/
article/10.1007/s10766-
024-00772-1.

Tekleyohannes:2021:DCH

[TRD21] Menbere Kina Tekleyohannes, Vladimir Rybalkin, and Andreas Dengel. iDocChip: A configurable hardware architecture for historical document image processing. International Journal of Parallel Programming, 49 (2):253–284, April 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00690-y.

Tierney:2009:SPC

[TRL09] Luke Tierney, A. J. Rossini, and Na Li. Snow: a parallel computing framework for the R system. International Journal of Parallel Programming, 37(1):78–90, February 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=1&spage=78.

Takahashi:2003:PEH

[TSB03] Daisuke Takahashi, Mitsuhisa Sato, and Taisuke Boku. Performance evaluation of the Hitachi SR8000 using SPEC OMP2001 benchmarks. International Journal of Parallel Programming, 31 (3):185–196, June 2003. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL /

ips/frames/Refs/referenceskapmain.
asp?J=4773&I=33&A=2&
LK=NM;   http://ipsapp007.
kluweronline.com/content/

getfile/4773/33/2/abstract.
htm;   http://ipsapp007.
kluweronline.com/content/
getfile/4773/33/2/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=31&issue=3&spage=
185.

Taylor:1986:PIF

[TSS86] S. Taylor, S. Safra, and E. Shapiro. A parallel implementation of Flat Concurrent Prolog. International Journal of Parallel Programming, 15(3):245– 275, June 1986. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=3&spage=245.

Tan:1999:PIB

[TSS99] Min Tan, Janet M.

Siegel, and Howard Jay Siegel. Parallel implementations of block-based motion vector estimation for video compression on four parallel processing systems. International Journal of Parallel Programming, 27(3):195–225, June 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        27&issue=3&spage=195.

Teodoro:2008:RTS

[TTF^{+}08] George Teodoro, Tulio Tavares, Renato Ferreira, Tahsin Kurc, Wagner Meira, Dorgival Guedes, Tony Pan, and Joel Saltz. A run-time system for efficient execution of scientific workflows on distributed environments. International Journal of Parallel Programming, 36(2):250– 266, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        36&issue=2&spage=250.

Thoman:2022:CHL

[TTF22] Peter Thoman, Florian Tischler, and Thomas Fahringer. The Celerity high-level API: C++20 for accelerator clusters. International Journal of Parallel Programming, 50(3– 4):341–359, August 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-022-00731-8.

Tonci:2023:DMF

[TTMD23] Nicolò Tonci, Massimo Torquati, Gabriele Mencagli, and Marco Danelutto. Distributed-memory Fast-Flow building blocks. International Journal of Parallel Programming, 51 (1):1–21, February 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00750-5.

Tousimojarad:2015:SLS

[TV15] Ashkan Tousimojarad and Wim Vanderbauwhede. Steal locally, share globally. International Journal of Parallel Programming, 43(5):894–917, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
015-0350-0.

Khan:2016:EOM

[uHKAMFM16a] Ayaz ul Hassan Khan, Mayez Al-Mouhamed, Allam Fatayer, and Nazeeruddin Mohammad. Erratum to: Optimizing the Matrix Multiplication Using Strassen and Winograd Algorithms with Limited Recursions on Many–Core. International Journal of Parallel Programming, 44 (4):831, August 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-015-0397-y.pdf.

See [uHKAMFM16b].

Khan:2016:OMM

[uHKAMFM16b] Ayaz ul Hassan Khan, Mayez Al-Mouhamed, Allam Fatayer, and Nazeeruddin Mohammad. Optimizing the matrix multiplication using Strassen and Winograd algorithms with limited recursions on many-core. International Journal of Parallel Programming, 44(4): 801–830, August 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0378-1. See

erratum [uHKAMFM16a].

Uchihira:2000:SBH

[UKT00] Naoshi Uchihira, Hideji Kawata, and Fumitaka Tamura. Scenario-based hypersequential programming. International Journal of Parallel Programming, 28(2):155–157, April 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        28&issue=2&spage=155.

urRehman:2014:PCS

[uRHH14] Zia ur Rehman, Omar Khadeer Hussain, and Farookh Khadeer Hussain. Parallel cloud service selection and ranking based on QoS history. International Journal of Parallel Programming, 42(5):820–852, October 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
article/10.1007/s10766-
013-0276-3.

Ugarte:2005:VES

[US05] Iñigo Ugarte and Pablo Sanchez. Verification of embedded systems based on interval analysis. International Journal of Parallel Programming, 33(6):697–720, December 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=6&spage=697.

Ullah:2020:LBS

[UWF^{+}20] Farhan Ullah, Junfeng Wang, Muhammad Farhan, Sohail Jabbar, Muhammad Kashif Naseer, and Muhammad Asif. LSA based smart assessment methodology for SDN infrastructure in IoT environment. International Journal of Parallel Programming, 48(2):162–177, April 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Vianna:2013:APM

[VCP^{+}13] Emanuel Vianna, Giovanni Comarela, Tatiana Pontes, Jussara Almeida, Virgílio Almeida, Kevin Wilkinson, Harumi Kuno, and Umeshwar Dayal. Analytical performance models for MapReduce workloads. International Journal of Parallel Programming, 41(4): 495–525, August 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0227-4.

Vassiliadis:2016:ESC

[VCP^{+}16] Vassilis Vassiliadis, Charalampos Chalios, Konstantinos Parasyris, Christos D. Antonopoulos, Spyros Lalis, Nikolaos Bellas, Hans Vandierendonck, and Dimitrios S. Nikolopoulos. Exploiting significance of computations for energyconstrained approximate computing. International Journal of Parallel Programming, 44(5): 1078–1098, October 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-016-0409-6.

vanderSpek:2008:CRT

[vdSGBW08] H. L. A. van der Spek, S. Groot, E. M. Bakker, and H. A. G. Wijshoff. A compile/run-time environment for the automatic transformation of linked list data structures. International Journal of Parallel Programming, 36(6):592–623, December 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=6&spage=592.

Veidenbaum:2001:GEI

[Vei01] Alex Veidenbaum. Guest Editor’s introduction. International Journal of Parallel Programming, 29 (5):461–462, October 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://ipsapp009.lwwonline.
com/content/getfile/
4773/23/1/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/23/1/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&

                                        volume=29&issue=5&spage=
                                        461.

Veidenbaum:2002:GEI

[Vei02] Alex Veidenbaum. Guest Editor’s introduction. International Journal of Parallel Programming, 30 (4):223–224, August 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/1/abstract.htm;
                                         http://ipsapp009.lwwonline.
                                        com/content/getfile/
                                        4773/28/1/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=30&issue=4&spage=
                                        223.

Venkatasubramanian:2012:TTT

[VFIN12] Girish Venkatasubramanian, Renato J. Figueiredo, Ramesh Illikkal, and Donald Newell. TMT: a TLB tag management framework for virtualized platforms. International Journal of Parallel Programming, 40(3):353–380, June 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        40&issue=3&spage=353.

Viitanen:2018:VLI

[VHK^{+}18] Timo Viitanen, Janne Helkala, Heikki Kultala, Pekka Jääskeläinen, Jarmo Takala, Tommi Zetterman, and Heikki Berg. Variable length instruction compression on transport triggered architectures. International Journal of Parallel Programming, 46 (6):1283–1303, December 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Valduriez:1988:PET

[VK88] Patrick Valduriez and Setrag Khoshfian. Parallel evaluation of the transitive closure of a database relation. International Journal of Parallel Programming, 17(1):19–42, February 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=1&spage=19.

Velasquez:2015:BBA

[VMS15] Ricardo A. Velásquez, Pierre Michaud, and André Seznec. BADCO: Behavioral applicationdependent superscalar core models. International Journal of Parallel Programming, 43(1):130–157, February 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0278-1.

vanNieuwpoort:2011:CRA

[vNR11] Rob V. van Nieuwpoort and John W. Romein. Correlating radio astronomy signals with many-core hardware. International Journal of Parallel Programming, 39(1):88–114, February 2011. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=1&spage=88.

Vanderbauwhede:2019:TDA

[VNU19] Wim Vanderbauwhede, Syed Waqar Nabi, and Cristian Urlea. Typedriven automated program transformations and cost modelling for optimising streaming programs on FPGAs. International Journal of Parallel Programming, 47(1):114–136, February 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0572-z.pdf.

Vandevoorde:1988:WAC

[VR88] Mark T. Vandevoorde and Eric S. Roberts. WorkCrews: an abstraction for controlling parallelism. International Journal of Parallel Programming, 17(4):347–366, August 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
17&issue=4&spage=347.

Vega-Rodriguez:2019:PPB

[VRGC19] Miguel A. Vega-Rodríguez and José M. Granado-Criado. Parallel programming in bioinformatics: Some interesting approaches. International Journal of Parallel Programming, 47(2):293–295, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-018-0605-7.pdf.

Vander-Swalmen:2009:CAM

[VSDK09] Pascal Vander-Swalmen, Gilles Dequen, and Michaël Krajecki. A collaborative approach for multithreaded SAT solving. International Journal of Parallel Programming, 37 (3):324–342, June 2009.

CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=37&issue=3&spage=
324.

Vallejo:2011:HTM

[VSH^{+}11] Enrique Vallejo, Sutirtha Sanyal, Tim Harris, Fernando Vallejo, Ramón Beivide, et al. Hybrid transactional memory with pessimistic concurrency control. International Journal of Parallel Programming, 39(3): 375–396, June 2011. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
39&issue=3&spage=375.

Velentzas:2023:GBA

[VVCA23] Polychronis Velentzas, Michael Vassilakopoulos, Antonio Corral, and Christos Antonopoulos. GPUbased algorithms for processing the k nearest– neighbor query on spatial data using partitioning and concurrent kernel execution. International Journal of Parallel Programming, 51(6): 275–308, December 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-023-00755-8.

Wainwright:1987:DPC

[Wai87] R. L. Wainwright. Deriving parallel computations from functional specifications: a seismic example on a hypercube. International Journal of Parallel Programming, 16(3): 243–260, June 1987. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=3&spage=243.

Wolfe:1987:DDA

[WB87] Michael Wolfe and Utpal Banerjee. Data dependence and its application to parallel processing. International Journal of Parallel Programming, 16(2):137–178, April 1987. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
16&issue=2&spage=137.

Wang:2016:CIK

[WCC16] Huang Wang, Xianglan Chen, and Huaping Chen. A cross-ISA kernelized high-performance parallel emulator. International Journal of Parallel Programming, 44(6):1118– 1141, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0379-0.

Wang:2017:GPH

[WdSAM^{+}17] Biao Wang, Diego F. de Souza, Mauricio Alvarez-Mesa, Chi Ching Chi, Ben Juurlink, Aleksandar Ilic, Nuno Roma, and Leonel Sousa. GPU parallelization of HEVC in-loop filters. International Journal of Parallel Programming, 45(6):1515–1535, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wrede:2018:SCG

[WE18] Fabian Wrede and Steffen Ernsting. Simultaneous CPU–GPU execution of data parallel algorithmic skeletons. International Journal of Parallel Programming, 46(1): 42–61, February 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wang:1994:DSP

[WEJS94] Jian Wang, Christine Eisenbeis, Martin Jourdan, and Bogong Su. Decomposed software pipelining: a new perspective and a new approach. International Journal of Parallel Programming, 22(3):351– 373, June 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Weis:2016:ASF

[WGF^{+}16] Sebastian Weis, Arne Garbade, Bernhard Fechner, Avi Mendelson, Roberto Giorgi, and Theo Ungerer. Architectural support for fault tolerance in a teradevice dataflow system. International Journal of Parallel Programming, 44 (2):208–232, April 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0312-y.

Woo:2004:AAJ

[WGW04] Jongwook Woo, Jean-Luc Gaudiot, and Andrew L. Wendelborn. Alias analysis in Java with reference-set representation for high-performance computing. International Journal of Parallel Programming, 32(1):39–76, February 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        32&issue=1&spage=39.

Wang:2017:IBM

[WHC^{+}17] Chonghua Wang, Zhiyu Hao, Lei Cui, Xiangyu Zhang, and Xiaochun Yun. Introspection-based memory pruning for live VM migration. International Journal of Parallel Programming, 45(6):1298– 1309, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wei:2024:ECH

[WHC^{+}24] Bing Wei, Qiang Huang, Hui Chen, Chenhao Zhang, and Limin Xiao. Erasurecoded hybrid writes based on data delta. International Journal of Parallel Programming, 52(4): 231–252, August 2024. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-024-00773-0.

Wing:1989:VAD

[Win89] Jeannette M. Wing. Verifying atomic data types. International Journal of Parallel Programming, 18 (5):315–357, October 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=18&issue=5&spage=
315.

Wrede:2020:THP

[WK20] Fabian Wrede and Herbert Kuchen. Towards highperformance code generation for multi-GPU clusters based on a domainspecific language for algorithmic skeletons. International Journal of Parallel Programming, 48(4): 713–728, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00659-x.

Wang:2016:GDS

[WL16] Ye Wang and Zhiyuan Li.

GridFOR: a domain specific language for parallel grid-based applications. International Journal of Parallel Programming, 44 (3):427–448, June 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0348-z.

Wang:2008:DIA

[WLL^{+}08] Hsiao-Hsi Wang, Kuan-Ching Li, Ssu-Hsuan Lu, Chun-Chieh Yang, and Jean-Luc Gaudiot. Design and implementation of an agent home scheme strategy for prefetch-based DSM systems. International Journal of Parallel Programming, 36(6):521– 542, December 2008. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
36&issue=6&spage=521.

Wang:2017:PSH

[WLL17] Yan Wang, Kenli Li, and Keqin Li. Partition scheduling on heterogeneous multicore processors for multi-dimensional loops applications. International Journal of Parallel Programming, 45(4): 827–852, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wu:2017:VPP

[WLW^{+}17] Song Wu, Yongchang Li, Xinhou Wang, Hai Jin, and Hanhua Chen. Vshadow: Promoting physical servers into virtualization world. International Journal of Parallel Programming, 45(1): 45–66, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

com/article/10.1007/
s10766-015-0385-2.

Wu:2015:CTC

[WLWZ15] Zhendong Wu, Kai Lu, Xiaoping Wang, and Xu Zhou. Collaborative technique for concurrency bug detection. International Journal of Parallel Programming, 43 (2):260–285, April 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0304-y.

Wolf:1998:CLT

[WMC98] Michael E. Wolf, Dror E.

Maydan, and Ding-Kai Chen. Combining loop transformations considering caches and scheduling. International Journal of Parallel Programming, 26(4):479–503, August 1998. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=4&spage=479.

Wrede:2019:FSS

[WMK19] Fabian Wrede, Breno Menezes, and Herbert Kuchen. Fish school search with algorithmic skeletons. International Journal of Parallel Programming, 47 (2):234–252, April 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Weis:2017:DHL

[WMN^{+}17] Christian Weis, Abdul Mutaal, Omar Naji, Matthias Jung, Andreas Hansson, and Norbert Wehn. DRAMSpec: a high-level DRAM timing, power and area exploration tool. International Journal of Parallel Programming, 45 (6):1566–1591, December 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wang:2016:PSI

[WNMW16] Chao Wang, Nadia Nedjah, Luiza M. Mourelle, and Aili Wang. Preface to the special issue on sequential code parallelization. International Journal of Parallel Programming, 44(6):1099– 1101, December 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //link.springer.com/
                                        content/pdf/10.1007/
                                        s10766-016-0447-0.pdf.

Wolfe:1986:LSW

[Wol86] Michael Wolfe. Loops skewing: The wavefront method revisited. International Journal of Parallel Programming, 15(4):279– 293, August 1986. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
15&issue=4&spage=279.

Wonnacott:2002:ASL

[Won02] David Wonnacott. Achieving scalable locality with time skewing. International Journal of Parallel Programming, 30(3): 181–221, June 2002. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//ipsapp009.lwwonline.
com/content/getfile/
4773/27/2/abstract.htm;
 http://ipsapp009.lwwonline.
com/content/getfile/
4773/27/2/fulltext.pdf;
 http://www.springerlink.
com/openurl.asp?genre=
article&issn=0885-7458&
volume=30&issue=3&spage=
181.

Wu:2000:CPG

[WP00] Peng Wu and David Padua. Containers on the parallelization of generalpurpose Java programs. International Journal of Parallel Programming, 28 (6):589–605, December 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

URL http://www.springerlink.

                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=28&issue=6&spage=
                                        589.

Weng:2007:OIS

[WPC07] Tien-Hsiung Weng, Ruey-Kuen Perng, and Barbara Chapman. OpenMP implementation of SPICE3 circuit simulator. International Journal of Parallel Programming, 35(5):493– 505, October 2007. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

                                        /www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        35&issue=5&spage=493.

Wang:2017:CIS

[WQJY17] Sheng Wang, Weizhong Qiang, Hai Jin, and Jinfeng Yuan. CovertInspector: Identification of shared memory covert timing channel in multitenanted cloud. International Journal of Parallel Programming, 45(1): 142–156, February 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0391-4.

Wen:2024:AMD

[WQZ^{+}24] Yingpeng Wen, Zhilin Qiu, Dongyu Zhang, Dan Huang, Nong Xiao, and Liang Lin. Accelerating massively distributed deep learning through efficient pseudo-synchronous update method. International Journal of Parallel Programming, 52(3): 125–146, June 2024. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-023-00759-4.

Wehr:2018:PTC

[WR18] David Wehr and Rafael Radkowski. Parallel kdtree construction on the GPU with an adaptive split and sort strategy. International Journal of Parallel Programming, 46 (6):1139–1156, December 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Warg:2008:DTS

[WS08] Fredrik Warg and Per Stenstrom. Dual-thread speculation: a simple approach to uncover threadlevel parallelism on a simultaneous multithreaded processor. International Journal of Parallel Programming, 36(2):166–183, April 2008. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        36&issue=2&spage=166.

Waliullah:2014:RCH

[WS14] M. M. Waliullah and Per Stenstrom. Removal of conflicts in hardware transactional memory systems. International Journal of Parallel Programming, 42(1): 198–218, February 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0210-0.

Williamson:2015:PIN

[WS15] Matthew Williamson and K. Subramani. A parallel implementation for the negative cost girth problem. International Journal of Parallel Programming, 43(2):240–259, April 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-013-0300-7.

Wang:2020:LMA

[WSC20] Wei Wang, Huansheng Song, and Hua Cui. Landslide multi-attitude data measurement of bedding rock slope model. International Journal of Parallel Programming, 48(5): 928–939, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00638-x.

Williams:2007:SCK

[WSO^{+}07] Samuel Williams, John Shalf, Leonid Oliker, Shoaib Kamil, Parry Husbands, and Katherine Yelick. Scientific computing kernels on the cell processor. International Journal of Parallel Programming, 35(3):263– 298, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=3&spage=263.

Wang:2018:HPM

[WSS18] Ke Wang, Elaheh Sadredini, and Kevin Skadron. Hierarchical pattern mining with the automata processor. International Journal of Parallel Programming, 46(2):376–411, April 2018. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wang:2023:SPF

[WTL^{+}23] Haoran Wang, Thibaut Tachon, Chong Li, Sophie Robert, and Sébastien Limet. SMSG: Profilingfree parallelism modeling for distributed training of DNN. International Journal of Parallel Programming, 51(2– 3):109–127, June 2023. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00741-6.

Wang:2021:TAF

[WTQ21] Bo Wang, Jie Tang, and Deyu Qi. A task-aware fine-grained storage selection mechanism for inmemory big data computing frameworks. International Journal of Parallel Programming, 49(1): 25–50, February 2021. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-020-00662-2.

Wang:2019:DAS

[WTZ^{+}19] Bo Wang, Jie Tang, Rui Zhang, Wei Ding, and Deyu Qi. A dependencyaware storage schema selection mechanism for inmemory big data computing frameworks. International Journal of Parallel Programming, 47(3):502– 519, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wu:2017:GLS

[WW17] Hsiang-Huang Wu and Chien-Min Wang. Generalization of large-scale data processing in one MapReduce job for coarsegrained parallelism. International Journal of Parallel Programming, 45(4): 797–826, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wang:2019:NAB

[WWG^{+}19] Xingwang Wang, Xiaohui Wei, Shang Gao, Yuanyuan Liu, and Zongpeng Li. A novel auction-based query pricing schema. International Journal of Parallel Programming, 47(4):759–780, August 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wolfson:1992:PPG

[WZB^{+}92] Ouri Wolfson, Weining Zhang, Harish Butani, Akira Kawaguchi, and Mok Kui. Parallel processing of graph reachability in databases. International Journal of Parallel Programming, 21(4):269– 302, August 1992. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=
                                        21&issue=4&spage=269.

Wang:2017:PCN

[WZG^{+}17] Qicong Wang, Jinhao Zhao, Dingxi Gong, Yehu Shen, Maozhen Li, and Yunqi Lei. Parallelizing convolutional neural networks for action event recognition in surveillance videos. International Journal of Parallel Programming, 45(4):734–759, August 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Wang:2013:BQE

[WZTH13] Junchang Wang, Kai Zhang, Xinan Tang, and Bei Hua. B-queue: Efficient and practical queuing for fast core-to-core communication. International Journal of Parallel Programming, 41(1): 137–159, February 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-012-0213-x.

Xue:1998:RDT

[XH98] Jingling Xue and Chua-Huang Huang. Reusedriven tiling for improving data locality. International Journal of Parallel Programming, 26(6):671– 696, December 1998. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:/

/www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
26&issue=6&spage=671.

Xu:2019:MCE

[XLWX19] Heyang Xu, Yang Liu, Wei Wei, and Ying Xue. Migration cost and energyaware virtual machine consolidation under cloud environments considering remaining runtime. International Journal of Parallel Programming, 47(3):481– 501, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Xavier:2009:MLP

[XOdFV^{+}09] Carolina Ribeiro Xavier, Rafael Sachetto Oliveira, Vinicius da Fonseca Vieira, Rodrigo Weber dos Santos, and Wagner Meira. Multilevel parallelism for the cardiac bidomain equations. International Journal of Parallel Programming, 37(6):572–592, December 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
37&issue=6&spage=572.

Xing:2021:AAA

[XWH21] Biao Xing, DanDan Wang, and Cuihua He. Accelerating DES and AES algorithms for a heterogeneous many-core processor. International Journal of Parallel Programming, 49(3):463–486, June 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00692-4.

Xiao:2020:FDO

[XZT20] Junmin Xiao, Guizhao Zhang, and Guangming Tan. Fast data-obtaining algorithm for data assimilation with large data set. International Journal of Parallel Programming, 48 (4):750–770, August 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-019-00653-y.

Xu:2015:YSE

[XZX^{+}15] Quanqing Xu, Liang Zhao, Mingzhong Xiao, Anna Liu, and Yafei Dai. YuruBackup: a spaceefficient and highly scalable incremental backup system in the cloud. International Journal of Parallel Programming, 43(3): 316–338, June 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0280-7.

Yang:1995:MDD

[YAI95] Yi-Qing Yang, Corinne Ancourt, and François Irigoin. Minimal data dependence abstractions for loop transformations: Extended version. International Journal of Parallel Programming, 23(4):359– 388, August 1995. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Ye:2017:REC

[YBDJ17] Chencheng Ye, Jacob Brock, Chen Ding, and Hai Jin. Rochester Elastic Cache Utility (RECU): Unequal cache sharing is good economics. International Journal of Parallel Programming, 45(1): 30–44, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0384-3.

Yzelman:2014:MCH

[YBRM14] A. N. Yzelman, R. H.

Bisseling, D. Roose, and K. Meerbergen. MulticoreBSP for C: a highperformance library for shared-memory parallel programming. International Journal of Parallel Programming, 42(4): 619–642, August 2014. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0262-9.

Yao:2019:CTC

[YDV19] Fan Yao, Miloš Doroslovački, and Guru Venkataramani. Covert timing channels exploiting cache coherence hardware: Characterization and defense. International Journal of Parallel Programming, 47(4):595– 620, August 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yu:2021:LBT

[YFC21] Mengshan Yu, Guisheng Fan, and Liang Chen. Location-based and timeaware service recommendation in mobile edge computing. International Journal of Parallel Programming, 49(5): 715–731, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00702-5.

Yang:2018:SPC

[YH18] Jixiang Yang and Qingbi He. Scheduling parallel computations by work stealing: a survey. International Journal of Parallel Programming, 46(2): 173–197, April 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yu:2016:CBL

[YHGW16] Yulong Yu, Xubin He, He Guo, and Yuxin Wang. A credit-based load-balance-aware CTA scheduling optimization scheme in GPGPU. International Journal of Parallel Programming, 44(1): 109–129, February 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0318-5.

Yang:2016:EBM

[YJY16] Chengcheng Yang, Peiquan Jin, and Lihua Yue. Efficient buffer management for tree indexes on solid state drives. International Journal of Parallel Programming, 44(1):5– 25, February 2016. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

                                        com/article/10.1007/
                                        s10766-014-0340-7.

YarKhan:2017:PPN

[YKLD17] Asim YarKhan, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Porting the PLASMA numerical library to the OpenMP standard. International Journal of Parallel Programming, 45(3):612–633, June 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yun:2003:TOS

[YKM03] Han-Saem Yun, Jihong Kim, and Soo-Mook Moon. Time optimal software pipelining of loops with control flows. International Journal of Parallel Programming, 31(5):339– 391, October 2003. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        ipsapp007.kluweronline.
                                        com/content/getfile/
                                        4773/36/1/abstract.htm;
                                        http://ipsapp007.kluweronline.
                                        com/content/getfile/
                                        4773/36/1/fulltext.pdf;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=
                                        article&issn=0885-7458&
                                        volume=31&issue=5&spage=
                                        339.

Yu:2019:AEH

[YLB19] Misun Yu, Joon-Sang Lee, and Doo-Hwan Bae. AdaptiveLock: Efficient hybrid data race detection based on real-world locking patterns. International Journal of Parallel Programming, 47(5–6): 805–837, December 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-018-0579-5.

Yao:2017:ONC

[YMW^{+}17] Jiansheng Yao, Chunguang Ma, Peng Wu, Gang Du, and Qi Yuan. An opportunistic network coding routing for opportunistic networks. International Journal of Parallel Programming, 45(1): 157–171, February 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0392-3.

Yin:2022:CSB

[YS22] Fei Yin and Feng Shi. A comparative survey of big data computing and HPC: From a parallel programming model to a cluster architecture. International Journal of Parallel Programming, 50(1): 27–64, February 2022. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/
                                        s10766-021-00717-y.

Yao:2019:IDF

[YWW^{+}19] Haipeng Yao, Qiyi Wang, Luyao Wang, Peiying Zhang, Maozhen Li, and Yunjie Liu. An intrusion detection framework based on hybrid multilevel data mining. International Journal of Parallel Programming, 47(4):740– 758, August 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yang:2018:IIS

[YWY^{+}18] Xiaomin Yang, Wei Wu, Binyu Yan, Huiqian Wang, Kai Zhou, and Kai Liu. Infrared image superresolution with parallel random forest. International Journal of Parallel Programming, 46(5):838– 858, October 2018. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yang:2020:EMD

[YYYX20] Wensi Yang, Qingfeng Yao, Kejiang Ye, and Cheng-Zhong Xu. Empirical mode decomposition and temporal convolutional networks for remaining useful life estimation. International Journal of Parallel Programming, 48 (1):61–79, February 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yang:2013:IHP

[YZ13] Yi Yang and Huiyang Zhou. The implementation of a high performance GPGPU compiler. International Journal of Parallel Programming, 41(6): 768–781, December 2013. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-012-0228-3.

Yu:2019:BLC

[YZZ^{+}19] Yong Yu, Tian Zhi, Xuda Zhou, Shaoli Liu, Yunji Chen, and Shuyao Cheng. BSHIFT: a low cost deep neural networks accelerator. International Journal of Parallel Programming, 47(3):360–372, June 2019. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Yu:2020:CCB

[YZZ20] Zhen Yu, Yu Zuo, and Yong Zhao. Convoider: a concurrency bug avoider based on transparent software transactional memory. International Journal of Parallel Programming, 48(1):32–60, February 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Zhao:2009:LTL

[ZC09] Yang Zhao and Krishnendu Chakrabarty. Online testing of lab-on-chip using reconfigurable digital-microfluidic compactors. International Journal of Parallel Programming, 37(4):370–388, August 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=4&spage=370.

Zhang:2017:DDM

[ZC17] Yu Zhang and Huifang Cao. DMR: a deterministic MapReduce for multicore systems. International Journal of Parallel Programming, 45(1): 128–141, February 2017. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

                                        http://link.springer.
                                        com/article/10.1007/
                                        s10766-015-0390-5.

Zins:2019:TPM

[ZD19] Pierre Zins and Michel Dagenais. Tracing and profiling machine learning dataflow applications on GPU. International Journal of Parallel Programming, 47(5–6):973– 1013, December 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-019-00630-5.

Zeyao:2005:CAP

[Zey05] Mo Zeyao. Concatenation algorithms for parallel numerical simulation of radiation hydrodynamics coupled with neutron transport. International Journal of Parallel Programming, 33(1):57–71, February 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=1&spage=57.

Zhang:2015:HTP

[ZGH^{+}15] Jianxun Zhang, Zhimin Gu, Yan Huang, Ninghan Zheng, and Xiaohan Hu. Helper thread prefetching control framework on chip multi-processor. International Journal of Parallel Programming, 43(2):180– 202, April 2015. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0299-9.

Zhang:1989:PAM

[Zha89] Yi Xin Zhang. Parallel algorithms for minimal spanning trees of directed graphs. International Journal of Parallel Programming, 18(3):205– 221, June 1989. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        18&issue=3&spage=205.

Zhang:2010:COP

[Zha10]  Nan Zhang.  Comput-

ing optimised parallel speeded-up robust features (P-SURF) on multicore processors. International Journal of Parallel Programming, 38(2):138– 158, April 2010. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

                                        www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        38&issue=2&spage=138.

Zhu:2019:WAI

[ZHF^{+}19] Huanzhou Zhu, Ligang He, Songling Fu, Rui Li, Xie Han, Zhangjie Fu, Yongjian Hu, and Chang-Tsun Li. Wolf-Path: Accelerating iterative traversing-based graph processing algorithms on GPU. International Journal of Parallel Programming, 47(4):644– 667, August 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-017-0533-y.pdf.

Zheng:2017:LLB

[ZJG17] Xinnian Zheng, Lizy K.

John, and Andreas Gerstlauer. LACross: Learningbased analytical crossplatform performance and power prediction. International Journal of Parallel Programming, 45(6):1488– 1514, December 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Zurstrassen:2022:ADG

[ZJL22] Niko Zurstraßen, Lukas Jünger, and Rainer Leupers. AMAIX in-depth: a generic analytical model for deep learning accelerators. International Journal of Parallel Programming, 50(2):295–318, April 2022. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-022-00728-3.

Zhang:2007:RCM

[ZK07] Chunhui Zhang and Fadi Kurdahi. Reducing offchip memory access via stream-conscious tiling on multimedia applications. International Journal of Parallel Programming, 35(1):63–98, February 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
35&issue=1&spage=63.

Zhang:2021:RBA

[ZLA21] Ziyu Zhang, Zitan Liu, and Hong An. RDMAbased Apache Storm for high-performance stream data processing. International Journal of Parallel Programming, 49(5): 671–684, October 2021. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.
com/article/10.1007/
s10766-021-00696-0.

Zalamea:2004:SHT

[ZLAV04] Javier Zalamea, Josep Llosa, Eduard Ayguadé, and Mateo Valero. Software and hardware techniques to optimize register file utilization in VLIW architectures. International Journal of Parallel Programming, 32(6):447–474, December 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
32&issue=6&spage=447.

Zhao:2019:EAS

[ZLC^{+}19] Peng Zhao, Lei Liu, Wei Cao, Xiao Dong, Jiansong Li, and Xiaobing Feng. ElasticActor: an actor system with automatic granularity adjustment. International Journal of Parallel Programming, 47(3): 520–534, June 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Zhang:2015:QBA

[ZLD15] Deli Zhang, Brendan Lynch, and Damian Dechev. Queue-based and adaptive lock algorithms for scalable resource allocation on shared–memory multiprocessors. International Journal of Parallel Programming, 43(5): 721–751, October 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-014-0317-6.

Zhang:2017:OPL

[ZLJ^{+}17] Jinbao Zhang, Xiaofei Liao, Hai Jin, Dong Liu, Li Lin, and Kao Zhao. An optimal page-level power management strategy in PCM–DRAM hybrid memory. International Journal of Parallel Programming, 45(1):4– 16, February 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-015-0382-5.

Zhang:2012:DDA

[ZLJA12] Yun Zhang, Jae W. Lee, Nick P. Johnson, and David I. August. DAFT: Decoupled Acyclic Fault Tolerance. International Journal of Parallel Programming, 40(1):118–140, February 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=1&spage=118.

Zhang:2020:VBM

[ZQT20] Han Zhang, Yurong Qian, and Chenwei Tian. A ViBe based moving targets edge detection algorithm and its parallel implementation. International Journal of Parallel Programming, 48(5): 890–908, October 2020. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

https://link.springer.

com/article/10.1007/
s10766-019-00628-z.

Zyulkyarov:2012:POT

[ZSH^{+}12] Ferad Zyulkyarov, Srdjan Stipic, Tim Harris, Osman S. Unsal, Adrián Cristal, Ibrahim Hur, and Mateo Valero. Profiling and optimizing transactional memory applications. International Journal of Parallel Programming, 40(1):25–56, February 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
40&issue=1&spage=25.

Zou:2019:AAS

[ZTY^{+}19] Huihui Zou, Shanjiang Tang, Ce Yu, Hao Fu, Yusen Li, and Wenjie Tang. ASW: Accelerating Smith–Waterman algorithm on coupled CPU– GPU architecture. International Journal of Parallel Programming, 47(3): 388–402, June 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Zheng:2005:SBP

[ZWJK05] Gengbin Zheng, Terry Wilmarth, Praveen Jagadishprasad, and Laxmikant V. Kalé. Simulation-based performance prediction for large parallel machines. International Journal of Parallel Programming, 33(2–3):183– 207, June 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=
33&issue=2&spage=183.

Zhong:2015:VBM

[ZXY^{+}15] Xianming Zhong, Chengcheng Xiang, Miao Yu, Zhengwei Qi, and Haibing Guan. A virtualization based monitoring system for miniintrusive live forensics. International Journal of Parallel Programming, 43 (3):455–471, June 2015. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.
com/article/10.1007/
s10766-013-0285-2.

Zhang:2013:KPM

[ZYOY13] Shixun Zhang, Shinichi Yamagiwa, Masahiko Okumura, and Seiji Yunoki. Kernel polynomial method on GPU. International Journal of Parallel Programming, 41(1):59–88, February 2013. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL

http://link.springer.

com/article/10.1007/
s10766-012-0204-y.

Zhang:2019:GES

[ZZS^{+}19] Feng Zhang, Jidong Zhai, Marc Snir, Hai Jin, Hironori Kasahara, and Mateo Valero. Guest editorial: Special issue on network and parallel computing for emerging architectures and applications. International Journal of Parallel Programming, 47(3): 343–344, June 2019. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//link.springer.com/
content/pdf/10.1007/
s10766-019-00634-1.pdf.